A Selected Bibliography of Publications by, and about, Hans Bethe

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

18 October 2022 Version 1.144

Title word cross-reference

+ [AS70]. 0 [AS70]. $1 [Duf46]. 1 [AS70]. 1/2 [Gün64, PcPZ90]. $12.95 [Edg91]. 15M_{sun} [Bar88]. $19.50 [Gol68]. $24.95 [Fra01]. $32.50 [Edg91]. $35.00 [Hod13, Hub14, Wal15a, Wal15b]. $39.50 [Edg91]. $5.00 [Sch68]. $50.00 [Wal15a, Wal15b]. $75.00 [Wal15a, Wal15b]. $8.95 [Edg91]. $87.00 [Wal15a, Wal15b]. $9.50 [Bet54b]. ^{12} [HB40]. ^{15} [HB40]. ^{27} [BH39]. ^{7} [Bet38c]. ^{∘} [Bet42a]. _{1−x} [BB76]. _{2} [BS37, SBR^{+}86]. _{3} [BB76, BS37]. _{4} [BS37]. _{x} [BB76]. A_{n−1} [HSY04]. α [Bet38c, HB40]. αβγ [AWCT09, Tur08]. δ [YLGC02]. Δ = −1/2 [dGBNM02]. E_{τ,η}(A_{2}^{(2)}) [MN07]. f [FBE^{+}34]. $h [Bet81e]. J [AAHS99]. J = 0 [FT75]. K^{−} [Bet70b]. λ_{l} [CWL^{+}04]. osp(1|2) [KM01]. N [CWL^{+}04, WG68, BA52]. O(4) [AS70]. O(5) [BB70, BB71]. p [BA52]. P_{μ} = 0 [Flo76]. π [Bet55d, CB60]. π^{+} [BA52]. r [App88, CFR82, Skr07, Skr09]. s [Ong86, Ong88]. SL(N) [Sco94]. SO(5) [Flo76]. SU(4) [YLGC02]. SU(n) [LYS01]. t [AAHS99]. × [Sch67b, Sch67a]. XXZ [LFK99]. Z [HB51a, Por03].

-decay [FBE^{+}34]. -Dependence [HB51a]. -field [CFR82]. -function [YLGC02]. -matrices [Skr07, Skr09]. -meson [Bet55d]. -Nucleon [CB60]. -Particle [Bet38c]. -process [App88]. -wave [Ong86, Ong88].

0 [Bet91b]. 0-465-03678-3 [Bet91b]. 0-8047-1713-3 [Edg91]. 0-8047-1714-1 [Edg91]. 0-8047-1721-4 [Edg91]. 0-8047-1722-2 [Edg91]. 0-85964-226-7 [Hen94]. ’09 [ACM09].

1 [Hod13, Hub14]. 1-55655-217-3 [Hen94]. 100th [KRW05]. 11 [HD72]. 11th [AP77]. 12th [AP77]. 1655-40 [BLLB00a, BLLB00b]. 1930/41 [Fer68, Fer71]. 1930s [Stu79]. 1933 [CCJ^{+}34]. 1937 [Ano37, For87]. 1950 [Ano50]. 1955 [Bet55c]. 1956 [Bet56b, Bet56c, Bet94a]. 1959 [Wal15a, Wal15b]. 1960s [Mla98]. 1966 [Bec67, MB66, Gol68]. 1967 [Bet67d, Bet67b, Bet97b, RB67a]. 1969 [WH72]. 1972 [CC73]. 1980s [BBG85b]. 1981 [RS82]. 1983 [BS83]. 1985 [BH88, CJK92]. 1986 [KS87]. 1987 [Blu87, DS88]. 1987A [BP90, Bet93h]. 1993 [Bet93e, Bet94b]. 1999 [Bet99a, GSS^{+}03].

2005 [BL09, Dre07, Gib05, LB07]. 2009 [ACM09]. 20th [APB96, ABP98]. 22 [CCJ^{+}34]. 25.95 [Nav14]. 2nd [Bet54b].

3 [Bet91b].

4 [Bet70c]. 40th [MKR87].

5-Year [BG82b]. 51 [RB37a].

6 [Tho99]. 6.7-Minute [BH39]. 60th [Gol68, MB66].

70th [HBD71]. 76 [FBB14].

80 [Sch67b, Sch67a]. 80th [DS88].

90th [Hof96]. 978 [Hod13, Hub14]. 978-0-226-61063-4 [Wal15a, Wal15b]. 978-0-230-35487-6 [Wal15a, Wal15b]. 978-0-674-06587-1 [Nav14, Wal15a, Wal15b, Hod13]. 978-0674-06587-1 [Hub14]. 978-981-4307-81-9 [Wal15a, Wal15b]. 981 [Tho99]. 981-02-2692-6 [Tho99]. 981-02-2876-7 [Tho99].

= [Uns00, Wei95].

A. [Sch67a]. AAPT [BSF93]. Abbildungen [HH14, Sch67b, Sch67a]. Aberdeen [Bet98a]. ablation [BA59]. ABM [Bet69a, Bet84b, Bet84a, CWB^{+}69]. Abnormal [BB71, BB70]. Absence [RB39]. absolute [CSB^{+}31]. Absorption [BS37, KMB57, LB62, WLB36, BBD^{+}95]. Absorptive [NR66]. Academy [WH72]. Accelerator [JLBF66, SSB^{+}98]. Accelerators [Bet97a]. acceptance [Bet93e]. Accident [SB86, SBR^{+}86]. Accidental [BB69]. Accretion [BLB00]. accuse [AA54]. ACM [ACM09, Blu87]. Acoustic [Gur01]. acoustics [Gur01]. Addenda [IIT78]. Address [Bet77b]. Administration [ABB^{+}80]. Administrators [STS^{+}49]. Advanced [RS82]. Advection [BLB00]. Advisers [BB86]. Advisors [Edg91, Yor89]. Advisory [GBG^{+}86]. AEC [Bet63b]. AEI [Bet75a]. Affairs [BBG^{+}85a, GR63]. affine [AAHS99]. Affinity [Bet29a]. Affirm [ACU^{+}54, BC00]. Again [Rai85]. Against [Bet84b]. Age [GR63, Stu18, Ano05b, Bra05, Bet62a, Dow83]. Ages [Jea24]. ago [Geh78]. Air [Bet42a, Bet46b, Bet50a, Bet65c, KMB57, Bet20]. Alamogordo [LW71]. Alamos [MKR87, Bet91j, Bet93k, Bet93i, Inc92, vB92]. Alan [Eks94]. Albrecht [BL09, Gib05, Hod81, LB07]. algebra [FV86, LFK99]. Algebraic [AMS11, Kyr72, LC71, MN07, Jur89, Liu72, RM06, Skr07, Skr09, ZLGM03]. Algorithm [Yui01, WW88]. Algorithms [PR97]. Alliance [She75]. Alloys [Bet38f]. Alpha [Bet50a, Bet49c, Rig07, Rig06]. alpha-Centaury [Rig07, Rig06]. Alpha-Particles [Bet50a, Bet49c]. Alpher [AH73, Mar08, Rig06, Rig07]. Alternative [BAA^{+}75, Gir75]. Always [Bet67e]. Amateur [Har01]. amazing [Tur08]. America [Hen94]. American [Bet55c, Bet75e, Bet93e, WH72, BBM78, Bet63a, Bet82e, Bet99a, Bof76, CHW91, EBU^{+}52, LBC^{+}75, Ros93, McC20]. Americans [ABB^{+}80]. Among [Bet95b, Van32, Pan05]. Amplitude [ROB63]. Amplitudes [SL64]. Amsterdam [Bet56b, Bet56c]. Analog [BS68]. Analysis [BdH54a, BdH54b, dHMAB54, Bet55b, Fle73]. Analytic [Bet93h]. analytical [BN68a]. Analyzed [Gou47]. Andrew [Dow83, Bet98b]. Angle [BR52]. Angular [BA52, Gün64]. Anisotropy [Van32]. Anja [Wal15a, Wal15b]. Annihilation [Bet35e, BH56]. Anniversary [MKR87]. Annotation [FBB14]. Announced [Bet63d]. Annual [Bet63b]. Anomalous [Ben38, BBD^{+}95]. ansatz [AFF01, AMS11, BK93, DL83, FCC09, Gar02, HSY04, Jur89, KW03, KZJ02, LYS01, LFK99, MIT06, MN07, Ray97, Sco94, Skr07, Skr09, YLGC02, ZLGM03, AAHS99, CH02, KMGT97, KHM98, RS05, Sut03]. Anti [BH56, Bet69b, GB68, WA94]. Anti-Ballistic-Missile [GB68]. Anti-Missiles [Bet69b]. Anti-Proton [BH56]. anti-quark [WA94]. antiballistic [CWB^{+}69]. antifermion [Kea71, Zum75]. Any [GB89, RR07, Roj09]. apart [Kai05b]. appeal [Ano82]. Appendix [BBAL80, BG85]. Application [vdLB47]. applications [Fur85]. Applied [MB40]. Appraises [Bet60b]. appreciation [Har01, Hol03]. Approach [BFK49, Bet93h, Nak63b, CH02, FCC09, Jur89]. Approximate [Vos60]. Approximation [CCG^{+}11, DRS^{+}54, SG73, WF11, BW87, Frü03, HM83]. April [Bet55c, Bet93e, Bet97b, HPA97]. Arbeiten [Hou30]. arbitrary [Bet43b, Bet98a, Lai74]. Argentine [ABB^{+}77]. Armor [BW39]. Arms [Bet82a, Bet8xa, BKD82, Bet83b, Bet69a]. Army [Bet98a, BS47, Gou47]. Arnold [Ano38, Sch67b, PS69, Sch67b]. arsenals [Bet89c]. art [KMN^{+}96]. Arthur [Duf46]. Article [Bof76, McC20]. Articles [For87, GR63, BBL86]. Arts [WH72]. Ascent [Bro73a]. assesses [Bet86c]. Assessing [BG82a]. Assignment [BH39]. Assistant [Kil77]. associated [HSY04]. Association [Bet93e, Ros93]. assumptions [Dud76]. astounding [Haw11]. Astronomers [Huf81, WGO^{+}91]. Astronomical [Kra17]. astronomy [FF91]. Astrophysical [Bro05b, Bro06a]. Astrophysics [Bet97a, Bet06, BCSF82, Bet03, MR85, ST86]. Asymptotic [WB51]. Atmosphere [KMT46]. atmospheres [BG91a]. Atmospheric [Bre53]. Atom [Bet70c, Duf46, LR47, Sul75, Bas71, BM66, CHW91]. Atomic [Ano95, Bet51, Bet54d, Bet62a, Bet77e, CP50, Duf46, FR13, Gas93, GR63, Pal00, PS69, STS^{+}49, WM15, Bet31b, Bet49c, Bet91j, BGP^{+}93, Bet94a, CCJ^{+}34, Hou30, Jun58, Jun82, LW71, MW46, MW07, Opp46, She75, Van03, Edg91]. atomiques [CCJ^{+}34]. Atomkette [Bet31b]. Atoms [Bet35a, Bet66b, Cho01, FSB71, Ino71, BS57a, BS57b, BS77, BS08, IIT78, Sch78, Bet01a]. Atomskerns [Hou30]. attached [Roj09, RM09]. attack [Bie03]. Attempt [Bet36a]. attempts [Nav06]. Attention [Ken54]. Aufbau [Bet33a]. August [BH88, CJK92]. auspices [CCJ^{+}34]. autobiography [Dys18]. autour [SB55b]. Award [Ano61, Ano12b, Ano85b, Ano12a, Bet63b, Bet63d, WB64]. awarded [Ano92a]. awards [BSF93]. away [Ano05b].

B [Bet98a, Ano93, Bet37a, CP93]. Ba [BB76]. Bacher [For87]. Back [BB86, Bet01b, Sal02]. background [WA94]. bad [Mad83]. Bakerian [Bet94b]. ball [MIT06]. Ballistic [BGGK84, Bet98a, BGGK12, GB68]. Ban [GSB^{+}83, McK86]. Bardeen [Skr09]. Barkas [Por04a]. Barkas-Effect [Por04a]. Barriers [BFK49]. Baryon [BJ74a]. Based [BGGK84, BGGK12]. Basic [BBL86, Bet90b, For87]. Basis [BMG87, CTSOS99, KB38]. BCS [AFF01]. be [Bet46b, Bet20, MB85]. Beam [CB48b, HM83]. Beams [SSB^{+}97d]. Bedrohung [SB55a]. before [BBM78]. Befuddles [Bet99b]. Behandlung [Bet30a]. Behavior [BK39, BK62]. Behind [Bet93a, Bet94a]. Being [Bet89b]. Beiträge [KRW05]. Bela [Bet93a]. Bell [Whi95]. Bemerkung [BBR31]. Bemerkungen [Bet48]. Berechnung [Bet34b, Bet29a]. Berichte [KRW05]. Berlin [Sch67b, Sch67a]. Berlin-Heidelberg-New [Sch67b, Sch67a]. Bernstein [Huf82]. Bet [Bet77d]. Beta [BHP39, BHP97]. Beta-Disintegration [BHP39, BHP97]. Betatron [CB48b]. Bethe [Aft02, Ano50, Ano61, Ano12a, Ano12b, BB71, BL09, Dow82, Dow83, Dre07, Dys55, Dys56, Dys05a, Edg91, For57, For87, Fra01, GvH05, Gib05, Gol68, Got98, GS05, Har07, Hof96, HD72, Huf82, Jeh56, Kut05b, Lan00, LB07, Mar08, Mer68, Olw01, Pei57, Pri56, RS08, Sar01, Sch14a, Sch78, Sch67b, Sch67a, Sch15, Sei01, Tho99, Wie14, vB92, Ada06, All59, AH73, AAHS99, AFF01, Bla82, Ano85b, Ano92a, Ano00a, Ano01, Ano05b, Ano05c, Ano05a, Ano12c, Anoxx, AMS11, Aui02, BB69, BB70, Bas71, Ber96a, Ber80, Ber81, Ber06, Bet65a, Bet80a, BWMH81, Bet84b, BBL86, Bet93j, BGP^{+}93, Bet93c, Bet93e, Bet97e, Bet99d, Bha05, Bis67, BG91b, Bof76, BK93, BSS05, Bra05, Bre75, Bri95b, Bro82, BT69, Bro73b, Bro71, Bro05b, Bro05a, Bro06a]. Bethe [BL06a, Bro06b, BL06b, BJMK92, BW87, Bur68, CTSOS99, Cab78, CF12, CWL^{+}04, CCG^{+}11, CH02, Cli97, CFR82, DL83, Dre05, Dys05b, Dys06, Eks94, Eva84, Ewa58, FV86, FCC09, Fle73, FT75, Flo76, FBB14, Fri99, Frü03, Fun81, Fur85, Gar02, GvH06, Gas93, Gau83, Gau14, Gol05, Goo99, Got05, Got06a, Got06b, Gün64, Ham73, Hod81, Hof96, HM83, HB06, HHD73, HSY04, HD70, Inc92, Ino71, IIT78, Iof05, Iof06, Jac09, Jur89, Kai05a, KMGT97, KHM98, KW03, Kea66, Kea68, Kea69, Kea70, Kea71, KZJ02, KM01, Kut05b, Kut05a, Kyr72, Lee05, Lee07, Ler71, LYS01, LFK99, MIT06, Mad86, MP91, Mai03, Mai04, MN07, ML83, MB66, Mar70, McD62, MTG05, MA06, MR85, Mor02, Moy97, Nak63a, Nak63b, Nak66, Nak71, Nav14]. Bethe [Neg05, Neg06, NR66, Ong86, Ong88, Orb05, PcPZ90, Pan05, Pep05, Per84a, Per84b, Pev78, Por97, Por98, Por99, Por03, Por04a, Por04b, PR97, Ran48, Ray97, RS05, Rig06, Rig07, RJ10, RM06, RR07, Roj07, Roj08, Roj09, RM09, RJ11, Ros66, Ros93, Run67, Sae67, Sal05, Sal06, Sat63, Saz87, Sch14a, Sch34, Sch67a, Sch95, Sch96, Sch00b, Sch05a, Sch08, Sch12, Sch14b, Sco94, SG73, Skr07, Skr09, Sut03, SL64, Tel87, Tom83, Tuc05, Vos60, Wal15a, Wal15b, WF11, WA94, WG68, Whi95, Wie14, Wie81, Wil73, Wri10, YG06, YLGC02, Yui01, ZLGM03, Zum75, dGBNM02, Hod13, Hub14, Kra13, Mer68, Low13]. Bethe-ansatz [DL83]. Bethesda [ACM09, Blu87, Hen94, Sch68]. Betrachtungen [HBD71]. Betrayed [Bet99f]. Betrt [BW87]. Better [Bet89d]. Between [SBB^{+}97, Stu18, BDSB94, MIT06, dHMAB54]. Beugung [Bet28a]. Bewußtseinswandel [vW88, vW91]. Big [Hil15]. Biggest [OM18]. Binaries [BB99, BLB99, BLZB01, BHL^{+}01, BBL07]. Binary [BB98]. Binding [Bet38a, Bet91a, DSB91]. Biographies [Wal15a, Wal15b]. Biography [Bet93a, Bet98b, Sch96, Sch14b, KMN^{+}96]. Bird [Pei85]. Birth [Hil15]. Birthday [HBD71, KRW05, Wri10, BCSF82, DS88, Gol68, Hof96, MB66]. bis [HH14]. Black [Bet94d, BB99, BBL06, BB94a, BB94b, BLB99, BLW^{+}00b, BHL^{+}01, BLB01, BBL03]. Black-Holes [BB94a, BB94b]. Blackwood [Dow82, Dow83]. Blaker [Mer68]. Blast [BFvN^{+}44, BFH^{+}47, Sul75, PPvNF44]. Bloch [Por97, Por98, Por99, Por03, Por04a, Por04b]. Blocks [SSB^{+}99b]. Blunder [OM18]. BMD [BBG85b, BG85]. Body [Bet36b, Bet56a, Bet65b, Bet67a, Bet73b, HB06, Kai05a, KB38, KS87, MJBA75, RB67b, BR67, CC73, Mai03, Mai04, MHFW72, Per63, Tom83]. Bohr [Duf46, Ano95, Bet85a, BGS95b, FK85, OBxx, Sch00a]. Bomb [Aft02, All56, ABB^{+}50, Ano93, Bac50, BBSS50, B^{+}50, Bet50d, Bet50c, Bet54e, Bet63a, Bet71e, Bet82b, Bet82d, Bet82f, Bet89e, Bet91c, Bet93a, Bet93c, Bet94a, Bet98b, Duf46, Fra01, Gas93, Ken54, Lan00, Olw01, Rid50, Sar01, Sei01, STS^{+}49, TB82, WM15, Ano50, Bet50b, Bet91h, BGP^{+}93, Bro82, Eli50, MW46, MW07, McC20, Sch00b, SB54, SB55b, SB55a, SB71, She75, Van03, WG17, Bet90d, Bet93a, Duf46, Bet95a]. Bombs [Bet54d, KMT46]. Book [Bet54b, Bet79a, Bet89b, Bet90b, Bet91b, Bet93b, Bet93a, Bet94a, Bet95a, Bet95b, Bet98b, Bet01a, Bet01b, Dys55, Dys56, Edg91, Ewa58, For57, For87, Fra01, Gol68, Got98, Har07, Hen94, Hod13, Hub14, Huf82, Inc92, Jeh56, Kra13, Lan00, Low13, Mer68, Nav14, Olw01, Pei57, Pri56, RS08, Sar01, Sch78, Sch68, Sei01, Tho99, Wal15a, Wal15b, vB92, Sch67b]. Books [ABB^{+}80, Bet58a, Bet90b, Sch68]. born [ML83, CF12]. Bose [ZLGM03]. Boson [FT75]. Both [BG82b]. Bound [Bis67, SB51, Kea71, Mai03, Mai04, MB82, WA94]. Bound-State [Bis67, SB51, Mai03, Mai04, MB82]. Boundaries [Sch95, Wei95]. boundary [Bet42e, Fur85, KZJ02, LYS01, Mai04]. bounds [RJ10, RM06, RR07]. box [MIT06]. box-ball [MIT06]. boycott [BBM^{+}76]. Boyle [Hen94]. Braking [Bet14, FBB14]. Breadth [DS88]. Breakout [Bet96a]. Breeder [Ham73]. Breeders [Bet75c]. Bremsformel [Bet32]. bremsstrahlung [Frü03, BML53, BM54, Bet72a, DB52, DBM54, MB52]. Briefe [KRW05]. Brien [Ano50]. Brighter [Jun58, Jun82, Bet58a]. British [Wal15a, Wal15b, Fak83]. Broad [MWBB02]. Broadway [Sch00a]. Broglie [Whi95]. Brook [GSS^{+}03, KS87]. Brown [Bet98b, Har07, RS08, Lee11]. Brueckner [Bet56c, Bet63c, Bro71]. Brussels [CCJ^{+}34]. Bruxelles [CCJ^{+}34]. BS [AS70]. Buchbesprechung [Sch67b]. Build [Bet93a, GB89]. Building [SSB^{+}99b]. bulk [OBxx, Rabxx]. Bulletin [GR63, BFG^{+}85]. Bursts [BLW^{+}00a, MWBB02, BLW^{+}03, Eps88]. Bush [Ano85b].

C [Bet42a, Hol03, HB40, LB37]. C. [Geh73]. C.B.E. [AP77]. ca [Ano00a, Kra16]. Calculate [Bet36a]. Calculated [Bet29d]. Calculation [Bet34b, CTSOS99, NB68, Cab78, KW03, Liu72, Bet29a]. Calculations [BJ74a, Bar88]. California [Bet76g]. Called [Cor12]. Calls [Ham73]. Cambridge [Hen94, Hod13, Hub14, Nav14, RS82, Wal15a, Wal15b]. campus [Ber96a]. Can [Bet46b, YB85, MB85, Bet20]. Can’t [Bet91d, SB86]. Capacity [Maw78]. Capture [Bet70b, HB51b]. Carbon [Bet40d, BLZB01, Hoy54]. Carbon-Oxygen [BLZB01]. career [NS15, Sch15]. Carl [HH14, Wie14]. Carnegie [HPA97]. Carson [PBB97]. Case [AAB^{+}88, Bet60a, Bet76b, Bet80b, Bet84b, Gou47, Ken54, Nak63a]. Cases [Cle48]. Catalogue [Hen94, Sch68]. Catastrophe [Dud75, Bet76h]. Cathode [Bet29c]. Causal [Kea66]. Cavendish [Nav06]. cavities [Bet43a, BS43]. Cease [Bet95c]. celebrate [Gol05]. Celebration [KRW05]. Centaury [Rig06, Rig07]. Centenary [Bet99a, FK85]. Centennial [PS69]. Centuries [Sch68]. Century [ABP98, Bet01b, Sch68, APB96, MB85, Har10, Wal15a, Wal15b]. cervical [HB69]. Chadwick [Bet98b]. Chadwyck [Hen94]. Chadwyck-Healey [Hen94]. chain [Bet31b, CWL^{+}04, LFK99, Sco94, dGBNM02]. chairman [BBG^{+}85a]. Chandrasekhar [BS83, Bet05b]. Chang [Har07, RS08]. Chang-Hwan [Har07, RS08]. Change [Bet31a, Bet75d, vW88, vW91]. Charge [BE68]. Charged [Ino71, IIT78]. Charles [Bet78a, Hol03].

Chemical [ABG48, Gam66]. Chernobyl [Bet86c, Bet91d, Bet91i, SB86, SBR^{+}86]. Chernobyl-Type [SB86]. Chicago [Wal15a, Wal15b]. China [Bet69a, DeN76]. chiral [DL83]. chiral-invariant [DL83]. choices [BBM78]. Chop [Bet89c]. Chown [Bet01a]. Christoph [Wal15a, Wal15b]. CI [PR97]. Circular [Mol89]. Citizen [McD62]. City [Bet69b]. Classical [Sco94, Skr09]. Clerical [Wei75]. Climbing [Kai05a]. Clinton [Bet97c]. Close [SB47]. closed [LFK99, Ong86]. closer [Har10]. cloth [Fra01]. club [Ber96b]. cm [Sch67b, Sch67a]. Coal [Bet80b]. Cold [AK10]. Collaboration [Ber10]. Collapse [Bar88, BBAL79, Bet81d, Blu88, Myr88, WM88]. Colleagues [BC00, DK91]. Collected [Hen94]. Collections [Hen94]. Collective [BDSB94, SSB^{+}99a, TBDB98, Bet33a]. College [Tho99]. Colliding [SSBB75]. Collisions [Bet51, BA52, BJBP93, Ino71, BW87, IIT78]. Column [BV74, BV75]. Combination [BC38a, BC38b]. combinations [RJ11]. Come [STS^{+}49]. commentaries [Yan13, Bet95b]. Commentary [BWH80, BBL06, Bey80, Bet97e, BBL03, DP97, Tho99, Got98]. Comments [Bet82b, Bet91c, RYBB73, Moy97]. Committee [Bet75e, BBG^{+}85a, GBG^{+}86]. Communication [Bet71e]. Community [Bet95d, Hab69, Bet81e]. Compact [BB98, BB99, BBL07]. Company [Pei57, Tho99]. Compared [Maw78]. Comparison [SBR^{+}86, Bet29d]. competition [SB55b, SB55b]. Compiled [Sch68]. Complex [Bet82e, Hil15, BG57]. Composition [CB34]. Compound [Bet40a]. Compton [Duf46]. Computational [ML83]. computer [Fle73]. Computers [FT70]. Computing [ACM09]. Concept [CBB84]. conception [SB55b]. Concepts [BBG85b, KS87, ST86]. Concerning [Bet48]. condensates [ZLGM03]. Condition [Sat63]. conditions [KZJ02, LYS01, Mai04]. Conference [Bec67, BM55, Bet56b, Bet56c, Blu87, KMN^{+}96, KS87, KMP98]. conferences [WH72]. configurational [Per84b]. conflict [SB55a]. Congress [BG83b]. connection [Saz87]. connectivity [RM06]. Conscience [BBH^{+}73, Bra05]. consciousness [vW88, vW91]. conseil [CCJ^{+}34]. conserved [MIT06]. Conserving [Fes77]. Considérations [Hei34, Hei34]. Consistency [JB78]. Constant [BL49]. constants [Bet43c]. Constitution [Edd20, Gam30]. Constraints [BB95, Bet96b]. Construction [Wal15a, Wal15b, Per84a, SB55a, Bet33a]. Contextual [Sch14b]. continent [Ano85a]. Continuity [CWL^{+}04]. Continuum [Bet40a]. Contribution [BM71, BM72, BB99]. Contributions [KRW05, MA06, Pan05]. control [Bet8xa, Opp46]. controversies [ST86]. controversy [SB76]. Convection [BBC87b, BL88, SSB^{+}97a, WM88]. convective [WW88]. Conventional [Usd09]. Conversation [BM04, Goo99, BM05, Goo01]. conversations [Bri95a]. Cooper [Skr09]. Cooperative [Bet37b]. Coordinate [LYS01, MP91]. Copenhagen [Sch00a]. copies [Roj07, Roj09]. core [BG57, Blu88, Ong86, Ong88, WM88]. Cornell [Ano37, AK10, Bet78a, Bet93k, Dys89, Gol05, Pev78]. corpuscular [Bet30b].

Correction [BBF53, Por04a]. Corrections [Ano93, BB51, WW88]. Correlation [PR97]. Correlations [Bet65b, Bet67a, JB78, BJ76]. Correspondence [Mor02, Lee07, Lee09, MIT06]. Corvallis [KMN^{+}96]. Cosmic [Bet35d, BKP40, BKB92, CB34, Mon49]. Cosmogony [Gam49]. Cosmological [Gam66]. cosmology [Tur08]. Could [Sul75]. Coulomb [Bet38b, BR52, CH02, Kyr72, WW88]. Counters [BB47]. Counting [CCG^{+}11]. Coupling [SBB^{+}97]. Courant [Bet54b]. Course [Bet47b]. Courting [Gla76]. Creation [Bet34a, BH34, KM01]. Credibility [Sau75]. Credulity [vHWD76]. Critical [BK39, BSS05]. criticism [BF33]. critics [Cor12]. critique [Bet34c]. Cross [BM54, BBC56a, BBC56b, BBC57a, BBC57b, DB52, DBM54, HB51a, HB40, HB51b, MB52, Bet43b, BW87, ML83]. Cross-Sections [BBC56a, BBC56b, BBC57a, BBC57b]. Crossing [Sch95, Wei95]. crowd [Fri99]. crystal [Bet42e]. Crystals [Bet42c, Bec29, BB76, Bet27, Bet28b, Bet28a, Bet29b, Bet42d, Bet49d, Bet50f, Ewa58]. Culture [Wal15a, Wal15b]. current [RS82]. curtain [Bet94a]. Curve [Gam30]. curves [Bet49c, Bet49e]. Cutkosky [MP91]. Cycle [Bet40d, Wie14]. Cyclotron [BR37b, Hil15, SSB^{+}97a]. Czech [Bet71b].

D [Bet54b, Bet78a]. Dalitz [Tho99]. Damping [BO46]. Dancoff [DRS^{+}54]. Danger [SB47]. Dangerous [BGG^{+}85b]. dans [Dem15, NBP^{+}34]. Dante [BBG^{+}85a]. Dark [Bet95a]. Data [Bet35a, BHP39, BHP97]. Daughter [Pal00]. David [Ano50, Bet94a, OBxx]. Davisson [Geh73]. day [HHW99]. Days [WM15]. DC [KMP98]. Deaths [Bha05]. Debate [Bet77a]. Decay [BN40, Aui02, FBE^{+}34]. Decision [Dow83, GB89, CWB^{+}69]. decisions [Str63]. Defect [Gam30]. Defended [WB64]. Defends [Pal76]. Defense [Bet69b, BGGK84, Bet8xb, BCMR92, BGGK12, TB84, Tel87]. Defining [Mla98]. Deflection [CB48b]. Deformation [BW39]. deformed [AMS11]. Degeneracy [BB69]. delta [Fun81]. delta-function [Fun81]. Denaturing [ABB^{+}46]. denounce [Bie03]. Dense [BJ74a, BDSB94, Han74, PB73, SBB^{+}97, SSB^{+}97b, BBD^{+}95, Bro88a]. Densities [Bet74]. Density [Bet74, BBCW83, BLMB41, HB51b, MJB75]. Department [Sch68]. Dependence [HB51a, Por03]. deploy [CWB^{+}69]. Dep’t [ABB^{+}46]. Derivation [Bet63c, MP91, Wie81]. Description [BBD^{+}96]. Desist [Bet95c]. destroyed [She75]. Destruction [Gla76]. detect [Nav06]. Detection [All59]. Detectors [ATP82]. determinants [dGBNM02]. Determination [WLB36]. Deuteron [Bet38a, Bet40c, BL50, GB51, RYBB73]. Deuterons [BC38a, BC38b]. deutsche [Bet05a]. Develop [Bet98a]. Development [Bet54e, Bet82f, Bet84b, Bet89e, HB69, HBE87]. Developments [CC73]. Deviations [BT41, BT45, BT51a, BT51b, BT53]. devices [Ula66]. DEW [Bet87a, CBR87]. DeWolf [Edg91]. Dexer [Duf46]. Diagram [DeS73]. Diagrams [Kai05a, Kai05b]. Dialyse [BBT24]. dialysis [BBT24]. Did [BGS95b, STS^{+}49]. Dielectric [ATP82]. Dies [Ano05a, Bra05]. Different [Bet29d, Van32]. Differential [BM54, MB52, Wil73]. Diffraction [Bet44, CJK92, Geh73, Bet28a, Geh78, HM83]. Diffused [BB76]. Diffusion [BFK49, BV74, BJ77, Bet82g, App88, Bet49b, BV75]. Digit [LG78, Rai85]. dimension [Lai74, MB82]. dimensional [LYS01, YLGC02]. Dinner [Bet77b]. Diplon [BP34c, BP35a]. Dipole [LB50]. Dirac [Bet48, CH02]. Diracschen [Bet48]. Direct [HLB37, KW03]. Directions [MKR87]. Director [Orb05]. Disarmament [Bet62b, Bet64a, Bet70c, Bet71a, B^{+}62]. Disclaimer [Nov76]. discourse [KMN^{+}96]. Discovery [Geh73, Hah58, CHW91]. Discusses [Bet65a]. Discussion [FBE^{+}34, SG73, Ano50]. discussions [CCJ^{+}34]. Disintegration [BP34c, Bet35b, BHP39, BL50, BHP97]. Disorder [Bet38f, BK39]. Dispersion [Nak63b, Kai05b]. Dispute [Bey80]. Disputes [Ken54]. Distances [BTH50a, Bet49b, BTH50b, BTH50c]. distributed [Dow83]. Distribution [Bet29d, BA52, BE68, BLMB41, MYZ67, SB59]. Disturbing [Dys79, Bet79a]. DM [Sch67b, Sch67a]. Do [BB93, BG83a, Ano50, Pol19]. documentary [CHW91]. dollar [CP93]. domain [KZJ02]. Dominated [BLB00]. d’onde [Gau83]. Don’t [BBL06]. Doomesday [BG82a]. Doppler [Bet57]. Double [Nak63b, Yui01]. Double-Loop [Yui01]. Doubt [Sul75]. Douse [Bro97]. Down [BTH50a, BTH50b, BTH50c, Bet89c]. Dr. [Bet93j, Ken54, Orb05, Wri10]. Drama [Ken54]. Drawing [Ber03, Kai05b]. dreams [Haw11]. durch [Bet29c, Bet30b]. Durchgang [Bet29c]. Durchgangs [Bet30b]. during [Ber96a]. Durling [Sch68]. Dwarf [BLZB01]. Dynamical [CJK92]. Dynamics [Bet37a, LB37, FCC09]. Dynamo [Ler71]. Dyson [Bet79a, Bet79a, Ler71].

Eagles [Bro01]. Early [Bet97a, Nav06, BL88, Sch15]. Earth [Bet30c, Bet70a]. economic [BKD82]. Ed [Gol68, Bet54b, Hen94]. Edited [Duf46, Bet90b]. Editor [GT38, Bet60b, Mar08]. Editorial [Ale05, Bet81a, BL83, Bof76]. Editors [Mer68, RS08]. Education [Ano06, Sch68, Bet83b]. Edward [Bet01b, Bro82, Har10]. Effect [Bet30c, Bet57, BG57, LB50, Por04a, Bet30a]. Effective [Bet49a, BL50, Bet73b, HM83, MJBA75]. Effects [BP37, BJ76]. Eigenfunction [Bet34b]. Eigenfunctions [Bet48, vdLB47, Bet31b]. Eigenfunktion [Bet34b]. Eigenfunktionen [Bet31b, Bet48]. eigenvalue [RR07]. Eigenvalues [HSY04, Wil73, vdLB47, RR07, Bet31b, RJ11]. Eigenwerte [Bet31b]. Einleitung [Bec29]. Einstein [Ano92a, BGA^{+}13, OM18, ZLGM03]. Einsteins [Ale05]. Eisenhower [Kil77]. Elastic [BM71, BM72, RYBB73, DeS73, Gur01]. Eleanor [Ano50]. Elect [ABB^{+}80]. Elections [Bet32]. Electric [Bet29c, BJ77]. Electrodynamics [Dys05b, Dys06, Sch58, Sch03]. electrokinetic [Rig06, Rig07]. Electromagnetic [Bet47a, Bet72a]. Electron [Bet33b, BO46, BM71, BM72, Geh73, Geh78, RYBB73, RB39, Sch78, SSB^{+}97a, SB53, SB59, BS57a, BS57b, BS77, BS08, BW87, CF12, CH02, Frü03, HBE87, SSB^{+}98, BZ38, Bet67c, Sch67b, Sch67a, SB33, SB34, SB67e].

Electron-Cyclotron-Harmonic [SSB^{+}97a]. Electron-Deuteron [RYBB73]. Electron-Diffraction [Geh73]. electron-hydrogen [BW87]. Electronic [Bet29a, vdLB47]. Electrons [Bet29d, Bet34a, Bet34b, BH34, BRS38, BBW50, BBAL80, FBB14, WB51, Bet27, Bet28b, Bet28a, BF32, BF33, BJ77, Bet14, NBP^{+}34, SB67c, SB53, SB67d, NBP^{+}34]. Electrostatic [CB48b]. elektrische [Bet29c]. Elektronen [Bet27, Bet28b, Bet28a, Bet32, BF32, SB67d, SB67c]. Elektronenaffinität [Bet29a]. Elektronenprobleme [Bet33b]. Elektronentheorie [Sch67a, Bet67c, Sch67b, SB33, SB34, SB67e]. Elektronenverteilung [Bet29d]. Elektronnaja [BZ38]. element [vW38a, vW38b]. Elemental [Wal15a, Wal15b]. Elementary [Bet47b, BM56, BM06, For57, Pri56]. Elements [ABG48, BBFH57, Hoy54, Gam66, Hoy46, Pen79]. Elementumwandlungen [vW38a, vW38b]. elite [Fin06]. elliptic [MN07]. Elusive [BG83a]. Émigré [Wal15a, Wal15b]. Eminent [EBU^{+}52]. Emission [BAB80, MWBB02]. Empirical [Bet93h]. Employment [BS47, Gou47]. End [Sul75]. Ending [Bet60a, BMG87]. energie [Bet71b]. Energieerzeugung [Bet68a]. Energies [BML53, Bet91a, DSB91]. Energy [Anoxx, BS05, BS06, BB47, Bet36a, BR37a, BR37b, Bet38a, Bet38b, Bet39c, Bet39a, Bet42b, Bet47a, Bet50a, BK62, Bet64c, Bet65a, Bet66a, Bet68g, Bet68d, Bet68e, Bet70d, Bet71c, Bet75a, Bet76c, Bet77c, Bet77e, Bet80b, Bet80c, Bet81e, Bet82c, Bet83a, BP90, Bet94a, CB48a, CB60, CP50, Edg91, GB51, Huf81, Iof05, Iof06, MB41, PR97, Ran48, Seg85, SBR^{+}86, Yui01, Bla82, BBM78, Ber80, Ber81, Bet49c, Bet49e, Bet53d, Bet68c, BN68b, Bet70a, BV75, Bet76a, CF12, Frü03, KW03, Kra16, KMP98, LW71, Mai04, Men27, Opp46, RS05, RJ10, RJ11, SB76, Sel77, Bet68a, Bet71b, Dow82, Dow83, Huf82]. Engineer [Bet78a, Pev78]. Engineers [SBR^{+}86]. Enough [Dow83, Gul75]. Enrichment [Bet82i]. Enrico [Bet55c, Ano61, BB02, Ore04]. Entgegnung [Sch34]. Entwicklung [HB69]. entzweite [Bet01d]. Environment [KMP98, KMP98]. Episode [Pan05]. Equation [BB69, BB71, Bet74, BBAL79, BBAL80, BBCW83, Bis67, BT69, FT75, Gün64, HD72, Kea66, Kea68, Kea69, Kea70, Ler71, Nak63a, Nak63b, NR66, Ros66, Sae67, SB51, Sat63, Vos60, AS70, BB70, Bet98a, Blu88, Bre75, Bro88a, DSB56, FV86, Fle73, Flo76, Fur85, HHD73, HD70, KW03, Kea71, MP91, Mai03, MB82, Nak71, Ong86, Ong88, Saz87, WA94, Wie81, Zum75]. Equations [MJB75, Bas71, BK93, Gar02, Kyr72, MIT06, Mai04]. Equilibrium [BT41, BT45, BFK49, BT51a, BT53, BT51b]. equivalence [Bas71]. Equivalent [Bet73b]. Erden [Bet30c]. Eric [Bet90b]. Erinnerungen [Bet00b]. Ernest [Hil15]. Errata [HD72]. Erratum [BB71, Bet83a, RB37a]. Error [Wei75]. errors [Bro82]. Erscheinungen [Bec29]. Espionage [Usd09]. essay [Edg91]. Essays [APB96, ABP98, BCSF82, Gol68, Mer68, dSFvLH66, MB66]. est [Dem15]. Estimate [Bet50e]. etc [Sch67b, Sch67a]. Ethical [BSP^{+}83]. Europe [Fer68, Fer71]. evaluate [BW87]. evaluation [CWB^{+}69]. Evanston [Pei57]. even [RS05]. Evening [BDH70, Gas93, BGP^{+}93]. Events [Ano05c].

Evidence [BH39, Bet40d, HLB37]. Evolution [BM39, Bet91a, BB98, BBL07, BLW^{+}00b, BLZB01, DSB91, Huf06, BBL03, BL88, NaH88]. Evolutionary [FvHW79]. Ewald [BH88, CJK92, CJK92, JMWB86]. Exact [Nak63a, ROB63, MIT06, Zum75]. Exactly [Eva84]. Examined [Bet60c]. Exchange [Bet84a, FT75, Ros66]. Excitation [Bet43a, KB38, RS05]. Excitement [Bet67e]. Exiled [RB02]. Expansions [PR97]. Expected [STS^{+}49]. experience [Bet93e]. experiment [CP93]. Experimental [BL49, LB37, Seg59a, Seg53, Seg59b]. Experiments [BB93, Bet89a, BBT24]. Experts [Sul75]. Explanation [Bet86a, Bet94c]. explicit [RR07]. Exploded [Bet46b, Bet20]. explodes [BB85, BB91b, BB12]. exploring [MR85, WH72]. Explosion [Bet50e, Bet81d]. Explosions [All59, BBB^{+}87]. Extension [YLGC02]. extensions [AAHS99]. extracted [Por04b].

F [Bet89b, Dys56, Edg91, For87, dSFvLH66]. F. [Bet79a]. F.R.S [AP77]. Face [TB84]. Face-Off [TB84]. Faces [Iri06]. factor [CF12, HB69]. factorization [Bas71]. factors [HM83]. Facts [BBSS50, B^{+}50]. Faith [ACU^{+}54]. Farm [Bet93b, Ber96b]. FAS [Ano12a]. Fascell [BBG^{+}85a]. Fashioned [Duf46]. Fast [BB47, BH34, Bet57, Bet99b, Ino71, BJ77, IIT78]. Fast-Moving [Bet99b]. faster [Bet30b]. Fateev [Ray97, RS05]. Fault [Bof76]. FBI [BLM19]. Fear [WM15]. February [Bet97b, KMN^{+}96, Ano50]. Felder [HBD71, Bet29c]. Fermi [Ano61, Bet55c, KS87, Bet63b, Bet68b, BB02, Bro71, MB40, NB68, Ore04, WB64]. fermion [Kea71, Zum75]. fermion-antifermion [Kea71, Zum75]. fermions [YLGC02]. Ferromagnetism [Bet33d, Cab78]. Ferromagnetismus [Bet33d]. Festschrift [KRW05, Lee11]. Fetter [Ano12b]. Fewer [BG83a]. Feynman [Bet88d, Bet91g, Bet93k, BR93, Dys89, Hal17, Kai05a, Kai05b]. Field [Bal62, BB51, BR52, Bet72a, Kut05b, Kut05a, Ler71, Pei57, BJ77, CFR82, SB53, SB67d]. Fields [Bet29c, Bet31a, Dys55, Dys56, Jeh56, Pei57, SSB^{+}97d, BdH55, SBdH56, CH02, HBD71, SBdH55, SBdH55, Dys56, Jeh56, Pei57]. fifty [Geh78]. fight [SB76]. Figures [Maw78]. Files [BLM19]. Film [Dow83, Dow82]. Final [BB54]. Fine [BL49, LR47]. finite [Mai04]. finite-energy [Mai04]. Fire [Bro97]. Fireball [Bet64d, Bet65c]. Fireworks [BKB92]. First [ABB^{+}50, Ano50, Bet93a, Bet95b, Ken54, LG78, Rai85, WM15, ML83, MB85, Kil77, Stu18]. first-born [ML83]. Fission [BBC56a, BBC56b, BBC57a, BBC57b, Bet76e, Bet76f, FvHW79, Hah58, Bet01d, CHW91]. Fission-Spectrum [BBC56a, BBC56b, BBC57a, BBC57b]. Fissionable [ABB^{+}46]. five [Ano85a, LW71]. fizike [BDH70]. flare [BG91a]. Flow [BLB00]. Fluctuation [Ler71]. Fluxes [SBB^{+}97, SSB^{+}97b]. Fly [Bro01]. Focus [Bar08]. Focuses [Ken54]. focusing [Rig06]. Foes [Bof76]. fonction [Gau83]. Force [BBD^{+}96]. Forces [Bet39b, Bet40b, Bet40c, BL50, Bet54a, Bet55a, Hod13, Hub14, Kra13, Low13, Nav14, Wal15a, Wal15b, Bet54c, Sch12]. Foreign [BBG^{+}85a]. Foreword [Bet75b, Bet76d, Bet01c, Bro88b, Duf46, Edg91]. Forget [BBL06]. Form [ROB63, Ong86]. Formal [Bas71, Bet43b, Sat63]. Formalism [Nak66, HSY04]. Forman [Sch14b]. Format [Sch67b, Sch67a]. Formation [BC38a, BC38b, BBL03, BLB99, BHL^{+}01]. formed [Bet29c]. Formula [FBB14, WB51, WG68, Bet14, RR07]. Forum [Bet69c, Ano50]. Forward [NR66]. Four [Aui02]. Four-particle [Aui02]. Fowler [BCSF82, BS83]. Fragments [Pal00]. Franco [Goo01]. Frederic [Dys55, Jeh56, Pei57]. Free [Yui01, Hen94, KW03, RS05, SB67c]. Freeman [Bet89b, Bet79a]. Freeze [BT82, BL82, BL83]. freien [SB67c]. French [CCJ^{+}34, Dem15, Gau83, Hei34, NBP^{+}34, SB55b]. Frenkel [BF33]. Frequency [Bet42c, Bet42d]. Friedensforschung [HH14]. Friedrich [HH14, Wie14]. Friend [Sch05a]. Friends [Bet95b, DK91]. Frisch [BW80, Dit80]. Fuchs [Wal15a, Wal15b, Ber10]. Fuel [Bet75c]. Fulfilling [Bet80c]. Full [Duf46, MW46, MW07, PR97]. Full-CI [PR97]. Function [Kea66, Nak66, Fun81, YLGC02, WF11]. functional [Sco94]. Functions [HD72, KB38, HHD73, HD70, Kyr72, PcPZ90, dGBNM02]. Fundamental [Bet59, Bet75d]. Furnace [Bet01a, Cho01]. Further [Por03]. Fusion [Bet78b, Bet80d, Bet81b, Bet79b]. Future [Bet93i, BGA^{+}13, Iri06, BT61].

GaAs [Por98]. Galaxy [Bet94d, BB94a, BB94b, BLW^{+}00b, BBL03]. Gamma [BLW^{+}00a, Eps88, MWBB02, BLW^{+}03]. Gamma-Ray [BLW^{+}00a, MWBB02, Eps88, BLW^{+}03]. Gamow [HPA97, AH73, Bet97a, Har01, Mar08, Rig06, Rig07]. Ganzes [SB67b]. Gap [BLB01]. Gas [BBAL80, Run67, Lai74]. Gatlinburg [Bec67]. Gaudin [AMS11, Gar02, KM01, Sch15]. Gaussian [Frü03]. Gaussian-mixture [Frü03]. Geburtstag [HBD71, Hof96, KRW05]. General [Bet40b, Bet56b, Blu88, Nak63a, Hei34]. General-Mass [Nak63a]. générales [Hei34]. Generalist [Sal02]. Generalization [SG73]. Generalized [CTSOS99, MB40, Roj07, Roj08, Roj09, RM09, RJ11]. Generation [BS05, BS06, SSB^{+}97d, SSB^{+}97b, Wal15a, Wal15b]. Geneva [MB85]. Genius [Bri95a, Bet93a]. Genocidal [Bet90b]. George [Har01]. Gerald [Har07, Lee11, RS08]. German [Bet54b, Bet05a, Ano93, BBR31, Bec29, BBT24, Bet27, Bet28b, Bet28a, Bet29a, Bet29c, Bet29b, Bet29d, Bet30a, Bet30b, Bet30c, Bet31b, Bet32, BF32, Bet33a, BF33, Bet33c, Bet33b, Bet33d, Bet34c, Bet34b, BS47, Bet48, Bet55d, Bet67c, Bet68a, BGP^{+}93, Bet93c, Bet00a, Bet01d, Bet05a, CSB^{+}31, Gas93, Gou47, HBD71, HB69, Hof96, Hou30, KRW05, Kut05b, Sch14a, Sch32, Sch34, Sch67b, Sch67a, SB55a, SB33, SB34, SB67d, SB67e, SB67c, SB67b, Wie14, vW35, vW38a, vW38b, vW88, vW91]. Germans [Wal15a, Wal15b]. Germer [Geh73]. Geschwindigkeit [Bet32]. giant [MB85]. Giants [Bet95b, Dre05, Mat98]. gitterförmige [Bet29c]. Give [EBU^{+}52]. given [Hen94]. Glasgow [BM55]. glassy [BA59]. glitching [Eps88]. Global [Iof05, Iof06]. Gluon [BBP91, BJBP93]. go [Mat98]. Goes [Bet87a, CBR87]. Goldstein [Kea70]. Goldstone [Ong86, Ong88, PR97]. Goldstone-Type [PR97]. Good [BR93]. gradients [Rig07]. Grand [She75]. Graph [WF11, Roj07, Roj09, RJ11]. Gravitational [BBAL79]. GRB [BLLB00a, BLLB00b]. great [Fri99]. Greater [Bet74]. Green [HD72, HHD73, HD70, Kea66, Kyr72]. Grenzüberschreitungen [Wei95]. grid [BK93, DHB^{+}06]. Grids [Bet29c]. GRO [BLLB00a, BLLB00b]. Gross [DL83]. Ground [Bet29d, Bet98a, BJ76, Ray97, Ken54]. Group [ABB^{+}80, Bet75e, MWBB02, Van32, AAHS99, MN07]. guerre [SB55b]. Guide [Ano00a, Hen94]. guides [Bet43b, Bet43d]. Guth [Eks94].

H [BS37, Duf46, Dys56, Eks94, Ewa58, Geh73, Sch67a, Tho99, ABB^{+}50, Bet50b, Bet54e, Bet82b, Bet82f, Bet89e, Bet90d, Bet91c, Bet91h, Bro82, Eli50, Ken54, McC20]. H-bomb [Bet90d, ABB^{+}50, Bet54e, Bet82b, Bet82f, Bet89e, Bet91c, Ken54, Bet91h, Bro82, McC20]. H. [Bet65a, Bet95b, MR85, Sch78, Sch95, Tho99]. Hackinsack [Wal15a, Wal15b]. Hadron [BBP91, BJBP93]. Half [BH39]. Half-Life [BH39]. Hall [Bet93b, Ber96b]. Halle [HH14]. Hamiltonian [Per84a, Per84b]. Hans [Ano50, Ano61, Ano05b, Ano12a, Ano12c, Bof76, Bur68, Cli97, Edg91, For87, Gol68, Got98, Hod13, Hub14, Inc92, Jeh56, Kra13, Low13, Mer68, Nav14, Orb05, Pei57, Ros93, Sch67b, Tho99, Wal15a, Wal15b, Ada06, Bla82, Ano85b, Ano92a, Ano00a, Ano01, Ano05c, Ano05a, Ano12b, Anoxx, Ber96a, Ber80, Ber81, Ber06, Bet80a, BWMH81, Bet84b, BGP^{+}93, Bet93e, Bet97e, Bet99d, Bha05, Bra05, Bri95b, Bro82, Bro73b, Bro05b, Bro05a, Bro06a, BL06a, Bro06b, BL06b, BL09, Dow82, Dre05, Dre07, Dys55, Dys05a, Dys05b, Dys06, Eks94, FBB14, For57, Fri99, GG05, GvH05, GvH06, Gas93, Gib05, Gol05, Goo99, Got98, GS05, Got05, Got06a, Got06b, Har07, Hod81, Hof96, HB06, Iof05, Iof06, Jac09, Kut05b, Kut05a, Lee05, LB07, Mad86, MB66]. Hans [Mar70, McD62, MTG05, MA06, Mer68, Neg05, Neg06, Pan05, Pep05, Pev78, Pri56, RS08, Sal05, Sal06, Sch96, Sch05a, Sch08, Sch12, Sch14b, Tel87, Tuc05, vB92, Dow83, Huf82]. Happen [Bet91d, SB86]. Happy [Bet79c, Sch05b, Sch06, Wri10]. Hard [BFK49, Bet69b, Run67, Ong88]. hardback [Hub14, Nav14, Tho99]. hardcover [Edg91, Wal15a, Wal15b]. Harmonic [SSB^{+}97a, Bas71]. Harmonics [BB71, BB70, Flo76]. Hartree [RB37a, BR37a, RB37b]. Harvard [Hod13, Hub14, Nav14, Wal15a, Wal15b]. Healey [Hen94]. Health [Sch68]. hears [Fri99]. Heat [Bet42a]. Heated [BV74, BV75]. Heating [BW85]. Heavy [Bet36a, BTH50a, BJBP93, JLBF66, SB67a, BTH50b, BTH50c]. Heidelberg [Sch67b, Sch67a]. Heisenberg [KRW05, Sch00a, HBD71, KW03, KRW05, LFK99]. Heitler [BO46, Frü03]. Held [Bet55c, AP77, Bec67, CCJ^{+}34, KS87, KMP98, RS82]. Helium [Bet29d]. Heliumgrundzustand [Bet29d]. Henrik [OBxx]. Henry [Bet95b, Edg91]. Herbert [Edg91]. Here [Bet91d, SB86]. Herman [BKJ84]. Herru [Sch34]. High [Bet42c, BML53, Bet53d, Bet64c, Bet65a, BB99, BLB99, BHL^{+}01, BLB01, CB48a, CB60, GB51, MJB75, Aui02, Bet42d, Bet84d]. High-Density [MJB75]. High-Energy [Bet64c, CB60]. High-Mass [BB99, BLB99]. high-temperature [Aui02]. Higher [CFR82]. Highlights [ST86]. Hilbert [Bet54b]. Himself [RB02]. hindered [Bro82]. historians [Fri99]. Historical [ABP98, Seg85, APB96, Edg91]. histories [BWMH81]. History [Ber03, Bet52a, Bet82b, Bet82d, Bet90d, Bet91c, Blu87, Bro82, Hod81, Sch14b, TB82, AWCT09, CHW91, Fin06, Jun58, Jun82, KHFA67, WH72]. Hitler [Ber96b]. Hoerlin [BKJ84]. Hoffmann [Dys55, Dys56, Jeh56, Pei57]. Holds [Bet53c, Bet84c]. Hole [BHL^{+}01, BLB01]. Holes [Bet44, Bet94d, BB99, BBL06, BB94a, BB94b, BLB99, BLW^{+}00b, BLB01, BBL03]. Holloway [Bet94a]. Honor [Bet55c, Bet99a, Mer68, Gol68, Lee11, MB66, dSFvLH66]. Hot [BBAL80, Hoy54, KMB57]. House [BBG^{+}85a, Duf46]. Hoyle [AWCT09, Mad83]. Hubbard [LYS01]. Human [Duf46]. Humanitarian [Sch05a]. Hundred [Bet95b]. Hunter [Hen94]. hvezdach [Bet71b]. Hwan [Har07, RS08]. Hybrid [Bet78b, Bet80d, Bet81b, Bet79b]. Hydrodynamics [BFvN^{+}44, PPvNF44]. Hydrogen [All56, Bac50, Bet29a, Bet48, BBSS50, B^{+}50, Bet50d, Bet50c, Bet63a, Bet82d, Bet95a, Ken54, LR47, LB62, Rid50, Sal52, TB82, dHMAB54, Ano50, Bas71, BJMK92, BW87, Hoy46, SB54, SB55b, SB55a, SB71, WG17]. Hypercritical [BLB00]. Hypernova [MWBB02]. Hypothese [SB67c]. Hypothesis [MB47, YG06, Fun81, Kra16, SB67c, Wie81].

I. [Rabxx]. i.e [HPA97]. IEF [Rig07]. Ignition [Bre53, KMT46]. II [Dys55, BBB^{+}87, Ber79b, Bet40c, BS47, Bet49c, BL50, Bet50d, BA53, BdH55, BBC56b, Bet67a, BYB82, BL88, DBM54, DHB^{+}06, HHD73, KHM98, Per84b, Rig07, SB53, Yan13, vW38b]. III [Bac50, Ber79c, BBC57a]. Ill [Pei57]. illus [Gol68, Wal15a, Wal15b]. illustrative [Cab78]. Illustrious [Fer68, Fer71]. im [Bet29d, SB67d]. immigrants [Fer68, Fer71]. Immobilized [Rig07]. impact [BKD82, LW71]. imperative [Bet85c]. Imperial [Tho99]. Imply [BB93]. Imprisoned [ABB^{+}77]. Improved [RJ10]. Incomparable [Sch05a]. Incorrect [BH39]. Independence [Bet82c]. Independent [CCG^{+}11]. index [Wal15a, Wal15b]. Indian [Bey80]. Individual [Cle48]. Industrial [Hil15, Usd09]. Industrial-Scale [Usd09]. Inelastic [BBC56a, BBC56b, BBC57a, BBC57b, Ino71, IIT78, RYBB73]. Inferior [Bet82j, Bet83c]. Inferiority [Bet82e]. infinite [NB68, Ong86]. Influence [Bet34a, Bet97a]. influential [Har10]. Informatics [Blu87]. Ingoing [BB54]. Inhomogeneous [KZJ02]. initiative [Ano85a, Bet76g, Bet8xb, SAB^{+}8x]. Innern [vW38a, vW38b]. Innocence [Stu18]. Insight [Bet91b, Wei91]. institut [CCJ^{+}34]. Institute [CCJ^{+}34, RS82, WH72, Bet85a]. Institution [HPA97]. integrable [Jur89, Skr07, Skr09]. Integral [DB52, DBM54, DSB56]. intellectual [Fer68, Fer71]. Intense [SBB^{+}97, SSB^{+}97b]. Intensity [Ler71]. Interaction [Bet73b, CB60, MJBA75, BF32, BF33, Fun81, Lai74, YLGC02]. Interactions [BDSB94, CB60, SSB^{+}99a, SBDB94, SBDB95, TBDB98, Gur01]. Interest [Bet95d]. Interiors [BM39, vW38a, vW38b]. Intermediate [BTH50a, Bet64b, BJ68, BJ73, BJ74b, BJ80, BJ86, BTH50b, BTH50c]. Internal [BLMB41, Edd20]. International [ACM09, CC73, KS87, KMP98, Opp46, CCJ^{+}34, Bec67]. Interpretation [BHP39, BKP40, BHP97]. Interrogating [OM18]. Interscience [Bet54b, Gol68]. Interview [Ano01, Bet80a, Hof96, McD62, Pev78, Bet78a]. Intrigues [Ano95, BGS95a]. Introduction [Bal62, BdH54c, Bet56b, Bet56c, Bet70e, Bet99a, Duf46, KMGT97, KHM98, Bec29]. Invariant [Nak63a, AAHS99, DL83]. inventory [KHFA67]. Inverse [Bet72a, Wie81]. Investigation [Ano95, Sch32]. Ion [BJBP93, BJMK92]. Ionization [Bet35c]. Ions [JLBF66]. Iraq [Bie03]. irises [Bet43c]. Iron [MWBB02, Van32]. Iron-Group [MWBB02]. irreversible [Eva84]. Isaac [Hen94]. ISBN [Bet91b, Edg91, Hen94, Hod13, Hub14, Nav14, Tho99, Wal15a, Wal15b]. Ising [Aui02, BG91b, WF11]. isoelectric [Rig06]. Issue [Bet99a, Got05]. Issues [Sal05]. Istochniki [Bet68c]. Ithaca [Ano37]. IV [BBC57b, BBH^{+}73]. ix [Hub14].

J [Ano50, Bet90b, Bet97c, Geh73, HD72, Mer68, Sch68]. J. [Bet63b, Bet67d, BSK67, Bet68f, Bet91e, Bet97b, Goo80, Goo81, KW69, OBxx, PC06, PBB97, RB67a]. Jacobsen [Wal15a, Wal15b]. jadernoj [Bet76f]. jadernye [Bet55a]. James [Bet98b]. January [BH88, CJK92]. Jasons [Fin06]. jenergetiki [Bet76f]. jenergii [Bet68c]. Jenergija [Bet70a]. Jeremy [Huf82]. Jo [Wal15a, Wal15b]. John [Ber10, Hal17, Réd05]. joined [Roj08]. Jones [Hen94, WG68]. Joseph [Ano92a]. Journey [Bet01b, FR13]. Joy [Bet91b, Wei91]. Judging [Har10]. Judgment [Bof76]. July [AP77, Ano37, Bet56b, Bet56c, BL09, Dre07, Gol68, LB07, RS82, MB66]. June [ACM09, DS88, RS82]. Jungk [Bet58a]. Juni [HH14].

K-Electrons [BBW50, WB51]. kapton [Por97, Por99]. Kathlerine [Duf46]. Kathodenstrahlen [Bet29c]. keepers [Van03]. kernel [Aui02]. kernels [CFR82]. Kernmassen [vW35]. Kernspaltung [Bet01d]. Key [Ken54]. Kidder [Mor02]. Kinematic [Ler71]. Kinetic [BR37a]. Klaus [Wal15a, Wal15b, Ber10]. Klein [Eks94]. Knizhnik [Gar02]. Konflikt [SB55a]. Konstruktion [SB55a]. Korpuskularstrahlen [Bet30b]. Kristallen [Bec29, Bet28a, Bet29b]. Kritik [BF33, Bet34c]. Krystallen [Bet27, Bet28b].

L [Orb05]. L. [Geh73]. Laboratory [Bet98a, Nav06]. Labs [BCMR92, Bet93f]. Labyrinth [Hal17]. Lack [Bet71e]. Ladder [Kai05a]. Lamb [BBF53, BBS50, BJMK92]. Landau [Bro71]. landed [McC20]. Landmarks [Bar08]. Lanouette [Bet93a]. Laplacian [RJ10]. Large [Bet38d, BK62, Bet94d, BB94a, BB94b, Ler71, Bet49b]. Large-Scale [Ler71]. largest [RR07]. Laser [BV74, Bet84b, Bet84a, BV75]. laser-heated [BV75]. Last [Dre05]. Lattice [BSS05, Lai74, Run67]. Lattices [BG91b, Eva84, Per84a, Per84b]. Laucht [Wal15a, Wal15b]. Laureate [Ano05b, Bof76, Bro82]. laureates [Bie03, Hoi01]. Lauritsen [Hol03]. Law [Ben38]. Lawrence [Hil15]. layer [Bet42e]. Leadership [BC00]. Leap [BCMR92]. Lecture [Bet56b, Bet56c, Bet94b, BCC^{+}46, Ber96a, Bet68g]. Lectures [BFvN^{+}44, Bet99d, BDH70, Eks94]. Lee [Har07, RS08]. left [Ano50]. Legacy [Iri06, Tur08]. Legend [OM18]. Legendre [CFR82]. lekcii [BDH70]. length [CWL^{+}04]. Lennard [WG68]. Lennard-Jones [WG68]. Leo [Bet93a]. Léon [Wal15a, Wal15b]. Leopoldina [HH14]. Leopoldina-Symposium [HH14]. less [Bar88]. Let [ABB^{+}50]. Letter [Bet60b, Bet89d, BCMR92, Bet93c, Bet95c, Bet95d, Bet97c, Bla87, Cli97, DeN76, Fes77, GT38, Gla76, Mar08, Maw78, McK86, Mol89, Nov76, Pal76, Tel87, vHWD76, BBG^{+}85a, Dem15, Moy97]. Letters [Ger75, Gir75, Gul75, Hen94, KRW05, Sau75, Smi75, Wei75, Dys18, Réd05, CSB^{+}31]. Level [HB51b, Bet84d, Sch15]. Levels [Bet36a, Bet47a, BN68a]. Li [Bet38c]. Library [Blu87, Sch68]. Life [BH39, Bet67e, Bet81c, Bet93d, Bet06, KMN^{+}96, Sch95, Sul75, Bet88b, BDH^{+}89, Bet03, Gol05, PC06, BDH70]. Lifton [Bet90b]. Ligand [Bal62, Kut05b, Kut05a]. Ligandenfeldtheorie [Kut05b]. Light [Bet35a, Bet38b, BR52, Bet75e, GB89, LBC^{+}75, RV88]. light-water [LBC^{+}75]. Lilienthal [Ano50]. limit [AAHS99]. Line [RJ11]. linear [Bet31b]. lineare [Bet31b]. Lines [MWBB02]. Linus [KMN^{+}96]. Liquid [LB62]. List [BL06b]. Lit [Bro97]. Lives [Bet80e, Bet93f]. Living [Sch96]. Livingston [For87]. Lloyd [BKHS89]. logarithm [BJMK92]. London [Bet54b, Nav14, Wal15a, Wal15b]. Long [Bet01b]. Look [Bet01b, Har10]. Looks [Sal02]. Loop [Yui01]. Loser [BG82b]. loss [CF12, Frü03]. Lost [Sau75]. Low [Bet94d, BB94a, BB94b, BLB99]. Low-Mass [BB94a, BB94b, BLB99]. luminaries [Bri95a]. Luminosity [BYB82]. Lumped [Bet43c]. LWR [Bet75c].

M [BGM91b, For87]. M.I.T. [Bet69c]. MA [Hod13, Hub14, Nav14, Wal15a, Wal15b]. Macmillan [Wal15a, Wal15b]. Made [Bet99d, Haw11]. Magic [Bet01a, Cho01]. Magnetic [Bet31a, Bet35c, Bet38c, HLB37, RYBB73, RB37a, RB37b, SSB^{+}97d, CH02, Sco94, BF33]. Magnetische [BF33]. Magnetized [SSB^{+}97b, SSB^{+}97c]. magneto [Bec29]. magneto-optical [Bec29]. magneto-optischen [Bec29]. Magnetoplasmas [SBB^{+}97, SSB^{+}97a]. magnon [MB82]. Make [BT82, YB85]. Maker [Dys18]. Makes [Bar08]. Making [Bet95a, Dow83, Hod13, Hub14, Kra13, Low13, Nav14, Wal15a, Wal15b, Sch12]. Man [Bet93a, Bro73a, Eds68, Hol03]. Management [Bet84d]. Manhattan [Ber96a, Bet94g, BC00, HHW99]. Manuscripts [Hen94]. Many [Bet36b, Bet56a, HB06, Kai05a, KB38, KS87, Pan05, Tel55, CC73, HM83, MHFW72, Per63]. many-beam [HM83]. Many-Body [Bet36b, Bet56a, HB06, Kai05a, KB38, KS87, CC73, Per63]. March [Bet99a, BL09, Dre07, KMN^{+}96, LB07]. Marcus [Bet01a]. mark [AP77, PBB97]. Marks [Sch00a]. Markusen [Bet90b]. Marshak [Gol68, Mer68]. Marvin [BGM91b]. Mary [Wal15a, Wal15b]. Maryland [ACM09, Blu87, Hen94, Sch68]. Mass [Bet46a, BP90, Bet94d, BB95, BB99, BB94a, BB94b, BLB99, BHL^{+}01, BLB01, Gam30, Kea68, Nak63a, WG68, Bar88, Bre75, vW35]. Mass-Energy [BP90]. Masses [Bet35a, BLB01, Jea24]. Massive [Ros66, NaH88]. Master [Ber79a, Ber79b, Ber79c, Ore04]. Masters [Duf46]. Materials [ABB^{+}46, BA59]. Materie [Bet30b, Bet33a]. Math [HD72]. Mathematical [Bet54b]. Mathematics [Ber03, FF91]. matrices [Skr07, Skr09]. Matrix [SB61, BW87]. Matter [BBP71, BA53, BBP63, Bet65b, Bet67a, BR67, Bet67f, BBS70, Bet71d, BJ74a, BBC87a, Dys66, Han74, Neg05, Neg06, RB67b, Bet30b, Bet33a, BN68b, Bro88a, KW69]. Maximum [BR37b, Bet50e, BB95]. May [ACM09, GSS^{+}03]. McMahon [Ano50]. MD [Bet98a]. Meaning [Duf46, MW46, MW07]. Meaningless [Bet81a]. Measurements [BKP40, Por04b]. measures [Per84a, Per84b]. mechanical [HBE87, Sch32]. Mechanics [Bet33b, Sch78, BS57a, BS57b, Bet64b, BJ68, BJ73, BJ74b, BS77, BJ80, BJ86, BS08, Saz87]. Mechanism [Bet94b, SB54, SB71, All56]. mechanisms [Bet90a]. Medal [Bet93e]. Medalist [Ros93]. media [DeS73]. Medical [Blu87]. Medicine [Blu87, Sch68]. Medium [BTH50a, SSB^{+}97c, BTH50b, BTH50c]. Meet [PS69]. Meeting [Bet55c, CCJ^{+}34, MR85]. memoir [Kil77]. Memoirs [Bet01b]. Memorandum [Bet35d, Bet52a]. Memorial [CJK92, Eks94, Bet55c]. Memoriam [GvH06]. Memories [BR93, MTG05]. Memorium [GvH05]. Men [All56, Ano92b, Str63, SB54, SB71, Van03]. Menace [All56, SB54, SB71]. Mentality [Bet90b]. Mentor [Sch05a]. Merely [Ber01]. Merge [BB98]. Mergers [BB99]. merging [BBL07]. Meson [BB51, BM55, Bet39b, Bet40b, Bet40c, BN40, Bet46a, BW51, Bet53b, BdH54a, BdH54b, DRS^{+}54, LB62, MB47, Pei57, BdH54c, Bet55d, DSB56]. meson-nucleon [DSB56]. Mesonen [Bet55d]. Mesons [Bet54c, Bet54a, BdH55, SBdH56, Bet70b, SBdH55, Bet49c, WA94, BdH55, Bet55a, Dys55, Dys56, Jeh56, Pei57, Dys55, Pei57]. Messages [BBH^{+}73]. metal [BF33, SB67b]. Metall [SB67b]. Metalle [Bet31b, Sch67a, Bet67c, Sch67b, SB33, SB34, SB67e]. Metallelektronen [BF33, Bet34b]. Metallic [Ano37]. metallow [BZ38]. Metals [Bet34b, BZ38, Bet31b, Bet67c, HBE87, NBP^{+}34, Sch67a, SB33, SB53, SB67e, SB34, Sch67b]. métaux [NBP^{+}34]. meteoric [Kra16]. Method [Bet38d, BBP63, LR47, MB40, PB73, WLB36, vdLB47, BK93, Cab78]. Methoden [Bet29d]. methodologies [Rig06, Rig07]. Methods [Bet29d, Dir69, DeS73, Dir70, Bet54b]. Mev [dHMAB54]. Mezony [Bet55a]. Michael [Dow83, Dow82, Hen94, Wal15a, Wal15b]. Michel [Sch15]. Microfilm [Hen94]. microfilms [Hen94]. Microwave [LR47]. middle [McC20]. Migdal [Bro71]. migration [Fer68, Fer71, Hol84]. Mikheyev [BB91a]. Military [Bet70c, Edg91, Hil15]. Miller [BGM91b]. Minimize [Yui01]. Minute [BH39]. Missile [BGGK84, BGGK12, GB68, CWB^{+}69]. Missiles [Bet69b, Bet89d]. Mission [Fak83]. Mixed [BB51]. mixture [Frü03]. Mnemonic [Ran48]. Model [BR37a, Bet38c, BB52, BG91b, KB38, RB37a, RB37b, AAHS99, AFF01, AMS11, Aui02, DL83, FCC09, Frü03, Gar02, KW03, KZJ02, KM01, LYS01, MP91, PcPZ90, Ray97, RS05]. Models [MJB75, AFF01, HSY04, Jur89, Per84a, Per84b, WW88, ZLGM03]. Modern [Gol68, Mer68, Tho99, GSS^{+}03, MB66, Opp66, ST86, Uns00]. moderne [Uns00]. Modes [SSB^{+}97a]. Modified [Por99, Por03]. molecular [BJMK92]. Molecules [Ino71, Run67, IIT78]. Molière [Bet53a]. Moment [Ano00b, Bet35c, Bet38c, HLB37]. Moments [RB37a, RB37b]. Momentum [Sae67, SB59]. mondial [SB55b]. monograph [Sch15]. Moral [Aft02, Fra01, Lan00, Olw01, Sar01, Sei01, Sch00b]. Mordechai [AAB^{+}88]. Morrison [Edg91, For57, Pri56]. Most [BR93, Har10, Haw11]. Motion [SBSW60]. movement [BFG^{+}85]. Moving [Bet99b]. Moynihan [Moy97]. Mr. [Sch34]. Mrs [Ano50]. Mrs. [RMR^{+}50]. MRT [Per84a]. Much [Dow83, Ger75]. multi [BK93]. multi-grid [BK93]. Multiple [BRS38, Bet46a, Bet53a, YLGC02]. Muonic [BN68a]. must [Pol19]. My [Bet81c, Bet93e, Bet03, Bet06, Jac09, Sal06, NS15]. mylar [Por97]. Mystery [FR13].

N [HB40]. nach [Bet29d]. Named [Bet63b]. National [Blu87, GR63, Sch68, Tel87]. NATO [RS82]. Natural [FBE^{+}34]. NDRC [Bet98a]. NDRC-B-237 [Bet98a]. Near [Bet75c, BBCW83]. Near-Breeders [Bet75c]. nearest [Lai74]. necessary [BBG^{+}85a]. Necessity [Bet76e, Bet76f]. Need [Bet77e, BG83a, Bet91f, Bet98c, Smi75, Bet83b]. Needed [Bet88c]. Negative [LB62]. Negotiations [Bet54d]. neighbor [Lai74]. Neobhodimost [Bet76f]. Net [CBB84]. Neuere [Hou30]. Neumann [Ber10, Réd05]. Neutrino [BB90, BB93, BP34a, BP34b, Bet35c, BAB80, BYB82, Bet82g, BW85, Bet86a, Bet94c, BP97a, BP97b, Myr88, SSB^{+}97d, SBB^{+}97, SSB^{+}97b, SSB^{+}99a, SBDB94, SBDB95, TBDB98, Bet89a, Nav06, RV88, SSB^{+}98]. Neutrinos [BS05, BS06, BBAL80, Bet92, Bet94f, BDSB94, BBD^{+}96, Coo88, SSB^{+}97c, SSB^{+}97a, SSB^{+}99b, BB91a, BBD^{+}95, Mad86]. Neutron [App88, BB47, BBP71, BB51, BKP40, Bet49b, BTH50b, BTH50c, BTH50a, BS68, BBS70, BB95, Bet98b, Bet01e, BLZB01, GB51, HLB37, HB51b, LB52, MJB75, Bet52b, BL88, Eps88]. Neutron-Star [BLZB01]. Neutrons [BP35b, Bet35b, BBC56a, BBC56b, BBC57a, BBC57b, BP97c, CB48a, WLB36]. Neveu [DL83]. Newton [Hen94]. Next [Ale05, BL83]. nichtstationäre [Bet30a]. Nickel [Hoy54]. Niels [Duf46, Ano95, Bet85a, FK85, OBxx]. nine [Van03]. Nitric [KMB57]. No [BAA^{+}75, BBM^{+}76, Bet85d, Bet98a, Gir75, Hen94, BG83b, Bet75d]. Nobel [Ano12c, BS83, Ano05b, Bet63d, Bet67b, Bet68g, Bie03, Bof76, Bri95a, Bro82, Hoi01, Kra17, Mad83, Sch14a]. Nobelpreis [Sch14a]. Non [Bet73c, Skr09, Bet30a, Skr07]. Non-skew-symmetric [Skr09, Skr07]. non-stationary [Bet30a]. None [Duf46, MW46, MW47, MW07]. Nonlinear [BBD^{+}96, SBB^{+}97, SSB^{+}97b, SSB^{+}97c, SBDB94, SBDB95, Jur89, PcPZ90, Wie81]. Nonstatic [BM71, BM72]. Nonuniform [SSB^{+}97d].

Normalization [DL83, Sat63, Tom83]. Norms [PcPZ90]. Note [Bet64c, Bet72a]. Notes [Bre53, RS08]. nova [Bet93j, BLLB00a, BLLB00b]. Novel [CC73]. November [Blu87]. noyau [Hei34]. noyaux [CCJ^{+}34]. Nuclear [Anoxx, Bec67, BM55, BB36, Bet36b, Bet37a, BP37, Bet39b, Bet40b, Bet40c, Bet40d, Bet47b, Bet49a, BL50, Bet50e, BB52, Bet54a, Bet55a, Bet56b, Bet56c, Bet56a, Bet60a, BBP63, Bet65b, Bet67a, BR67, Bet67f, Bet70b, Bet71d, Bet74, Bet75d, BAA^{+}75, Bet76b, Bet76g, Bet77a, Bet77d, Bet77e, Bet80b, Bet82e, BT82, Bet83b, BBCW83, BMG87, Bet88c, Bet90d, Bet91f, BGM91a, BGM91b, Bet93f, BGS95b, BGS95a, Bet97a, Bet98c, Bet99c, BCC^{+}46, Bof76, Bro97, Bro05a, Dow83, For57, For87, Gam30, HB06, Hoy54, KMT46, LB50, LB52, LB37, McC20, MB85, MB86, Mla98, MJBA75, MHFW72, Neg05, Neg06, Pei91a, Pei91b, Pri56, RB67b, RB37b, Sal52, Sch12, Seg59a, Seg53, Seg59b, Seg85, Stu18, TB84, Usd09, Wal15a, Wal15b, BCSF82, Bet54c, BM56, BT61, BN68b, BJ76, BBL86, Bet89c, BG91a, Bet91j]. nuclear [BM06, Blu88, CC73, CHW91, KMP98, MR85, Pol19, SB76, WH72, vW35, Bet37a, Kra13, LB37, Nav14, RB37a, Stu79, Wal15a, Wal15b, Hod13, Hub14, Low13]. Nuclei [ATP82, Bet35b, BB36, BR37a, Bet38b, BTH50a, Bet58b, BE68, Bet68b, BBS70, Bet73c, Bet73a, Bet77f, FSB71, HB51a, NB68, SB67a, SSBB75, Bet49c, BTH50b, BTH50c, BG57, CCJ^{+}34, Lee11, MR85]. Nucleon [Bet53b, CB60, DRS^{+}54, FT75, SB61, DSB56, Fle73]. Nucleon-Nucleon [SB61, Fle73]. nucleosynthesis [AWCT09, App88]. Nucleus [Bet36a, Bet40a, Bet53c, BM71, BM72, Bet84c, FR13, Hei34, Hou30]. Nuke [Ano05a]. Nukes [BG83a]. Nullpunktstemperatur [BBR31]. Number [Bet36a, Bet94d, BB94a, BB94b]. Numbers [Ben38, CWL^{+}04]. Numerical [BBS50]. Nurturing [Ale05]. NY [Ano37]. Nye [Wal15a, Wal15b].

O [BS37]. Obituary [BKJ84, Bet88d, Bet05b, GS05, Got06b, PBB97, Tuc05, Dit80]. Objects [BB98, BBL07]. Observation [RV88]. Observational [BB95]. Observations [Ber01, Bet54e, Bet82f, Bet89e, BP90, Por03]. observed [Por04b]. Obtainable [BR37b]. obtained [Roj07]. Obtaining [vdLB47]. occasion [AP77, BCSF82, Gol68, MB66]. occupancies [YLGC02]. Occurring [Hoy54]. ocean [Gur01]. October [CCJ^{+}34, KMP98]. octobre [CCJ^{+}34]. Oersted [Bet93e, Ros93]. Off [TB84, AFF01]. off-shell [AFF01]. Office [Bet98a, Orb05]. Old [Bet95b]. One [Bet33b, FT75, MW46, MW47, MW07, Sch78, BS57a, BS57b, BS77, BS08, Fri99, Har10, Lai74, LYS01, MB82, YLGC02, Duf46]. One- [Bet33b, Sch78, BS57a, BS57b, BS08]. one-and [BS77]. One-Boson [FT75]. one-dimensional [LYS01, YLGC02]. Open [Tel87, LYS01, BBG^{+}85a]. opened [MR85]. Opening [Sch00a]. operators [KM01]. Oppenheimer [Aft02, Bet67d, Bet97b, Edg91, Fra01, Lan00, Olw01, RB67a, Sar01, Sei01, ACU^{+}54, Ano50, Bet38e, Bet63b, Bet67e, BSK67, Bet68f, Bet91e, Bro82, Goo80, Goo81, Ken54, KW69, MS15, OBxx, PC06, Sch00b, Yor89]. Oppie [BC00]. Oppose [Bet93a]. opposition [Bro82]. optical [Bec29]. optics [Jur89]. optischen [Bec29]. Oral [Hod81, BWMH81]. Order [Bet38f, BK39, BM71, BM72, SG73]. Order-Disorder [BK39]. Oregon [KMN^{+}96]. Origin [ABG48, Bet01a, Gam66, Pen79]. Origins [Cho01, Wal15a, Wal15b]. Oscillator [Bet56d, CTSOS99, Bas71]. Oskar [Eks94]. Osteochondrosen [HB69]. osteochondroses [HB69]. other [Bri95a]. Otto [BW80]. overflow [Fri99]. overlooked [Mad83]. Own [BCMR92]. Oxford [AP77, DS88]. Oxide [KMB57]. Oxygen [BLZB01].

P [BKHS89, Bet91g, CJK92, JMWB86, HB40]. P. [CJK92]. Pacific [Ken54]. pages [Fra01]. Pair [BM54, DB52, DBM54, MB52, SB59]. Palgrave [Wal15a, Wal15b]. Pandora [Van03]. Panel [SBR^{+}86, Bet86c, Ham73]. paper [FBB14, Tur08]. paperback [Edg91]. Papers [Ano00a, Bet95b, Hen94, Sch58, Sch03, BBL03, Blu87, DP97, Haw11, OBxx, Rabxx, Tho99, Yan13]. Paramagnetic [Van32]. Parameters [Por98, Por97, Por99, Por03, Por04b]. Part [AK10, Gur01, WGO^{+}91, Bet40c, BA53, BBC56a, BBC56b, BBC57a, BBC57b, Bet70c, KMB57, Rig06, Rig07]. Partial [Fle73]. Participating [Ano50]. Particle [Bet38c, Aui02, MHFW72, Saz87]. Particles [BH34, Bet50a, Bet59, Bet73b, Gün64, Ino71, Ros66, Bet49c, BV75, IIT78, PcPZ90]. Partition [WF11]. Parts [NR66]. Passage [Pei85, Bet30b, Bet29c, BA53]. passes [Ano05b]. Passions [Bet91b, Wei91]. path [MR85, RM09]. patterns [Dys18]. Paul [BH88, CJK92, JMWB86, Sch14b]. Pauli [Uns00, Bet72b, Bet00b, Uns00]. Pauling [KMN^{+}96]. Pb [BS68]. Peace [Ano92a, GG05, Ano85a, Ano92b]. peers [Fri99]. Peierls [AP77, Cab78, DS88, Tho99, Wal15a, Wal15b, Bet88a, BGM91b, DP97, Lee07, Lee09]. Penetration [BFK49, BTH50a, BTH50b, BTH50c]. People [Ano05c, Tel55]. Percentages [DF79, MYZ67]. Percolation [BSS05]. periodic [MIT06, SB53, SB67d]. periodischen [SB67d]. Periphery [Bet70b]. Personal [Ber01, Bet91j, Jun58, Jun82]. Perspective [GSB^{+}83, Seg85, Bet91j]. Perspectives [MB66, Opp66, Sch95, Mer68, Gol68]. Perturbation [BS43, Kea71, Liu72]. Perturbations [Bet38d]. Peter [BH88, Hen94]. Peterson [Pei57]. Pevsner [Bet78a]. pH [Rig07]. Phase [Bet53b, BdH54a, BdH54b, Bet55b, Run67, dHMAB54, BB76, Cab78]. Phase-shift [Bet55b]. Phases [Ano37, BG91b]. Phenomena [Bet37b, Bec29, Bet53d]. Phenomenon [LG78, Rai85]. Philip [Edg91, For57, Pri56]. Phillips [Bet88d, Bet38e]. philosophical [HBD71]. Philosophie [HH14]. philosophische [HBD71]. Philosophy [Wal15a, Wal15b]. Photo [BL50, LB50]. Photo-Disintegration [BL50]. Photo-Effect [LB50]. Photocapture [HB51a]. Photoeffect [LB52]. Photoeffekts [Bet30a]. Photoelectric [BP34c, Bet30a]. Phys [HD72, RB37a]. Physical [Bet55c, Bet75e, Bet95b, Bet99a, LBC^{+}75, FT70, HBD71]. Physicist [Bet81c, Hod13, Hub14, Iri06, Kra13, Low13, Nav14, Pei85, Wal15a, Wal15b, Wei91, Sch12, Bet89b, Bet91b]. Physicists [PS69, Bet01d, Hol84]. Physics [ABP98, Ano12c, Anoxx, BB93, Bec67, BM55, BB36, Bet37a, BM39, Bet65a, Bet67b, BDH70, BS83, Bet88c, Bet93d, Bet95b, Bet97a, Bet99d, BM04, BGA^{+}13, BCC^{+}46, Bro05a, Bro06b, DS88, For87, Gol68, Han74, LB37, MA06, Mer68, MKR87, Mla98, Mon49, Pan05, Ros93, Seg59a, Seg53, Seg59b, SSB^{+}99a, Stu18, Tho99, Wal15a, Wal15b, AP77, APB96, BBL86, BDH^{+}89, Bet93e, Bet99c, BM05, Bra05, BL06a, Dir69, Dir70, FF91, Fri99, GSS^{+}03, Haw11, Kai05b, KHFA67, MB66, MR85, MHFW72, Opp66, Stu79, Uns00, WH72, dSFvLH66, Bet55d, CCJ^{+}34, RS08, Bet54b, Har07]. Physik [HH14, Bet55d, Uns00, FBB14]. physikalische [HBD71]. Physiker [Bet01d]. physique [CCJ^{+}34]. pi [Bet55d]. pi-Mesonen [Bet55d]. picture [SBSW60]. Pilgrimages [Bet04]. Pion [Bet72b, Bet73a, JB78, dHMAB54, Bet55b, BJ76]. pion-proton [Bet55b]. Pioneer [Ano05a]. Pioneers [Tho99]. Pions [Bet77f]. Plan [Bet60c, BG82b]. Planar [CH02]. planet [Ano85a]. Plasma [BV74, BBD^{+}96, SSB^{+}99a, TBDB98, BV75]. Plasmas [BDSB94, SSB^{+}97b, SBDB94, SBDB95, BBD^{+}95]. Plate [BW39]. Pledge [ABB^{+}50]. Plenum [Sch78]. Plenum/Rosetta [Sch78]. plus [LW71]. Point [Bet69b, Bet94g, Bey80]. Polanyi [Wal15a, Wal15b]. Polarization [Bet58b, RB39, SB61]. Polarized [SB61]. Pole [ROB63]. Poles [BK62, Nak66]. policies [CHW91]. Policy [Bet75a, Bet82e, Bet84b, Dre06, EBU^{+}52]. political [Bro82]. Politics [Hab69, Wal15a, Wal15b, LW71, Bet01b]. Polymath [Har01]. Polynomials [WF11]. polystyrene [Por97]. Ponderomotive [BBD^{+}96]. Positive [BH34]. Positrons [Bet35e]. possibility [BG91a]. Possible [Bet50e, Bet82c, Bet86a, Bet94c]. postgraduate [Sch15]. postwar [Fin06, Kai05b]. Potential [BK62, Kea69, Ong88, SB53, SB67d]. Potentialfeld [SB67d]. Potentials [BJ74a, WG68]. Power [Bet35c, BBW50, BAA^{+}75, Bet76b, Bet76e, Bet76f, Bet77a, Bet77d, Bet77e, Bet91f, Bet98c, FvHW79, Por98, Por04a, WB51, Bet91j, Por97, Por99, Por04b, SB76]. Pp [Edg91, Hen94, Nav14, Pei57, Sch68, Tho99, Bet54b, Duf46, Gol68, Hod13, Hub14, Wal15a, Wal15b]. Practice [Pei57]. Preceded [Ken54]. precision [Tur08]. Predictions [Bet99b]. Preis [Sch78, Sch67b, Sch67a]. Preludes [dSFvLH66]. presence [BJ77]. Present [CC73, RYBB73, CHW91]. presented [BCSF82, Bet93e, Blu87]. Preservation [Bet93g]. President [Kil77, ABB^{+}80, Bet97c, GBG^{+}86]. President-Elect [ABB^{+}80]. Press [Edg91, Fra01, Hod13, Hub14, Nav14, Tho99, Wal15a, Wal15b]. Pressure [CBB84]. Presupernova [NaH88, WW88]. prevent [Pol19]. price [Hen94]. primordial [App88]. Princeton [Fra01]. Principle [Bet72b]. Printed [Sch68]. private [Lee09]. Privilege [Bet89b]. Prize [Ano12c, Bet67b, BS83, Mad83, Ano92a, Bet63b, Kra17, Sch14a]. Probability [Per84a, Per84b]. Problem [BB90, Bet36b, Bet56a, BR67, Bis67, HB06, Huf81, Kai05a, Kea70, RB67b, Bet76a, CC73, Fun81, Per63]. Problems [Bet33b, B^{+}62, Bet70c, BB54, Iof05, Iof06, KS87, SB51, Bet71a, Bet8xb, Per84b]. procedure [BW87]. Procedures [Cle48]. Proceedings [BM55, DS88, KMP98, RS82, ACM09, AP77, BBM78, Blu87, CC73, KMN^{+}96, KS87, Per63, Stu79, WH72]. Process [Bet38e, App88]. Processes [BP37, BG91a, Eva84]. produced [Dow83]. Producer [Dow82]. Production [Anoxx, Bet39c, Bet39a, Bet42b, BA52, BM54, Bet66a, Bet68e, Bet70d, Bet71b, Bet83a, DB52, DBM54, MB52, SB59, Bet68g, Bet68d, Bet68a]. Produkce [Bet71b]. Prof [Pei57]. Professor [AP77, Ano38, Bur68, Cli97]. Profiles [Ber79a, Ber79b, Ber79c]. Profound [Bet91j]. Program [Ano93, Bet52a, Bet93c, Ham73]. project [BGP^{+}93, Ber96a, Bet94g, Bet00a, BC00, Bet05a, Gas93, HHW99]. projectile [Por03]. projectile- [Por03]. projectors [CFR82]. prompt [BBB^{+}87]. Propagation [CBB84, SSB^{+}97c, Fur85, Gur01]. Proper [Nak66]. Properties [Kea66, Bet49c, CCJ^{+}34, ML83, Per84a]. Prophet [Ber81, Dow82, Dow83, Huf82, Bla82, Ber80]. Proposal [Bet60b, BGM91a, Pei91a, Pei91b]. Proposed [BB52]. propriétés [CCJ^{+}34]. Protest [BS47]. Proton [BB51, BC38a, BC38b, BdH54a, BdH54b, BH56, Bet64c, Bet55b]. Protons [BP35b, Bet50a, Bet58b, BP97c, CB48a, Bet49c, Tom83]. provided [Hen94]. Proving [Bet98a, Ken54]. Public [Bet87a, Bet95d, CBR87, Dre06, Duf46, Sch68, MW46, MW07]. Publications [BL06b, Hen94, Lee05]. Published [Hen94, Hub14]. Publishing [Bof76, Tho99]. Pure [BG91b]. Purposes [Edg91]. Puzzle [Bet86a, Bet94c].

Quake [DHB^{+}06]. Qualitative [Bet73c, Bet73a]. Quanta [HBD71]. Quanten [HBD71]. Quantenmechanik [Bet33b]. Quantentheorie [BBR31, Hou30, Bet33c]. Quantitative [Bet34b]. quantities [MIT06]. Quantum [BP35a, BS57a, BS57b, BS77, Bet99d, Bet99e, BS08, Dys05b, Dys06, Fer32, Hal17, Sch78, Sch58, Sch03, Skr07, AAHS99, BBR31, Bet64b, BJ68, BJ73, BJ74b, BJ80, BJ86, CSB^{+}31, Haw11, HBE87, Hou30, Jur89, KHFA67, LFK99, MN07, Saz87, Wie81, Bet33c, Bet33b]. quantum-mechanical [HBE87]. Quark [BBP91, BJBP93, WA94]. Quark/Gluon [BBP91]. quarreled [Bet01d]. Quasars [Tel66]. quench [FCC09]. Questions [Pal00]. Qui [Dem15].

R [Bet34c, Bet54b, Gol68, Mer68, Tho99]. R. [Bet34c]. Rabi [Rabxx]. Race [Bet82a, Duf46, WM15, BKD82, Bet69a]. Radiation [Bet35e, BKP40, BO46, Fer32, KMB57, SSB^{+}97b]. Radiations [BA53, RCE30]. radioactive [Bet84d, Kra16, RCE30]. Radioactivity [BH39]. Radius [Bet36b, BS68]. Ram [CBB84]. Random [Web73, DeS73]. Range [Bet49e, Bet49a, BL50, Bet50a, Bet49c]. Range-energy [Bet49c]. Rapports [CCJ^{+}34]. Rare [Bet30c]. Rasetti [Goo01]. Rate [GT38]. Ray [BLW^{+}00a, BHL^{+}01, BLB01, MWBB02, Mon49, BN68a, BLB99, BLW^{+}03, CJK92, Eps88, HM83]. Rays [Bet29c, Bet35d, BFK49, BKB92, CB34, Bet30b]. RBMK [Bla87]. reactance [BW87]. Reaction [BO46, HB40, McC20]. Reactions [Bet40d, Bet56b, Bet56c, GT38, Hoy54, Ran48, Sal52]. Reactor [Bet56d, Bet75d, Bet75e, Bet80d, Bet81b, WPBW76, LBC^{+}75]. Reactors [Bet57, SBR^{+}86]. Reagan [ABB^{+}80, BG82a, GKB^{+}84]. Realistic [BJ74a, MJB75]. Reality [Hal17]. Realization [LC71, Liu72]. Really [BT82]. rearrangement [BN68b]. reasonable [Hol03]. reasoning [Mol89]. Rebuttal [Sel77]. Receive [Bet63b]. receives [Ano85b]. reciprocity [Bet52b]. Recoil [SB59]. Recollections [ABP98, Bet80f, Pei85, APB96]. Recommendations [BBF^{+}82]. Reconstruction [SB61]. Record [Ano06, Mat98]. recording [Bet81e, KW69]. recordings [Ber96b]. recounts [Ber96a]. Rectification [Bet42c, Bet42d]. rectifiers [Bet42e]. Red [AK10, McC20]. Reducing [MB85, MB86]. Reduction [Kea68]. Reference [BBP63]. referendum [BL83]. Reflection [CF12]. Refugee [Seg85]. Regge [BK62, ROB63]. Regions [WLB36]. related [Bet8xb, Jur89]. Relation [Bet50a, BP90, Gar02]. relationship [CFR82]. Relatively [Bet99d]. Relativistic [ATP82, BBF53, Bet32, Bis67, BJBP93, FBB14, Gam49, SB51, Bet14, Blu88, Saz87]. relativistischer [Bet32]. reliability [McK86]. Relic [BLLB00a, BLLB00b]. Remark [BBR31]. Remarks [CSB^{+}31, Dys66, Nak63b, Bet48]. remembered [DK91]. Remembering [Bet00b]. Remind [Duf46]. Reminiscences [Bet88a, NS15]. Reorientation [Ham73]. Rep [Bet98a]. Repercussions [ABB^{+}46]. Reply [Bet84a, Bet87a, Bet91a, BGM91b, BGS95a, Sch34]. Report [ABB^{+}46, Bet75e, Bet95d, Duf46, LBC^{+}75, KHFA67, MW46, MW07]. Reports [KRW05, CCJ^{+}34]. Representatives [BBG^{+}85a]. reprints [Eks94]. repulsive [BG57, Ong86]. Res [Bet98a]. Research [Bet98a, RS82]. researcher [Sch15]. Resistance [Bet31a]. resolution [BK93]. Resonance [BP37]. respect [CWL^{+}04]. Response [Bet77d]. Responsibilities [BSP^{+}83]. Responsibility [Aft02, Fra01, Lan00, Olw01, Sar01, Sei01, Sch00b]. Restraint [Smi75]. Retirement [GSS^{+}03, AP77]. retrospect [Stu79]. Rev [RB37a]. Revealed [Bet90c, Usd09]. Reverse [DeN76]. Review [All56, Bet54b, Bet75e, Bet79a, Bet89b, Bet90b, Bet91b, Bet93b, Bet93a, Bet94a, Bet95a, Bet95b, Bet98b, Bet01a, Bet01b, Duf46, Dys55, Dys56, Ewa58, For57, For87, Fra01, Gol68, Got98, Har07, Hod13, Hub14, Huf82, Inc92, Jeh56, Kra13, Lan00, Low13, Mer68, Olw01, Pei57, Pri56, Sar01, Sch78, Sch68, Sch67b, Sei01, SBR^{+}86, Wal15a, Wal15b, vB92, Aft02, Tho99]. Reviews [Dow83, Edg91, Hen94, Nav14]. Revisited [BSS05, GBG^{+}86, Ino71, IIT78]. Revival [BW85, Bet90c]. Revolutionized [Hal17]. Rewriting [Bro82]. Richard [Sch68, Bet88d, Bet91g, BR93, Hal17]. Richardson [FCC09]. right [Ano50]. Ring [Usd09]. Risk [MB85, MB86, SB55b]. risque [SB55b]. Road [Inc92, vB92, Bet91j]. Robert [Ano50, Bet58a, Bet63b, Bet67d, BSK67, Bet68f, BW80, Bet90b, Bet91e, Bet97b, For87, Goo80, Goo81, Hen94, KW69, RB67a, Bro82]. Role [CB60, Ken54, CF12, FT70, Mai04]. Roma [CC73]. Rome [BB02, Bet04].

Roosevelt [Ano50, Ano50, Dem15, RMR^{+}50]. root [HSY04, Roj08]. roots [dGBNM02]. Rosenberg [Usd09]. Rosenfeld [Wal15a, Wal15b]. Rosetta [Sch78]. rotated [Zum75]. Rotblat [Ano92a]. rough [DeS73, Fur85]. Rounded [MYZ67]. Rounding [DF79]. Row [Pei57]. Royal [Hen94]. Rudi [Bet88a]. Rudolf [Lee09, Tho99, Wal15a, Wal15b, AP77, BGM91b]. Rudolph [AP77]. Ruijsenaars [HSY04]. Rule [CTSOS99]. Running [DHB^{+}06]. Russian [BZ38, Bet55a, Bet60b, Bet68c, Bet70a, BDH70, Bet76f, Mar70].

S [Bet98a, Dys56, Fra01, Hod13, Hub14, Jeh56, Kra13, Lan00, Nav14, Sar01, Sei01, Wal15a, Wal15b]. S. [Dys56, Fra01, Sar01]. Saale [HH14]. sadly [Mad83]. Safe [Bet77d, Gul75, YB85]. Safer [BT82]. Safety [Bet56d, Bet75d, Bla87, LBC^{+}75, WPBW76, Bet75e]. Sakharov [Bet90d, Bet91h, DK91]. Salpeter [BB71, HD72, Sch78, Aui02, BB69, BB70, Bas71, Bis67, Bre75, BT69, CFR82, FV86, Fle73, FT75, Flo76, Fur85, Gün64, HHD73, HD70, Kea66, Kea68, Kea69, Kea70, Kea71, Kyr72, Ler71, MP91, Mai03, Mai04, Nak63a, Nak63b, Nak66, Nak71, NR66, Ros66, Sae67, Sat63, Saz87, SL64, Tom83, Vos60, WA94, Zum75]. Salts [Bet30c, Van32]. Salzen [Bet30c]. Save [BG83b]. says [Bro82]. scalar [Nak71]. Scale [Ler71, Usd09]. Scare [McC20]. Scattering [BB51, BP35b, BRS38, Bet46a, BO46, Bet49a, BW51, BR52, Bet53a, Bet53b, BdH54a, BdH54b, Bet58b, Bet64c, Bet66b, BM71, BM72, Bet72b, Bet73c, Bet73a, Bet77f, BP97c, BBD^{+}96, BB54, CB48a, DeS73, DRS^{+}54, FT75, GB51, JB78, NR66, RYBB73, ROB63, RB39, SB61, SSB^{+}97a, SL64, Bet52b, Bet27, Bet28b, Bet55b, BJ76, BBD^{+}95, DSB56, Fur85, Gur01, ML83, Wie81, Bet32]. Scenario [BG82a, Bet94d, BB94a, BB94b, MWBB02]. Schachenmeier [Bet34c]. scheme [ZLGM03]. Schneider [HSY04]. schneller [Bet30b]. Schrieffer [Skr09]. Schrieffer-type [Skr09]. Schrödinger [PcPZ90, Wie81]. Schweber [Dys56, Fra01, Hod13, Hub14, Jeh56, Kra13, Nav14, Pei57, Sar01, Wal15a, Wal15b, Lan00, Sei01]. Sci [Bet98a]. Science [Ber01, Bet69c, BB86, Bet95d, Bet99b, Bet01b, GBG^{+}86, Hab69, Hil15, Kil77, McD62, RB02, Web73, Bet91j, Bra05, Fin06, LW71, Wei95, Orb05, Wal15a, Wal15b]. Sciences [Kra17, WH72, FT70]. Scientific [Eds68, Har01, RS08, Tho99, Wal15a, Wal15b, DP97, Haw11, Lee09, Bof76, McC20]. Scientist [ABB^{+}77, Ore04, Sch96, Sch05a, BFG^{+}85, Bet94g, Sch15, Sch00b, Fra01, Lan00, Sar01, Sei01, Aft02, Olw01]. Scientists [ACU^{+}54, Ano06, Ber01, BS47, BAA^{+}75, BWH80, BSP^{+}83, BGG^{+}85a, BLM19, Duf46, EBU^{+}52, GB89, Gou47, GR63, Seg85, SBR^{+}86, STS^{+}49, Tel87, Bri95a, Har10, Jun58, Jun82, Kil77, GR63]. Sco [BLLB00a, BLLB00b]. Scope [Sch15]. Screening [Bet34a]. SE [SBR^{+}86]. Sea [KS87]. Search [Bet01a, Cho01]. Seasons [Eds68]. Second [BM71, BM72, SG73, HM83, Stu18]. Second-Order [SG73]. Seconds [Ano05a]. secret [Ber96b, Fin06, WG17]. secrète [SB55b]. Secrets [BGS95b]. Section [BM54, DB52, DBM54, HB51a, HB40, MB52, Bet43b, BW87]. Sections [BBC56a, BBC56b, BBC57a, BBC57b, HB51b, ML83]. sector [Nak71]. Secure [SBR^{+}86, MB85]. Security [BG83a, Ken54]. seinem [KRW05]. Seiten [Sch78, Sch67b, Sch67a]. Select [ABB^{+}80, SBR^{+}86]. Selected [Bet47b, Bet97e, DP97, Sch58, Sch03, Yan13, BBL03, Lee09, Réd05, Got98, Tho99]. Selection [ABP98, Bet95b, CTSOS99, APB96]. Selective [GT38, WLB36]. Self [JB78]. Self-Consistency [JB78]. seltenen [Bet30c]. Semi [BN68a, NB68]. Semi-analytical [BN68a]. Semi-infinite [NB68]. Seminal [Bet95b, For87, BBL86]. seminar [Bet00c]. Senator [Ano50, Moy97]. SENSE [BWH80]. Sensitivity [BB47, WW88]. separates [MP91]. Separation [Gün64, Sco94]. September [Bec67, CC73, KS87]. septième [CCJ^{+}34]. Serial [Bet98a]. Series [BCC^{+}46]. Service [Sch68]. Set [CTSOS99]. Sets [CCG^{+}11]. Seventeenth [Sch68]. Seventh [CCJ^{+}34]. seventieth [BCSF82]. Shadow [Aft02, Bet66b, Fra01, Lan00, Olw01, Sar01, Sei01, Sch00b]. Shadows [Bet93a]. Shape [SB67a]. Shaping [Dre06]. Share [BGS95b]. shatterer [Goo80, Goo81]. Shell [BB52, AFF01]. Shielding [BW39]. Shift [BBF53, Bet47a, BBS50, BdH54a, BdH54b, dHMAB54, Bet55b, BJMK92]. Shifted [MWBB02]. Shifts [Bet53b]. Shine [Bar08]. Shock [BT41, BFvN^{+}44, BT45, BT51a, BT53, BAB80, BW85, Bet90c, Bet93g, Bet95e, Bet96a, Bet96b, Bet96c, Bet97d, CBB84, PPvNF44, SSBB75, BT51b, Bet98a]. Shocks [Bet82g]. shook [Haw11]. Short [Bet47b]. Shoulders [Bet95b]. Si [BH39]. side [Bet43d]. Sign [Bet53b]. Silicon [Bet42c, Bet42d]. Silvan [Hod13, Hub14, Pei57, Wal15a, Wal15b, Jeh56, Kra13, Nav14]. sily [Bet55a]. Simple [Bet99d, MJBA75, NB68]. Simplified [BBCW83]. Singapore [RS08]. single [BB76]. Singular [SL64, DSB56]. Sir [AP77, Bet98b, Hen94, Tho99, Lee09]. Situation [Bet75d]. Six [Bet69b, Bet77e, KZJ02]. six-vertex [KZJ02]. Sixteenth [Sch68, Sch68]. Sixteenth-Century [Sch68]. Sixty [FR13, Sal06]. Skaar [Wal15a, Wal15b]. skew [Skr07, Skr09]. Slow [Bet50a, WLB36]. Slowing [BTH50a, BTH50b, BTH50c]. Small [Bet44, BR52, SBSW60, Bet43c]. Smirnov [BB91a]. Smith [BKHS89]. Smyth [Edg91]. SN [BP90]. SN1987A [RV88]. sneered [Cor12]. Social [Wal15a, Wal15b]. Societal [Sal05]. Society [Bet55c, Bet75e, Bet99a, Hen94, Pan05, LBC^{+}75]. Sodium [vdLB47]. Soft [BLB01, Blu88]. Solar [BB90, BB93, BS05, BS06, Bet86a, Bet89a, BB91a, Bet92, Bet94c, Ran48, Kra16, Mad86, Men27]. Solar-Energy [Ran48]. Solar-Neutrino [BB90, Bet86a, Bet94c, Bet89a]. Solid [BK39, Bet80f, BM04, MA06, BM05]. Solid-State [Bet80f, BM04, MA06]. Solids [vdLB47]. soliton [WA94]. Solution [BB90, Bet85d, Kea69, Sat63, LFK99, Ong86, Ong88, Per84b]. Solutions [BB71, BK39, Nak63a, Vos60, BB70, CWL^{+}04, MIT06, Nak71, Zum75]. Solvable [ZLGM03, Eva84]. Solvay [CCJ^{+}34, CCJ^{+}34]. Solving [Mai03, Mai04]. Some [Cle48, HLB37, NS15]. Sommerfeld [Sch67b, Sch67a, Ano38, Bet00c, PS69]. Sooner [STS^{+}49]. Sound [KW69, Bet81e]. source [Bet49b, Kra16, Men27]. Sources [KHFA67, MB41, Bet68c]. sous [CCJ^{+}34]. Soviet [Bet94a, Bet60c, SBR^{+}86, STS^{+}49, WG17]. Soviets [Bet61, Bet82j, Bet83c, Cor12]. Space [Bet84a, BGGK84, BG85, BGGK12, Sae67, BG83b, GSB^{+}83]. Space-Based [BGGK84, BGGK12]. Speak [BAA^{+}75]. Special [Ano95, Bet99a, Got05, Kil77]. specialist [Sch15]. species [Ano85a]. Specific [Bet42a]. Spectra [MJBA75, Roj08, Roj09, RM09, Roj07]. Spectral [BFK49, ML83]. spectroscopy [CF12]. Spectrum [BS37, BBC56a, BBC56b, BBC57a, BBC57b, BBP63, RS05]. speculations [Kra16]. speech [Bet93e]. Speed [FBB14, Bet14]. spherical [Flo76]. Spin [Gün64, LC71, AS70, LFK99, Liu72, PcPZ90, RS05, dGBNM02]. Spin- [Gün64, AS70, PcPZ90]. Spin-Wave [LC71]. spinor [Bre75, Flo76, Nak71]. spinor-spinor [Bre75]. Spins [RB37a, RB37b]. Splitting [Bet49d, Bet50f, Bet29b, Ewa58]. sponsored [HPA97]. Springer [Sch67b, Sch67a]. Springer-Verlag [Sch67b, Sch67a]. Sputnik [Kil77]. Square [Kea69]. Square-Well [Kea69]. Sr [BB76]. Stability [Bet64a, Dys66]. Stalin [Bet94a]. Stalled [BW85]. standard [Ong88]. Stands [Bet95b]. Stanford [Edg91]. Stanley [For87]. Star [BBP71, BB95, BLZB01, GKB^{+}84, MJB75, BGG^{+}85a, BGG^{+}85b, Cor12, YB85]. Stardust [WGO^{+}91]. Stars [Anoxx, Bar08, Bet39c, Bet39a, Bet42b, Bet68e, Bet70d, Bet71b, Bet76c, BBAL79, Bet80e, Bet83a, BBC87a, BBFH57, Edd20, Hoy54, Jea24, MB40, Sal52, Bar88, Bet68g, Bet68d, Bet68a, Bet70a, Bet88b, BG91a, BL88, Eps88, Lee11, MR85, NaH88]. State [Bet29d, BS68, Bet74, BBAL79, BBAL80, Bet80f, BBCW83, BM04, Bis67, BB54, GSS^{+}03, KS87, MJB75, MA06, SB51, BJ76, Bet98a, BM05, Blu88, Bro88a, Mai03, Mai04, MB82, Ray97, ABB^{+}46]. statement [Ano85a]. States [BB36, DL83, WA94, Hol84]. stationary [Bet30a, BB36]. Statistical [Bet35f, Bet37b, Fur85]. statistics [Lai74]. Status [CC73, RYBB73]. Steering [Bet75e]. Stellar [BS05, BM39, Huf81, Huf06, MB41, Bet68c, Myr88, vW38a, vW38b]. Stellar-Energy [Huf81]. Steller [BS06]. step [MB85]. Stern [Dow83]. Sterne [vW38b, vW38a]. Sternen [Bet68a]. Steve [Ano12b]. Stilled [BBH^{+}73]. Stimulated [SSB^{+}97a]. STOC [ACM09]. Stony [GSS^{+}03, KS87]. Stopping [Bet34a, BH34, BBW50, Por98, Por04a, WB51, Por97, Por99, Por04b]. Stopping-Power [Por98]. Strange [BBC87a]. Strategic [Bet64a, Bet8xb]. Strategy [Bet62b, Bet77c, Bet80c, FvHW79]. Strengths [CTSOS99]. Streuung [Bet27, Bet28b]. string [MIT06]. Stroke [Bet95b]. Strong [Bet72a]. Strongly [SSB^{+}97c]. Structure [Ano37, BL49, CCJ^{+}34, Huf06, LR47, Hei34, HM83, CCJ^{+}34]. structures [Gur01]. Strutinsky [Bet71c]. Studies [ATP82, Por04b, ML83]. Study [Bet75e, Bet87a, RS82, Ray97, RS05, WPBW76, CBR87]. Stuff [BR93, Haw11]. submerged [Gur01]. Subramanyan [Bet05b]. substances [RCE30]. Successful [Ken54]. suggest [Mad83]. Suggested [ABB^{+}80]. Sum [CTSOS99]. Sums [MYZ67]. Sun [Bet66a, Bet95a, BLMB41, HHW99, Bet94f]. suns [Jun58, Jun82, Bet58a]. Super [Bet93j, SSB^{+}97b]. Superbomb [Edg91, GB89, Bro82, Yor89]. superbombe [SB55b]. superconductivity [BF33, Bet34c, Sch32, Sch34]. superiority [Bet81a]. Superlative [Sch05a]. Superlattices [Bet35f]. Supernova [BAB80, Bet81d, BYB82, Bet82g, Bet82h, BW85, Bet85b, Bet86b, BBC87b, Bet87b, Bet90c, Bet90a, Bet91a, Bet93g, Bet93h, Bet94e, Bet95e, Bet96a, Bet96b, Bet96c, Bet96d, Bet97d, BBB82, DSB91, BB85, BB91b, BB12]. Supernovae [Bet87c, Bet88c, Bet90e, Bet94b, CBB84, BBB^{+}87, BL88, Coo88, WM88, RS82]. support [Ano85a]. Supposed [BH39]. Supraleitung [BF33, Bet34c, Sch34, Sch32]. sure [Cor12, Hei34]. surface [DeS73]. surfaces [ML83]. Surveillance [BLM19]. survey [RS82]. symmetric [Skr07, Skr09, dGBNM02]. Symmetries [Bis67]. Symmetry [BN68b, AS70, LFK99, GSS^{+}03]. Symposium [ACM09, Bet55c, CC73, GSS^{+}03, HPA97, HH14, KS87, AP77, DS88, KMN^{+}96, Stu79, Ano37, Sch00a]. Synchrotron [CB48b]. Syndrome [DeN76]. Synthesis [BBFH57, Hoy54, Hoy46]. System [Kea68, Kra17, Wil73, AS70, CWB^{+}69, HSY04, MIT06, MP91]. Systems [Bro71, GB68, PB73, Kea71, Skr07, Skr09]. Szilard [Bet93a, Bet93a, SBSW60].

Tabelle [HH14]. Take [Huf81]. Talk [Bre53, Bri95a]. Tamm [DRS^{+}54]. target [Por03]. Targets [SB61]. taste [Mad83]. Taxes [vHWD76]. Teacher [Jac09]. Teachers [Bet93e, Ros93]. teaching [Bet93e, NS15]. Team [Ano93, CP93]. Team-B [CP93]. Technical [BMG87, Bro82]. Technological [Bet85d, Bet8xb, Bet85c]. Technologies [BBG85b, BG85]. Technology [Bet82i, Bet84b, LW71, Kil77]. telecast [Ano50]. Teller [Bet01b, Edg91, Bet01b, Bro82, Har10, RB02, Yor89]. Temperature [BLMB41, Aui02, BBR31, CSB^{+}31]. Temperature-Density [BLMB41]. tenable [Dud76]. Tennessee [Bec67]. Tensor [SB61]. tenu [CCJ^{+}34]. teorija [BZ38]. Term [Bet29b]. Termaufspaltung [Bet29b]. terms [Bet49d, Bet50f, Ewa58]. Terrible [Duf46]. Test [BB52, Bet60c, Ken54, McK86]. Testimony [Bet76g]. Tests [Bet56d, Bet60a, Bet61, BBG^{+}85a, BMG87, MWBB02, BT61]. Their [EBU^{+}52, HHW99, CFR82, FT70]. Theorem [Bet71c, Bet52b]. Theoretical [Bet37a, Bet50e, Tho99, AP77, Dir69, Dir70, Hei34, dSFvLH66]. Theorie [Bec29, BBT24, Bet30b, Bet30c, Bet31b, BF33, Bet34c, Sch32, Sch34, vW35, NBP^{+}34, Bet28a, Bet33d, Bet48]. Theories [Bro71, Gam66, Kai05b]. théoriques [Hei34]. Theory [ACM09, Bal62, BB51, Bet33c, BP35a, Bet35f, Bet35b, Bet37b, Bet39b, Bet40a, Bet40b, Bet40c, BN40, Bet42e, Bet42d, Bet42c, Bet43d, Bet44, BO46, Bet47b, Bet48, Bet49a, Bet51, Bet53a, BM54, Bet56c, Bet63c, Bet64d, Bet67f, BN68a, Bet68b, Bet71d, Bet73c, Bet73a, Bet80f, Bet82h, Bet85b, Bet86b, BBC87b, Bet87b, Bet87c, Bet88c, Bet94e, Bet94f, Bet96d, BBB82, BLW^{+}00a, Bro05b, Bro06a, CJK92, DBM54, Fer32, Ino71, KB38, Ler71, LC71, Neg05, Neg06, Pei57, ROB63, Tel66, Van32, BBR31, Bec29, BZ38, BBT24, Bet30b, Bet31b, BF33, Bet34c, Bet43b, BS43, BdH54c, BM56, BG57, BA59, Bet67c, BB91a, Bet98a, Bet99e, BM06, BLW^{+}03, CSB^{+}31, DSB56, Fur85, HBE87, Hou30, IIT78, NBP^{+}34, Por03, Sch32, Sch34, Sch67b, Sch67a, SB33, SB34, SB53, SB67e, vW35, Bet28a, Bet30c, Bet33d]. Theory [KMB57, Kut05b, Kut05a, For57, Pri56]. There [Bet67e, Bet85d]. Thermal [BT41, BT45, BT51a, BT53, BT51b, Per84b]. thermodynamic [AAHS99, BK93]. Thermonuclear [Bet52a, GT38, Ula66]. Thick [BFK49]. thinking [Sel77]. Thinks [Bet61]. Thirties [Bet79c, Sch05b, Sch06]. Thomas [Bet68b, MB40, NB68]. Thousand [Bet58a, WM15, Jun58, Jun82]. Threat [BGM91a, BGM91b, Duf46, Pei91a, Pei91b, SB55a]. Threatening [Iri06]. Three [Ada06, Bet65b, Bet67a, BR67, BSK67, Bet99d, RB67b, MB82, Tom83]. Three-Body [Bet65b, Bet67a, RB67b, BR67, Tom83]. Tight [RM06]. tightly [Kea71]. Time [Ano00b, Hal17, HB69]. times [Bet79a]. TiO [BB76]. titan [Bra05]. Tm [BS37]. Today [RMR^{+}50]. Together [Bet53c, Bet84c]. Too [Ger75]. Topics [Bet47b]. Toroidal [FSB71]. Towering [Iri06]. Track [ATP82]. Trade [Ber79a, Ber79b, Ber79c]. Trans [Bet54b]. Transcript [Hod81]. Transcripts [Bet93b]. Transfer [SB61]. Transformation [BK39, Wie81]. transformations [vW38a, vW38b]. transforms [CFR82]. Transients [BLB01]. Transition [BBP91, BJBP93, Run67, Cab78]. Transitions [LB50, BB76]. translated [Eks94]. Transmutation [Bet35a]. transport [Myr88]. Transverse [BV75]. treasury [FF91]. treaties [BG83b]. Treating [Bet38d]. treatment [Bet30a, Kea71]. Treaty [Bet99f]. tree [RR07, Roj07, Roj09]. trees [RJ10, RM06, RR07, Roj08, RM09, RJ11]. trends [Por04b]. Tribute [Orb05, DK91]. tributes [BSK67]. Trieste [BDH70, BBM^{+}76]. trillion [CP93]. Truth [Bet91i]. Twentieth [Bet01b, Wal15a, Wal15b, Har10]. Twentieth-Century [Bet01b]. twenty [LW71, MB85]. twenty-first [MB85]. twenty-five [LW71]. Two [Bet33b, Bet73b, Gün64, MB47, MJBA75, Sch78, Tho99, BF32, BS57a, BS57b, BS77, BS08, Mai03, Mai04, Saz87]. Two-Body [Bet73b, MJBA75, Mai03, Mai04]. Two-Electron [Bet33b, Sch78, BS57a, BS57b, BS77, BS08]. Two-Meson [MB47]. two-particle [Saz87]. Type [JLBF66, PR97, SB86, BL88, Skr09, BBB^{+}87, BYB82].

U [Bet98a, Hod13]. U.S [BBG^{+}85a, Sch68]. U.S. [Bet86c, SBR^{+}86]. Ultimate [Bet76h, Dud75]. Underground [All59]. Unequal [Bre75]. Unified [KS87]. Union [Bet94a]. United [Hol84]. Universe [SSB^{+}99b, Dys79, Bet79a]. University [Ano37, DS88, Edg91, Fra01, GSS^{+}03, HPA97, Hen94, Hub14, KS87, Nav14, Wal15a, Wal15b]. Unlikely [Ber10]. unratified [BG83b]. Unravelling [FR13]. Unreal [Sel77]. unreasonable [Hol03]. upper [RR07]. Uranium [ATP82, Bet82i, Bet00a, Bet05a, Fes77, Ber96b]. Uranprojekt [Bet05a]. urgent [Bet83b]. US$8.95 [Sch78]. USA [ACM09, Ano37]. USC [BBF^{+}82]. Use [ABB^{+}50, HM83]. Usefulness [SB61]. Using [CCG^{+}11, PR97].

v [Bet70a, BDH70]. V. [dSFvLH66]. Validity [Fun81]. Value [BL49, BBS50, BL82]. values [Por04b]. Vannevar [Ano85b]. Vanunu [AAB^{+}88]. Variables [Gün64, Sco94]. Variational [PB73]. Variations [Van32]. ve [Bet71b]. Vechernie [BDH70]. Vector [Nak71]. Vector-scalar [Nak71]. vectors [KM01]. Velocity [Bet32]. Vergleich [Bet29d]. verification [Bet8xa]. Verlag [Sch67b, Sch67a]. verschiedenen [Bet29d]. Versuche [BBT24]. Versus [Duf46]. Vertex [Nak66, AFF01, KZJ02]. Very [BBAL80, Hoy54]. VI [Bet96c]. Victor [Bet89b, Bet91b]. View [Sul75, Bet79a, Bet94g]. viewpoints [Bet91j]. Views [Bet69b, Bet77e, Bet84b, EBU^{+}52, HBD71]. viii [Nav14, Tho99, Wal15a, Wal15b, Bet97d]. Viki [Bet93d]. Visa [EBU^{+}52]. Voice [BBH^{+}73]. Vol [Bet54b, Dys56, Pei57, Pei57]. Volterra [FV86]. Volume [CJK92, Dys55, Jeh56, FK85, MKR87]. vom [HH14]. Vorstudien [Sch32]. vs [Bet69b].

W [BKJ84, Mer68]. Walk [Web73]. wall [KZJ02]. Want [Bet61]. War [AK10, BG82b, GG05, MB85, MB86, Pol19]. Wars [BGG^{+}85a, BGG^{+}85b, Cor12, GKB^{+}84, Stu18, YB85]. Warsaw [Ano82]. Wartime [STS^{+}49]. Was [Bet67e, Bet90c, Mad83]. Washington [Bet55c, HPA97, KMP98]. Wasserstoff [Bet48]. Wasserstoff-Eigenfunktionen [Bet48]. Wasserstoffbombe [SB55a]. Wasserstoffs [Bet29a]. waste [Bet84d]. Water [Bet46b, Bet75e, Bet20, LBC^{+}75]. Wave [BFH^{+}47, LC71, Bet43b, Bet43d, Fle73, Ong86, Ong88, PcPZ90, Sch32]. Wave-mechanical [Sch32]. Wavefunction [Sat63, Gau83, Gau14, Tom83, Sch15]. Waves [BT41, BFvN^{+}44, BT45, BT51a, BT53, BBD^{+}96, BB54, SSBB75, BT51b, Bet98a, Liu72, PPvNF44]. Way [Duf46]. Ways [BG82b]. Weaponry [Bet84a]. Weapons [BG85, GSB^{+}83, BG83b, Bet91j]. Wechselwirkung [BF32, BF33]. weekly [Ano50]. Weeks [Ada06]. weighted [Roj08]. Weiss [Cab78]. Weisskopf [Bet89b, Bet91b, Bet93d, dSFvLH66]. Weizsäcker [HH14, Wie14, Sch14a, WG68, Wie14]. Weizsäcker-Zyklus [Wie14]. Welfare [Sch68]. Well [Kea69]. Wellenmechanische [Sch32]. were [Ano50]. weren’t [Cor12]. Werner [HBD71, KRW05]. Wheeler [Hal17]. Where [Bet67e, Mad83]. White [BLZB01]. White-Dwarf [BLZB01]. Whittlesey [Duf46]. Who [Duf46, Dem15, Dem15]. whole [SB67b]. Whose [Bet95b]. Wick [MP91, Zum75]. Wick-rotated [Zum75]. Wie [Bet01d]. Wiesner [SBSW60]. Wiley [Gol68]. Will [BT82]. William [Bet97c, BCSF82, Bet93a]. winding [CWL^{+}04]. windows [Bet43a, Bet43d]. Windsurfing [KS87]. Winners [Bet63d]. Wissenschaft [Wei95]. within [FSB71]. without [Sal52]. Wolfenstein [BB91a]. Wolfgang [Bet00b, Bet00b, Uns00]. women [HHW99]. Won’t [BGG^{+}85b]. Work [BGG^{+}85b, KMN^{+}96, Tel55, Bro82, Hou30]. Worked [Bet93a]. Works [Got98, Bet97e, Tho99]. World [Duf46, GR63, RS08, Stu18, Tho99, Wal15a, Wal15b, WM15, Bet81e, BM66, FF91, Haw11, Hol03, LW71, MW46, MW47, MW07, SB55b, She75, SBSW60]. worlds [Goo80, Goo81]. Would [Bro97]. Writing [Sch96, Sch14b]. Writings [ABP98, APB96]. wrong [Tur08].

x [Duf46, BFK49, BN68a, BLB99, BHL^{+}01, BLB01, CJK92, HM83]. X-Ray [BHL^{+}01, BLB01, BN68a, BLB99, CJK92, HM83]. X-Rays [BFK49]. xi [Hen94, Wal15a, Wal15b]. xii [Gol68, Sch68, Wal15a, Wal15b]. xiii [Pei57]. xiv [Edg91, Wal15a, Wal15b]. xlix [Hen94]. xvi [Edg91, Pei57]. xviii [Fra01]. xxiii [Tho99]. XXZ [dGBNM02].

Yalta [Bet71e]. Yang [GSS^{+}03, Fun81]. Year [BG82b, FR13, BBM78]. Years [Bet95b, Mla98, Sal06, Geh78, Hoi01, LW71]. Yield [LB52]. York [Bet54b, Bet89b, Bet90b, Duf46, Edg91, Gol68, GSS^{+}03, KS87, Sch78, Sch67b, Sch67a]. York-London [Bet54b]. Yorker [Ber06].

Zamolodchikov [Gar02, Ray97, RS05]. Zeeman [Bet30c]. Zeemaneffektes [Bet30c]. Zeitfaktor [HB69]. Zeitschrift [FBB14]. Zemle [Bet70a]. Zero [Kea68, BBR31, CSB^{+}31]. Zero-Mass [Kea68]. zervikaler [HB69]. Zhizn [BDH70]. Zirconium [Pal76]. ZnSe [Por98]. zum [HBD71, Hof96]. Zur [BF33, BBR31, BBT24, Bet30b, Bet30c, Bet31b, Bet34c, Sch34, vW35]. zusammenhängenden [Bet33a]. Zuschriften [CSB^{+}31]. zvezd [Bet68c]. zvezdah [Bet70a]. zwei [BF32, Bet33b]. Zwei-Elektronenprobleme [Bet33b]. Zyklus [Wie14].

References

Alsop:1954:WA

[AA54] Joseph Alsop and Stewart Alsop. We accuse! Harper’s Magazine, ??(??):25–45, October 1954. CODEN HAMAA3. ISSN 1045-7143.

           URL http://harpers.org/archive/1954/10/we-accuse/.

Alfven:1988:CMV

[AAB^{+}88] Hannes Alfven, Edoardo Amaldi, Paul Beeson, Hans A. Bethe, Owen Chamberlain, et al. The case of Mordechai Vanunu. New York Review of Books, 33(10):56, June 16, 1988. ISSN 0028-7504 (print), 1944-7744 (electronic). URL http://www.

nybooks.com/articles/archives/1988/jun/16/the-case-of-

mordechai-vanunu/. Appeal by twenty-seven scientists on behalf of Mordechai Vanunu, the Israeli nuclear technician who, on March 27, 1988, was sentenced by a military tribunal to eighteen years in prison for having made public information about Israel’s nuclear capacity.

Ambjorn:1999:BAT

[AAHS99] J. Ambjørn, A. Avakyan, T. Hakobyan, and A. Sedrakyan. Bethe Ansatz and thermodynamic limit of affine quantum group invariant extensions of the t – J model. Journal of Mathematical Physics, 40 (11):5687–5701, November 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Alvarez:1946:RSD

[ABB^{+}46] L. W. Alvarez, R. F. Bacher, M. Benedict, H. A. Bethe, A. H. Compton, Farrington Daniels, J. R. Oppenheimer, J. R. Ruhoff, G. T. Seaborg, S. H. Spedding, C. A. Thomas, and W. H. Zinn. Repercussions of the State Dep’t report: Denaturing fissionable materials. Bulletin of the Atomic Scientists, 1(9):11, April 15, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Allison:1950:LUP

[ABB^{+}50] Samuel K. Allison, Kenneth T. Bainbridge, Hans Bethe, Robert B. Brode, Charles C. Lauritsen, F. Wheeler Loomis, George B. Pegram, Bruno Rossi, Frederick Seitz, Merle A. Tuve, Victor F. Weisskopf, and Milton G. White. Let us pledge not to use H-bomb first! Bulletin of the Atomic Scientists, 6(3):75, March 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Adler:1977:IAS

[ABB^{+}77] Kraig Adler, Simon H. Bauer, Hans A. Bethe, Thomas Eisner, Roger H. Farrell, Michael E. Fisher, Peter J. Gierasch, Kenneth Greisen, Donald F. Holcomb, Peter J. Kahn, Harry Kesten, Jack Kiefer, James A. Krumhansl, Stephen Lichtenbaum, G. R. Livesay, Franklin A. Long, Thomas R. Podleski, Carl Sagan, Harold A. Scheraga, Frank Spitzer, Moss E. Sweedler, Yervant Terzian, James E. West, and Benjamin Widom. Imprisoned Argentine scientist. Science, 197(4307):938, September 2, 1977. CODEN SCI-EAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http:

//www.sciencemag.org/content/197/4307/938.2.full.pdf.

Anderson:1980:PEB

[ABB^{+}80] R. O. Anderson, D. Bell, H. Bethe, N. Glazer, T. M. Hesburgh, V. Jordan, H. H. Kay, E. I. Koch, R. W. Lyman, F. F. Piven, N. Podhoretz, D. Shalala, W. Simon, R. E. Tyrrell, and C. Yost. For the President-Elect + books suggested for Reagan Administration by select group of Americans. New York Times Book Review, 85(48):9–??, ???? 1980. ISSN 0028-7806.

Alpher:1948:OCE

[ABG48] Ralph A. Alpher, Hans Bethe, and George Gamow. The origin of chemical elements. Physical Review, 73(7):803–804, April 1, 1948. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://prola.aps.org/abstract/PR/

v73/i7/p803_1. See historical retrospective [Tur08] and comments [AWCT09].

Amaldi:1998:CPE

[ABP98] Edoardo Amaldi, Giovanni Battimelli, and Giovanni Paoloni, editors. 20th Century Physics: Essays and Recollections: a Selection of Historical Writings, volume 3 of Edoardo Amaldi Foundation series. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1998. ISBN 981-02-2369-2. xiii + 747 pp.

LCCN Q127.I8 A53 1998. URL http://www.worldscientific.
com/worldscibooks/10.1142/2852.

ACM:2009:SPA

[ACM09] ACM, editor. STOC ’09: proceedings of the 2009 ACM International Symposium on Theory of Computing, Bethesda, Maryland, USA, May 31–June 2, 2009. ACM Press, New York, NY, USA, 2009. ISBN 1-60558-613-7. LCCN QA75.5 .A22 2009.

Allison:1954:SAF

[ACU^{+}54] Samuel K. Allison, Edward Condon, Harold C. Urey, Farrington Daniels, F. W. Loomis, Edward Shils, Linus Pauling, S. A. Goudsmit, Julian Schwinger, Albert Einstein, Hugh C. Wolfe, Cyril S. Smith, Victor F. Weisskopf, and Hans Bethe. Scientists affirm faith in Oppenheimer. Bulletin of the Atomic Scientists, 10 (5):188–190, May 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Adami:2006:TWH

[Ada06] Christoph Adami. Three weeks with Hans Bethe. In Brown and Lee [BL06a], pages 45–110. ISBN 981-256-609-0, 981-256-610-4 (paperback), 981-277-450-5 (e-book). LCCN QC16.B46 H36 2006.

URL http://www.worldscibooks.com/physics/5989.html;
http://www.worldscientific.com/worldscibooks/10.1142/5989.

Amico:2001:BMS

[AFF01] Luigi Amico, G. Falci, and Rosario Fazio. The BCS model and the off-shell Bethe ansatz for vertex models. Journal of Physics A (Mathematical and General), 34(33):6425–6434, 2001. CODEN JPHAC5. ISSN 0305-4470 (print), 1361-6447 (electronic).

Aftergood:2002:SBO

[Aft02] Steven Aftergood. In the Shadow of the Bomb: Oppenheimer, Bethe, and the Moral Responsibility of the Scientist (review). Journal of Cold War Studies, 4(1):95–96, Winter 2002. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (elec-

tronic). URL http://muse.jhu.edu/article/9218;   http:/
/muse.jhu.edu/article/9218/pdf;   http://muse.jhu.edu/
journals/journal_of_cold_war_studies/v004/4.1aftergood.
html.

Alpher:1973:ABG

[AH73] R. A. Alpher and R. Herman. Alpher, Bethe and Gamow. In Weber [Web73], page ?? ISBN 0-85498-027-X. LCCN Q167.W42. Compiled by R. L. Weber, edited by E. Mendoza, with a foreword by William Cooper.

Altschuler:2010:RCC

[AK10] Glenn Altschuler and Isaac Kramnick. Red Cornell: Cornell in the Cold War (Part I). Cornell Alumni Magazine, 113(1):??, July

8, 2010. URL http://cornellalumnimagazine.com/index.php?
option=com_content&task=view&id=808&Itemid=9. Discusses

role of Hans Bethe in a committee affiliated with the Federation of American Scientists that was formed to “to investigate the spy investigators” during the ‘Red Scares’ of the late 1940s.

Alexander:2005:ENN

[Ale05] Lamar Alexander. Editorial: Nurturing the next Einsteins. Science, 307(5712):1013, February 18, 2005. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/307/5712/1013.full.pdf.

Allan:1956:RBH

[All56] H. R. Allan. Review of The Hydrogen Bomb: The Men, The Menace, the Mechanism. International Affairs (Royal Institute of International Affairs 1944–), 32(1):132–133, ???? 1956. ISSN 0020-5850 (print), 1468-2346 (electronic). URL http://www.jstor.

org/stable/2607946.

Allison:1959:BDU

[All59] Helen C. Allison. Bethe on detection of underground explosions. Bulletin of the Atomic Scientists, 15(6):257, 259, June 1959. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Antonio:2011:ABA

[AMS11] N. Cirilo António, N. Manojlović, and A. Stolin. Algebraic Bethe ansatz for deformed Gaudin model. Journal of Mathematical Physics, 52(10):103501, October 2011. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:

//jmp.aip.org/resource/1/jmapaq/v52/i10/p103501_s1.

Anonymous:1937:SSM

[Ano37] Anonymous, editor. Symposium on the Structure of Metallic Phases, Cornell University, Ithaca, NY, USA, 1–3 July 1937. Cornell University, Ithaca, NY, USA, 1937. LCCN ????

Anonymous:1938:PAS

[Ano38] Anonymous. Professor Arnold Sommerfeld. Journal of Applied Physics, 9(12):754–755, December 1938. CODEN JAPIAU. ISSN 0021-8979 (print), 1089-7550 (electronic), 1520-8850. Note for the 70th birthday of Arnold Sommerfeld.

Anonymous:1950:FTE

[Ano50] Anonymous. First telecast of Eleanor Roosevelt’s weekly forum on February 11, 1950. Participating in a discussion of “what to do with the hydrogen bomb” were, left to right: Senator Brien McMahon, Hans A. Bethe, Mrs. Roosevelt, David E. Lilienthal and J. Robert Oppenheimer, February 11, 1950. URL http://hdl.loc.gov/ loc.pnp/cph.3b3842. Photograph by Leonard McCombe.

Anonymous:1961:EFA

[Ano61] Anonymous. Enrico Fermi Award [to Hans A. Bethe]. Physics Today, 14(10):74, October 1961. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.physicstoday.

org/resource/1/phtoad/v14/i10/p74_s1.

Anonymous:1982:WWN

[Ano82] Anonymous. Warsaw: We now appeal. Bulletin of the Atomic Scientists, 38(9):3–4, November 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1985:SPS

[Ano85a] Anonymous. For the species and the planet: a statement in support of the five continent peace initiative. Bulletin of the Atomic Scientists, 41(10):4–5, November 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1985:HBR

[Ano85b] Anonymous. Hans A. Bethe receives Vannevar Bush Award. Bulletin of the Atomic Scientists, 41(6):55, June/July 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1992:EPP

[Ano92a] Anonymous. Einstein Peace Prize awarded: [to Hans Bethe and Joseph Rotblat]. Bulletin of the Atomic Scientists, 48(10):5, December 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1992:MP

[Ano92b] Anonymous. Men of peace. New Scientist, ??(??):??, July 11, 1992. CODEN NWSCAL. ISSN 0262-4079 (print), 1364-8500

(electronic). URL https://www.newscientist.com/article/
mg13518292-400-men-of-peace/. Brief report of the award of

the 1992 Einstein Peace Prize (US$50,000) to Joseph Rotblat and Hans Bethe.

Anonymous:1993:CGB

[Ano93] Anonymous. Corrections: [German bomb program and Team B]. Bulletin of the Atomic Scientists, 49(4):6, May 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bet93c, CP93].

Anonymous:1995:SIA

[Ano95] Anonymous. Special investigation: The atomic intrigues of Niels Bohr. Scientific American, 272(5):83, May 1995. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.nature.com/scientificamerican/
journal/v272/n5/pdf/scientificamerican0595-83.pdf.

Anonymous:2000:GHB

[Ano00a] Anonymous. Guide to the Hans Bethe papers, ca. 1931–1995. Cornell University Library archives., 2000. URL http://rmc.

library.cornell.edu/EAD/htmldocs/RMA00976.html.

Anonymous:2000:MT

[Ano00b] Anonymous. The moment in time. [Los Alamos National] Laboratory produced documentary on the Manhattan Project., April

20, 2000. URL http://www.governmentattic.org/11docs/
NNSAfilmVideoList_2014.pdf;   https://www.youtube.com/

watch?v=xwpgmEvlRpM. The hour-long program features interviews with some of those who worked on the Project Hans Bethe, Norman Ramsey, Phil Morrison, Edward Teller and others. The program profiles the creation of the Manhattan Project the obstacles that were faced, life in the “secret city,’ and events that led up to and followed the Trinity Test.

Anonymous:2001:IHB

[Ano01] Anonymous. Interview with Hans A. Bethe. Web document,

2001. URL http://oralhistories.library.caltech.edu/96/
;   http://resolver.caltech.edu/CaltechOH:OH_Bethe_H.

Anonymous:2005:SNP

[Ano05a] Anonymous. 60 seconds: Nuke pioneer [Hans Bethe] dies. New Scientist, ??(??):??, March 9, 2005. CODEN NWSCAL. ISSN 0262-4079 (print), 1364-8500 (electronic). URL https://www.

newscientist.com/article/mg18524904-200-60-seconds/.

Anonymous:2005:NLH

[Ano05b] Anonymous. Nobel Laureate Hans Bethe passes away at age of 98. Web news story., March 7, 2005. URL http://en.wikinews.

org/wiki/Nobel_Laureate_Hans_Bethe_passes_away_at_age_
of_98.

Anonymous:2005:PEH

[Ano05c] Anonymous. People & events: Hans A. Bethe, (1906–2005). PBS

broadcast., 2005. URL http://www.pbs.org/wgbh/amex/bomb/
peopleevents/pandeAMEX50.html.

Anonymous:2006:ESR

[Ano06] Anonymous. Education: Scientists on the record. Science, 312 (5775):821, May 12, 2006. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/312/5775/821.4.full.pdf.

Anonymous:2012:FHB

[Ano12a] Anonymous. FAS Hans Bethe Award. Web page., August 5, 2012.

URL http://www.fas.org/about/awards.html.

Anonymous:2012:HBA

[Ano12b] Anonymous. Hans Bethe Award to Steve Fetter. Web page.,

August 5, 2012. URL http://www.fas.org/press/events/
2005nov30_hansbetheaward.html.

Anonymous:2012:NPP

[Ano12c] Anonymous. The Nobel Prize in Physics 1967: Hans Bethe. Web

site., 2012. URL http://www.nobelprize.org/nobel_prizes/
physics/laureates/1967/.

Anonymous:20xx:HBE

[Anoxx] Anonymous. Hans Bethe, energy production in stars, and nuclear physics. Web site, 20xx. URL http://www.osti.gov/

accomplishments/bethe.html.

Aitchison:1977:RPT

[AP77] Ian Johnston Rhind Aitchison and J. E. Paton, editors. Rudolf Peierls and theoretical physics: proceedings of the symposium held in Oxford on July 11th and 12th, 1974, to mark the occasion of the retirement of Professor Sir Rudolph E. Peierls, F.R.S., C.B.E., volume 13 of Progress in nuclear physics. Pergamon, New York, NY, USA, 1977. ISBN 0-08-020606-9. LCCN QC1 .P3 1974.

Amaldi:1996:ERC

[APB96] Edoardo Amaldi, Giovanni Paoloni, and Giovanni Battimelli, editors. Essays and recollections on 20th century physics: a selection of historical writings, volume 2 of Edoardo Amaldi Foundation series. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1996. ISBN 981-02-2369-2. xiii + 747 pp. LCCN Q127.I8 A53 1996.

Applegate:1988:NDP

[App88] James H. Applegate. Neutron diffusion, primordial nucleosynthesis, and the r-process. Physics Reports, 163(1–3):141–154, 1988. CODEN PRPLCM. ISSN 0370-1573 (print), 1873-6270

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0370157388900415.

Abo-Shaeer:1970:SBE

[AS70] Muhsin A. R. Abo-Shaeer. O(4) symmetry and the BS equation for a spin-1 + spin-0 system. Journal of Mathematical Physics, 11 (8):2272–2279, August 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v11/i8/p2272_s1.

Ahlen:1982:SRU

[ATP82] S. P. Ahlen, G. Tarlé, and P. B. Price. Studies of relativistic uranium nuclei with dielectric track detectors. Science, 217 (4565):1139–1140, September 17, 1982. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/217/4565/1139.full.pdf.

Auil:2002:FPD

[Aui02] F. Auil. Four-particle decay of the Bethe–Salpeter kernel in the high-temperature Ising model. Journal of Mathematical Physics, 43(12):6209–6223, December 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Amett:2009:HHN

[AWCT09] David Amett, George Wallerstein, Ken Croswell, and Michael S. Turner. αβγ, Hoyle, and the history of nucleosynthesis. Physics Today, 62(5):10–11, May 2009. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.

physicstoday.org/resource/1/phtoad/v62/i5/p10_s1. See

[ABG48, Tur08].

Bethe:1950:FAHb

[B^{+}50] Hans A. Bethe et al. The Facts about the Hydrogen Bomb. University of Chicago Press, Chicago, IL, USA, 1950. ???? pp. LCCN ????

Bethe:1962:PD

[B^{+}62] Hans A. (Hans Albrecht) Bethe et al. Problems of disarmament, volume 19 of Monthly review pamphlet series. Monthly Review Press, New York, NY, USA, 1962. 31 pp. LCCN UA11 .C6.

Bethe:1952:ADP

[BA52] H. A. Bethe and N. Austern. Angular distribution of π^{+} production in n–p collisions. Physical Review, 86(1):121–122, April 1, 1952. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://prola.aps.org/abstract/PR/
v86/i1/p121_2.

Bethe:1953:PIP

[BA53] Hans A. Bethe and Julius Ashkin. Part II. Passage of radiations through matter. In Segrè [Seg59a], pages 166–357. LCCN QC173

.S313 1953. URL https://en.wikipedia.org/wiki/Julius_
Ashkin.

Bethe:1959:TAG

[BA59] Hans A. Bethe and Mac C. Adams. A theory for the ablation of glassy materials. Journal of the Aero/Space Sciences, 26(6): 321–328, 350, 1959. ISSN 0095-9820.

Bethe:1975:SSN

[BAA^{+}75] Hans A. Bethe, Luis Alvarez, Peter Auer, William O. Baker, John Bardeen, Robert F. Bacher, Felix Bloch, Norris E. Bradbury, Harold Brown, Richard H. Chamberlain, Cyril L. Comar, Arthur Kantrowitz, Ralph E. Lapp, Joshua Lederberg, Willard F. Libby, Franklin A. Long, Edwin M. McMillan, Kenneth S. Pitzer, Edward M. Purcell, I. I. Rabi, Norman Rasmussen, Roger Revelle, Glenn T. Seaborg, Frederick Seitz, Edward Teller, James A. Van Allen, Warren Weaver, Alvin Weinberg, Victor F. Weisskopf, Edward Wenk, Jr., Eugene Wigner, and Richard Wilson. 32 scientists speak out: “no alternative to nuclear power”. Bulletin of the Atomic Scientists, 31(3):4–5, March 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See correction [Wei75] about the incorrect inclusion of Victor Weisskopf in the authors. See comments [Gir75, Smi75, Sau75, Gul75].

Bethe:1980:NES

[BAB80] H. A. Bethe, J. H. Applegate, and G. E. Brown. Neutrino emission from a supernova shock. Astrophysical Journal, 241(??):343–354, ???? 1980. CODEN ASJOAB. ISSN 0004-637X (print), 1538-4357 (electronic).

Bacher:1950:HBI

[Bac50] Robert F. Bacher. The hydrogen bomb: III. Scientific American, 182(5):11–15, May 1950. CODEN SCAMAC.

ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.nature.com/scientificamerican/journal/v182/n5/pdf/
scientificamerican0550-11.pdf. See also Parts I [Rid50] and

II [Bet50d].

Ballhausen:1962:ILF

[Bal62] Carl J. Ballhausen. Introduction to Ligand Field Theory. McGraw-Hill, New York, NY, USA, 1962. ix + 298 pp. LCCN QD475 .B3.

Baron:1988:CCS

[Bar88] E. Baron. Collapse calculations in stars of mass less than 15M_{sun}. Physics Reports, 163(1–3):37–46, 1988. CODEN PRPLCM. ISSN 0370-1573 (print), 1873-6270 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0370157388900336.

Bardi:2008:FLW

[Bar08] Jason Socrates Bardi. Focus: Landmarks: What makes the stars shine? Physical Review Focus, 21(??):3–??, January 23, 2008. CODEN PRFHAQ. ISSN 1539-0748. URL http://physics.aps.

org/story/v21/st3.

Basu:1971:FEH

[Bas71] Debabrata Basu. Formal equivalence of the hydrogen atom and harmonic oscillator and factorization of the Bethe–Salpeter equations. Journal of Mathematical Physics, 12(8):1474–1482, August 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v12/i8/p1474_s1.

Bethe:1936:NPS

[BB36] H. A. Bethe and R. F. Bacher. Nuclear physics A. Stationary states of nuclei. Reviews of Modern Physics, 8(2):82–229, April 1, 1936. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.8.82;   http://rmp.aps.org/abstract/RMP/v8/
i2/p82_1.

Barschall:1947:ESF

[BB47] H. H. Barschall and H. A. Bethe. Energy sensitivity of fast neutron counters. Review of Scientific Instruments, 18(??):147–149, 1947. CODEN RSINAK. ISSN 1089-7623, 0034-6748.

Beard:1951:FCN

[BB51] D. B. Beard and H. A. Bethe. Field corrections to neutron–proton scattering in a new mixed meson theory. Physical Review (2), 83(6):1106–1114, September 15, 1951. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.

aps.org/abstract/PR/v83/i6/p1106_1.

Bethe:1952:PTN

[BB52] H. A. Bethe and S. T. Butler. A proposed test of the nuclear shell model. Physical Review, 85(6):1045–1046, March 15, 1952. CO-DEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

URL http://prola.aps.org/abstract/PR/v85/i6/p1045_2.

Breit:1954:IWF

[BB54] G. Breit and H. A. Bethe. Ingoing waves in final state of scattering problems. Physical Review (2), 93(4):888–890, February 15, 1954. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://prola.aps.org/abstract/PR/
v93/i4/p888_1.

Basu:1969:ADB

[BB69] Debabrata Basu and S. N. Biswas. Accidental degeneracy in the Bethe–Salpeter equation. Journal of Mathematical Physics, 10 (11):2104–2113, November 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v10/i11/p2104_s1.

Basu:1970:HAS

[BB70] Debabrata Basu and S. N. Biswas. O(5) harmonics and abnormal solutions in the Bethe–Salpeter equation. Journal of Mathematical Physics, 11(4):1204–1209, April 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v11/i4/p1204_s1. See erra-

tum [BB71].

Basu:1971:EHA

[BB71] Debabrata Basu and S. N. Biswas. Erratum: “O(5) Harmonics and Abnormal Solutions in the Bethe–Salpeter Equation”. Journal of Mathematical Physics, 12(7):1446, July 1971. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v12/i7/p1446_

s2. See [BB70].

Benguigui:1976:DPT

[BB76] L. Benguigui and K. Bethe. Diffused phase transitions in Ba_{x}Sr_{1−x}TiO_{3} single crystals. Journal of Applied Physics, 47(??): 2787–??, 1976. CODEN JAPIAU. ISSN 0021-8979 (print), 1089-7550 (electronic), 1520-8850.

Bethe:1985:HSE

[BB85] Hans A. Bethe and Gerald Brown. How a supernova explodes. Scientific American, 252(5):40–48, May 1985. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). See summary in [BB12].

Bethe:1986:BSA

[BB86] H. A. Bethe and John Bardeen. Back to science advisers. New York Times, page 27, May 17, 1986. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Bahcall:1990:SSN

[BB90] John N. Bahcall and H. A. Bethe. A solution of the solar-neutrino problem. Physical Review Letters, 65(18):2233–2235, October 29, 1990. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Bethe:1991:SNM

[BB91a] Hans A. Bethe and John N. Bahcall. Solar neutrinos and the Mikheyev–Smirnov–Wolfenstein theory. Physical Review D (Particles and Fields), 44(10):2962–2969, November 15, 1991. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM).

Bethe:1991:HSE

[BB91b] Hans A. Bethe and Gerald Brown. How a supernova explodes. In Ferris and Fadiman [FF91], pages 277–291. ISBN 0-316-28129-8. LCCN QC71 .W67 1991. Foreword by Clifton Fadiman.

Bahcall:1993:DSN

[BB93] John N. Bahcall and Hans A. Bethe. Do solar neutrino experiments imply new physics? Physical Review D (Particles and Fields), 47 (4):1298–1301, February 15, 1993. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM).

Brown:1994:SLNa

[BB94a] G. E. Brown and H. A. Bethe. A scenario for a large number of low-mass black-holes in the galaxy. Abstracts of Papers of the American Chemical Society, 207(??):31–??, March 13, 1994. CO-DEN ACSRAL. ISSN 0065-7727.

Brown:1994:SLNb

[BB94b] G. E. Brown and H. A. Bethe. A scenario for a large number of low-mass black-holes in the galaxy. Astrophysical Journal, 423(2): 659–664, ???? 1994. CODEN ASJOAB. ISSN 0004-637x (print), 1538-4357 (electronic).

Bethe:1995:OCM

[BB95] H. A. Bethe and G. E. Brown. Observational constraints on the maximum neutron star mass. Astrophysical Journal, 445(2):L129– L132, ???? 1995. CODEN ASJOAB. ISSN 0004-637X (print), 1538-4357 (electronic).

Bethe:1998:EBC

[BB98] Hans A. Bethe and G. E. Brown. Evolution of binary compact objects that merge. Astrophysical Journal, 506(2):780–789, ???? 1998. CODEN ASJOAB. ISSN 0004-637X (print), 1538-4357 (electronic).

Bethe:1999:CHM

[BB99] Hans A. Bethe and G. E. Brown. Contribution of high-mass black holes to mergers of compact binaries. Astrophysical Journal, 517 (1):318–327, ???? 1999. CODEN ASJOAB. ISSN 0004-637X (print), 1538-4357 (electronic).

Bethe:2002:EFR

[BB02] Hans A. Bethe and Henry Bethe. Enrico Fermi in Rome, 1931– 32. Physics Today, 55(6):28–29, June 2002. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.

physicstoday.org/resource/1/phtoad/v55/i6/p28_s1.

Bethe:2012:HSE

[BB12] H. A. Bethe and G. Brown. How a supernova explodes. Scientific American, 307(1):66–67, July 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). Summary of [BB85].

Bethe:1979:ESG

[BBAL79] H. A. Bethe, G. E. Brown, J. Applegate, and J. M. Lattimer. Equation of state in the gravitational collapse of stars. Nuclear Physics A, 324(2–3):487–533, July/August 1979. CODEN NUPABL. ISSN 0375-9474 (print), 1873-1554 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0375947479905967.

Bethe:1980:AES

[BBAL80] H. A. Bethe, G. E. Brown, J. Applegate, and J. M. Lattimer. Appendix: Equation of state of a very hot gas of electrons and neutrinos. Astrophysical Journal, 241(??):350–354, ???? 1980. CODEN ASJOAB. ISSN 0004-637X (print), 1538-4357 (electronic).

Brown:1982:ST

[BBB82] G. E. Brown, H. A. Bethe, and G. Baym. Supernova theory. Nuclear Physics A, 378(3):481–532, February 15, 1982. CO-DEN NUPABL. ISSN 0375-9474 (print), 1873-1554 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0375947482900252.

Baron:1987:TIS

[BBB^{+}87] E. Baron, H. A. Bethe, G. E. Brown, J. Cooperstein, and S. Kahana. Type II supernovae from prompt explosions. Physical Review Letters, 59(6):736–739, August 10, 1987. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Bethe:1956:ICSa

[BBC56a] Hans A. Bethe, J. R. Beyster, and R. E. Carter. Inelastic cross-sections for fission-spectrum neutrons. Part I. Journal of Nuclear Energy (1954), 3(3):207–223, October 1956. ISSN 0891-3919 (print), 1878-2051 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0891391956900730.

Bethe:1956:ICSb

[BBC56b] Hans A. Bethe, J. R. Beyster, and R. E. Carter. Inelastic cross-sections for fission-spectrum neutrons. Part II. Journal of Nuclear Energy (1954), 3(4):273–300, November 1956. ISSN 0891-3919 (print), 1878-2051 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0891391956900560.

Bethe:1957:ICSa

[BBC57a] Hans A. Bethe, J. R. Beyster, and R. E. Carter. Inelastic crosssections for fission-spectrum neutrons. Part III. Journal of Nuclear Energy (1954), 4(1):3–25, January 1957. ISSN 0891-3919 (print),

1878-2051 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0891391957901109.

Bethe:1957:ICSb

[BBC57b] Hans A. Bethe, J. R. Beyster, and R. E. Carter. Inelastic cross-sections for fission-spectrum neutrons. Part IV. Journal of Nuclear Energy (1954), 4(2):147–163, February 1957. ISSN 0891-3919 (print), 1878-2051 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0891391957900517.

Bethe:1987:SSM

[BBC87a] H. A. Bethe, G. E. Brown, and J. Cooperstein. Stars of strange matter. Nuclear Physics A, 462(4):791–802, February 9, 1987. CODEN NUPABL. ISSN 0375-9474 (print), 1873-1554

(electronic). URL http://www.sciencedirect.com/science/
article/pii/037594748790577X.

Bethe:1987:CST

[BBC87b] Hans A. Bethe, G. E. Brown, and J. Cooperstein. Convection in supernova theory. Astrophysical Journal, 322(1):201–205, ???? 1987. CODEN ASJOAB. ISSN 0004-637X (print), 1538-4357 (electronic).

Bethe:1983:SES

[BBCW83] H. A. Bethe, G. E. Brown, J. Cooperstein, and J. R. Wilson. A simplified equation of state near nuclear density. Nuclear Physics A, 403(3):625–648, July 25, 1983. CODEN NUPABL. ISSN 0375-9474 (print), 1873-1554 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0375947483906267.

Bingham:1995:ASA

[BBD^{+}95] R. Bingham, H. A. Bethe, J. M. Dawson, J. J. Su, and V. N. Tsytovich. Anomalous scattering and absorption of neutrinos in dense plasmas. AIP Conference Proceedings, 345(??):406–415, 1995. CO-DEN APCPCS. ISSN 0094-243X (print), 1551-7616 (electronic), 1935-0465.

Bingham:1996:NSN

[BBD^{+}96] R. Bingham, H. A. Bethe, J. M. Dawson, P. K. Shukla, and J. J. Su. Nonlinear scattering of neutrinos by plasma waves: a ponderomotive force description. Physics Letters A, 220(1–3):107–110, September 2, 1996. CODEN PYLAAG. ISSN 0375-9601 (print),

1873-2429 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0375960196005038.

Baranger:1953:RCL

[BBF53] M. Baranger, H. A. Bethe, and R. P. Feynman. Relativistic correction to the Lamb shift. Physical Review (2), 92(2):482–501, October 15, 1953. CODEN PHRVAO. ISSN 0031-899X (print),

1536-6065 (electronic). URL http://prola.aps.org/abstract/
PR/v92/i2/p482_1.

Bacher:1982:UR

[BBF^{+}82] Robert F. Bacher, Hans A. Bethe, Adrian Fisher, Richard L. Garwin, Noel Gayler, John M. Lee, Norman F. Ramsey, Herbert Scoville, Jr., Paul C. Warnke, Victor F. Weisskopf, Jerome B. Wiesner, Robert R. Wilson, and Herbert F. York. USC recommendations. Bulletin of the Atomic Scientists, 38(6):5, June 1982. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Burbidge:1957:SES

[BBFH57] E. Margaret Burbidge, G. R. Burbidge, William A. Fowler, and F. Hoyle. Synthesis of the elements in stars. Reviews of Modern Physics, 29(4):547–650, October 1957. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.29.547;
 http://rmp.aps.org/abstract/RMP/v29/i4/p547_1.

Bethe:1985:OLD

[BBG^{+}85a] Hans Bethe, Norris Bradbury, Richard Garwin, Spurgeon M. Keeny, Jr., Wolfgang Panofsky, George Rathjens, Herbert Scoville, Jr., and Paul Warnke. [Open letter to Dante Fascell, chairman, Committee on Foreign Affairs, U.S. House of Representatives]: Tests not necessary. Bulletin of the Atomic Scientists, 41(10):11, November 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bethe:1985:BTC

[BBG85b] Hans A. Bethe, Jeffrey Boutwell, and Richard L. Garwin. BMD technologies and concepts in the 1980s. Dædalus, 114(2):53– 71, Spring 1985. CODEN DAEDAU. ISSN 0011-5266 (print),

1548-6192 (electronic). URL http://www.jstor.org/stable/

20024978. Special issue on the Strategic Defense Initiative. Volume 1. Concepts and technologies. Volume 2. Implications for security.

Bethe:1973:VCS

[BBH^{+}73] Hans A. Bethe, Robert F. Bacher, Philip Handler, Harvey Brooks, Bentley Glass, Alvin Weinberg, Frederick Seitz, Joseph Rotblat, Julian Schwinger, Cyril Stanley Smith, Walter Orr Roberts, Glenn T. Seaborg, John A. Simpson, Detlev W. Bronk, Harold C. Urey, Sewall Wright, Philip M. Morse, and I. I. Rabi. A voice of conscience is stilled. IV. messages. Bulletin of the Atomic Scientists, 29(6):8–12, June 1973. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Eulogy for Eugene Rabinowitch (1901–1973), co-founder of the Bulletin.

Bethe:1986:BBS

[BBL86] Hans A. (Hans Albrecht) Bethe, Robert F. (Robert Fox) Bacher, and M. Stanley (Milton Stanley) Livingston. Basic Bethe: seminal articles on nuclear physics, 1936–1937, volume 6 of The History of modern physics, 1800-1950. Tomash Publishers, Los Angeles, CA, USA, 1986. ISBN 0-88318-495-8. xi + 476 pp. LCCN QC776 .B47 1986. US$37.50. Preface by Hans A. Bethe. Introduction by Roger H. Stuewer.

Bethe:2003:FEB

[BBL03] Hans A. (Hans Albrecht) Bethe, G. E. (Gerald Edward) Brown, and Chang-Hwan Lee, editors. Formation and evolution of black holes in the galaxy: selected papers with commentary, volume 33 of World Scientific series in 20th century physics. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2003. ISBN 981-238-211-9, 981-238-250-X (paperback). xi + 506 pp. LCCN QB843.B55 F67 2003.

Bethe:2006:DFB

[BBL06] Hans A. Bethe, Gerald E. Brown, and Chang-Hwan Lee. And don’t forget the black holes (with commentary). In Brown and Lee [BL06a], pages 239–249. ISBN 981-256-609-0, 981-256-610-4 (paperback), 981-277-450-5 (e-book). LCCN QC16.B46 H36 2006.

URL http://www.worldscibooks.com/physics/5989.html;
http://www.worldscientific.com/worldscibooks/10.1142/5989.

Bethe:2007:EMB

[BBL07] Hans A. Bethe, Gerald E. Brown, and Chang-Hwan Lee. Evolution and merging of binaries with compact objects. Physics Reports, 442(1–6):5–22, ???? 2007. CODEN PRPLCM. ISSN 0370-1573 (print), 1873-6270 (electronic).

Bethe:1976:NBT

[BBM^{+}76] Hans Bethe, Aage Bohr, Ben Mottelson, Victor Weisskopf, and John Ziman. No boycott for Trieste. Physics Today, 29(6):9–??, June 1976. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://www.physicstoday.org/resource/1/
phtoad/v29/i6/p9_s2.

Bergman:1978:AEC

[BBM78] Elihu Bergman, Hans Albrecht Bethe, and Robert Eugene Marshak, editors. American energy choices before the year 2000: [proceedings]. Lexington Books, Lexington, MA, USA, 1978. ISBN 0-669-02398-1. viii + 150 pp. LCCN TJ163.25.U6 C63 1978.

Bethe:1963:RSM

[BBP63] H. A. Bethe, B. H. Brandow, and A. G. Petschek. Reference spectrum method for nuclear matter. Physical Review, 129(1):225–264, January 1, 1963. CODEN PHRVAO. ISSN 0031-899X (print),

1536-6065 (electronic). URL http://prola.aps.org/abstract/
PR/v129/i1/p225_1.

Baym:1971:NSM

[BBP71] Gordon Baym, Hans A. Bethe, and Christopher J. Pethick. Neutron star matter. Nuclear Physics A, 175(2):225–240, November 8, 1971. CODEN NUPABL. ISSN 0375-9474 (print), 1873-1554

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0375947471902818.

Brown:1991:HQG

[BBP91] G. B. Brown, H. A. Bethe, and P. M. Pizzochero. The hadron to quark/gluon transition. Physics Letters B, 263(3–4):337–341, ???? 1991. CODEN PYLBAJ. ISSN 0031-9163 (print), 1873-2410 (electronic).

Beck:1931:BQN

[BBR31] G. Beck, H. Bethe, and W. Riezler. Bemerkung zur Quantentheorie der Nullpunktstemperatur (German) [Remark on the quantum theory of zero temperature]. Naturwissenschaften, 1:39, 1931. CO-DEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Bethe:1950:NVL

[BBS50] H. A. Bethe, L. M. Brown, and J. R. Stehn. Numerical value of the Lamb shift. Physical Review, 77(3):370–374, February 1, 1950. CO-DEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

URL http://prola.aps.org/abstract/PR/v77/i3/p370_1.

Bethe:1970:NNM

[BBS70] Hans A. Bethe, G. Bömer, and Katsuhiko Sato. Nuclei in neutron matter. Astronomy and Astrophysics, 7(??):279–288, ???? 1970. CODEN AAEJAF. ISSN 0004-6361 (print), 1432-0746 (elec-

tronic). URL http://adsabs.harvard.edu/abs/1970A%26A....
.7..279B.

Bethe:1950:FAHa

[BBSS50] Hans A. Bethe, Harrison S. Brown, Frederick Seitz, and Leo Szilard. The facts about the hydrogen bomb. Bulletin of the Atomic Scientists, 6(4):106–109, 126–127, April 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bethe:1924:VTD

[BBT24] Albrech J. Bethe, Hans A. Bethe, and Y. Terada. Versuche zur Theorie der Dialyse. (German) [Experiments on the theory of dialysis]. Z. physik. Chem., 112(??):250–269, ???? 1924. This is Hans Bethe’s first published paper, co-authored with his father and his father’s colleague.

Bethe:1950:SPK

[BBW50] H. A. Bethe, L. M. Brown, and M. C. Walske. Stopping power of K-electrons. Physical Review, 79(2):413, July 15, 1950. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://prola.aps.org/abstract/PR/v79/i2/p413_1.

Bethe:1938:FDPa

[BC38a] H. A. Bethe and C. L. Critchfield. The formation of deuterons by proton combination. Physical Review, 54(4):248–254, August 15, 1938. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://prola.aps.org/abstract/PR/
v54/i4/p248_1.

Bethe:1938:FDPb

[BC38b] H. A. Bethe and C. L. Critchfield. On the formation of deuterons by proton combination. Physical Review, 54(10):862, November 15, 1938. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://prola.aps.org/abstract/PR/
v54/i10/p862_2.

Bethe:2000:OCA

[BC00] Hans A. Bethe and Robert Christy. Oppie’s colleagues affirm his leadership in Manhattan Project. Physics Today, 53(5):15, June 2000. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://link.aip.org/link/phtoad/v53/i6/
p15/s1.

Bloch:1946:LSN

[BCC^{+}46] Felix Bloch, Robert F. Christy, C. Louis Critchfield, Edwin M. McMillan, Emilio Segrè, Edward Teller, Victor Frederick Weisskopf, John H. Williams, Hans A. Bethe, Robert E. Marshak, Lyman G. Parratt, Bruno Rossi, John W. Trischa, Rudolf Peierls, Klaus Fuchs, John Manley, Bernard T. Feld, Julian Schwinger, Joseph Keller, Julius Askin, Chaim Richman, George B. Kistiakowsky, Leonard Schiff, Eugene M. Baroody, and Emil J. Konopinski. Lecture series on nuclear physics, 1943–1944. Report LA-24, Los Alamos Scientific Laboratory, Los Alamos,

NM, USA, 1946. 329 pp. URL https://www.pbagalleries.
com/view-auctions/catalog/id/352/lot/106762/LA-Report-
24-Lecture-Series-on-Nuclear-Physics.

Bethe:1992:LLL

[BCMR92] Hans A. Bethe, George Cowan, Nicholas Metropolis, and Louis Rosen. Letter: Labs leap to own defense. Bulletin of the Atomic Scientists, 48(9):45–46, November 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Barnes:1982:ENA

[BCSF82] C. A. Barnes, Donald D. Clayton, David N. Schramm, and William A. Fowler, editors. Essays in nuclear astrophysics: presented to William A. Fowler, on the occasion of his seventieth birthday. Cambridge University Press, Cambridge, UK, 1982. ISBN 0-521-24410-2, 0-521-28876-2 (paperback). xiii + 562 pp. LCCN QB464 .E87 1982.

Bethe:1954:MPSa

[BdH54a] H. A. Bethe and F. de Hoffmann. Meson–proton scattering phase shift analysis. Physical Review, 95(4):1100–1101, August 15, 1954. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://prola.aps.org/abstract/PR/
v95/i4/p1100_1.

Bethe:1954:MPSb

[BdH54b] H. A. Bethe and F. de Hoffmann. Meson–proton scattering phase shift analysis. Physical Review, 96(6):1714, December 15, 1954. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065

(electronic). URL http://prola.aps.org/abstract/PR/v96/
i6/p1714_3.

Bethe:1954:IMT

[BdH54c] Hans A. (Hans Albrecht) Bethe and F. de Hoffmann. Introduction to meson theory. Technical report, Cornell University, Ithaca, NY, USA, 1954. ???? pp. Lectures given 1953 at Cornell University.

Bethe:1955:MFI

[BdH55] Hans A. (Hans Albrecht) Bethe and F. (Frederic) de Hoffmann. Mesons and fields. II. Mesons, volume 2. Row, Peterson, Evanston, IL, USA, 1955. xiii + 446 pp. LCCN QC173 .S288.

Bethe:1970:ZVF

[BDH70] H. A. Bethe, P. A. M. Dirac, and V. Heisenberg. Zhizn’ v fizike. Vechernie lekcii v Trieste. (Russian) [Life in physics: Evening lectures at Trieste]. Uspekhi Fizicheskikh Nauk, 102(10):279–312, October 1970. CODEN UFNAAG. ISSN 0042-1294 (print), 1996-

6652 (electronic). URL http://ufn.ru/ru/articles/1970/10/

f/. From a Life of Physics, Evening Lectures the International Centre of Theoretical Physics in Trieste, Vienna, IAIA, 1969 [BDH^{+}89]. Translated by V. K. Ignatovich.

Bethe:1989:LP

[BDH^{+}89] Hans A. (Hans Albrecht) Bethe, P. A. M. Dirac, W. Heisenberg, E. P. Wigner, O. Klein, L. D. Landau, and E. M. Lifshitz, editors. From a life of physics. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1989. ISBN 9971-5-0937-7. ix + 92 pp. LCCN QC71 .F74 1989. URL http://

adsabs.harvard.edu/abs/1989liph.book.....B.

Bingham:1994:CIB

[BDSB94] R. Bingham, J. M. Dawson, J. J. Su, and H. A. Bethe. Collective interactions between neutrinos and dense plasmas. Physics Letters A, 193(3):279–284, October 3, 1994. CODEN PYLAAG. ISSN 0375-9601 (print), 1873-2429 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0375960194905975.

Bethe:1968:CDN

[BE68] H. A. Bethe and L. R. B. Elton. Charge distribution in nuclei. Physical Review Letters, 20(??):745–747, April 1, 1968. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Becquerel:1929:ETM

[Bec29] Jean Becquerel. Einleitung in eine Theorie der magneto-optischen Erscheinungen in Kristallen. (German) [Introduction to the theory of magneto-optical phenomena in crystals]. Zeitschrift für Physik, 58(3–4):205–216, March 1929. CODEN ZEPYAA.

ISSN 0044-3328. URL http://www.springerlink.com/content/

m58v616271m62886/. This paper, together with [Bet29b], is the basis for Crystal Field Theory, which John H. Van Vleck later applied to chemistry [Van32].

Becker:1967:INP

[Bec67] Richard L. Becker, editor. International Nuclear Physics Conference, held at Gatlinburg, Tennessee, September 12–17, 1966. Academic Press, New York, NY, USA, 1967. LCCN QC771 1967.

Benford:1938:LAN

[Ben38] Frank Benford. The law of anomalous numbers. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 78(4):551–572, March 1938. CODEN PAP-CAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http:

//links.jstor.org/sici?sici=0003-049X%2819380331%2978%
3A4%3C551%3ATLOAN%3E2.0.CO%3B2-G. See comments about Ben-

ford’s biased rounding practices [DF79].

Bernstein:1979:PMTa

[Ber79a] Jeremy Bernstein. Profiles: Master of the trade — I. The New Yorker, 55(??):50–??, December 3, 1979. CODEN ???? ISSN 0028-

792X. URL http://www.newyorker.com/archive/1979/12/03/
1979_12_03_050_TNY_CARDS_000330752.

Bernstein:1979:PMTb

[Ber79b] Jeremy Bernstein. Profiles: Master of the trade — II. The New Yorker, 55(??):52–??, December 10, 1979. CODEN ???? ISSN

0028-792X. URL http://www.newyorker.com/archive/1979/
12/10/1979_12_10_052_TNY_CARDS_000328924.

Bernstein:1979:PMTc

[Ber79c] Jeremy Bernstein. Profiles: Master of the trade — III. The New Yorker, 55(??):48–??, December 17, 1979. CODEN ????

ISSN 0028-792X. URL http://www.newyorker.com/archive/
1979/12/17/1979_12_17_048_TNY_CARDS_000333121.

Bernstein:1980:HBP

[Ber80] Jeremy Bernstein. Hans Bethe, prophet of energy. Basic Books, New York, NY, USA, 1980. ISBN 0-465-02903-5. xii + 212 pp. LCCN QC16.B46 B47. US$11.95.

Bernstein:1981:PEH

[Ber81] Jeremy Bernstein. Prophet of energy, Hans Bethe. Dutton, New York, NY, USA, 1981. ISBN 0-525-47677-6 (paperback). xii + 212 pp. LCCN QC16.B4 B47. US$7.25.

Bernard:1996:HBR

[Ber96a] Larry Bernard. Hans Bethe recounts the Manhattan Project during campus lecture. Cornell Chronicle Web document., Novem-

ber 7, 1996. URL http://www.news.cornell.edu/Chronicle/
96/11.7.96/Bethe.html.

Bernstein:1996:HUC

[Ber96b] Jeremy Bernstein. Hitler’s uranium club: the secret recordings at Farm Hall. American Institute of Physics, Woodbury, NY, USA, 1996. ISBN 1-56396-258-6. xxx + 427 + 4 pp. LCCN QC773.3.G3 B47 1995. US$34.95. Introduction by David Cassidy.

Bernstein:2001:MPO

[Ber01] Jeremy Bernstein. The Merely Personal: Observations on Science and Scientists. Ivan R. Dee, Chicago, IL, USA, 2001. ISBN 1-56663-344-3. 245 pp. LCCN Q171 .B5375 2001.

Bernstein:2003:DWH

[Ber03] Jeremy Bernstein. The drawing, or why history is not mathematics. Physics in Perspective (PIP), 5(3):243–261, 2003. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Bernstein:2006:HBN

[Ber06] Jeremy Bernstein. Hans Bethe at The New Yorker. In Brown and Lee [BL06a], pages 111–114. ISBN 981-256-609-0, 981-256-610-4 (paperback), 981-277-450-5 (e-book). LCCN QC16.B46 H36 2006.

URL http://www.worldscibooks.com/physics/5989.html;
http://www.worldscientific.com/worldscibooks/10.1142/5989.

Bernstein:2010:JNK

[Ber10] Jeremy Bernstein. John von Neumann and Klaus Fuchs: an unlikely collaboration. Physics in Perspective (PIP), 12(1):36–50, March 2010. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.
1007/s00016-009-0001-1;   http://www.springerlink.com/
content/y54804274h88v546/.

Bethe:1927:SEK

[Bet27] Hans Bethe. Über die Streuung von Elektronen an Krystallen. (German) [On the scattering of electrons by crystals]. Naturwissenschaften, 15(38):786–788, ???? 1927. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL http://www.

springerlink.com/content/q8160663j6475318/fulltext.pdf.

Bethe:1928:TBE

[Bet28a] H. Bethe. Theorie der Beugung von Elektronen an Kristallen. (German) [Theory of diffraction of electrons by crystals]. Annalen der Physik (1900), 392(17):55–129, 1928. ISSN 1521-3889.

Bethe:1928:SEK

[Bet28b] Hans Bethe. Die Streuung von Elektronen an Krystallen. (German) [The scattering of electrons by crystals]. Naturwissenschaften, 16(19):333–334, ???? 1928. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL http://www.

springerlink.com/content/u614786357707752/fulltext.pdf.

Bethe:1929:BEW

[Bet29a] Hans A. Bethe. Berechnung der Elektronenaffinität des Wasserstoffs. (German) [Calculation of electronic affinity of hydrogen]. Zeitschrift für Physik, 57(??):815–821, ???? 1929. CODEN ZEPYAA. ISSN 0044-3328.

Bethe:1929:TKG

[Bet29b] Hans A. Bethe. Termaufspaltung in Kristallen. (German) [Term splitting in crystals]. Annalen der Physik (1900), 395(2):133–208, ???? 1929. ISSN 1521-3889. This paper, together with [Bec29], is the basis for Crystal Field Theory, which John H. Van Vleck later applied to chemistry [Van32].

Bethe:1929:DKD

[Bet29c] Hans A. Bethe. Über den Durchgang von Kathodenstrahlen durch gitterförmige elektrische Felder. (German) [Passage of cathode rays through electric fields formed by grids]. Zeitschrift für Physik, 54 (??):703–710, ???? 1929. CODEN ZEPYAA. ISSN 0044-3328.

Bethe:1929:VEI

[Bet29d] Hans A. Bethe. Vergleich der Elektronenverteilung im Heliumgrundzustand nach verschiedenen Methoden. (German) [Comparison of the distribution of electrons in the helium ground state as calculated by different methods]. Zeitschrift für Physik, 55(??): 431–436, ???? 1929. CODEN ZEPYAA. ISSN 0044-3328.

Bethe:1930:NBP

[Bet30a] H. Bethe. Über die nichtstationäre Behandlung des Photoeffekts. (German) [On the non-stationary treatment of the photoelectric effect]. Annalen der Physik (1900), 396(4):443–449, 1930. ISSN 1521-3889.

Bethe:1930:TDS

[Bet30b] H. Bethe. Zur Theorie des Durchgangs schneller Korpuskularstrahlen durch Materie. (German) [On the theory of passage of faster corpuscular rays through matter]. Annalen der Physik (1900), 397(3):325–400, 1930. ISSN 1521-3889.

Bethe:1930:TZS

[Bet30c] Hans A. Bethe. Zur Theorie des Zeemaneffektes an den Salzen der seltenen Erden. (German) [Theory of the Zeeman effect in the salts of rare earth]. Zeitschrift für Physik, 60(??):218–233, ???? 1930. CODEN ZEPYAA. ISSN 0044-3328.

Bethe:1931:CRM

[Bet31a] Hans A. Bethe. Change of resistance in magnetic fields. Nature, 127 (3201):336–337, March 7, 1931. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v127/n3201/pdf/127336a0.pdf.

Bethe:1931:TME

[Bet31b] Hans A. Bethe. Zur Theorie der Metalle. I. Eigenwerte und Eigenfunktionen der lineare Atomkette. (German) [On the theory of metals. I. Eigenvalues and eigenfunctions of the linear atomic chain]. Zeitschrift für Physik, 71(3–4):205–226, March 1931. CO-

DEN ZEPYAA. ISSN ???? URL http://www.springerlink.
com/content/g86x6317566uh2x6/.

Bethe:1932:BER

[Bet32] Hans A. Bethe. Bremsformel für Elektronen relativistischer Geschwindigkeit. (German) [Scattering of elections of relativistic velocity]. Zeitschrift für Physik, 76(5–6):293–299, ???? 1932. CO-DEN ZEPYAA. ISSN 0044-3328. See annotation [FBB14].

Bethe:1933:AZM

[Bet33a] Hans Bethe. Aufbau der zusammenhängenden Materie. (German) [Construction of collective matter]. In Adolf Gustav Smekal, editor, Handbuch der Physik, volume 24-2, pages xiii + 1203. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1933. LCCN ????

Bethe:1933:QZE

[Bet33b] Hans Bethe. Quantenmechanik der Ein und Zwei-Elektronenprobleme. (German) [Quantum mechanics of one- and two-electron problems]. In Adolf Gustav Smekal, editor, Handbuch der Physik, volume 24-1, pages 273–560. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1933. LCCN ????

Bethe:1933:QGQ

[Bet33c] Hans Bethe. Quantentheorie. (German) [Quantum theory]. In Adolf Gustav Smekal, editor, Handbuch der Physik, volume 24-1, pages ix + 853. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1933. LCCN ????

Bethe:1933:TFG

[Bet33d] Hans A. Bethe. Theorie des Ferromagnetismus. (German) [Theory of ferromagnetism]. In P. J. W. Debye, editor, Magnetismus, pages 74–81. Hirzel, Leipzig, Germany, 1933.

Bethe:1934:ISC

[Bet34a] Hans A. Bethe. The influence of screening on the creation and stopping of electrons. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 30:524–539, ???? 1934. CODEN PCPSA4. ISSN 0008-1981.

Bethe:1934:QBE

[Bet34b] Hans A. Bethe. Quantitative Berechnung der Eigenfunktion von Metallelektronen. (German) [Quantitative calculation of the eigenfunction of electrons in metals]. Helvetica Physica Acta, 7(2):18–23, ???? 1934. CODEN HPACAK. ISSN 0018-0238 (print), 2297-1971 (electronic).

Bethe:1934:KTS

[Bet34c] Hans A. Bethe. Zur Kritik der Theorie der Supraleitung von R. Schachenmeier. (German) [A critique of the theory of superconductivity by R. Schachenmeier]. Zeitschrift für Physik, 90(9–10):674–679, September 1934. CODEN ZEPYAA.

ISSN ????  URL http://www.springerlink.com/content/

v825g1x1jv5812l2/. See [Sch32] and response [Sch34].

Bethe:1935:MLA

[Bet35a] H. Bethe. Masses of light atoms from transmutation data. Physical Review, 47(8):633–634, April 15, 1935. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.

aps.org/abstract/PR/v47/i8/p633_1.

Bethe:1935:TDN

[Bet35b] H. A. Bethe. Theory of disintegration of nuclei by neutrons. Physical Review, 47(10):747–759, May 15, 1935. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://

prola.aps.org/abstract/PR/v47/i10/p747_1.

Bethe:1935:IPN

[Bet35c] Hans A. Bethe. Ionization power of a neutrino with magnetic moment. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 31(1):108–115, January 1935. CO-DEN PCPSA4. ISSN 0008-1981.

Bethe:1935:MCR

[Bet35d] Hans A. Bethe. Memorandum of cosmic rays. In Carnegie Institution of Washington Yearbook, volume 35, pages 333–335. Carnegie Institution, Washington, DC, USA, 1935.

Bethe:1935:ARP

[Bet35e] Hans A. Bethe. On the annihilation radiation of positrons. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 150(869):129–141, May 1, 1935.

CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-9169 (elec-

tronic).  URL http://rspa.royalsocietypublishing.org/
content/150/869/129.

Bethe:1935:STS

[Bet35f] Hans A. Bethe. Statistical theory of superlattices. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 150(??):552–575, ???? 1935. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-9169 (electronic).

Bethe:1936:ACN

[Bet36a] H. A. Bethe. An attempt to calculate the number of energy levels of a heavy nucleus. Physical Review, 50(4):332–341, August 15, 1936. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://prola.aps.org/abstract/PR/
v50/i4/p332_1.

Bethe:1936:NRM

[Bet36b] H. A. Bethe. Nuclear radius and many-body problem. Physical Review, 50(11):977–979, December 1, 1936. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://

prola.aps.org/abstract/PR/v50/i11/p977_1.

Bethe:1937:NPB

[Bet37a] H. A. Bethe. Nuclear physics B. Nuclear dynamics, theoretical. Reviews of Modern Physics, 9(2):69–244, April 1, 1937. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.9.
69;   http://rmp.aps.org/abstract/RMP/v9/i2/p69_1.

Bethe:1937:STC

[Bet37b] Hans A. Bethe. Statistical theory of cooperative phenomena. In Anonymous [Ano37], page ?? LCCN ????

Bethe:1938:BED

[Bet38a] H. A. Bethe. The binding energy of the deuteron. Physical Review, 53(4):313–314, February 15, 1938. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.aps.

org/abstract/PR/v53/i4/p313_1.

Bethe:1938:CEL

[Bet38b] H. A. Bethe. Coulomb energy of light nuclei. Physical Review, 54(6):436–439, September 15, 1938. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.

aps.org/abstract/PR/v54/i6/p436_1.

Bethe:1938:MML

[Bet38c] H. A. Bethe. Magnetic moment of Li^{7} in the α-particle model. Physical Review, 53(10):842, May 15, 1938. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://

prola.aps.org/abstract/PR/v53/i10/p842_1.

Bethe:1938:MTL

[Bet38d] H. A. Bethe. A method for treating large perturbations. Physical Review, 54(11):955–967, December 1, 1938. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://

prola.aps.org/abstract/PR/v54/i11/p955_1.

Bethe:1938:OPP

[Bet38e] H. A. Bethe. The Oppenheimer–Phillips process. Physical Review, 53(1):39–50, January 1, 1938. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.aps.

org/abstract/PR/v53/i1/p39_1.

Bethe:1938:ODA

[Bet38f] H. A. Bethe. Order and disorder in alloys. Journal of Applied Physics, 9(??):244–251, 1938. CODEN JAPIAU. ISSN 0021-8979 (print), 1089-7550 (electronic), 1520-8850.

Bethe:1939:EPSb

[Bet39a] H. A. Bethe. Energy production in stars. Physical Review, 55(1): 103, January 1, 1939. CODEN PHRVAO. ISSN 0031-899X (print),

1536-6065 (electronic). URL http://prola.aps.org/abstract/
PR/v55/i1/p103_1.

Bethe:1939:MTN

[Bet39b] H. A. Bethe. The meson theory of nuclear forces. Physical Review, 55(12):1261–1263, June 15, 1939. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.aps.

org/abstract/PR/v55/i12/p1261_1.

Bethe:1939:EPSa

[Bet39c] Hans A. Bethe. Energy production in stars. Physical Review, 55 (5):434–456, March 1, 1939. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://link.aps.org/doi/

10.1103/PhysRev.55.434;   http://prola.aps.org/abstract/

PR/v55/i5/p434_1. See further work [Hoy46, Hoy54]. According to [BL09, page 41], this paper won Bethe the A. Cressy Morrison Prize of the New York Academy of Sciences, and almost three decades later, the 1967 Nobel Prize in Physics.

Bethe:1940:CTC

[Bet40a] H. A. Bethe. A continuum theory of the compound nucleus. Physical Review, 57(12):1125–1144, June 15, 1940. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://

prola.aps.org/abstract/PR/v57/i12/p1125_1.

Bethe:1940:MTNa

[Bet40b] H. A. Bethe. The meson theory of nuclear forces I. general theory. Physical Review, 57(4):260–272, February 15, 1940. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://prola.aps.org/abstract/PR/v57/i4/p260_1.

Bethe:1940:MTNb

[Bet40c] H. A. Bethe. The meson theory of nuclear forces. Part II. theory of the deuteron. Physical Review, 57(5):390–413, March 1, 1940. CO-DEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

URL http://prola.aps.org/abstract/PR/v57/i5/p390_1.

Bethe:1940:REN

[Bet40d] Hans A. Bethe. Recent evidence on the nuclear reactions in the carbon cycle. Astrophysical Journal, 92(??):118–121, ???? 1940. CO-DEN ASJOAB. ISSN 0004-637X (print), 1538-4357 (electronic).

Bethe:1942:SHA

[Bet42a] H. A. Bethe. The specific heat of air up to 25,000^{∘}c. Report 1942 OSRD-369 (NDRC B-171), Los Alamos Scientific Laboratory, Los Alamos, NM, USA, February 9, 1942.

Bethe:1942:EPS

[Bet42b] Hans A. Bethe. Energy production in stars. American Scientist, 30(??):243–264, ???? 1942. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (electronic).

Bethe:1942:THFb

[Bet42c] Hans A. Bethe. Theory of high frequency rectification by silicon crystals. Report RL-184, US Atomic Energy Commission, Washington, DC, USA, October 29, 1942. URL http://www.osti.gov/

accomplishments/documents/fullText/ACC0099.pdf.

Bethe:1942:THFa

[Bet42d] Hans A. (Hans Albrecht) Bethe. Theory of high frequency rectification by silicon crystals. Technical report, Radiation Laboratory, Massachusetts Institute of Technology, Cambridge, MA, USA, 1942. 12 pp.

Bethe:1942:TBL

[Bet42e] Hans A. (Hans Albrecht) Bethe. Theory of the boundary layer of crystal rectifiers. Technical report, Radiation Laboratory, Massachusetts Institute of Technology, Cambridge, MA, USA, 1942. 23 + 1 pp.

Bethe:1943:ECT

[Bet43a] Hans A. (Hans Albrecht) Bethe. Excitation of cavities through windows. Technical report, Radiation Laboratory, Massachusetts Institute of Technology, Cambridge, MA, USA, 1943. 18 pp.

Bethe:1943:FTW

[Bet43b] Hans A. (Hans Albrecht) Bethe. Formal theory of wave guides of arbitrary cross section. Technical report, Radiation Laboratory, Massachusetts Institute of Technology, Cambridge, MA, USA, 1943. 16 pp.

Bethe:1943:LCS

[Bet43c] Hans A. (Hans Albrecht) Bethe. Lumped constants for small irises. Technical report, Radiation Laboratory, Massachusetts Institute of Technology, Cambridge, MA, USA, 1943. 40 pp.

Bethe:1943:TSW

[Bet43d] Hans A. (Hans Albrecht) Bethe. Theory of side windows in wave guides. Technical report, Radiation Laboratory, Massachusetts Institute of Technology, Cambridge, MA, USA, 1943. 26 pp.

Bethe:1944:TDS

[Bet44] H. A. Bethe. Theory of diffraction by small holes. Physical Review (2), 66(7):163–182, October 1, 1944. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.

aps.org/abstract/PR/v66/i7-8/p163_1. This paper, delayed in

publication by two and a half years because of war-time security considerations, describes the ‘Bethe coupler’, a device that makes bidirectional power flow measurements possible in radar waveguides.

Bethe:1946:MSM

[Bet46a] H. A. Bethe. Multiple scattering and the mass of the meson. Physical Review, 70(11):821–831, December 1, 1946. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://

prola.aps.org/abstract/PR/v70/i11-12/p821_1.

Bethe:1946:CAW

[Bet46b] Hans A. Bethe. Can air or water be exploded? Bulletin of the Atomic Scientists, 1(7):2, 14, March 15, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bethe:1947:ESE

[Bet47a] H. A. Bethe. The electromagnetic shift of energy levels. Physical Review, 72(4):339–341, August 15, 1947. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.

aps.org/abstract/PR/v72/i4/p339_1.

Bethe:1947:ENT

[Bet47b] Hans A. (Hans Albrecht) Bethe. Elementary Nuclear Theory: a Short Course on Selected Topics. Wiley, New York, NY, USA, 1947. vi + 147 pp. LCCN QC173 .B48 1947. URL https://

archive.org/details/Bethe-ElementaryNuclearTheory. Given

at the Research Laboratory of the General Electric Company at Schenectady.

Bethe:1948:BWE

[Bet48] Hans A. Bethe. Bemerkungen über die Wasserstoff-Eigenfunktionen in der Diracschen Theorie. (German) [Remarks concerning the hydrogen eigenfunctions in Dirac’s theory.]. Zeitschrift für Naturforschung, 3A(??):470–477, ???? 1948. CODEN ZNTFA2. ISSN 0372-9516.

Bethe:1949:TER

[Bet49a] H. A. Bethe. Theory of the effective range in nuclear scattering. Physical Review, 76(1):38–50, July 1, 1949. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://

prola.aps.org/abstract/PR/v76/i1/p38_1.

Bethe:1949:NDL

[Bet49b] Hans A. (Hans Albrecht) Bethe. Neutron diffusion at large distances from the source. Technical Report AECD 2105, U.S. Atomic Energy Commission, Technical Information Branch, Oak Ridge,

TN, USA, 1949. 15 pp. URL http://catalog.hathitrust.org/
Record/011571709;   http://pi.lib.uchicago.edu/1001/cat/

bib/ocn702212666. Declassified 9 July 1948.

Bethe:1949:PAN

[Bet49c] Hans A. (Hans Albrecht) Bethe. The properties of atomic nuclei. II. Range-energy curves, alpha-particles, protons, and mesons. Technical Report BNL-T-7 and AECU 374, Brookhaven National Laboratory, Upton, NY, USA, June 1, 1949. 25

pp. URL http://hdl.handle.net/2027/mdp.39015086572024;
 http://pi.lib.uchicago.edu/1001/cat/bib/ocm25358007.

Bethe:1949:STC

[Bet49d] Hans A. (Hans Albrecht) Bethe. Splitting of terms in crystals. Consultants Bureau, New York, NY, USA, 1949. 69 pp. LCCN QD339 .B4.

Bethe:1949:REC

[Bet49e] Hans Albrecht Bethe. Range–energy curves. Report, Brookhaven National Laboratory, Upton, NY, USA, 1949. 13 pp.

Bethe:1950:RER

[Bet50a] H. A. Bethe. The range–energy relation for slow alphaparticles and protons in air. Reviews of Modern Physics, 22 (2):213–219, April 1, 1950. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.22.213;   http://
rmp.aps.org/abstract/RMP/v22/i2/p213_1.

Bethe:1950:HBb

[Bet50b] Hans A. Bethe. The H bomb. In Eliot [Eli50], page ?? LCCN UF767 .H2. Introduction by Albert Einstein and commentary by George Fielding Eliot.

Bethe:1950:HBa

[Bet50c] Hans A. Bethe. The hydrogen bomb. Bulletin of the Atomic Scientists, 6(4):99–104, 125, April 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprinted in [Bet63a].

Bethe:1950:HBI

[Bet50d] Hans A. Bethe. The hydrogen bomb: II. Scientific American, 182(4):18–23, April 1950. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.nature.com/scientificamerican/journal/v182/n4/pdf/
scientificamerican0450-18.pdf. See also Parts I [Rid50] and

III [Bac50]. See [Hol03] for a comment about the delay in publishing this paper, due to confiscation by the Atomic Energy Commission of the printing plates.

Bethe:1950:TEM

[Bet50e] Hans A. Bethe. Theoretical estimate of maximum possible nuclear explosion. Report KAPL-294, US Atomic Energy Commission, Washington, DC, USA, January 31, 1950. URL http://www.

osti.gov/accomplishments/documents/fullText/ACC0100.pdf.

Declassified 22 February 1975.

Bethe:1950:STC

[Bet50f] Hans A. (Hans Albrecht) Bethe. Splitting of terms in crystals. Consultants Bureau, New York, NY, USA, 1950. 69 pp. LCCN QC174.2 .B5621; QD939.B56. URL http://pi.lib.uchicago. edu/1001/cat/bib/797721. English translation of [Bet29b].

Bethe:1951:TAC

[Bet51] Hans A. Bethe. Theory of atomic collisions. Nature, 167(4240): 165, February 3, 1951. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v167/n4240/pdf/167165a0.pdf.

Bethe:1952:MHT

[Bet52a] Hans A. Bethe. Memorandum on the history of thermonuclear program. Assembled by Chuck Hansen, editor of Swords of Armaged-

don, May 28, 1952. URL http://www.fas.org/nuke/guide/usa/
nuclear/bethe-52.htm.

Albrecht:1952:RTN

[Bet52b] Hans A. (Hans Albrecht) Bethe. The reciprocity theorem in neutron scattering. Report LA-1428, U.S. Atomic Energy Commission, Technical Information Service, Oak Ridge, TN, USA, June

20, 1952. 18 pp. URL http://pi.lib.uchicago.edu/1001/cat/
bib/ocn771930833.

Bethe:1953:MTM

[Bet53a] H. A. Bethe. Molière’s theory of multiple scattering. Physical Review, 89(6):1256–1266, March 15, 1953. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.

aps.org/abstract/PR/v89/i6/p1256_1.

Bethe:1953:SPS

[Bet53b] H. A. Bethe. The sign of the phase shifts in meson–nucleon scattering. Physical Review, 90(5):994–995, June 1, 1953. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://prola.aps.org/abstract/PR/v90/i5/p994_2.

Bethe:1953:WHN

[Bet53c] Hans A. Bethe. What holds the nucleus together? Scientific American, 189(3):56–63, September 1953. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.nature.com/scientificamerican/journal/v189/n3/pdf/
scientificamerican0953-58.pdf.

Bethe:1953:HEP

[Bet53d] Hans A. (Hans Albrecht) Bethe. High energy phenomena. Technical report, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, January 8, 1953. 100 pp. A course of lectures given at Los Alamos in the spring and summer of 1952, concerning phenomena involving particles with energy in the range of hundreds of MeV. Main emphasis is placed on the properties of π-mesons and on a relativistic treatment of the nucleon-nucleon interaction.

Bethe:1954:MNFb

[Bet54a] H. A. Bethe. Mesons and nuclear forces. Journal of the Washington Academy of Sciences, 44(4):97–105, April 1954. CODEN JWASA3.

ISSN 0043-0439. URL http://www.washacadsci.org/Journal/
journalscans/v.44-n.4.pdf.

Bethe:1954:BRB

[Bet54b] Hans A. Bethe. Book review: Methods of Mathematical Physics, Vol. I. Trans. from 2nd German ed. R. Courant and D. Hilbert. New York-London: Interscience, 1953. 561 pp. $9.50. Science, 119 (3080):75–76, January 8, 1954. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/119/3080/75.full.pdf.

Bethe:1954:MNFa

[Bet54c] Hans A. Bethe. Mesons and nuclear forces. Physics Today, 7 (2):5–11, February 1954. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.aip.org/

link/phtoad/v7/i2/p5/s1.

Bethe:1954:NAB

[Bet54d] Hans A. Bethe. Negotiations and atomic bombs. Bulletin of the Atomic Scientists, 10(1):9–10, January 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bethe:1954:ODH

[Bet54e] Hans A. Bethe. Observations on the development of the H-bomb. Unpublished Los Alamos text. Finally printed, with slight edits, in [Bet82f]., 1954.

Bethe:1955:MJS

[Bet55a] H. Bethe. Mezony i jadernye sily. (Russian) [Mesons and nuclear forces]. Uspekhi Fizicheskikh Nauk, 56(6):149–166, June 1955. CO-DEN UFNAAG. ISSN 0042-1294 (print), 1996-6652 (electronic).

URL http://ufn.ru/ru/articles/1955/6/a/. Translation of

[Bet54a].

Bethe:1955:PSA

[Bet55b] H. A. Bethe. Phase-shift analysis of pion-proton scattering. In Bellamy and Moorehouse [BM55], pages 265–267. LCCN ????

Bethe:1955:MSH

[Bet55c] Hans A. Bethe. Memorial Symposium Held in Honor of Enrico Fermi at the Washington Meeting of the American Physical Society, April 29, 1955. Reviews of Modern Physics, 27(3):249, July 1, 1955. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.27.249;   http://rmp.aps.org/abstract/
RMP/v27/i3/p249_1.

Bethe:1955:PPM

[Bet55d] Hans A. Bethe. Physik der pi-Mesonen. (German) [Physics of the π-meson]. In E. Brüche and H. Franke, editors, Physikertagung Wiesbaden 1955. Hautvortrage der Jahrestagung 1955, page ?? Physikverlag, Mosbach/Baden, West Germany, 1955. LCCN ????

Bethe:1956:NMB

[Bet56a] H. A. Bethe. Nuclear many-body problem. Physical Review, 103 (5):1353–1390, September 1, 1956. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.aps.

org/abstract/PR/v103/i5/p1353_1.

Bethe:1956:GIL

[Bet56b] Hans A. Bethe. General introduction. (lecture at Amsterdam Nuclear Reactions Conference, July 2–7, 1956). Physica, 22(??):941– 951, ???? 1956. CODEN PHYSAG. ISSN 0031-8914 (print), 1873-1767 (electronic).

Bethe:1956:IBT

[Bet56c] Hans A. Bethe. Introduction to the Brueckner theory. (lecture at Amsterdam Nuclear Reactions Conference, July 2–7, 1956). Physica, 22(??):987–993, ???? 1956. CODEN PHYSAG. ISSN 0031-8914 (print), 1873-1767 (electronic).

Bethe:1956:RSO

[Bet56d] Hans A. Bethe. Reactor safety and oscillator tests. Report 117, APDA, ????, October 15, 1956.

Bethe:1957:DEF

[Bet57] Hans A. Bethe. On the Doppler effect in fast reactors. Report 119, APDA, ????, March 11, 1957.

Bethe:1958:BBT

[Bet58a] Hans A. Bethe. Books: Brighter Than a Thousand Suns, By Robert Jungk. Bulletin of the Atomic Scientists, 14(10):426–428, December 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bethe:1958:SPP

[Bet58b] Hans A. Bethe. Scattering and polarization of protons by nuclei. Annals of Physics, 3(2):190–240, February 1958. CO-DEN APNYA6. ISSN 0003-4916 (print), 1096-035X (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0003491658900174.

Bethe:1959:FP

[Bet59] Hans A. Bethe. Fundamental particles. Journal of The Franklin Institute, 268(??):501–507, ???? 1959. CODEN JFINAB. ISSN 0016-0032 (print), 1879-2693 (electronic).

Bethe:1960:CEN

[Bet60a] Hans A. Bethe. The case for ending nuclear tests. Atlantic Monthly, ??(??):43–51, August 1960. ISSN 1072-7825 (print), 2151-9463 (electronic).

Bethe:1960:LEA

[Bet60b] Hans A. Bethe. Letter to the Editor: Appraises Russian proposal. Bulletin of the Atomic Scientists, 16(5):144, May 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bethe:1960:STP

[Bet60c] Hans A. Bethe. Soviet test plan examined. New York Times, page 36, March 29, 1960. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Bethe:1961:TSW

[Bet61] Hans A. Bethe. Thinks Soviets want tests. New York Times, page E6, July 2, 1961. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Bethe:1962:AA

[Bet62a] Hans Bethe. [The Atomic Age]. Bulletin of the Atomic Scientists, 18(10):19, December 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bethe:1962:DS

[Bet62b] Hans A. Bethe. Disarmament and strategy. Bulletin of the Atomic Scientists, 18(7):14–22, September 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bethe:1963:AHB

[Bet63a] Hans Bethe. The American hydrogen bomb. In Grodzins and Rabinowitch [GR63], pages 144–155. LCCN D842 .B78. Reprint of [Bet50c].

Bethe:1963:FPJ

[Bet63b] Hans Bethe. Fermi Prize: J. Robert Oppenheimer named to receive annual AEC Award. Science, 140(3563):161–163, April 12, 1963. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Bethe:1963:DBT

[Bet63c] Hans A. Bethe. Derivation of the Brueckner theory. In Percus [Per63], page 31. LCCN ????

Bethe:1963:NAW

[Bet63d] Hans A. Bethe. Nobel Award winners announced. Science, 142 (3594):937–938, November 15, 1963. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/142/3594/937.full.pdf.

Bethe:1964:DSS

[Bet64a] Hans A. Bethe. Disarmament and strategic stability. Fourth Annual Dewey F. Fagerburg Memorial Lecture, University of Michigan, 1964.

Bethe:1964:IQM

[Bet64b] Hans A. Bethe. Intermediate quantum mechanics. Notes by R. W. Jackiw. Benjamin/Cummings Pub. Co., Menlo Park, CA, USA, 1964. x + 276 pp. LCCN QC174.1 .B416.

Bethe:1964:NHE

[Bet64c] Hans A. Bethe. Note on high-energy proton–proton scattering. Il Nuovo Cimento (10), 33(??):1167–1172, ???? 1964. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (electronic).

Bethe:1964:TF

[Bet64d] Hans A. Bethe. Theory of the fireball. Report LA-3064, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, 1964.

Bethe:1965:HBD

[Bet65a] H. A. Bethe. H. A. Bethe discusses high energy physics. Science, 147(3665):1551–1552, March 26, 1965. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/147/3665/1551.full.pdf.

Bethe:1965:TBC

[Bet65b] H. A. Bethe. Three-body correlations in nuclear matter. Physical Review, 138(4B):B804–B822, May 24, 1965. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://

prola.aps.org/abstract/PR/v138/i4B/pB804_1.

Bethe:1965:FA

[Bet65c] Hans A. Bethe. The fireball in air. Journal of Quantitative Spectroscopy and Radiative Transfer, 5(??):9–12, ???? 1965. CODEN JQSRAE. ISSN 0022-4073 (print), 1879-1352 (electronic).

Bethe:1966:EPS

[Bet66a] Hans A. Bethe. Energy production in the sun. In Th. Page and L. Williams Page, editors, The Origin of the Solar System, pages 30–39. Macmillan Publishing Company, New York, NY, USA, 1966. LCCN ????

Bethe:1966:SSA

[Bet66b] Hans A. Bethe. Shadow scattering by atoms. In de Shalit et al. [dSFvLH66], pages 240–249. LCCN ????

Bethe:1967:TBC

[Bet67a] H. A. Bethe. Three-body correlations in nuclear matter. II. Physical Review, 158(4):941–947, June 20, 1967. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://

prola.aps.org/abstract/PR/v158/i4/p941_1.

Bethe:1967:NPP

[Bet67b] Hans Bethe. The Nobel Prize in Physics 1967. Nobelprize.org, 1967. The Nobel Prize in Physics 1967 was awarded to Hans Bethe “for his contributions to the theory of nuclear reactions, especially his discoveries concerning the energy production in stars”. Reprinted in [Bet97e, pages 380–395].

Bethe:1967:EM

[Bet67c] Hans A. Bethe. Elektronentheorie der Metalle. (German) [Electron theory of metals]. ????, ????, 1967. ???? pp. LCCN ????

Bethe:1967:JRO

[Bet67d] Hans A. Bethe. J. Robert Oppenheimer, 1904–1967. Bulletin of the Atomic Scientists, 23(8):3–6, October 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bethe:1967:OWW

[Bet67e] Hans A. Bethe. Oppenheimer: “Where He Was There Was Always Life and Excitement”. Science, 155(3766):1080–1084, March 3, 1967. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203

(electronic). URL http://www.sciencemag.org/content/155/
3766/1080.full.pdf.

Bethe:1967:TNM

[Bet67f] Hans A. Bethe. Theory of nuclear matter. In ????, editor, Proceedings of the International Conference on Nuclear Matter, Tokyo, Japan (Sept 1967), pages 56–62. ????, ????, 1967. LCCN ????

Bethe:1968:ESG

[Bet68a] H. A. Bethe. Energieerzeugung in Sternen. (German) [Energy production in stars]. Naturwissenschaften, 55(9):405–413, ???? 1968. CODEN NATWAY. ISSN 0028-1042 (print),

1432-1904 (electronic). URL http://www.springerlink.com/
content/l348q530lx321843/fulltext.pdf.

Bethe:1968:TFT

[Bet68b] H. A. Bethe. Thomas–Fermi theory of nuclei. Physical Review, 167(4):879–907, March 20, 1968. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.aps.

org/abstract/PR/v167/i4/p879_1.

Bethe:1968:IJZ

[Bet68c] Hans Bethe. Istochniki jenergii zvezd. (Russian) [Sources of stellar energy]. Uspekhi Fizicheskikh Nauk, 96(3):393–408, November 1968. CODEN UFNAAG. ISSN 0042-1294 (print), 1996-6652

(electronic). URL http://ufn.ru/ru/articles/1968/11/a/.

Bethe:1968:EPSb

[Bet68d] Hans A. Bethe. Energy production in stars. Physics Today, 21(9): 36–44, September 1968. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://scitation.aip.

org/content/aip/magazine/physicstoday/article/21/9/10.
1063/1.3035151.

Bethe:1968:EPSc

[Bet68e] Hans A. Bethe. Energy production in stars. Science, 161(3841): 541–547, August 9, 1968. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/161/3841/541.full.pdf. See erratum [Bet83a].

Bethe:1968:JRO

[Bet68f] Hans A. Bethe. J. Robert Oppenheimer. 1904–1967. Biographical Memoirs of Fellows of the Royal Society, 14:390–416, November 1968. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494

(electronic). URL http://www.jstor.org/stable/769451.

Bethe:1968:EPSa

[Bet68g] Hans A. (Hans Albrecht) Bethe. Energy production in stars. Nobel lecture. Norstedt & Söner, Stockholm, Sweden, December 11, 1968. 16 pp. LCCN QB464 .B46.

Bethe:1969:ACA

[Bet69a] Hans A. Bethe. The ABM, China and the Arms Race. Bulletin of the Atomic Scientists, 25(5):41–44, May 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bethe:1969:MAM

[Bet69b] Hans A. Bethe. Missiles and anti-missiles … six views: Hard point vs. city defense. Bulletin of the Atomic Scientists, 25(6):25–26, June 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Washington Scene: Missiles and Antimissiles.

Bethe:1969:MSF

[Bet69c] Hans A. Bethe. M.I.T. science forum. New York Times, page 28, February 15, 1969. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Bethe:1970:JZV

[Bet70a] Hans Bethe. Jenergija na zemle i v zvezdah. (Russian) [The energy on Earth and in the stars]. Uspekhi Fizicheskikh Nauk, 102(2):279– 290, October 1970. CODEN UFNAAG. ISSN 0042-1294 (print),

1996-6652 (electronic). URL http://ufn.ru/ru/articles/1970/

10/i/. From a Life of Physics, Evening Lectures the International Centre of Theoretical Physics in Trieste, Vienna, IAIA, 1969 [BDH^{+}89]. Translated by V. K. Ignatovich.

Bethe:1970:CMN

[Bet70b] Hans A. Bethe. Capture of K^{−} mesons in the nuclear periphery. Nuclear Physics B, 21(??):589–605, ???? 1970. CODEN NUPBBO. ISSN 0550-3213 (print), 1873-1562 (electronic).

Bethe:1970:DPP

[Bet70c] Hans A. Bethe. Disarmament problems. Part 4 of “The Military Atom”. Bulletin of the Atomic Scientists, 26(6):99–102, June 1970. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bethe:1970:EPS

[Bet70d] Hans A. Bethe. Energy production in stars. Proceedings of the Physical Society of Japan, 25(??):??, May 1970.

Bethe:1970:I

[Bet70e] Hans A. Bethe. Introduction. In Fernbach and Taub [FT70], page ?? ISBN 0-677-14030-4. LCCN Q183.9 .C65.

Bethe:1971:DP

[Bet71a] Hans A. Bethe. Disarmament problems. In Lewis and Wilson [LW71], page ?? ISBN 0-670-11151-1. LCCN QC792 .A43 1971. With Eugene Rabinowitch.

Bethe:1971:PEV

[Bet71b] Hans A. Bethe. Produkce energie ve hvezdach. (czech) [Energy production in stars]. Geskoslovensky Gasopis Pro Fysiku [Sekce A], 21(??):35–??, ???? 1971.

Bethe:1971:SET

[Bet71c] Hans A. Bethe. Strutinsky’s energy theorem. In Heisenberg et al. [HBD71], pages 237–?? LCCN QC174.45 .Q35.

Bethe:1971:TNM

[Bet71d] Hans A. Bethe. Theory of nuclear matter. Annual Review of Nuclear Science, 21:93–244, December 1971. CODEN ARNUA8. ISSN 0066-4243.

Bethe:1971:YLC

[Bet71e] Hans A. Bethe. Yalta: Lack of communication on bomb. New York Times, page E12, February 28, 1971. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Bethe:1972:NIB

[Bet72a] Hans A. Bethe. Note on inverse Bremsstrahlung in a strong electromagnetic field. Report LA-5031, US Atomic Energy Commission, Washington, DC, USA, September 1972. URL http://www.osti.

gov/accomplishments/documents/fullText/ACC0102.pdf.

Bethe:1972:PPP

[Bet72b] Hans A. Bethe. Pauli Principle and pion scattering. Report LA-5052, US Atomic Energy Commission, Washington, DC, USA,

October 1972. URL http://www.osti.gov/accomplishments/
documents/fullText/ACC0103.pdf.

Bethe:1973:QTP

[Bet73a] H. A. Bethe. Qualitative theory of pion scattering by nuclei. Physical Review Letters, 30(??):105–108, January 15, 1973. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Bethe:1973:ETB

[Bet73b] Hans A. Bethe. Effective two-body interaction for equivalent particles. In Calogero and Ciofi Degli Atti [CC73], pages 828–?? ISBN 020. LCCN QC793.9 .I73 1972.

Bethe:1973:QTN

[Bet73c] Hans A. Bethe. Qualitative theory of non scattering by nuclei. Physical Review Letters, 30(3):105–108, January 15, 1973. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-

0145. URL http://prl.aps.org/abstract/PRL/v30/i3/p105_
1.

Bethe:1974:ESD

[Bet74] Hans A. Bethe. Equation of state at densities greater than nuclear density. In Hansen [Han74], pages 27–46. ISBN 90-277-0406-6, 90-277-0407-4 (paperback). LCCN QB806 .P48.

Bethe:1975:AEP

[Bet75a] Hans A. Bethe. AEI energy policy. Science, 190(4220):1152, December 19, 1975. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Bethe:1975:F

[Bet75b] Hans A. Bethe. Foreword. In A world destroyed: the atomic bomb and the Grand Alliance [She75], page ?? ISBN 0-394-49794-5. LCCN D753 .S48 1975. US$10.00.

Bethe:1975:FLB

[Bet75c] Hans A. Bethe. Fuel for LWR, breeders and near-breeders. Annals of Nuclear Energy, 2(??):763–766, ???? 1975. CODEN ANENDJ. ISSN 0306-4549 (print), 1873-2100 (electronic).

Bethe:1975:NFC

[Bet75d] Hans A. Bethe. “No fundamental change in the situation” [nuclear reactor safety]. Bulletin of the Atomic Scientists, 31(7):40– 41, September 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bethe:1975:RSR

[Bet75e] Hans A. Bethe. Report of Steering Review Committee of the American Physical Society’s Study Group on Light Water Reactor Safety. Bulletin of the Atomic Scientists, 31(9):35–37, November 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bethe:1976:EP

[Bet76a] Hans Bethe. The energy problem. AIP Conference Proceedings, 28(??):7–??, 1976. CODEN APCPCS. ISSN 0094-243X (print), 1551-7616 (electronic), 1935-0465.

Bethe:1976:CNP

[Bet76b] Hans A. Bethe. The case for nuclear power. Skeptic, 14(??):22, July/August 1976.

Bethe:1976:ES

[Bet76c] Hans A. Bethe. The energy of the stars. Technology Review (M.I.T.), 78(7):58–63, June 1976. CODEN TEREAU. ISSN 0040-1692.

Bethe:1976:F

[Bet76d] Hans A. Bethe. Foreword. In The fight over nuclear power: the energy controversy [SB76], pages vii–?? ISBN 0-87843-002-4. LCCN ???? Foreword by Hans A. Bethe.

Bethe:1976:NFPa

[Bet76e] Hans A. Bethe. The necessity of fission power. Scientific American, 234(1):21–31, January 1976. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://alsos.wlu.

edu/information.aspx?id=3441;   http://www.nature.com/
scientificamerican/journal/v234/n1/pdf/scientificamerican0176-
21.pdf.

Bethe:1976:NJJ

[Bet76f] Hans A. Bethe. Neobhodimost’ jadernoj jenergetiki. (Russian) [Necessity of fission power]. Uspekhi Fizicheskikh Nauk, 120(3):455– 475, November 1976. CODEN UFNAAG. ISSN 0042-1294 (print), 1996-6652 (electronic).

Bethe:1976:TCN

[Bet76g] Hans A. Bethe. Testimony on the California Nuclear Initiative. Energy, 1(??):223–228, ???? 1976.

Bethe:1976:UC

[Bet76h] Hans A. Bethe. Ultimate catastrophe? Bulletin of the Atomic Scientists, 32(6):36–37, June 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Dud75].

Bethe:1977:DNP

[Bet77a] Hans A. Bethe. The debate on nuclear power. In ????, editor, Nuclear Power and Its Fuel Cycle, volume 7, pages 3–17. International Atomic Energy Agency, Vienna, Austria, 1977.

Bethe:1977:DA

[Bet77b] Hans A. Bethe. Dinner address. Annals of the New York Academy of Sciences, 294(1):98–102, November 1977. CODEN ANYAA9. ISSN 0077-8923 (print), 1749-6632 (electronic).

Bethe:1977:ES

[Bet77c] Hans A. Bethe. Energy strategy. Foreign Affairs, 55(3):636–637, ???? 1977. CODEN FRNAA3. ISSN 0015-7120.

Bethe:1977:RNP

[Bet77d] Hans A. Bethe. Response: Nuclear power: a safe bet. Bulletin of the Atomic Scientists, 33(7):55, September 1977. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bet77e, Sel77, Fes77].

Bethe:1977:SVA

[Bet77e] Hans A. Bethe. Six views on atomic energy: The need for nuclear power. Bulletin of the Atomic Scientists, 33(3):59–63, March 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Sel77, Bet77d].

Bethe:1977:SPN

[Bet77f] Hans Albrecht Bethe. The scattering of pions by nuclei. In Aitchison and Paton [AP77], pages 1–17. ISBN 0-08-020606-9. LCCN QC1 .P3 1974.

Bethe:1978:CEI

[Bet78a] Hans A. Bethe. Cornell Engineer (interview with Charles D. Pevsner). Cornell Engineer, 44(??):14–20, November 1978. CODEN CNLEAC. ISSN 0010-8790.

Bethe:1978:FH

[Bet78b] Hans A. Bethe. The fusion hybrid. Nuclear News, 21(7):41, May 1978. CODEN NUNWA8. ISSN 0029-5574.

Bethe:1979:BRF

[Bet79a] Hans Bethe. Book review: Freeman Dyson’s view of his times: F. Dyson, Disturbing the Universe. Physics Today, 32(12):51–52, December 1979. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-

0699 (electronic). URL http://link.aip.org/link/phtoad/
v32/i12/p51/s1.

Bethe:1979:FH

[Bet79b] Hans A. Bethe. The fusion hybrid. Physics Today, 32(5):44–51, May 1979. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://link.aip.org/link/phtoad/v32/i5/
p44/s1.

Bethe:1979:HT

[Bet79c] Hans A. Bethe. The happy thirties. In Stuewer [Stu79], pages 9–31. ISBN 0-8166-0869-5. LCCN QC773 .S95 1977.

Bethe:1980:IHB

[Bet80a] H. Bethe. An interview with Hans Bethe. Center Magazine, 13 (3):23–27, ???? 1980. ISSN 0008-9125.

Bethe:1980:CCN

[Bet80b] Hans A. Bethe. The case for coal and nuclear energy. The Center Magazine, 13(??):14–22, May/June 1980. CODEN ???? ISSN ????

Bethe:1980:FES

[Bet80c] Hans A. Bethe. Fulfilling energy strategy. New York Times, page A19, July 1, 1980. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Bethe:1980:FHR

[Bet80d] Hans A. Bethe. The fusion hybrid reactor. Report SAND 81-1265, Sandia National Laboratories, Albuquerque, NM, August 19, 1980.

Colloquium at Sandia National Laboratories, Albuquerque, New Mexico, September 19, 1980.

Bethe:1980:LS

[Bet80e] Hans A. Bethe. The lives of the stars. The Sciences (New York), 20(8):6–9, ???? 1980. CODEN SCNCAD. ISSN 0036-861X.

Bethe:1980:RSS

[Bet80f] Hans A. Bethe. Recollections of solid-state theory, 1926–33. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 371(1744):49–51, ???? 1980. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-9169 (electronic).

Bethe:1981:EMS

[Bet81a] Hans A. Bethe. Editorial: Meaningless superiority. Bulletin of the Atomic Scientists, 37(8):1, 4, October 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bethe:1981:FHR

[Bet81b] Hans A. Bethe. Fusion hybrid reactor. Report SAND-81-1265, Sandia National Laboratories, Albuquerque, NM, USA,

August 1981. URL http://www.osti.gov/accomplishments/
documents/fullText/ACC0104.pdf.

Bethe:1981:MLP

[Bet81c] Hans A. Bethe. My life as a physicist. Talk at Erice, Sicily, Italy., April 1981.

Bethe:1981:SCE

[Bet81d] Hans A. Bethe. Supernova collapse and explosion. Bulletin of the Atomic Scientists, 26(1):17, January 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bethe:1981:EWC

[Bet81e] Hans A. (Hans Albrecht) Bethe. Energy and the world community $h [sound recording]. 2 sound cassettes, 1981.

Bethe:1982:AR

[Bet82a] Hans A. Bethe. The arms race. Colloquium at Sandia National Laboratories, Albuquerque, New Mexico, July 28, 1982.

Bethe:1982:CHH

[Bet82b] Hans A. Bethe. Comments on the history of the H-bomb. Los Alamos Science, 6:42–53, Fall 1982. CODEN LASCDI. ISSN

0273-7116. URL http://library.lanl.gov/cgi-bin/getfile?

06-03.pdf. This article corrects a few errors in one book [Goo80, Goo81], and points out numerous errors, fallacies, misrepresentations, and misunderstandings, in an earlier book [SB54]. It also contains a substantial discussion of the role of Edward Teller in the work as Los Alamos, and later, at Lawrence Livermore National Laboratory.

Bethe:1982:EIP

[Bet82c] Hans A. Bethe. Energy independence is possible. E. L. Miller Lecture on Public Policy, Texas A&M University, College Station, TX, USA., October 8, 1982.

Bethe:1982:HBH

[Bet82d] Hans A. Bethe. Hydrogen bomb history. Science, 218(4579): 1270, December 24, 1982. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/218/4579/1270.2.full.pdf.

Bethe:1982:ICA

[Bet82e] Hans A. Bethe. The inferiority complex and American nuclear policy. New York Review of Books, 29(10):3, June 10, 1982. ISSN 0028-7504 (print), 1944-7744 (electronic).

URL http://www.nybooks.com/articles/archives/1982/jun/
10/the-inferiority-complex/. Testimony to the US Senate For-

eign Relations Committee on May 13, 1982.

Bethe:1982:ODH

[Bet82f] Hans A. Bethe. Observations on the development of the H-bomb. Los Alamos Science, ??(??):43–??, Fall 1982. CODEN LASCDI. ISSN 0273-7116. Edited publication of unpublished work [Bet54e].

Bethe:1982:SSN

[Bet82g] Hans A. Bethe. Supernova shocks and neutrino diffusion. In Rees and Stoneham [RS82], pages 35–52. ISBN 90-277-1442-8. LCCN QB843.S95 N37 1981.

Bethe:1982:ST

[Bet82h] Hans A. Bethe. Supernova theory. In Barnes et al. [BCSF82], pages 439–466. ISBN 0-521-24410-2, 0-521-28876-2 (paperback).

LCCN QB464 .E87 1982. Presented to William A. Fowler, on the occasion of his 70th birthday.

Bethe:1982:UET

[Bet82i] Hans A. Bethe. Uranium enrichment technology. Science, 217 (4558):398, July 30, 1982. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/217/4558/398.1.full.pdf.

Bethe:1982:WIS

[Bet82j] Hans A. Bethe. We are not inferior to the Soviets. In Proceedings of the Symposium of the American Physical Society’s Forum on Physics and Society, April, 1982, page ?? American Physical Society, Ridge, NY 11961, USA, 1982. URL www.aps.org/units/

fps/newsletters/upload/july83.pdf.

Bethe:1983:EEP

[Bet83a] H. A. Bethe. Erratum: Energy production in stars. Science, 219 (4582):238, January 21, 1983. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/219/4582/238.3.full.pdf. See [Bet68e]. A figure

of 1500 billion years in that article should be 15,000 billion.

Bethe:1983:NAE

[Bet83b] Hans A. Bethe. Nuclear arms education: the urgent need. Physics Today, 36(3):120, March 1983. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.aip.org/

link/phtoad/v36/i3/p120/s1.

Bethe:1983:WIS

[Bet83c] Hans A. Bethe. We are not inferior to the Soviets. In Kenneth W. Ford, editor, Nuclear Weapons and Nuclear War, pages 9–18. American Association of Physics Teachers, SUNY, Stony Brook, NY, USA, 1983. Papers based on a symposium of the Forum on Physics and Society of the American Physical Society held in Washington, D.C., in April 1982.

Bethe:1984:ELA

[Bet84a] H. Bethe. Exchange on laser ABM — laser space weaponry — reply. Laser Focus/Electro-Optics Magazine, 20(8):16–??, ???? 1984. ISSN 0275-1399.

Bethe:1984:CAL

[Bet84b] Hans A. Bethe. The case against laser ABM development technology and policy — views of Hans Bethe. Laser Focus/Electro-Optics Magazine, 20(3):10–??, March 20, 1984. ISSN 0275-1399.

Bethe:1984:WHN

[Bet84c] Hans A. Bethe. What holds the nucleus together? Scientific American, 251(4):39–49, October 1984. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.nature.com/scientificamerican/journal/v251/n4/pdf/
scientificamerican1084-39.pdf.

Bethe:1984:MHL

[Bet84d] Hans A. (Hans Albrecht) Bethe. Management of high level radioactive waste. 2 sound cassettes (95 min.), 1984.

Bethe:1985:NBH

[Bet85a] Hans Bethe. Niels Bohr and his institute. In French and Kennedy [FK85], pages 232–234. ISBN 0-674-62415-7, 0-674-62416-5 (paperback). LCCN QC16.B63 N49 1985. US$27.50.

Bethe:1985:ST

[Bet85b] Hans A. Bethe. Supernova theory. In Molinari and Ricci [MR85], pages 181–209. ISBN 0-444-86988-3 (U.S.). LCCN QC770 .I645 1984.

Bethe:1985:TI

[Bet85c] Hans A. Bethe. The technological imperative. Bulletin of the Atomic Scientists, 41(7):34–36, August 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bethe:1985:TNT

[Bet85d] Hans A. Bethe. There is no technological solution. Quest for Peace, ??(??):??, ???? 1985. URL http://www.lib.uci.edu/

quest/index.php?page=bethe.

Bethe:1986:PES

[Bet86a] H. A. Bethe. Possible explanation of the solar-neutrino puzzle. Physical Review Letters, 56(12):1305–1308, March 24, 1986. CO-DEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Bethe:1986:ST

[Bet86b] Hans A. Bethe. Supernova theory. In Shapiro and Teukolsky [ST86], pages 45–83. ISBN 0-471-82421-6. LCCN QB461 .H54 1986. US$35.95 (est.). Derived from a symposium held at Cornell University in October 1984, in honor of E. S. Salpeter.

Bethe:1986:UPA

[Bet86c] Hans A. Bethe. U.S. panel assesses Chernobyl. Bulletin of the Atomic Scientists, 42(10):45–46, December 1986. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Bla87].

Bethe:1987:DSG

[Bet87a] H. Bethe. The DEW study goes public — reply. Physics Today, 40(10):13, October 1987. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Bethe:1987:ST

[Bet87b] Hans A. Bethe. Supernova theory. In Metropolis et al. [MKR87], pages 235–256. ISBN 0-12-492155-8. LCCN QC44 .N49

1987. URL http://www.osti.gov/energycitations/product.
biblio.jsp?osti_id=6120718&query_id=0.

Bethe:1987:TS

[Bet87c] Hans A. Bethe. The theory of supernovae. In Kuo and Speth [KS87], pages 3–12. ISBN 0-444-87064-4 (set), 0-444-87065-2 (v. 1), 0-444-87066-0 (v. 2). LCCN QC174.17.P7 I56 1986.

Bethe:1988:RRP

[Bet88a] H. A. Bethe. Reminiscences of Rudi Peierls. In A breadth of physics (Oxford, 1987), pages 223–228. World Sci. Publ., Teaneck, NJ, 1988.

Bethe:1988:LS

[Bet88b] Hans A. Bethe. The life of the stars. AIP Conference Proceedings, 169(??):49–??, 1988. CODEN APCPCS. ISSN 0094-243X (print), 1551-7616 (electronic), 1935-0465.

Bethe:1988:NPN

[Bet88c] Hans A. Bethe. Nuclear physics needed for the theory of supernovae. Annual Review of Nuclear and Particle Science, 38:1–29, December 1988. CODEN ARPSDF. ISSN 0163-8998 (print), 1545-4134 (electronic). Reprinted in Autobiographies by Nobel Laureates in Physics Vol. 1 (2009): 1–29.

Bethe:1988:ORP

[Bet88d] Hans A. Bethe. Obituary: Richard Phillips Feynman (1918– 1988). Nature, 332(6165):588, April 14, 1988. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.

nature.com/nature/journal/v332/n6165/332588a0.

Bethe:1989:SNE

[Bet89a] H. A. Bethe. Solar-neutrino experiments. Physical Review Letters, 63(8):837–839, August 21, 1989. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Bethe:1989:BRP

[Bet89b] Hans A. Bethe. Book review: The Privilege of Being a Physicist, by Victor F. Weisskopf, (New York: Freeman, 1989). American Journal of Physics, 57(10):957, October 1989. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://

aapt.scitation.org/doi/10.1119/1.15808;   http://link.
aip.org/link/ajpias/v57/i10/p957/s1.

Bethe:1989:CNA

[Bet89c] Hans A. Bethe. Chop down the nuclear arsenals. Bulletin of the Atomic Scientists, 45(2):11–15, March 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Mol89].

Bethe:1989:LBM

[Bet89d] Hans A. Bethe. Letter: Better missiles. Bulletin of the Atomic Scientists, 45(7):44, September 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bethe:1989:ODH

[Bet89e] Hans A. Bethe. Observations on the development of the Hbomb. In The advisors: Oppenheimer, Teller, and the superbomb [Yor89], pages 163–181. ISBN 0-8047-1713-3, 0-8047-1714-1 (paperback). LCCN UG1282.A8 Y67 1989. URL http://www.

fas.org/nuke/guide/usa/nuclear/bethe-54.htm. Appendix II.

Written in 1954, but unpublished until the 1989 edition of this book.

Bethe:198x:ACV

[Bet8xa] Hans A. Bethe. Arms control verification. Found in USGPO Congressional Record document list; publication year and status unknown, 198x.

Bethe:198x:TPR

[Bet8xb] Hans A. Bethe. Technological problems related to Strategic Defense Initiative. Found in USGPO Congressional Record document list; publication year and status unknown, 198x.

Bethe:1990:SM

[Bet90a] H. A. Bethe. Supernova mechanisms. Reviews of Modern Physics, 62(4):801–866, October 1, 1990. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.62.801;
 http://rmp.aps.org/abstract/RMP/v62/i4/p801_1.

Bethe:1990:GM

[Bet90b] Hans A. Bethe. Book review: The Genocidal Mentality, edited by Robert J. Lifton & Eric Markusen (New York: Basic Books, 1990). Issues in Science and Technology, 7(1):97–98, Fall 1990. CODEN ???? ISSN 0748-5492 (print), 1938-1557 (electronic).

Bethe:1990:HSS

[Bet90c] Hans A. Bethe. How supernova shock revival was revealed. Physics Today, 43(12):91, December 1990. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.aip.org/

link/phtoad/v43/i12/p91/s1.

Bethe:1990:NHS

[Bet90d] Hans A. Bethe. Nuclear history: Sakharov’s H-bomb. Bulletin of the Atomic Scientists, 46(8):8–9, October 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

alsos.wlu.edu/information.aspx?id=2432.

Bethe:1990:S

[Bet90e] Hans A. Bethe. Supernovae. Physics Today, 43(9):24–27, September 1990. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://link.aip.org/link/phtoad/v43/i9/

p24/s1. See comments and reply [DSB91].

Bethe:1991:BES

[Bet91a] H. A. Bethe. Binding energies and supernova evolution — reply. Physics Today, 44(4):15–??, April 1991. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Bethe:1991:BRB

[Bet91b] Hans Bethe. Book review: The Joy of Insight: Passions of a Physicist, by Victor Weisskopf, ISBN 0-465-03678-3. Physics Today, 44 (10):111–112, October 1991. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.physicstoday.

org/resource/1/phtoad/v44/i10/p111_s1.

Bethe:1991:CHH

[Bet91c] Hans Bethe. Comments on the history of the H-bomb, 1954. In Cantelon et al. [CHW91], pages 129–138. ISBN 0-8122-3096-5 (hardcover), 0-8122-1354-8 (paper). LCCN UA23 .A597 1991.

Bethe:1991:CIC

[Bet91d] Hans A. Bethe. Chernobyl: It can’t happen here. New York Times, page A25, May 2, 1991. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Bethe:1991:JRO

[Bet91e] Hans A. Bethe. J. Robert Oppenheimer. In The road from Los Alamos: Profound perspective and personal viewpoints on atomic weapons, nuclear power, and science [Bet91j], pages 221– 230. ISBN 0-88318-707-8, 0-671-74012-1. LCCN U264 .B455 1991. US$24.95.

Bethe:1991:NNP

[Bet91f] Hans A. Bethe. The need for nuclear power. Engineering & Science, California Institute of Technology, 54(3):28–35, Spring 1991. CODEN ENSCAT. ISSN 0013-7812. URL http://resolver.

caltech.edu/CaltechES:54.3.Bethe.

Bethe:1991:RPF

[Bet91g] Hans A. Bethe. Richard P. Feynman. In The road from Los Alamos: Profound perspective and personal viewpoints on atomic weapons, nuclear power, and science [Bet91j], pages 239–244. ISBN 0-88318-707-8, 0-671-74012-1. LCCN U264 .B455 1991. US$24.95.

Bethe:1991:SHB

[Bet91h] Hans A. Bethe. Sakharov’s H-bomb. In Drell and Kapitsa [DK91], pages 149–152. ISBN 0-88318-853-8 (hardcover), 0-88318-852-X (paperback). LCCN QC16.S255 S25 1991.

Bethe:1991:TAC

[Bet91i] Hans A. Bethe. The truth about Chernobyl. New York Times, page BR34, May 5, 1991. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Bethe:1991:RAP

[Bet91j] Hans Albrecht Bethe. The road from Los Alamos: Profound perspective and personal viewpoints on atomic weapons, nuclear power, and science. Masters of modern physics. American Institute of Physics, Woodbury, NY, USA, 1991. ISBN 0-88318-707-8, 0-671-74012-1. xvii + 286 pp. LCCN U264 .B455 1991. US$24.95.

Bethe:1992:SN

[Bet92] Hans A. Bethe. Solar neutrinos. International Journal of Modern Physics. D, Gravitation, Astrophysics, Cosmology, 1(1):1–12, ???? 1992. CODEN IMPDEO. ISSN 0218-2718. URL http://www.

worldscientific.com/doi/abs/10.1142/S0218271892000021.

Bethe:1993:BRS

[Bet93a] Hans Bethe. Book review: Szilard worked first to build the bomb and then to oppose it: William Lanouette and Bela Szilard, Genius in the Shadows: a Biography of Leo Szilard, the Man Behind the Bomb. Physics Today, 46(9):63–64, September 1993. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

URL http://www.physicstoday.org/resource/1/phtoad/v46/
i9/p63_s1.

Bethe:1993:BRF

[Bet93b] Hans A. Bethe. Book review: The Farm Hall Transcripts. American Journal of Physics, 61(3):??, March 1993. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Bethe:1993:LBG

[Bet93c] Hans A. Bethe. Letter: Bethe on the German bomb program. Bulletin of the Atomic Scientists, 49(1):53–54, January/February 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See corrections [Ano93].

Bethe:1993:LPV

[Bet93d] Hans A. Bethe. Life and physics with Viki [Weisskopf]. In ????, editor, Achievements in Physics: The International Community of Physicists (Honoring the lifetime achievements of V. F. Weisskopf), page ?? American Institute of Physics, Woodbury, NY, USA, 1993.

Bethe:1993:MET

[Bet93e] Hans A. Bethe. “my experience in teaching physics,” Hans Bethe’s acceptance speech for the 1993 Oersted Medal presented by the American Association of Physics Teachers, 13 April 1993. American Journal of Physics, 61(11):972–973, November 1993. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL

http://link.aip.org/link/ajpias/v61/i11/p972/s1.

Bethe:1993:NLN

[Bet93f] Hans A. Bethe. New lives for nuclear labs. New York Times, page 17, December 6, 1993. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Bethe:1993:PSS

[Bet93g] Hans A. Bethe. Preservation of the supernova shock. Astrophysical Journal, 419(1):197–199, ???? 1993. CODEN ASJOAB. ISSN 0004-637X (print), 1538-4357 (electronic).

Bethe:1993:SEA

[Bet93h] Hans A. Bethe. Supernova 1987A: An empirical and analytic approach. Astrophysical Journal, 412(1):192–202, ???? 1993. CO-DEN ASJOAB. ISSN 0004-637X (print), 1538-4357 (electronic).

Bethe:1993:WFA

[Bet93i] Hans A. Bethe. What is the future of Los Alamos? Los Alamos Science, 21:4–30, 1993. CODEN LASCDI. ISSN 0273-7116. URL

http://library.lanl.gov/cgi-bin/getfile?21-02.pdf.

Bethe:1993:DHB

[Bet93j] Hans A. (Hans Albrecht) Bethe. Dr. Hans Bethe, super nova. 1 videocassette (69 min.)., 1993.

Bethe:1993:FAC

[Bet93k] Hans Albrecht Bethe. Feynman in Los Alamos and Cornell. In Brown and Rigden [BR93], pages 33–36. ISBN 0-88318-870-8. LCCN QC16.F49 A3 1993. Commentary by Joan Feynman, John Wheeler, Hans Bethe, Julian Schwinger, Murray Gell-Mann, Daniel Hillis, David Goodstein, Freeman Dyson, and Laurie Brown.

Bethe:1994:BAC

[Bet94a] Hans A. Bethe. Behind the atomic curtain: Book review: Stalin and the Bomb: the Soviet Union and Atomic Energy 1939–1956, by David Holloway. Nature, 372(6503):281–283, November 17, 1994. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v372/
n6503/pdf/372281a0.pdf.

Bethe:1994:MSB

[Bet94b] Hans A. Bethe. Mechanism of supernovae: The Bakerian Lecture, 1993. Philosophical transactions of the Royal Society of London, A346(1680):251–258, February 15, 1994. CODEN PTRSAV. ISSN 0370-2316.

Bethe:1994:PES

[Bet94c] Hans A. Bethe. Possible explanation of the solar-neutrino puzzle. In John N. Bahcall, Raymond Davis, Jr., Peter Parker, Alexei Smimov, and Roger Ulrich, editors, Solar Neutrinos. The First Thirty Years, pages 65–68. Addison-Wesley, Reading, MA, USA, 1994.

Bethe:1994:SLN

[Bet94d] Hans A. Bethe. A scenario for a large number of low mass black holes in the galaxy. Astrophysical Journal, 423(??):659–664, ???? 1994. CODEN ASJOAB. ISSN 0004-637X (print), 1538-4357 (electronic).

Bethe:1994:ST

[Bet94e] Hans A. Bethe. Supernova theory. In Ekspong [Eks94], pages 17–26. ISBN 981-02-1450-2 (hardcover). LCCN QC71 .O84 1991.

Bethe:1994:TNS

[Bet94f] Hans A. Bethe. Theory of neutrinos from the Sun. In Ekspong [Eks94], pages 1–16. ISBN 981-02-1450-2 (hardcover). LCCN QC71 .O84 1991.

Bethe:1994:MPS

[Bet94g] Hans A. (Hans Albrecht) Bethe. Manhattan Project: a scientist’s point of view. 1 sound cassette (76 min.), 1994.

Bethe:1995:BRB

[Bet95a] Hans A. Bethe. Book review: Dark Sun: The Making of the Hydrogen Bomb. Science, 269(5225):1455–1457, September 8, 1995. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Bethe:1995:BRO

[Bet95b] Hans A. Bethe. Book review: Old friends among the giants on whose shoulders physics stands: H. Henry Stroke, The Physical Review: The First Hundred Years. A Selection of Seminal Papers and Commentaries. Physics Today, 48(11):77, November 1995. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

URL http://link.aip.org/link/phtoad/v48/i11/p77/s1.

Bethe:1995:LCD

[Bet95c] Hans A. Bethe. Letter: Cease and desist. Bulletin of the Atomic Scientists, 51(6):3, November/December 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bethe:1995:PIR

[Bet95d] Hans A. Bethe. Public interest report: Letter to the science community. Journal of the Federation of American Scientists, 48 (5):8, September/October 1995. URL http://www.pugwash.org/

about/bethe.htm.

Bethe:1995:SS

[Bet95e] Hans A. Bethe. The supernova shock. Astrophysical Journal, 449 (2):714–726, ???? 1995. CODEN ASJOAB. ISSN 0004-637X (print), 1538-4357 (electronic).

Bethe:1996:BSS

[Bet96a] Hans A. Bethe. Breakout of the supernova shock. Astrophysical Journal, 469(2):737–739, ???? 1996. CODEN ASJOAB. ISSN 0004-637X (print), 1538-4357 (electronic).

Bethe:1996:CSS

[Bet96b] Hans A. Bethe. Constraints on the supernova shock. Astrophysical Journal, ??(??):??, August 1996. CODEN ASJOAB. ISSN 0004-637X (print), 1538-4357 (electronic).

Bethe:1996:SSV

[Bet96c] Hans A. Bethe. The supernova shock VI. Astrophysical Journal, 473(1):343–346, ???? 1996. CODEN ASJOAB. ISSN 0004-637X (print), 1538-4357 (electronic).

Bethe:1996:ST

[Bet96d] Hans A. Bethe. Supernova theory. Nuclear Physics A, 606(1–2):95– 117, August 26, 1996. CODEN NUPABL. ISSN 0375-9474 (print),

1873-1554 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0375947496002345.

Bethe:1997:IGE

[Bet97a] Hans Bethe. Influence of Gamow on early astrophysics and on early accelerators in nuclear physics. In Harper et al. [HPA97], pages 44–48. ISBN 1-886733-49-X. LCCN QC16.G37 G46 1996.

URL http://adsabs.harvard.edu/abs/1997ASPC..129...44B.

Bethe:1997:JRO

[Bet97b] Hans A. Bethe. J. Robert Oppenheimer: April 22, 1904– February 18, 1967. Biographical memoirs — National Academy of Sciences of the United States of America, 71:175–218, 1997. CODEN BMNSAC. ISBN 0-309-59031-0. ISSN 0077-2933.

URL http://alsos.wlu.edu/information.aspx?id=3181;
http://www.nap.edu/catalog.php?record_id=5737;   http:
//www.nap.edu/catalog/5737.html.

Bethe:1997:LPW

[Bet97c] Hans A. Bethe. Letter to President William J. Clinton. On behalf of the Federation of Atomic Scientists. See response [Cli97]., April

25, 1997. URL http://www.fas.org/bethecr.htm.

Bethe:1997:SSV

[Bet97d] Hans A. Bethe. Supernova shock VIII. Astrophysical Journal, 490(2):765–771, ???? 1997. CODEN ASJOAB. ISSN 0004-637X (print), 1538-4357 (electronic).

Bethe:1997:SWH

[Bet97e] Hans Albrecht Bethe. Selected works of Hans A. Bethe: with commentary, volume 18 of World Scientific series in 20th century physics. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1997. ISBN 981-02-2876-7. viii + 605 pp. LCCN QC780 .B452 1996.

URL http://www.zentralblattmath.org/zmath/en/search/?

an=0916.01027. English translations (when originals are in German) of 28 of the author’s papers, accompanied by a list of 290 of his publications.

Bethe:1998:TSW

[Bet98a] H. A. Bethe. On the theory of shock waves for an arbitrary equation of state [Rep. No. 545, Serial No. NDRC-B-237, Office Sci. Res. Develop., U. S. Army Ballistic Research Laboratory, Aberdeen Proving Ground, MD, 1942]. In Classic papers in shock compression science, High-press. Shock Compression Condens. Matter, pages 421–492. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998.

Bethe:1998:BRA

[Bet98b] Hans A. Bethe. Book review: Andrew Brown, The Neutron and the Bomb: a Biography of Sir James Chadwick. American Journal of Physics, 66(5):454–??, May 1998. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Bethe:1998:NNP

[Bet98c] Hans A. Bethe. The need for nuclear power. In Kursunoğlu et al. [KMP98], pages 39–41. ISBN 0-306-45921-3.

LCCN TK9006 .E54 1998. URL http://www.loc.gov/catdir/
enhancements/fy0821/98041249-d.html;   http://www.loc.
gov/catdir/enhancements/fy0821/98041249-t.html.

Bethe:1999:SIH

[Bet99a] H. A. Bethe. Special issue in honor of the centenary of the American Physical Society — March, 1999 — introduction. Reviews of Modern Physics, 71(2):U6, ???? 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

Bethe:1999:FMS

[Bet99b] Hans A. Bethe. Fast-moving science befuddles predictions. New Perspectives Quarterly: NPQ, 16(4):43–44, Fall 1999. CODEN ???? ISSN 0893-7850 (print), 1540-5842 (electronic).

Bethe:1999:NP

[Bet99c] Hans A. Bethe. Nuclear physics. Reviews of Modern Physics, 71(2):S6–S15, March 1, 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.71.S6;   http://
rmp.aps.org/abstract/RMP/v71/i2/pS6_1. Special Issue in

Honor of the Centenary of the American Physical Society.

Bethe:1999:QPM

[Bet99d] Hans A. Bethe. Quantum physics made relatively simple: Three lectures by Hans Bethe. Web document and 173 minutes of video

recording, 1999. URL http://bethe.cornell.edu/index.html.

Bethe:1999:QT

[Bet99e] Hans A. Bethe. Quantum theory. Reviews of Modern Physics, 71(2):S1–S5, March 1, 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.71.S1;   http://
rmp.aps.org/abstract/RMP/v71/i2/pS1_1.

Bethe:1999:TB

[Bet99f] Hans A. Bethe. The treaty betrayed. New York Review of Books, 46(18):6, November 18, 1999. ISSN 0028-7504 (print), 1944-

7744 (electronic). URL http://www.nybooks.com/articles/
archives/1999/nov/18/the-treaty-betrayed/. Comments on

the US Senate’s failure to ratify the Comprehensive Test Ban Treaty.

Bethe:2000:GUP

[Bet00a] Hans A. Bethe. The German Uranium Project. Physics Today, 53(7):34–36, July 2000. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://alsos.wlu.edu/

information.aspx?id=2657;   http://www.physicstoday.org/
resource/1/phtoad/v53/i7/p34_s1.

Bethe:2000:RWP

[Bet00b] Hans A. Bethe. Remembering Wolfgang Pauli / Erinnerungen an Wolfgang Pauli. In Unser [Uns00], pages 85–90. ISBN 3-9521386-2-2. LCCN QC16.P37 W635 2000.

Bethe:2000:SS

[Bet00c] Hans A. Bethe. Sommerfeld’s seminar. Physics in Perspective (PIP), 2(1):3–5, 2000. CODEN PHPEF2. ISSN 1422-6944 (print),

1422-6960 (electronic). URL http://www.springerlink.com/
content/2h5uxxgpn6pvngx7/fulltext.pdf.

Bethe:2001:BRBa

[Bet01a] Hans A. Bethe. Book review: The Magic Furnace: The Search for the Origin of Atoms, by Marcus Chown. Nature, 411(6840): 888, June 21, 2001. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v411/n6840/full/411888a0.html.

Bethe:2001:ETL

[Bet01b] Hans A. Bethe. Edward Teller: A long look back: Book review: Edward Teller, Memoirs: A Twentieth-Century Journey in Science and Politics. Physics Today, 54(11):55–56, November 2001. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

URL http://link.aip.org/link/phtoad/v54/i11/p55/s1.

Bethe:2001:F

[Bet01c] Hans A. Bethe. Foreword. In Guran [Gur01], page ?? ISBN 981-02-2964-X. LCCN ???? Foreword by Hans A. Bethe.

Bethe:2001:WKP

[Bet01d] Hans A. Bethe. Wie die Kernspaltung die Physiker entzweite. (German) [As the fission physicists quarreled]. Süddeutsche Zeitung, ?? (??):??, December 11, 2001.

Bethe:2001:N

[Bet01e] Hans Albrecht Bethe. Neutron. In Hoiberg [Hoi01], pages 84–96. ISBN 81-88237-00-0. LCCN AS911.N9 A15 2001.

Bethe:2003:MLA

[Bet03] Hans A. Bethe. My life in astrophysics. Annual Review of Astronomy and Astrophysics, 41:1–14, 2003. CODEN ARAAAJ. ISSN 0066-4146 (print), 1545-4282 (electronic). URL http://adsabs.

harvard.edu/abs/2003ARA%26A..41....1B. Reprinted in book-

titleAutobiographies by Nobel Laureates in Physics, Vol. 1 (2009): 1–14, and in [Bet06].

Bethe:2004:PR

[Bet04] Hans Bethe. Pilgrimages to Rome. In Orear [Ore04], chapter 18, pages 75–82. ISBN ???? LCCN ???? URL http://dspace.

library.cornell.edu/handle/1813/62.

Bethe:2005:DU

[Bet05a] Hans A. Bethe. Das deutsche Uranprojekt. (German) [The German Uranium Project]. In Kleint et al. [KRW05], pages 201– 203. ISBN 3-7776-1402-5. ISSN 0365-6470. LCCN QC16.H35

W474 2005. URL http://www.loc.gov/catdir/toc/fy0613/
2006402966.html.

Bethe:2005:OSC

[Bet05b] Hans A. Bethe. Obituary: Subramanyan Chandrasekhar (1910– 95). Nature, 377(6549):484, October 12, 2005. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.

nature.com/nature/journal/v377/n6549/pdf/377484a0.pdf.

Bethe:2006:MLA

[Bet06] Hans A. Bethe. My life in astrophysics. In Brown and Lee [BL06a], pages 27–43. ISBN 981-256-609-0, 981-256-610-4 (paperback), 981-277-450-5 (e-book). LCCN QC16.B46 H36 2006. URL http://

www.worldscibooks.com/physics/5989.html;   http://www.
worldscientific.com/worldscibooks/10.1142/5989.

Bethe:2014:BFE

[Bet14] Hans Bethe. Braking formula for electrons of relativistic speed. European Physical Journal H, 39(5):537–542, December 2014. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (elec-

tronic). URL http://link.springer.com/article/10.1140/

epjh/e2014-50028-5. See annotation [FBB14].

Bethe:2020:CAW

[Bet20] H. A. Bethe. 1946: Can air or water be exploded? Bulletin of the Atomic Scientists, 76(6):341–343, 2020. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/00963402.2020.1847466.

Beyea:1980:CDI

[Bey80] Jan Beyea. Commentary: Dispute at Indian Point. Bulletin of the Atomic Scientists, 36(5):63–64, May 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [BWH80].

Bethe:1932:WZE

[BF32] H. Bethe and E. Fermi. Über die Wechselwirkung von zwei Elektronen. (German) [On the interaction of two electrons]. Zeitschrift für Physik, 77(5–6):296–306, May 1932. CODEN ZEPYAA.

ISSN ????  URL http://www.springerlink.com/content/
l561955527vv4513/.

Bethe:1933:MWM

[BF33] H. Bethe and H. Fröhlich. Magnetische Wechselwirkung der Metallelektronen. Zur Kritik der Theorie der Supraleitung von Frenkel. (German) [Magnetic interaction of the metal electrons.

On a criticism of the theory of superconductivity by Frenkel]. Zeitschrift für Physik, 85(5–6):389–397, May 1933. CODEN

ZEPYAA. ISSN ???? URL http://www.springerlink.com/
content/m5w484m248g01q45/.

Bethe:1985:BSM

[BFG^{+}85] H. Bethe, R. Forsberg, R. Garwin, W. A. Higinbotham, P. Morrison, J. A. Simpson, C. Smith, and R. Wilson. The Bulletin and the scientists’ movement. Bulletin of the Atomic Scientists, 41(11): 19–31, December 1985. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bethe:1947:BW

[BFH^{+}47] Hans A. Bethe, Klaus Fuchs, Joseph O. Hirschfelder, John L. Magee, Rudolf E. Peierls, and John von Neumann. Blast wave. U.S. Government Document LA-2000 (TID-4500, 13th edition, supplement), Los Alamos Scientific Laboratory, Los Alamos, NM, USA, August 1947. 301 pp. URL https://fas.org/sgp/

othergov/doe/lanl/docs1/00320773.pdf. First distributed 27

March 1958.

Bethe:1949:PDH

[BFK49] H. A. Bethe, U. Fano, and P. R. Karr. Penetration and diffusion of hard X-rays through thick barriers. I. the approach to spectral equilibrium. Physical Review, 76(4):538–540, August 15, 1949. CO-DEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

URL http://prola.aps.org/abstract/PR/v76/i4/p538_1.

Bethe:1944:LSH

[BFvN^{+}44] Hans A. Bethe, Klaus Fuchs, John von Neumann, Rudolf Peierls, and W. G. Penney. Lectures in Shock Hydrodynamics and Blast Waves. ????, October 28, 1944. Notes by J. O. Hirschfelder.

Bethe:1957:ERC

[BG57] H. A. Bethe and J. Goldstone. Effect of a repulsive core in the theory of complex nuclei. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 238:551–567, 1957. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-9169 (electronic).

Bethe:1982:ARD

[BG82a] Hans A. Bethe and K. Gottfried. Assessing Reagan’s Doomesday scenario. New York Times, page E17, April 11, 1982. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Bethe:1982:YWP

[BG82b] Hans A. Bethe and Kurt Gottfried. The 5-year war plan — a loser both ways. New York Times, page A31, June 10, 1982. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Bethe:1983:ESD

[BG83a] H. A. Bethe and K. Gottfried. Elusive security: Do we need more nukes or fewer? Washington Post, page ??, February 6, 1983. ISSN 0190-8286.

Bethe:1983:SUA

[BG83b] Hans A. Bethe and Kurt Gottfried. Save us all, Congress: no weapons in space, no unratified treaties. Washington Post, ??(??): ??, February 6, 1983. ISSN 0190-8286.

Bethe:1985:WSA

[BG85] Hans A. Bethe and Richard L. Garwin. Weapons in space: Appendix A: New BMD technologies. Dædalus, 114(3):331–368, Summer 1985. CODEN DAEDAU. ISSN 0011-5266 (print),

1548-6192 (electronic). URL http://www.jstor.org/stable/

20024997. Special issue on the Strategic Defense Initiative. Volume 1. Concepts and technologies. Volume 2. Implications for security.

Bethe:1991:PNP

[BG91a] Hans A. Bethe and Grigor A. Gurzadyan. On the possibility of nuclear processes in atmospheres of flare stars. Astrophysics and space science, 183(1):1–10, ???? 1991. CODEN APSSBE. ISSN 0004-640X (print), 1572-946X (electronic). URL http://www.

springerlink.com/content/j8120406v014048q/fulltext.pdf.

Blekher:1991:PPI

[BG91b] P. M. Blekher and N. N. Ganikhodzhaev. On pure phases of the Ising model on the Bethe lattices. Theory of Probability and its Applications, 35(2):216–227, ???? 1991. CODEN TPRBAU. ISSN 0040-585X (print), 1095-7219 (electronic). Original Russian article in Teor. Veroyatnost. i Primenen., 35(2), (1990), pp. 220–230.

Bethe:2013:EPF

[BGA^{+}13] Hans Bethe, Marvin L. Goldberger, Stephen L. Adler, Steven Weinberg, Freeman Dyson, and C. N. Yang. Einstein and the physics of the future. In Yang [Yan13], pages 21–34. ISBN 981-4449-00-8 (hardcover), 981-4449-01-6 (paperback), 981-4449-02-4

(e-book). LCCN QC21.3. URL http://www.worldscientific.
com/doi/abs/10.1142/9789814449021_0006.

Bethe:1985:SWS

[BGG^{+}85a] H. A. Bethe, R. L. Garwin, K. Gotthied, H. W. Kendall, C. Sagan, and V. F. Weisskopf. Star Wars and the scientists. Commentary, 79(3):6–11, ???? 1985.

Bethe:1985:WSW

[BGG^{+}85b] Hans A. Bethe, Richard L. Garwin, Kurt Gottfried, Henry W. Kendall, Carl Sagan, et al. Why Star Wars is dangerous and won’t work. New York Review of Books, 32(2):26, February 14, 1985. ISSN 0028-7504 (print), 1944-7744 (electronic).

URL http://www.nybooks.com/articles/archives/1985/feb/
14/why-star-wars-is-dangerous-and-wont-work/. Statement

by six prominent scientists on the dangers of Star Wars that appeared as part of a letter to The Wall Street Journal on January 2, 1985.

Bethe:1984:SBB

[BGGK84] Hans A. Bethe, Richard Garwin, Kurt Gottfried, and Henry W. Kendall. Space-based ballistic missile defense. Scientific American, 251(4):39–49, October 1984. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.nature.com/scientificamerican/journal/v251/n4/pdf/
scientificamerican1084-39.pdf. See summary in [BGGK12].

Bethe:2012:SBB

[BGGK12] Hans A. Bethe, Richard Garwin, Kurt Gottfried, and Henry W. Kendall. Space-based ballistic missile defense. Scientific American, 307(1):72, July 2012. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). Summary of [BGGK84].

Bethe:1991:NTP

[BGM91a] Hans A. Bethe, Kurt Gottfried, and Robert S. McNamara. The nuclear threat: a proposal. New York Review of Books, 38(12):48– 50, June 27, 1991. ISSN 0028-7504 (print), 1944-7744 (electronic).

URL http://www.nybooks.com/articles/archives/1991/jun/
27/the-nuclear-threat-a-proposal/.

Bethe:1991:NTR

[BGM91b] Hans A. Bethe, Kurt Gottfried, and Robert S. McNamara. The nuclear threat: A reply to Marvin M. Miller and Rudolf Peierls. New York Review of Books, 38(18):61, November 7, 1991. ISSN 0028-7504 (print), 1944-7744 (electronic). URL http://www.

nybooks.com/articles/archives/1991/nov/07/the-nuclear-

threat-2/. See [Pei91a].

Bethe:1993:EHB

[BGP^{+}93] Hans A. (Hans Albrecht) Bethe, Kurt Gottfried, Thomas Powers, Eric L. Gasteiger, and Mike E. Allmendinger. An evening with Hans Bethe: the German atomic bomb project. 1 videocassette (90 min.), 1993.

Bethe:1995:NIR

[BGS95a] H. A. Bethe, K. Gottfried, and R. Z. Sagdeev. Nuclear intrigues — a reply. Scientific American, 273(3):10, ???? 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Bethe:1995:DBS

[BGS95b] Hans A. Bethe, Kurt Gottfried, and Roald Z. Sagdeev. Did Bohr share nuclear secrets? Scientific American, 272(5):84–90 (Intl. ed. 64–??), May 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Bethe:1934:SFP

[BH34] H. A. Bethe and W. Heitler. On the stopping of fast particles and on the creation of positive electrons. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 146(856):83–112, August 1, 1934. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-9169 (electronic). URL http://www.jstor.

org/stable/2935479.

Bethe:1939:EIA

[BH39] H. A. Bethe and W. J. Henderson. Evidence for incorrect assignment of the supposed Si^{27} radioactivity of 6.7-minute half-life. Physical Review, 56(10):1060–1061, November 15, 1939. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://prola.aps.org/abstract/PR/v56/i10/p1060_2.

Bethe:1956:APA

[BH56] H. A. Bethe and J. Hamilton. Anti-proton annihilation. Il Nuovo Cimento (10), 4(??):1–22, ???? 1956. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (electronic).

Bethe:1988:PPE

[BH88] H. A. Bethe and G. Hildebrandt. Paul Peter Ewald: 23 January 1888–22 August 1985. Biographical Memoirs of Fellows of the Royal Society, 34:135–176, December 1988. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL http://www.

jstor.org/stable/770049.

Bhattacharjee:2005:DHB

[Bha05] Yudhijit Bhattacharjee. Deaths: Hans A. Bethe. Science, 307 (5716):1719, March 18, 2005. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/307/5716/1719.2.full.pdf.

Brown:2001:FHM

[BHL^{+}01] G. E. Brown, A. Heger, N. Langer, C. H. Lee, S. Welistein, and H. A. Bethe. Formation of high mass X-ray black hole binaries. New Astronomy, 6(7):457–470, ???? 2001. CODEN NEASFS. ISSN 1384-1076 (print), 1384-1092 (electronic).

Bethe:1939:IBD

[BHP39] H. A. Bethe, F. Hoyle, and R. Peierls. Interpretation of betadisintegration data. Nature, 143(3614):200–201, February 4, 1939. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v143/
n3614/pdf/143200a0.pdf;   https://ui.adsabs.harvard.edu/
#abs/1939Natur.143..200B.

Bethe:1997:IBD

[BHP97] H. A. Bethe, F. Hoyle, and R. Peierls. Interpretation of betadisintegration data. In Dalitz and Peierls [DP97], pages 270–272. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN QC21.2 .P42 1997. URL https://ui.

adsabs.harvard.edu/#abs/1997sspr.book..270B.

Biever:2003:NLD

[Bie03] Celeste Biever. Nobel laureates denounce a US attack on Iraq. New Scientist, ??(??):??, January 29, 2003. CODEN NWSCAL. ISSN 0262-4079 (print), 1364-8500 (electronic).

URL https://www.newscientist.com/article/dn3324-nobel-
laureates-denounce-a-us-attack-on-iraq/.

Biswas:1967:SBS

[Bis67] S. N. Biswas. Symmetries of the Bethe–Salpeter equation for relativistic bound-state problem. Journal of Mathematical Physics, 8(5):1109–1111, May 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v8/i5/p1109_s1.

Bethe:1968:IQM

[BJ68] Hans A. (Hans Albrecht) Bethe and Roman W. Jackiw. Intermediate quantum mechanics. Lecture notes and supplements in physics. Benjamin/Cummings Pub. Co., Menlo Park, CA, USA, second edition, 1968. ISBN 0-8053-0755-9. xvi + 393 pp. LCCN QC174.1 .B562 1968.

Bethe:1973:IQM

[BJ73] Hans A. (Hans Albrecht) Bethe and Roman W. Jackiw. Intermediate quantum mechanics. Lecture notes and supplements in physics. Benjamin/Cummings Pub. Co., Menlo Park, CA, USA, second edition, 1973. ISBN 0-8053-0755-9. xvi + 393 pp. LCCN QC174.12 .B47 1973.

Bethe:1974:DBM

[BJ74a] H. A. Bethe and M. B. Johnson. Dense baryon matter calculations with realistic potentials. Nuclear Physics A, 230(1):1–58, September 16, 1974. CODEN NUPABL. ISSN 0375-9474 (print),

1873-1554 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0375947474905284.

Bethe:1974:IQM

[BJ74b] Hans A. (Hans Albrecht) Bethe and Roman W. Jackiw. Intermediate quantum mechanics. Lecture notes and supplements in physics. W. A. Benjamin, New York, second edition, 1974. xvi + 393 pp. LCCN QC174.1 .B416 1968.

Bethe:1976:ENG

[BJ76] Hans A. (Hans Albrecht) Bethe and Mikkel B. Johnson. Effects of nuclear ground state correlations on pion scattering. Technical Report LA-UR-76-1844, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, 1976. 135 pp.

Bethe:1977:DFE

[BJ77] H. A. Bethe and J. H. Jacob. Diffusion of fast electrons in the presence of an electric field. Physical Review A (Atomic, Molecular, and Optical Physics), 16(5):1952–1963, November 1, 1977. CO-DEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-4519.

Bethe:1980:IQM

[BJ80] Hans A. (Hans Albrecht) Bethe and Roman W. Jackiw. Intermediate quantum mechanics. Lecture notes and supplements in physics. Benjamin/Cummings Pub. Co., Menlo Park, CA, USA, second edition, 1980. ISBN 0-8053-0755-9 (paperback). xvi + 393 pp. LCCN QC174.1 .B562 1980. Fourth printing, with revisions and additions.

Bethe:1986:IQM

[BJ86] Hans A. (Hans Albrecht) Bethe and Roman W. Jackiw. Intermediate quantum mechanics. Benjamin/Cummings Pub. Co., Menlo Park, CA, USA, third edition, 1986. ISBN 0-8053-0757-5 (paperback). xviii + 396 pp. LCCN QC174.12 .B47 1986.

Brown:1993:HQG

[BJBP93] G. E. Brown, A. D. Jackson, H. A. Bethe, and P. M. Pizzochero. The hadron to quark gluon transition in relativistic heavy ion collisions. Nuclear Physics A, 560(4):1035–1074, August 2, 1993. CODEN NUPABL. ISSN 0375-9474 (print),

1873-1554 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037594749390145N.

Bukowski:1992:BLL

[BJMK92] Robert Bukowski, Bogumił Jeziorski, Robert Moszyński, and Włodzimierz Kołos. Bethe logarithm and Lamb shift for the hydrogen molecular ion. International Journal of Quantum Chemistry, 42(2):287–319, April 20, 1992. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Bethe:1939:CBS

[BK39] H. A. Bethe and J. G. Kirkwood. Critical behavior of solid solutions in the order-disorder transformation. Journal of Chemical Physics, 7(??):578–582, 1939. CODEN JCPSA6. ISSN 0021-9606 (print), 1089-7690 (electronic).

Bethe:1962:BRP

[BK62] H. A. Bethe and Toichiro Kinoshita. Behavior of Regge poles in a potential at large energy. Physical Review (2), 128(3):1418–1424, November 1, 1962. CODEN PHRVAO. ISSN 0031-899X (print),

1536-6065 (electronic). URL http://prola.aps.org/abstract/
PR/v128/i3/p1418_1.

Borzi:1993:MGM

[BK93] Alfio Borzì and Anni Koubek. A multi-grid method for the resolution of thermodynamic Bethe ansatz equations. Computer Physics Communications, 75(1–2):118–125, April 1993. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
001046559390169D.

Bjorksten:1992:FCR

[BKB92] Johan Bjorksten, Serge A. Korff, and Hans Bethe. From fireworks to cosmic rays, 1906 to 1940. Annals of the New York Academy of Sciences, 655(1):15–??, June 1992. CODEN ANYAA9. ISSN 0077-8923 (print), 1749-6632 (electronic). Frontiers in Cosmic Physics: A Special Colloquium and Symposium in Memory of Serge Alexander Itorff.

Bethe:1982:EIA

[BKD82] Hans A. (Hans Albrecht) Bethe, Alfred E. (Alfred Edward) Kahn, and Robert F. Drinan. The economic impact of the arms race. 2 sound cassettes, 1982.

Bethe:1989:LPS

[BKHS89] Hans A. Bethe, James A. Krumhansl, Paul L. Hartman, and Robert L. Sproull. Lloyd P. Smith. Physics Today, 42(8):83–??, August 1989. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-

0699 (electronic). URL http://link.aip.org/link/phtoad/
v42/i8/p83/s1.

Bethe:1984:OHW

[BKJ84] Hans A. Bethe, Donald M. Kerr, and Robert A. Jeffries. Obituary: Herman W. Hoerlin. Physics Today, 37(12):82–83, December 1984. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-

0699 (electronic). URL http://link.aip.org/link/phtoad/
v37/i12/p82/s3.

Bethe:1940:INM

[BKP40] H. A. Bethe, S. A. Korff, and G. Placzek. On the interpretation of neutron measurements in cosmic radiation. Physical Review, 57(7):573–587, April 1, 1940. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.aps.

org/abstract/PR/v57/i7/p573_1.

Bethe:1949:EVF

[BL49] H. A. Bethe and C. Longmire. On the experimental value of the fine structure constant. Physical Review, 75(2):306–307, January 15, 1949. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://prola.aps.org/abstract/PR/
v75/i2/p306_1.

Bethe:1950:ERN

[BL50] H. A. Bethe and Conrad Longmire. The effective range of nuclear forces II. photo-disintegration of the deuteron. Physical Review, 77(5):647–654, March 1, 1950. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.aps.

org/abstract/PR/v77/i5/p647_1.

Bethe:1982:VF

[BL82] Hans A. Bethe and Franklin A. Long. The value of a freeze. New York Times, September 22, 1982. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Bethe:1983:EFR

[BL83] Hans A. Bethe and Franklin A. Long. Editorial: The freeze referendum: what next? Bulletin of the Atomic Scientists, 39(2): 2–3, February 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Burrows:1988:CTI

[BL88] Adam Burrows and James M. Lattimer. Convection, type II supernovae, and the early evolution of neutron stars. Physics Reports, 163(1–3):51–62, 1988. CODEN PRPLCM. ISSN 0370-1573 (print),

1873-6270 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037015738890035X.

Brown:2006:HBHa

[BL06a] G. E. (Gerald Edward) Brown and Chang-Hwan Lee, editors. Hans Bethe and his physics. World Scientific Publishing Co.

Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2006. ISBN 981-256-609-0, 981-256-610-4 (paperback), 981-277-450-5 (ebook). xiii + 314 pp. LCCN QC16.B46 H36 2006. URL http://

www.worldscibooks.com/physics/5989.html;   http://www.
worldscientific.com/worldscibooks/10.1142/5989.

Brown:2006:LPH

[BL06b] G. E. (Gerald Edward) Brown and Chang-Hwan Lee. List of publications of Hans A. Bethe. In Hans Bethe and his physics [BL06a], pages 283–314. ISBN 981-256-609-0, 981-256-610-4 (paperback), 981-277-450-5 (e-book). LCCN QC16.B46 H36 2006. URL http://

www.worldscibooks.com/physics/5989.html;   http://www.
worldscientific.com/worldscibooks/10.1142/5989.

Brown:2009:HAB

[BL09] Gerald E. Brown and Sabine Lee. Hans Albrecht Bethe: July 2, 1906–March 6, 2005. Biographical memoirs — National Academy of Sciences of the United States of America, 91:31–56, 2009. CO-DEN BMNSAC. ISBN 0-309-14560-0. ISSN 0077-2933. URL

http://www.nap.edu/catalog.php?record_id=12776.

Anonymous:1982:HBP

[Bla82] Michael Blackwood. Hans Bethe — prophet of energy. 1 videocassette (58 minutes)., 1982. Co-production of Blackwood Productions, Inc. and West German Television. Produced and directed by Michael Blackwood.

Blackith:1987:LRS

[Bla87] Robert Blackith. Letter: RBMK safety. Bulletin of the Atomic Scientists, 43(3):52, April 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bet86c].

Brown:1999:FHM

[BLB99] G. E. Brown, C. H. Lee, and Hans A. Bethe. The formation of highmass black holes in low-mass X-ray binaries. New Astronomy, 4(4): 313–323, ???? 1999. CODEN NEASFS. ISSN 1384-1076 (print), 1384-1092 (electronic).

Brown:2000:HAD

[BLB00] G. E. Brown, C. H. Lee, and H. A. Bethe. Hypercritical advection dominated accretion flow. Astrophysical Journal, 541(2):918–923, ???? 2000. CODEN ASJOAB. ISSN 0004-637X (print), 1538-4357 (electronic).

Brown:2001:HMB

[BLB01] G. E. Brown, C. H. Lee, and H. A. Bethe. High mass black holes in soft X-ray transients: Gap in black hole masses? In C. Kouveliotou, J. Ventura, and Ed van den Heuvel, editors, The neutron star black hole connection, pages 343–348. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2001. ISBN ???? LCCN ????

Brovm:2000:NSG

[BLLB00a] G. E. Brovm, C. H. Lee, H. K. Lee, and H. A. Bethe. Is Nova Sco 1994 (GRO 1655-40) a relic of a GRB? In R. Marc Kippen, Robert S. Mallozzi, and Gerald J. Fishman, editors, Gamma-Ray Bursts, AlP Conference Series, pages 628–?? American Institute of Physics, Woodbury, NY, USA, 2000. ISBN ???? LCCN ????

Brown:2000:NSG

[BLLB00b] G. E. Brown, C.-H. Lee, H. K. Lee, and H. A. Bethe. Is Nova Sco 1994 (GRO 1655-40) a relic of a GRB? AIP Conference Proceedings, 526:628–632, 2000. CODEN APCPCS. ISSN 0094-243X (print), 1551-7616 (electronic), 1935-0465.

Brown:2019:SUS

[BLM19] J. Patrick Brown, Beryl Lipton, and Michael Morisy. Scientists under Surveillance: the FBI Files. MIT Press, Cambridge, MA, USA, 2019. ISBN 0-262-53688-9 (paperback), 0-262-35302-4 (ebook). xviii + 413 pp. LCCN Q125 .S4366 2019. URL http:/

/mitpress.mit.edu/9780262536882. Foreword by Steven After-

good and introduction by Walter V. Robinson.

Blanch:1941:ITD

[BLMB41] G. Blanch, A. N. Lowan, R. E. Marshak, and H. A. Bethe. The internal temperature-density distribution of the sun. Astrophysical Journal, 94(??):37–45, ???? 1941. CODEN ASJOAB. ISSN 0004-637X (print), 1538-4357 (electronic).

Blum:1987:CPA

[Blu87] Bruce I. Blum, editor. Conference proceedings: ACM Conference on the History of Medical Informatics: papers presented at the Conference, National Library of Medicine, Bethesda, Maryland, November 5–6, 1987. ACM Press, New York, NY, USA, 1987. ISBN 0-89791-248-9. 205 pp. LCCN W 26.5 A184c 1987.

Bludman:1988:GRC

[Blu88] Sidney A. Bludman. General relativistic core collapse with a soft nuclear equation of state. Physics Reports, 163(1–3):47–50, 1988. CODEN PRPLCM. ISSN 0370-1573 (print), 1873-6270

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0370157388900348.

Brown:2000:TGR

[BLW^{+}00a] G. B. Brown, C. H. Lee, R. A. M. J. Wijers, H. K. Lee, G. Israelian, and H. A. Bethe. A theory of gamma-ray bursts. New Astronomy, 5(4):191–210, ???? 2000. CODEN NEASFS. ISSN 1384-1076 (print), 1384-1092 (electronic).

Brown:2000:EBH

[BLW^{+}00b] G. E. Brown, C. H. Lee, R. A. M. J. Wijers, G. Israelian, and H. A. Bethe. Evolution of black holes in the galaxy. Physics Reports, 333 (1–6):471–504, ???? 2000. CODEN PRPLCM. ISSN 0370-1573 (print), 1873-6270 (electronic).

Brown:2003:TGR

[BLW^{+}03] G. E. Brown, C.-H. Lee, R. A. M. J. Wijers, H. K. Lee, G. Israelian, and H. A. Bethe. A theory of gamma-ray bursts. In Goldhaber et al. [GSS^{+}03], pages 65–92. ISBN 981-238-503-7 (hardcover), 981-238-563-0 (paperback), 981-279-508-1 (e-book). LCCN

QC174.17.S9 S994 2003. URL http://www.worldscientific.
com/doi/pdf/10.1142/9789812795083_0005.

Brown:2001:ENS

[BLZB01] G. E. Brown, C. H. Lee, Simon F. Portegies Zwart, and H. A. Bethe. Evolution of neutron-star, carbon-oxygen white-dwarf binaries. Astrophysical Journal, 547(1):345–354, ???? 2001. CODEN ASJOAB. ISSN 0004-637X (print), 1538-4357 (electronic).

Bethe:1939:PSI

[BM39] Hans A. Bethe and Robert E. Marshak. The physics of stellar interiors and stellar evolution. Reports on Progress in Physics, 6 (??):1–15, ???? 1939. CODEN RPPHAG. ISSN 0034-4885 (print), 1361-6633 (electronic).

Bethe:1954:TBP

[BM54] H. A. Bethe and L. C. Maximon. Theory of Bremsstrahlung and pair production. I. differential cross section. Physical Review (2), 93(4):768–784, February 15, 1954. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.aps.

org/abstract/PR/v93/i4/p768_1.

Bellamy:1955:PGC

[BM55] E. H. Bellamy and R. G. Moorehouse, editors. Proceedings of the 1954 Glasgow Conference on Nuclear and Meson Physics. Pergamon Press, Oxford, UK, 1955. LCCN ????

Bethe:1956:ENT

[BM56] Hans A. (Hans Albrecht) Bethe and Philip Morrison. Elementary nuclear theory. Wiley, New York, NY, USA, second edition, 1956. xi + 274 pp. LCCN QC173 .B48 1956.

Boorse:1966:WA

[BM66] Henry A. (Henry Abraham) Boorse and Lloyd Motz, editors. The world of the atom. Basic Books, New York, NY, USA, 1966. xxvi + 1873 (two volumes) pp. LCCN QC173 .B56.

Bethe:1971:NSO

[BM71] H. A. Bethe and A. Molinari. Nonstatic second order contribution to the elastic electron–nucleus scattering. Annals of Physics, 63(2): 393–431, April 1971. CODEN APNYA6. ISSN 0003-4916 (print),

1096-035X (electronic). URL http://www.sciencedirect.com/
science/article/pii/0003491671900194. Reprinted in [BM72].

Bethe:1972:NSO

[BM72] H. A. Bethe and A. Molinari. Nonstatic second order contribution to the elastic electron–nucleus scattering. In Morse et al. [MHFW72], pages 393–431. ISBN 0-12-508201-0 (vol. 1), 0-323-14303-2 (e-book). LCCN QC173 .N8837 1972.

URL http://www.sciencedirect.com/science/article/pii/
B9780125082013500235. Reprint of [BM71].

Bethe:2004:CAS

[BM04] Hans A. Bethe and N. David Mermin. A conversation about solid-state physics. Physics Today, 57(6):53–56, June 2004. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

URL http://ifup.cit.cornell.edu/bethe/;   http://link.
aip.org/link/phtoad/v57/i6/p53/s1. Also available as a video-

recording, made 25 February 2003, archived with the Cornell University Library.

Bethe:2005:CAS

[BM05] Hans A. Bethe and N. David Mermin. A conversation about solid state physics. Resonance, 10(10):96–104, ???? 2005. CODEN RESOFE. ISSN 0971-8044 (print), 0973-712X

(electronic). URL http://www.springerlink.com/content/
x353562724514503/fulltext.pdf.

Bethe:2006:ENT

[BM06] Hans A. (Hans Albrecht) Bethe and Philip Morrison. Elementary nuclear theory. Dover, New York, NY, USA, second edition, 2006. ISBN 0-486-45048-1 (paperback). xi + 274 pp. LCCN QC173

.B48 2006. URL http://www.loc.gov/catdir/enhancements/
fy0625/2005055054-d.html;   http://www.loc.gov/catdir/
toc/fy0612/2005055054.html.

Bethe:1987:ENT

[BMG87] Hans A. Bethe, Carson Mark, and Richard Garwin. Ending nuclear tests: A technical basis. F.A.S. Public Interest Report, 40

(4):??, April 1987. URL http://www.fas.org/programs/ssp/
nukes/testing/kidderucrl53820.pdf.

Bethe:1953:BHE

[BML53] H. A. Bethe, L. Maximon, and F. Low. Bremsstrahlung at high energies. Physical Review, 91(2):417–418, July 15, 1953. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://prola.aps.org/abstract/PR/v91/i2/p417_2.

Bethe:1940:TMD

[BN40] H. A. Bethe and L. W. Nordheim. On the theory of meson decay. Physical Review, 57(11):998–1006, June 1, 1940. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://prola.aps.org/abstract/PR/v57/i11/p998_1.

Bethe:1968:SAT

[BN68a] H. A. Bethe and J. W. Negele. Semi-analytical theory of muonic X-ray levels. Nuclear Physics A, 117(3):575–585, September 23, 1968. CODEN NUPABL. ISSN 0375-9474 (print), 1873-1554

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0375947468908324.

Bethe:1968:SRE

[BN68b] Hans A. (Hans Albrecht) Bethe and J. Németh. Symmetry and rearrangement energy in nuclear matter. Technical report, Institute for Theoretical Physics, Roland Eötvös University, Budapest, Hungary, 1968. 11 pp.

Bethe:1946:RRE

[BO46] H. A. Bethe and J. Robert Oppenheimer. Reaction of radiation on electron scattering and Heitler’s theory of radiation damping. Physical Review, 70(7–8):451–458, October 1, 1946. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://prola.aps.org/abstract/PR/v70/i7-8/p451_1.

Boffey:1976:NFF

[Bof76] Philip M. Boffey. Nuclear foes fault Scientific American’s editorial judgment in publishing a recent article by Nobel Laureate Hans Bethe. Science, 191(4233):1248–1249, March 26, 1976. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (elec-

tronic). URL http://www.sciencemag.org/content/191/4233/
1248.full.pdf.

Bethe:1934:Na

[BP34a] H. Bethe and R. Peierls. The “neutrino”. Nature, 133(3362): 532, April 7, 1934. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v133/n3362/abs/133532a0.html;   http://
www.nature.com/nature/journal/v133/n3362/pdf/133532a0.
pdf;   https://ui.adsabs.harvard.edu/#abs/1934Natur.133.
.532B.

Bethe:1934:Nb

[BP34b] H. Bethe and R. Peierls. The “neutrino”. Nature, 133(3366): 689–690, May 5, 1934. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v133/n3366/pdf/133689b0.pdf;   https://
ui.adsabs.harvard.edu/#abs/1934Natur.133..689B.

Bethe:1934:PDD

[BP34c] Hans A. Bethe and Rudolf E. Peierls. Photoelectric disintegration of the diplon. In International Conference on Physics, London, pages 93–94. Cambridge University Press, Cambridge, UK, 1934.

Bethe:1935:QTD

[BP35a] Hans A. Bethe and Rudolf E. Peierls. Quantum theory of the diplon. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 148(863):146–156, January 1, 1935. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-9169 (elec-

tronic). URL http://www.jstor.org/stable/96363;   https:
//ui.adsabs.harvard.edu/#abs/1935RSPSA.148..146B.

Bethe:1935:SNP

[BP35b] Hans A. Bethe and Rudolf E. Peierls. The scattering of neutrons by protons. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 149(866):176–183, March 1, 1935. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-9169 (elec-

tronic). URL http://www.jstor.org/stable/96272;   https:
//ui.adsabs.harvard.edu/#abs/1935RSPSA.149..176B.

Bethe:1937:REN

[BP37] H. A. Bethe and G. Placzek. Resonance effects in nuclear processes. Physical Review, 51(6):450–484, March 15, 1937. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://prola.aps.org/abstract/PR/v51/i6/p450_1.

Bethe:1990:MER

[BP90] H. A. Bethe and P. Pizzochero. Mass-energy relation for SN 1987A from observations. Astrophysical Journal, 350(2):L33–L35, February 20, 1990. CODEN ASJOAB. ISSN 0004-637X (print), 1538-4357 (electronic).

Bethe:1997:Na

[BP97a] H. Bethe and R. Peierls. The “neutrino”. In Dalitz and Peierls [DP97], pages 139–140. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN QC21.2 .P42

1997. URL https://ui.adsabs.harvard.edu/#abs/1997sspr.
book..139B.

Bethe:1997:Nb

[BP97b] H. Bethe and R. Peierls. The neutrino. In Dalitz and Peierls [DP97], page 141. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). LCCN QC21.2 .P42 1997.

URL https://ui.adsabs.harvard.edu/#abs/1997sspr.book..
141B.

Bethe:1997:SNP

[BP97c] H. A. Bethe and R. Peierls. The scattering of neutrons by protons. In Dalitz and Peierls [DP97], pages 154–161. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book).

LCCN QC21.2 .P42 1997. URL https://ui.adsabs.harvard.
edu/#abs/1997sspr.book..154B.

Bethe:1937:KEN

[BR37a] H. A. Bethe and M. E. Rose. Kinetic energy of nuclei in the Hartree model. Physical Review, 51(4):283–285, February 15, 1937. CO-DEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

URL http://prola.aps.org/abstract/PR/v51/i4/p283_1.

Bethe:1937:MEO

[BR37b] H. A. Bethe and M. E. Rose. The maximum energy obtainable from the cyclotron. Physical Review, 52(12):1254–1255, December 15, 1937. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://prola.aps.org/abstract/PR/
v52/i12/p1254_2.

Bethe:1952:SAS

[BR52] H. A. Bethe and F. Rohrlich. Small angle scattering of light by a Coulomb field. Physical Review, 86(1):10–16, April 1, 1952. CO-DEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

URL http://prola.aps.org/abstract/PR/v86/i1/p10_1.

Bethe:1967:TBP

[BR67] H. A. Bethe and R. Rajaraman. The three-body problem in nuclear matter. Report TID-23736, US Atomic Energy Commission, Washington, DC, USA, March 16, 1967. 98

pp. URL http://www.osti.gov/accomplishments/documents/
fullText/ACC0101.pdf.

Brown:1993:MGS

[BR93] Laurie M. Brown and John S. Rigden, editors. Most of the Good Stuff: Memories of Richard Feynman. American Institute of Physics, Woodbury, NY, USA, 1993. ISBN 0-88318-870-8. 181 + 16 pp. LCCN QC16.F49 A3 1993. Commentary by Joan Feynman, John Wheeler, Hans Bethe, Julian Schwinger, Murray Gell-Mann, Daniel Hillis, David Goodstein, Freeman Dyson, and Laurie Brown.

Brand:2005:HBT

[Bra05] David Brand. Hans Bethe, a titan of physics and conscience of science, dies at age 98. Web document., March 7,

2005. URL http://www.news.cornell.edu/stories/March05/
Betheobit.deb.html.

Breit:1953:NTA

[Bre53] Gregory Breit. Notes for talk on atmospheric ignition. Report ????, Los Alamos Scientific Laboratory of the University of California, Los Alamos, NM, USA, 1953. This is a talk presented before the first hydrogen-bomb test on 1 November 1952. The work showed that it was impossible for the bomb explosion to ignite the atmosphere or oceans, as had earlier work by Bethe and by Konopinski, Marvin, and Teller, for the atomic bomb. According to [Har10, page 224], this report remained classified until about 1975, when it was mentioned in an newspaper story [Sul75].

Brennan:1975:UMS

[Bre75] B. J. Brennan. Unequal mass spinor-spinor Bethe–Salpeter equation. Journal of Mathematical Physics, 16(11):2241–2249, November 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v16/i11/p2241_s1.

Brian:1995:GTC

[Bri95a] Denis Brian. Genius talk: conversations with Nobel scientists and other luminaries. Plenum Press, New York, NY, USA; London, UK, 1995. ISBN 0-306-45089-5. xii + 423 pp. LCCN fa1941.

Brian:1995:HB

[Bri95b] Denis Brian. Hans Bethe. In Genius talk: conversations with Nobel scientists and other luminaries [Bri95a], pages 105–118. ISBN 0-306-45089-5. LCCN fa1941.

Brown:1971:LBB

[Bro71] G. E. Brown. Landau, Brueckner–Bethe, and Migdal theories of Fermi systems. Reviews of Modern Physics, 43(1):1–14, January 1971. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.43.1;   http://rmp.aps.org/
abstract/RMP/v43/i1/p1_1.

Bronowski:1973:AM

[Bro73a] Jacob Bronowski. The Ascent of Man. British Broadcasting Corporation, London, UK, 1973. ISBN 0-563-10498-8. 448 pp. LCCN Q175 .B7918 1973; CB151.

Bronowski:1973:HB

[Bro73b] Jacob Bronowski. Hans Bethe. In The Ascent of Man [Bro73a], pages 343–344. ISBN 0-563-10498-8. LCCN Q175 .B7918 1973; CB151.

Broad:1982:RHH

[Bro82] William J. Broad. Rewriting the history of the H-bomb: Nobel laureate Hans Bethe says technical errors by Edward Teller, not political opposition by Robert Oppenheimer, hindered work on the superbomb. Science, 218(4574):769–772, November 19, 1982. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

URL http://www.jstor.org/stable/1689737;   http://www.
sciencemag.org/content/218/4574/769.extract. See response

[TB82].

Brown:1988:ESD

[Bro88a] G. E. Brown. The equation of state of dense matter. Physics Reports, 163(1–3):167–203, 1988. CODEN PRPLCM. ISSN 0370-1573 (print), 1873-6270 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0370157388900439.

Brown:1988:F

[Bro88b] G. E. Brown. Foreword. Physics Reports, 163(1–3):3–12, 1988. CODEN PRPLCM. ISSN 0370-1573 (print), 1873-6270

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0370157388900312.

Broad:1997:LNF

[Bro97] William J. Broad. He lit nuclear fire; now he would douse it. New York Times, ??(??):??, June 17, 1997. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://www.fas.

org/bethecr.htm.

Brown:2001:FE

[Bro01] Gerald E. Brown. Fly with eagles. Annual Review of Nuclear and Particle Science, 51:1–22, December 2001. CODEN ARPSDF. ISSN 0163-8998 (print), 1545-4134 (electronic). URL http://

adsabs.harvard.edu/abs/2001ARNPS..51....1B.

Brown:2005:HBH

[Bro05a] Gerald E. Brown. Hans Bethe 1906–2005: Hans Bethe’s nuclear physics. Nuclear Physics A, 762(1–2):2–12, November 28, 2005. CODEN NUPABL. ISSN 0375-9474 (print), 1873-1554 (electronic).

Brown:2005:HBA

[Bro05b] Gerald E. Brown. Hans Bethe and astrophysical theory. Physics Today, 10(10):62–65, October 2005. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.

physicstoday.org/resource/1/phtoad/v58/i10/p62_s1.

Brown:2006:HBA

[Bro06a] Gerald E. Brown. Hans Bethe and astrophysical theory. In Brown and Lee [BL06a], pages 175–183. ISBN 981-256-609-0, 981-256-610-4 (paperback), 981-277-450-5 (e-book). LCCN QC16.B46 H36

2006. URL http://www.worldscibooks.com/physics/5989.
html;   http://www.worldscientific.com/worldscibooks/10.
1142/5989.

Brown:2006:HBHb

[Bro06b] Gerald E. Brown. Hans Bethe and his physics. In Brown and Lee [BL06a], pages 1–26. ISBN 981-256-609-0, 981-256-610-4 (paperback), 981-277-450-5 (e-book). LCCN QC16.B46 H36 2006.

URL http://www.worldscibooks.com/physics/5989.html;
http://www.worldscientific.com/worldscibooks/10.1142/5989.

Bethe:1938:MSE

[BRS38] H. A. Bethe, M. E. Rose, and L. P. Smith. The multiple scattering of electrons. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 78 (4):573–585, March 1938. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http://www.jstor.org/ stable/984803. This paper, which immediately follows Benford’s [Ben38] in this journal issue, is reported in [LG78, p. 197] to have been of considerable interest to scientists involved in secret nuclear physics work in World War II. That is how Benford’s paper “in a journal of rather limited circulation and not usually read by mathematicians” came to be noticed by physicists.

Bethe:1937:AST

[BS37] H. A. Bethe and F. H. Spedding. The absorption spectrum of Tm_{2} (SO_{4})_{3} : 8 H_{2}O. Physical Review, 52(5):454–455, September 1, 1937. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://prola.aps.org/abstract/PR/
v52/i5/p454_1.

Bethe:1943:PTC

[BS43] Hans A. (Hans Albrecht) Bethe and Julian Schwinger. Perturbation theory for cavities. Technical Report D1-117, Radiation Laboratory, Massachusetts Institute of Technology, Cambridge, MA, USA, 1943. 19 pp.

Bethe:1947:GSA

[BS47] H. A. Bethe and H. S. Sack. German scientists in Army employment: II. A protest. Bulletin of the Atomic Scientists, 3(2):65, 67, February 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bethe:1957:QMOa

[BS57a] Hans A. Bethe and Edwin E. Salpeter. Quantum mechanics of oneand two-electron atoms. In ????, editor, Handbuch der Physik, volume 35, pages 88–436. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1957. LCCN ????

Bethe:1957:QMOb

[BS57b] Hans A. Bethe and Edwin E. Salpeter. Quantum mechanics of one- and two-electron atoms. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1957. viii + 369 pp. LCCN QC174.1 .B563.

Bethe:1968:NRP

[BS68] H. A. Bethe and P. J. Siemens. Neutron radius in Pb from the analog state. Physics Letters, 27B(??):549–551, ???? 1968. CODEN PYLBAJ. ISSN 0031-9163 (print), 1873-2410 (electronic).

Bethe:1977:QMOa

[BS77] Hans A. (Hans Albrecht) Bethe and Edwin E. Salpeter. Quantum mechanics of one-and two-electron atoms. Plenum Press, New York, NY, USA; London, UK, 1977. ISBN 0-306-20022-8 (paperback). xii + 369 pp. LCCN QC174.17.P7 B47 1977. Reprint of the 1957 original.

Bethe:1983:NPP

[BS83] Hans A. Bethe and E. B. Salpeter. The 1983 Nobel Prize in Physics — Chandrasekhar and Fowler. Science, 222(4626):881– 883, November 25, 1983. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/222/4626/881.extract.

Bahcall:2005:SEG

[BS05] John N. Bahcall and Edwin E. Salpeter. Stellar energy generation and solar neutrinos. Physics Today, 58(10):44–??, October 2005. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-

0699 (electronic). URL http://link.aip.org/link/phtoad/
v58/i10/p44/s1.

Bahcall:2006:SEG

[BS06] John N. Bahcall and Edwin E. Salpeter. Steller energy generation and solar neutrinos. In Brown and Lee [BL06a], pages 147–155. ISBN 981-256-609-0, 981-256-610-4 (paperback), 981-277-450-5 (e-book). LCCN QC16.B46 H36 2006. URL http://

www.worldscibooks.com/physics/5989.html;   http://www.
worldscientific.com/worldscibooks/10.1142/5989.

Bethe:2008:QMO

[BS08] Hans A. (Hans Albrecht) Bethe and Edwin E. Salpeter. Quantum mechanics of one- and two-electron atoms. Dover books on physics. Dover, New York, NY, USA, 2008. ISBN 0-486-46667-1. xii + 369 pp. LCCN QC174.17.P7 B47 2008. URL http://www.loc.gov/

catdir/enhancements/fy0829/2008012290-d.html.

Bethe:1993:AA

[BSF93] Hans Bethe, Richard E. Smalley, and Anthony P. French. AAPT awards. The Physics Teacher, 31(??):266–??, 1993. CODEN PH-TEAH. ISSN 0031-921X (print), 1943-4928 (electronic).

Bethe:1967:TTJ

[BSK67] Hans A. Bethe, Henry DeWolfe [i.e., Wolf] Smyth, and George Kennan. Three tributes to J. Robert Oppenheimer. Institute for Advanced Study, Princeton, NJ, USA, 1967. 29 pp. LCCN QC16.O62 B45 1967.

Bethe:1983:ERS

[BSP^{+}83] H. Bethe, C. L. Schwartz, C. Pursell, W. J. Mead, T. Atwater, W. Kohn, and D. McDonald. The ethical responsibilities of scientists. Center Magazine, 16(5):2–11, ???? 1983. ISSN 0008-9125.

Braga:2005:CPB

[BSS05] Gastão A. Braga, Rémy Sanchis, and Tiago A. Schieber. Critical percolation on a Bethe lattice revisited. SIAM Review, 47(2):349– 365, June 2005. CODEN SIREAD. ISSN 0036-1445 (print), 1095-

7200 (electronic). URL http://epubs.siam.org/sam-bin/dbq/
article/43368.

Bethe:1941:DTE

[BT41] H. A. Bethe and E. Teller. Deviations from thermal equilibrium in shock waves. Memorandum ATI-18278 (NP-4898, BRL-X-117), Ballistic Research Laboratories, Aberdeen, MD, USA, ????

1941. iv + 46 pp. URL http://books.google.de/books?id=

2qPvAAAAMAAJ. Undated, perhaps 1941 or 1942. Classified until after World War II, and reprinted in [Bet97e, pages 296ff]. Discussed briefly in [Ber80, page 64], which says work in 1940. Bethe, in Goodstein’s interview [Goo99, pages 256–258], also says 1940, and describes it as arising from a suggestion given to the authors by aerodynamicist Theodore von Kármán.

Bethe:1945:DTE

[BT45] H. A. Bethe and E. Teller. Deviations from thermal equilibrium in shock waves. Report X-117, Ballistic Research Laboratories, Aberdeen, MD, USA, 1945.

Bethe:1951:DTEa

[BT51a] Hans A. Bethe and Edward Teller. Deviations from thermal equilibrium in shock waves. Report AM-562, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, 1951. iv + 77 pp.

Bethe:1951:DTEb

[BT51b] Hans A. (Hans Albrecht) Bethe and Edward Teller. Deviations from thermal equilibrium in shock waves. Technical report, Engineering Research Institute, University of Michigan, Ann Arbor,

MI, USA, 1951. iv + 77 pp. URL http://pi.lib.uchicago.edu/

1001/cat/bib/ocm23330822. First reproduced for limited distribution by the Ballistic Research Laboratories, Aberdeen Proving Ground, Aberdeen, Maryland.

Bethe:1953:DTE

[BT53] Hans A. Bethe and Edward Teller. Deviations from thermal equilibrium in shock waves. Report BRL-X-117 (NP-4898), United States Department of Energy, Oak Ridge, TN, USA, 1953. 82 pp.

URL http://www.osti.gov/servlets/purl/4420349/.

Bethe:1961:FNT

[BT61] Hans A. (Hans Albrecht) Bethe and Edward Teller. The future of nuclear tests: Nuclear tests, volume 145 of Headline series. Foreign Policy Association — World Affairs Center, New York, NY, USA, 1961. 61 pp. LCCN E744 .H43 no. 145.

Broido:1969:BSE

[BT69] Michael M. Broido and John G. Taylor. Bethe–Salpeter equation. Journal of Mathematical Physics, 10(1):184–209, January 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v10/i1/p184_s1.

Bethe:1982:NFW

[BT82] Hans A. Bethe and E. Teller. The nuclear freeze: Will it really make us safer? Los Angeles Times, October 17, 1982.

Bethe:1950:NPSc

[BTH50a] H. A. Bethe, L. Tonks, and H. Hurwitz, Jr. Neutron penetration and slowing down at intermediate distances through medium and heavy nuclei. Physical Review (2), 80(1):11–19, October 1, 1950. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://prola.aps.org/abstract/PR/
v80/i1/p11_1.

Bethe:1950:NPSa

[BTH50b] H. A. (Hans Albrecht) Bethe, L. Tonks, and H. (Henry) Hurwitz, Jr. Neutron penetration and slowing down at intermediate distances through medium and heavy nuclei. Technical Report AECD-2790, Knolls Atomic Power Laboratory, Schenectady, NY,

USA, 1950. 25 pp. URL http://hdl.handle.net/2027/mdp.
39015086448928;   http://pi.lib.uchicago.edu/1001/cat/

bib/ocn702216714. Published 29 September 1949. Declassified 16 January 1950.

Bethe:1950:NPSb

[BTH50c] H. A. (Hans Albrecht) Bethe, L. Tonks, and H. (Henry) Hurwitz, Jr. Neutron penetration and slowing down at intermediate distances through medium and heavy nuclei. Technical Report KAPL-P-450, Knolls Atomic Power Laboratory, Schenectady, New York, 1950. 9 pp. Reprinted from [BTH50a].

Burhop:1968:PHB

[Bur68] E. H. S. Burhop. Professor Hans Bethe. Physics Education, 3 (1):16, 1968. CODEN PHEDA7. ISSN 0031-9120 (print), 1361-

6552 (electronic). URL http://stacks.iop.org/0031-9120/3/
i=1/a=304.

Bethe:1974:DLH

[BV74] H. A. Bethe and G. C. Viases. Diffusion in a laser heated plasma column. Bulletin of the American Physical Society, 19(??):932, ???? 1974. CODEN BAPSA6. ISSN 0003-0503.

Bethe:1975:TDP

[BV75] H. A. Bethe and G. C. Vlases. Transverse diffusion of particles and energy in a laser-heated plasma column. Physics of Fluids, 18 (??):982–990, 1975. CODEN PHFLE6. ISSN 1070-6631.

Bethe:1939:APD

[BW39] Hans A. Bethe and George Winter. Armor plate deformation and shielding. Title and year uncertain. The memoir [BL09, page 42] says “The paper was never published and as a potentially significant contribution to the war effort it was soon classified and thereby put outside Bethe’s reach, who — as enemy alien — was not allowed to access such classified material! Only after the war Nevill Mott told him that it had been very useful.”, 1939.

Bethe:1951:MS

[BW51] H. A. Bethe and R. R. Wilson. Meson scattering. Physical Review, 83(3):690–692, August 1, 1951. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.aps.

org/abstract/PR/v83/i3/p690_1.

Bethe:1980:ORF

[BW80] H. A. Bethe and George Winter. Otto Robert Frisch. Physics Today, 33(1):99–100, January 1980. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://physicstoday.

scitation.org/doi/abs/10.1063/1.2913924. See correction

[Dit80].

Bethe:1985:RSS

[BW85] H. A. Bethe and J. R. Wilson. Revival of a stalled supernova shock by neutrino heating. Astrophysical Journal, 295(1):14–23, ???? 1985. CODEN ASJOAB. ISSN 0004-637X (print), 1538-4357 (electronic).

Burgess:1987:BPE

[BW87] Alan Burgess and Colm T. Whelan. Betrt — a procedure to evaluate the cross section for electron-hydrogen collisions in the Bethe approximation to the reactance matrix. Computer Physics Communications, 47(2–3):295–304, November/December 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587901159.

Bethe:1980:CSS

[BWH80] Hans A. Bethe, Richard Wilson, and Andrew Hull. Commentary: Scientists for SENSE. Bulletin of the Atomic Scientists, 36(7): 64, September 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bey80].

Bethe:1981:HAB

[BWMH81] Hans A. (Hans Albrecht) Bethe, Charles Wiener, Jagdish Mehra, and Lillian Hoddeson. Hans Albrecht Bethe oral histories: 1966– 1981, 1981.

Bethe:1982:NLT

[BYB82] H. A. Bethe, A. Yahil, and G. B. Brown. On the neutrino luminosity from a Type II supernova. Astrophysical Journal, 262(1): L7–L10, ???? 1982. CODEN ASJOAB. ISSN 0004-637X (print), 1538-4357 (electronic).

Bete:1938:ETM

[BZ38] G. Bete and A. Zommerfel’d. Elektronnaja teorija metallow. (Russian) [Electron theory of metals]. In Handbuch der Physik, volume 24 (part 2), page ?? ONTI, Moskva, USSR, 1938.

Cabrera:1978:BPW

[Cab78] G. G. Cabrera. On the Bethe–Peierls–Weiss method of ferromagnetism: An illustrative calculation of a phase transition. American Journal of Physics, 46(10):1062–??, October 1978. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL

http://link.aip.org/link/ajpias/v46/i10/p1062/s1.

Compton:1934:CCR

[CB34] A. H. Compton and H. A. Bethe. Composition of cosmic rays. Nature, 134(3393):734–735, November 10, 1934. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.

nature.com/nature/journal/v134/n3393/pdf/134734b0.pdf.

Camac:1948:SHE

[CB48a] M. Camac and H. A. Bethe. The scattering of high energy neutrons by protons. Physical Review, 73(3):191–196, February 1, 1948. CO-DEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

URL http://prola.aps.org/abstract/PR/v73/i3/p191_1.

Courant:1948:EDB

[CB48b] E. D. Courant and H. A. Bethe. Electrostatic deflection of a betatron or synchrotron beam. Review of Scientific Instruments, 19 (??):632–637, 1948. CODEN RSINAK. ISSN 1089-7623, 0034-6748.

Carruthers:1960:RIH

[CB60] P. Carruthers and H. A. Bethe. Role of the π–π interaction in highenergy π-nucleon interactions. Physical Review Letters, 4(??):536– 539, May 15, 1960. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Cooperstein:1984:SPS

[CBB84] J. Cooperstein, H. A. Bethe, and G. E. Brown. Shock propagation in supernovae: Concept of net ram pressure. Nuclear Physics A, 429(3):527–555, November 12, 1984. CODEN NUPABL. ISSN 0375-9474 (print), 1873-1554 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0375947484906961.

Cranberg:1987:DSG

[CBR87] Lawrence Cranberg, Hans Bethe, and William J. Rice. The DEW Study goes public. Physics Today, 40(10):11–??, October 1987. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-

0699 (electronic). URL http://link.aip.org/link/phtoad/
v40/i10/p11/s1.

Calogero:1973:NMB

[CC73] F. Calogero and C. Ciofi Degli Atti, editors. The nuclear manybody problem: proceedings of the International Symposium on Present Status and Novel Developments in the Nuclear Many-body Problem: Roma, September 19–23, 1972, volume 4 of International physics series. Compositori, Bologna, Italy, 1973. ISBN 020. LCCN QC793.9 .I73 1972.

Chandrasekaran:2011:CIS

[CCG^{+}11] Venkat Chandrasekaran, Misha Chertkov, David Gamarnik, Devavrat Shah, and Jinwoo Shin. Counting independent sets using the Bethe approximation. SIAM Journal on Discrete Mathematics, 25(2):1012–1034, ???? 2011. CODEN SJDMEC. ISSN 0895-4801 (print), 1095-7146 (electronic). URL http://epubs.siam.org/

sidma/resource/1/sjdmec/v25/i2/p1012_s1.

Cockcroft:1934:SPN

[CCJ^{+}34] J. Cockcroft, J. Chadwick, F. Joliot, J. Joliot, N. Bohr, G. Gamov, P. A. M. Dirac, and W. Heisenberg, editors. Structure et propriétés des noyaux atomiques. Rapports et discussions du septième conseil de physique tenu à Bruxelles du 22 au 29 octobre 1933 sous les auspices de l’institut international de physique Solvay. (French) [Structure and properties of atomic nuclei. Reports and discussions of the Seventh Meeting on Physics held in Brussels from 22 to 29 October 1933 under the auspices of the Solvay International Institute of Physics]. Gauthier-Villars et cie, Paris, France, 1934. LCCN ???? Publiés par la commission administrative de l’institut.

Calliari:2012:REE

[CF12] L. Calliari and S. Fanchenko. Reflection electron energy loss spectroscopy: role of the Bethe–Born factor. Surface and Interface Analysis, 44(8):1104–1109, August 2012. CODEN SIANDQ. ISSN 0142-2421 (print), 1096-9918 (electronic). URL http://

onlinelibrary.wiley.com/doi/10.1002/sia.4827/abstract.

Cooper:1982:HLT

[CFR82] Alan Cooper, Joel Feldman, and Lon Rosen. Higher Legendre transforms and their relationship to Bethe–Salpeter kernels and rfield projectors. Journal of Mathematical Physics, 23(5):846–868, May 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v23/i5/p846_s1.

Chiang:2002:PDE

[CH02] Chun-Ming Chiang and Choon-Lin Ho. Planar Dirac electron in Coulomb and magnetic fields: a Bethe Ansatz approach. Journal of Mathematical Physics, 43(1):43–51, January 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Chown:2001:MFS

[Cho01] Marcus Chown. The Magic Furnace: the Search for the Origins of Atoms. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2001. ISBN 0-19-514305-1, 0-19-803266-8 (e-book). LCCN

QC171.2 .C5 2001. URL http://lib.myilibrary.com?id=83521.

Cantelon:1991:AAD

[CHW91] Philip L. (Philip Louis) Cantelon, Richard G. Hewlett, and Robert Chadwell Williams, editors. The American atom: a documentary history of nuclear policies from the discovery of fission to the present. University of Pennsylvania Press, Philadelphia, second edition, 1991. ISBN 0-8122-3096-5 (hardcover), 0-8122-1354-8 (paper). xviii + 369 pp. LCCN UA23 .A597 1991.

Cruickshank:1992:PPE

[CJK92] D. W. J. (Durward W. J.) Cruickshank, Hellmut J. Juretschke, and N. (Norio) Katō, editors. P. P. Ewald and His Dynamical Theory of X-ray Diffraction: a Memorial Volume for Paul P. Ewald, 23 January 1888–22 August 1985, volume 2 of International Union of Crystallography monographs on crystallography. International Union of Crystallography, Chester, UK, 1992. ISBN 0-19-855379-X. x + 161 pp.

LCCN QD945 .P22 1992. URL http://www.loc.gov/catdir/
enhancements/fy0606/92009876-d.html;   http://www.loc.
gov/catdir/enhancements/fy0606/92009876-t.html.

CSCFAS:1948:SIC

[Cle48] Committee on Secrecy and Federation of American Scientists Clearance. Some individual cases of clearance procedures. Bulletin of the Atomic Scientists, 4(9):281–285, September 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Clinton:1997:LPH

[Cli97] Bill Clinton. Letter to Professor Hans Bethe. Response to [Bet97c],

June 2, 1997. URL http://www.fas.org/bethecr.htm.

Cooperstein:1988:NS

[Coo88] J. Cooperstein. Neutrinos in supernovae. Physics Reports, 163 (1–3):95–126, 1988. CODEN PRPLCM. ISSN 0370-1573 (print),

1873-6270 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0370157388900385.

Correll:2012:TCI

[Cor12] John T. Correll. They called it Star Wars: The critics sneered at it, but the Soviets weren’t so sure. Air Force magazine, 95(6):66–70, June 2012. CODEN AFORCO. ISSN 0730-6784 (print), 1943-4782 (electronic). URL http://www.

airforce-magazine.com/MagazineArchive/Pages/2012/June%
202012/0612starwars.aspx.

Crammer:1950:AE

[CP50] J. L. (John Lewis) Crammer and R. E. Peierls, editors. Atomic Energy. Penguin, London, UK and New York, NY, USA, 1950. 199 pp. LCCN TK9145 .C75. Revised and expanded version of Science News volume 2 (February 1947).

Cahn:1993:TBT

[CP93] Anne Hessing Cahn and John Prados. Team-B: the trillion dollar experiment. Bulletin of the Atomic Scientists, 49(3):22–31, April 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.highbeam.com/doc/1G1-13610752.

html. See corrections [Ano93].

Corlin:1931:ZGL

[CSB^{+}31] Axel Corlin, J. S. Stein, G. Beck, H. Bethe, and W. Riezler. Zuschriften. (German) [Letters: Remarks on the quantum theory of the absolute zero of temperature]. Naturwissenschaften, 19(2):38–39, January 1931. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic). URL http:

//www.math.tohoku.ac.jp/~kuroki/Sokal/misc/bethespoof.
html;   http://www.springerlink.com/content/x25n8m515025742u/

fulltext.pdf. This paper is a parody on Eddington’s numerology.

Cabrera-Trujillo:1999:BSR

[CTSOS99] R. Cabrera-Trujillo, John R. Sabin, J. Oddershede, and Stephan P. A. Sauer. The Bethe sum rule and basis set selection in the calculation of generalized oscillator strengths. Advances in Quantum Chemistry, 35:175–192, 1999. CODEN AQCHA9. ISSN 0065-

3276. URL http://www.sciencedirect.com/science/article/
pii/S0065327608604608.

Chayes:1969:AED

[CWB^{+}69] Abram Chayes, Jerome B. (Jerome Bert) Wiesner, Hans A. (Hans Albrecht) Bethe, et al., editors. ABM; an evaluation of the decision to deploy an antiballistic missile system, volume 7 of Signet broadside. Harper & Row, New York, NY, USA, 1969. xxii + 282 pp. LCCN UG633 .A714 1969; UG633 .C493 1969b; UG633 .C512 1969b.

Caspers:2004:CBS

[CWL^{+}04] W. J. Caspers, A. Wal, M. Labuz, M. Kuźma, and T. Lulek. Continuity of Bethe solutions with respect to chain length N and winding numbers λ_{l}. Journal of Mathematical Physics, 45(1):391– 399, January 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v45/i1/p391_s1.

Davies:1952:ICS

[DB52] Handel Davies and H. A. Bethe. Integral cross section for Bremsstrahlung and pair production. Physical Review, 87(1):156– 157, July 1, 1952. CODEN PHRVAO. ISSN 0031-899X (print),

1536-6065 (electronic). URL http://prola.aps.org/abstract/
PR/v87/i1/p156_2.

Davies:1954:TBP

[DBM54] Handel Davies, H. A. Bethe, and L. C. Maximon. Theory of Bremsstrahlung and pair production. II. integral cross section for pair production. Physical Review (2), 93(4):788–795, February 15, 1954. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://prola.aps.org/abstract/PR/
v93/i4/p788_1.

Demortier:2015:QEQ

[Dem15] Guy Demortier. Qui est qui dans ^{≪}the letter to Roosevelt ^{≫}?. (French) [Who is who in “the letter to Roosevelt”?]. Revue des Questions Scientifiques, 186(1):125–152, ???? 2015. CODEN RQS-

CAN. ISSN 0035-2160. URL https://www.unamur.be/sciences/
philosoc/revueqs/textes-en-ligne/rqs_186_1_demortier.

DeNike:1976:LRC

[DeN76] L. Douglas DeNike. Letter: Reverse China syndrome. Bulletin of the Atomic Scientists, 32(1):3, January 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Dud75].

DeSanto:1973:SRR

[DeS73] John A. DeSanto. Scattering from a random rough surface: Diagram methods for elastic media. Journal of Mathematical Physics, 14(11):1566–1573, November 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:

//jmp.aip.org/resource/1/jmapaq/v14/i11/p1566_s1.

Diaconis:1979:RP

[DF79] Persi Diaconis and David Freedman. On rounding percentages. Journal of the American Statistical Association, 74(366):359–364, June 1979. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X

(electronic). URL http://www.jstor.org/stable/2284288. The

authors extend prior work on correctness of sums of rounded percentages [MYZ67], and criticize biased rounding practices in [Ben38].

deGier:2002:XSC

[dGBNM02] J. de Gier, M. T. Batchelor, B. Nienhuis, and S. Mitra. The XXZ spin chain at Δ = −1/2: Bethe roots, symmetric functions, and determinants. Journal of Mathematical Physics, 43(8):4135–4146, August 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Deen:2006:RQI

[DHB^{+}06] G. Deen, M. Hammer, J. Bethencourt, I. Eiron, J. Thomas, and J. H. Kaufman. Running Quake II on a grid. IBM Systems Journal, 45(1):21–44, ???? 2006. CODEN IBMSA7. ISSN 0018-8670. URL

http://www.research.ibm.com/journal/sj/451/deen.html.

deHoffmann:1954:PHP

[dHMAB54] F. de Hoffmann, N. Metropolis, E. F. Alei, and H. A. Bethe. Pion–hydrogen phase shift analysis between 120 and 217 Mev. Physical Review, 95(6):1586–1605, September 15, 1954. CO-DEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

URL http://prola.aps.org/abstract/PR/v95/i6/p1586_1;
http://prola.aps.org/pdf/PR/v95/i6/p1586_1.

Dirac:1969:MTP

[Dir69] P. A. M. Dirac. Methods in theoretical physics. International Atomic Energy Agency Bulletin, ??(??):21–28, ???? 1969. CO-DEN IAEBAB. ISSN 0020-6067 (print), 1564-2690 (electronic). Reprinted from [BDH^{+}89].

Dirac:1970:MTP

[Dir70] P. A. M. Dirac. The methods of theoretical physics. Uspekhi Fizicheskikh Nauk, 102(10):291, October 1970. CODEN UFNAAG. ISSN 0042-1294 (print), 1996-6652 (electronic). URL http://ufn.ru/ ru/articles/1970/10/j/. From a Life of Physics, Evening Lectures the International Centre of Theoretical Physics in Trieste, Vienna, IAIA, 1969 [BDH^{+}89]. Translated by V. K. Ignatovich.

Ditchburn:1980:FO

[Dit80] R. W. Ditchburn. Frisch obituary. Physics Today, 33(5):86, May 1980. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). Correction to [BW80], noting that Frisch did not succeed to Rutherford’s chair, but rather, was Jacksonian Professor of Natural Philosophy.

Sakharov:1991:SRT

[DK91] Sidney D. (Sidney David) Drell and S. P. (Sergei Petrovich) Kapitsa, editors. Sakharov remembered: a tribute by friends and colleagues. American Institute of Physics, Woodbury, NY, USA, 1991. ISBN 0-88318-853-8 (hardcover), 0-88318-852-X (paperback). x + 303 pp. LCCN QC16.S255 S25 1991.

Destri:1983:NBA

[DL83] C. Destri and J. H. Lowenstein. Normalization of Bethe-ansatz states in the chiral-invariant Gross–Neveu model. Journal of Mathematical Physics, 24(3):677–682, March 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v24/i3/p677_s1.

Dowling:1982:MBP

[Dow82] John Dowling. Michael Blackwood, Producer, Hans Bethe: a Prophet of Energy (film). American Journal of Physics, 50(5):480– ??, May 1982. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-

2909 (electronic). URL http://link.aip.org/link/ajpias/
v50/i5/p480/s1.

Dowling:1983:FRBd

[Dow83] John Dowling. Film reviews: Hans Bethe: A Prophet of Energy, produced and distributed by Michael Blackwood; How Much is Enough: Decision Making in the Nuclear Age, produced by Andrew A. Stern. Bulletin of the Atomic Scientists, 39(5):44–45, May 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dalitz:1997:SSP

[DP97] Richard Henry Dalitz and Sir Rudolf Peierls, editors. Selected scientific papers of Sir Rudolf Peierls: with commentary, volume 19 of World Scientific series in 20th century physics. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1997. ISBN 981-02-2692-6 (hardcover), 981-02-2693-4 (paperback), 981-279-577-4 (e-book). xxiii + 805 pp. LCCN QC21.2

.P42 1997.  URL https://ui.adsabs.harvard.edu/#abs/
1997sspr.book.....D;   https://ui.adsabs.harvard.edu/#
abs/1997ssps.book.....P;   https://ui.adsabs.harvard.edu/
#abs/1997WSSP...19.....D;   https://www.worldscientific.
com/worldscibooks/10.1142/3128.

Drell:2005:HBL

[Dre05] Sidney Drell. Hans Bethe — last of the giants. American Physical Society Forum on Physics and Society Newsletter, April

2005. URL http://www.aps.org/units/fps/newsletters/
2005/april/articles.html#drell.

Drell:2006:SPP

[Dre06] Sidney D. Drell. Shaping public policy. In Brown and Lee [BL06a], pages 251–262. ISBN 981-256-609-0, 981-256-610-4 (paperback), 981-277-450-5 (e-book). LCCN QC16.B46 H36 2006. URL http://

www.worldscibooks.com/physics/5989.html;   http://www.
worldscientific.com/worldscibooks/10.1142/5989.

Drell:2007:HBJ

[Dre07] Sidney D. Drell. Hans Bethe, 2 July 1906–6 March 2005. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 151(3):337–344, September 2007. CO-DEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

URL http://www.jstor.org/stable/4599076.

Dyson:1954:MNS

[DRS^{+}54] F. J. Dyson, M. Ross, E. E. Salpeter, S. S. Schweber, M. K. Sundaresan, W. M. Visscher, and H. A. Bethe. Meson– nucleon scattering in the Tamm–Dancoff approximation. Physical Review, 95(6):1644–1658, September 15, 1954. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

URL http://link.aps.org/doi/10.1103/PhysRev.95.1644;
http://prola.aps.org/abstract/PR/v95/i6/p1644_1.

Dalitz:1988:BPP

[DS88] R. H. (Richard Henry) Dalitz and R. B. Stinchcombe, editors. A Breadth of Physics: the Proceedings of the Peierls 80th birthday symposium, Oxford University, 27 June 1987. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1988. ISBN 9971-5-0519-3, 9971-5-0520-7 (paperback). LCCN QC770.B67 1988.

Dalitz:1956:SIE

[DSB56] R. H. Dalitz, M. K. Sundaresan, and H. A. Bethe. A singular integral equation in the theory of meson-nucleon scattering. Proceedings of the Cambridge Philosophical Society. Mathematical and physical sciences, 52:251–272, 1956. CODEN PCPSA4. ISSN 0008-1981.

Derringh:1991:BES

[DSB91] Edward Derringh, Richard Shurtleff, and Hans A. Bethe. Binding energies and supernova evolution. Physics Today, 44(4):15, April 1991. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://link.aip.org/link/phtoad/v44/i4/

p15/s1. See [Bet90e].

de-Shalit:1966:PTP

[dSFvLH66] Amos de Shalit, Herman Feshbach, and L. van (Léon) Hove, editors. Preludes in theoretical physics, in honor of V. F. Weisskopf: essays. North-Holland Publishing Co., Amsterdam, The Netherlands, 1966. x + 351 pp. LCCN QC71 .S39 1966.

Dudley:1975:UCT

[Dud75] H. C. Dudley. The ultimate catastrophe. Bulletin of the Atomic Scientists, 31(9):21–24, November 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments and reply [vHWD76, DeN76, Gla76, Bet76h].

Dudley:1976:YAT

[Dud76] H. C. Dudley. Are your assumptions tenable? Bulletin of the Atomic Scientists, 32(6):38, June 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bet76h].

Duffus:1946:ABV

[Duf46] R. L. Duffus. Atom bomb versus ‘human race’: The scientists who fashioned it remind us of its terrible threat: Review of One World or None. A Report to the Public on the Full Meaning of the Atomic Bomb. Edited by Dexer Masters and Kathlerine Way. Foreword by Niels Bohr. Introduction by Arthur H. Compton. x + 79 pp. New York: Whittlesey House. $1. New York Times, ??(??): BR1–BR2, March 17, 1946. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.proquest.

com/docview/107455627.

Dyson:1955:BRH

[Dys55] Freeman J. Dyson. Book review: Hans A. Bethe and Frederic de Hoffmann, Mesons and Fields. Volume II: Mesons. Physics Today, 8(6):27, June 1955. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://link.aip.org/link/?PTO/
8/27/1;   http://link.aip.org/link/phtoad/v8/i6/p27/s1.

Dyson:1956:BRS

[Dys56] Freeman J. Dyson. Book review: S. S. Schweber, H. A. Bethe, and F. De Hoffmann, Mesons and Fields; Vol. 1: Fields. Physics Today, 9(5):32–??, May 1956. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.aip.org/

link/phtoad/v9/i5/p32/s1.

Dyson:1966:RSM

[Dys66] Freeman J. Dyson. Remarks on the stability of matter. Talk delivered at the sixtieth birthday celebration of Hans Bethe at Cornell University, Ithaca, October 1966., October 1966.

Dyson:1979:DU

[Dys79] Freeman J. Dyson. Disturbing the Universe. Harper & Row, New York, NY, USA, 1979. ISBN 0-06-011108-9. x + 283 pp. LCCN QC16.D95 A33 1979. US$12.95. Commissioned by the Science Book Program of the Alfred P. Sloan Foundation. Parts of the book were published in The New Yorker, August 6, 13, and 20, 1979, and in The Observer, October 28, 1979.

Dyson:1989:FC

[Dys89] Freeman J. Dyson. Feynman at Cornell. Physics Today, 42(2):32– 38, February 1989. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://link.aip.org/link/?PTO/

42/32/1. Special issue on Richard Feynman.

Dyson:2005:HB

[Dys05a] Freeman Dyson. Hans A. Bethe (1906–2005). Science, 308(5719): 219, April 8, 2005. CODEN SCIEAS. ISSN 0036-8075 (print),

1095-9203 (electronic). URL http://www.jstor.org/stable/
3841346;   http://www.sciencemag.org/content/308/5719/
219.full.pdf.

Dyson:2005:HBQ

[Dys05b] Freeman Dyson. Hans Bethe and quantum electrodynamics. Physics Today, 58(10):48–50, October 2005. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

link.aip.org/link/phtoad/v58/i10/p48/s1.

Dyson:2006:HBQ

[Dys06] Freeman J. Dyson. Hans Bethe and quantum electrodynamics. In Brown and Lee [BL06a], pages 157–163. ISBN 981-256-609-0, 981-256-610-4 (paperback), 981-277-450-5 (e-book). LCCN QC16.B46

H36 2006. URL http://www.worldscientific.com/doi/abs/
10.1142/9789812774507_0009.

Dyson:2018:MPA

[Dys18] Freeman J. Dyson. Maker of patterns: an autobiography through letters. Liveright Publishing Corporation, New York, NY, USA, 2018. ISBN 0-87140-386-2 (hardcover). xvi + 400 pp. LCCN QC16.D95 A3 2018.

Einstein:1952:ESG

[EBU^{+}52] Albert Einstein, Hans A. Bethe, Harold C. Urey, James Franck, Samuel Goudsmit, Cyril S. Smith, Arthur H. Compton, and William P. Murphy. Eminent scientists give their views on American visa policy. Bulletin of the Atomic Scientists, 8(7):217–220, October 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Issue devoted to the restrictions of the McCarran Acts — the Internal Security Act of 1950, and the Immigration and Nationality Act of 1952 — on freedom of travel by scientists across US borders.

Eddington:1920:ICS

[Edd20] A. S. Eddington. The internal constitution of the stars. Nature, 106(2653):14–20, September 1920. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Edgerton:1991:BRA

[Edg91] David Edgerton. Book reviews: Atomic Energy for Military Purposes, with a new foreword by Philip Morrison and an essay by Henry DeWolf Smyth. Stanford: Stanford University Press, 1990. Pp. xvi + 324. ISBN 0-8047-1721-4, $39.50 (hardcover); 0-8047-1722-2, $12.95 (paperback). Herbert F. York. The Advisors: Oppenheimer, Teller and the Superbomb, with a historical essay by Hans A. Bethe. Stanford: Stanford University Press, 1989. Pp. xiv + 201. ISBN 0-8047-1713-3, $32.50 (hardcover); 0-8047-1714-1, $8.95 (paperback). British Journal for the History of Science, 24(4):476–477, December 1991. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.jstor.

org/stable/4027137.

Edson:1968:SMA

[Eds68] Lee Edson. Scientific man for all seasons. New York Times, ??(??): 28, 29, 122, March 10, 1968. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL https://search.proquest.

com/hnpnewyorktimes/docview/118200068/.

Ekspong:1994:OKM

[Eks94] Gösta Ekspong, editor. The Oskar Klein memorial lectures: Lectures by Hans A. Bethe and Alan H. Guth with translated reprints by Oskar Klein, volume 2. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1994. ISBN 981-02-1450-2 (hardcover). ix + 149 pp. LCCN QC71 .O84 1991.

Eliot:1950:HB

[Eli50] George Fielding Eliot, editor. The H bomb. Didier, New York, NY, USA, 1950. 175 pp. LCCN UF767 .H2. Introduction by Albert Einstein and commentary by George Fielding Eliot.

Epstein:1988:GRB

[Eps88] Richard I. Epstein. Gamma-ray bursts and glitching neutron stars. Physics Reports, 163(1–3):155–166, 1988. CODEN PRPLCM. ISSN 0370-1573 (print), 1873-6270 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0370157388900427.

Evans:1984:ESI

[Eva84] J. W. Evans. Exactly solvable irreversible processes on Bethe lattices. Journal of Mathematical Physics, 25(8):2527–2532, August 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v25/i8/p2527_s1.

Ewald:1958:BRB

[Ewa58] P. P. Ewald. Book review: Splitting of terms in crystals by H. A. Bethe. Acta Crystallographica, 11(10):755, October 1958. CODEN ACCRA9. ISSN 0365-110X (print), 1600-8642 (electronic).

Fakley:1983:BM

[Fak83] Dennis C. Fakley. The British Mission. Los Alamos Science, 7: 186–189, Winter/Spring 1983. CODEN LASCDI. ISSN 0273-7116.

URL http://library.lanl.gov/cgi-bin/getfile?07-21.pdf.

Fontes:2014:AHB

[FBB14] Christopher J. Fontes, Christopher J. Bostock, and Klaus Bartschat. Annotation of Hans Bethe’s paper, Zeitschrift für Physik 76, 293 (1932), “Braking Formula for Electrons of Relativistic Speed”. European Physical Journal H, 39(5):517–536, December 2014. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-

6467 (electronic). URL http://link.springer.com/article/
10.1140/epjh/e2014-50027-1. See [Bet32, Bet14].

Fermi:1934:NDD

[FBE^{+}34] Enrico Fermi, H. A. Bethe, W. M. Elsasser, H. O. W. Richardson, and K. Sitte. Natural f-decay: Discussion. In International Conference on Physics, London, Vol. I. Nuclear Physics, pages 66–71. Physical Society, London, UK (??), 1934.

Faribault:2009:BAA

[FCC09] Alexandre Faribault, Pasquale Calabrese, and Jean-Sébastien Caux. Bethe ansatz approach to quench dynamics in the Richardson model. Journal of Mathematical Physics, 50(9):095212, September 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v50/i9/p095212_s1.

Fermi:1932:QTR

[Fer32] Enrico Fermi. Quantum theory of radiation. Reviews of Modern Physics, 4(1):87–132, January 1, 1932. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.4.87;
http://rmp.aps.org/abstract/RMP/v4/i1/p87_1.

Fermi:1968:III

[Fer68] Laura Fermi. Illustrious immigrants; the intellectual migration from Europe, 1930/41. University of Chicago Press, Chicago, IL, USA, 1968. xi + 440 pp. LCCN E184.A1 F47.

Fermi:1971:III

[Fer71] Laura Fermi. Illustrious immigrants; the intellectual migration from Europe, 1930/41. University of Chicago Press, Chicago, IL, USA, second edition, 1971. ISBN 0-226-24376-1, 0-226-24378-8 (paperback). xi + 431 pp. LCCN E184.A1 F47 1971.

Festag:1977:LCU

[Fes77] Eckhard Festag. Letter: Conserving uranium. Bulletin of the Atomic Scientists, 33(10):4–5, December 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bet77d, Maw78].

Ferris:1991:WTP

[FF91] Timothy Ferris and Clifton Fadiman, editors. The world treasury of physics, astronomy, and mathematics. Little, Brown and Co., Boston, MA, USA, 1991. ISBN 0-316-28129-8. xv + 859 pp. LCCN QC71 .W67 1991. Foreword by Clifton Fadiman.

Finkbeiner:2006:JSH

[Fin06] Ann K. Finkbeiner. The Jasons: the secret history of science’s postwar elite. Viking Press, New York, NY, USA, 2006. ISBN 0-670-03489-4. xxx + 304 pp. LCCN Q141

.F536 2006. URL http://en.wikipedia.org/wiki/JASON_%
28advisory_group%29;   http://www.aip.org/aip/writing/
books08.html;   http://www.loc.gov/catdir/enhancements/
fy0719/2005043471-b.html;   http://www.loc.gov/catdir/
enhancements/fy0719/2005043471-d.html.

French:1985:NBC

[FK85] A. P. (Anthony Philip) French and P. J. Kennedy, editors. Niels Bohr: a centenary volume. Harvard University Press, Cambridge, MA, USA, 1985. ISBN 0-674-62415-7, 0-674-62416-5 (paperback). xiv + 403 pp. LCCN QC16.B63 N49 1985. US$27.50.

Fleischer:1973:PWA

[Fle73] J. Fleischer. Partial wave analysis of nucleon-nucleon Bethe– Salpeter equation on the computer. Journal of Computational Physics, 12(1):112–123, May 1973. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999173901721.

Floratos:1976:BSS

[Flo76] E. G. Floratos. Bethe–Salpeter spinor equation at P_{μ} = 0 and SO(5) spinor spherical harmonics. Journal of Mathematical Physics, 17(5):706–709, May 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:

//jmp.aip.org/resource/1/jmapaq/v17/i5/p706_s1.

Ford:1957:BRH

[For57] Kenneth W. Ford. Book review: Hans A. Bethe and Philip Morrison, Elementary Nuclear Theory. American Journal of Physics, 25(1):48–??, January 1957. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://ajp.aapt.org/

resource/1/ajpias/v25/i1/p48_s1.

Forman:1987:BRB

[For87] Paul Forman. Book review: Basic Bethe: Seminal Articles on Nuclear Physics, 1936–1937 by Hans A. Bethe; Robert F. Bacher; M. Stanley Livingston. Isis, 78(3):453, September 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL

http://www.jstor.org/stable/232027.

Fernandez:2013:UMA

[FR13] Bernard Fernandez and Georges Ripka. Unravelling the Mystery of the Atomic Nucleus — a Sixty Year Journey 1896–1956. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 1-4614-4180-3 (hardcover), 1-4614-4181-1 (ebook). xviii + 522 pp. LCCN QC773 .F47 2013.

Frauenfelder:2001:BRS

[Fra01] Hans Frauenfelder. Book review: S. S. Schweber, In the Shadow of the Bomb: Bethe, Oppenheimer, and the Moral Responsibility of the Scientist. Princeton: Princeton University Press, 2000, xviii + 260 pages. $24.95 (cloth). Physics in Perspective (PIP), 3(4):493– 495, November 2001. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Friedlander:1999:OCH

[Fri99] Blaine P. Friedlander Jr. An overflow crowd of Hans Bethe’s peers hears one of the great historians of physics. Cornell University News Service Web document., March 1999. URL http://www.

news.cornell.edu/releases/March99/APS/Bethe.Manhattan.
html.

Fruhwirth:2003:GMA

[Frü03] R. Frühwirth. A Gaussian-mixture approximation of the Bethe– Heitler model of electron energy loss by bremsstrahlung. Computer Physics Communications, 154(2):131–142, August 1, 2003. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0010465503002923.

Fowler:1971:AWT

[FSB71] J. L. Fowler, P. J. Siemens, and H. A. Bethe. Atoms within toroidal nuclei. Bulletin of the American Physical Society, 16(??):1151, ???? 1971. CODEN BAPSA6. ISSN 0003-0503.

Fernbach:1970:CTR

[FT70] Sidney Fernbach and Abraham Haskel Taub, editors. Computers and their role in the physical sciences. Gordon and Breach, New York, 1970. ISBN 0-677-14030-4. xiv + 624 pp. LCCN Q183.9 .C65.

Fleischer:1975:BSE

[FT75] J. Fleischer and J. A. Tjon. Bethe–Salpeter equation for J = 0 nucleon–nucleon scattering with one-boson exchange. Nuclear Physics B, 84(2):375–396, January 13, 1975. CODEN NUPBBO. ISSN 0550-3213 (print), 1873-1562 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0550321375903132.

Fung:1981:VBY

[Fun81] M. K. Fung. Validity of the Bethe–Yang hypothesis in the deltafunction interaction problem. Journal of Mathematical Physics, 22 (9):2017–2019, September 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v22/i9/p2017_s1.

Furutsu:1985:STS

[Fur85] K. Furutsu. Statistical theory of scattering and propagation through a rough boundary—the Bethe–Salpeter equation and applications. Journal of Mathematical Physics, 26(9):2352–2369, September 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v26/i9/p2352_s1.

Faraut:1986:VAB

[FV86] J. Faraut and G. A. Viano. Volterra algebra and the Bethe– Salpeter equation. Journal of Mathematical Physics, 27(3):840– 848, March 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v27/i3/p840_s1.

Feiveson:1979:FPE

[FvHW79] Harold A. Feiveson, Frank von Hippel, and Robert H. Williams. Fission power: An evolutionary strategy. Science, 203(4378):330– 337, January 26, 1979. CODEN SCIEAS. ISSN 0036-8075 (print),

1095-9203 (electronic).  URL http://inflationdata.com/
Inflation/Inflation_Rate/Historical_Oil_Prices_Table.asp;
 http://world-nuclear.org/info/Nuclear-Fuel-Cycle/Uranium-
Resources/Supply-of-Uranium/;   http://www.infomine.com/
investment/metal-prices/uranium-oxide/;   http://www.jstor.
org/stable/1747758;   http://www.nuclearfaq.ca/cnf_sectionG.
htm#uranium_supply;   http://www.oil-price.net/;   http:
//www.sciencemag.org/content/203/4378/330.abstract.

Gamow:1930:MDC

[Gam30] George Gamow. Mass defect curve and nuclear constitution. Proceedings of the Royal Society of London. Series A, Containing Papers of a Mathematical and Physical Character, 126(803):632–644, March 3, 1930. ISSN 0950-1207 (print), 2053-9150 (electronic).

URL http://www.jstor.org/stable/pdfplus/95297.pdf.

Gamow:1949:RC

[Gam49] George Gamow. On relativistic cosmogony. Reviews of Modern Physics, 21(3):367–373, July 1949. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//adsabs.harvard.edu/abs/1949RvMP...21..367G;   http:/
/link.aps.org/doi/10.1103/RevModPhys.21.367;   http://
rmp.aps.org/abstract/RMP/v21/i3/p367_1.

Gamow:1966:CTO

[Gam66] George Gamow. Cosmological theories of the origin of chemical elements. In Robert E. Marshak and J. Warren Blaker, editors, Perspectives in modern physics: essays in honor of Hans A. Bethe on the occasion of his 60th birthday, July 1966, page ?? Wiley, New York, NY, USA, 1966. LCCN ????

Garajeu:2002:BAG

[Gar02] Daniela Garajeu. Bethe ansatz for the Gaudin model and its relation with Knizhnik–Zamolodchikov equations. Journal of Mathematical Physics, 43(11):5732–5756, November 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Gasieiger:1993:EHB

[Gas93] Eric L. Gasieiger. An evening with Hans Bethe: the German Atomic Bomb Project. Videorecording produced at Cornell University, Dept. of Physics, 1993. URL http://ifup.cit.cornell.

edu/bethe/.

Gaudin:1983:FOB

[Gau83] Michel Gaudin. La fonction d’onde de Bethe. (French) [The Bethe wavefunction]. Collection du Commissariat à l’énergie atomique. Série scientifique. Masson Editeur, Masson, France, 1983. ISBN 2-225-79607-6. ISSN 0294-0248. xvi + 330 pp. LCCN QC174.26.W3 G38 1983.

Gaudin:2014:BW

[Gau14] Michel Gaudin. The Bethe wavefunction. Cambridge University Press, Cambridge, UK, 2014. ISBN 1-107-04585-1. xv + 324 pp. LCCN QC174.26.W3 G3813 2014. URL http://

www.cambridge.org/9781107045859;   http://www.loc.gov/
catdir/enhancements/fy1408/2013036990-b.html;   http:/
/www.loc.gov/catdir/enhancements/fy1408/2013036990-d.html;
 http://www.loc.gov/catdir/enhancements/fy1408/2013036990-
t.html.

Gluckstern:1951:NDS

[GB51] R. L. Gluckstern and H. A. Bethe. Neutron–deuteron scattering at high energy. Physical Review, 81(5):761–781, March 1, 1951. CO-DEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

URL http://prola.aps.org/abstract/PR/v81/i5/p761_1.

Garwin:1968:ABM

[GB68] Richard L. Garwin and Hans A. Bethe. Anti-ballistic-missile systems. Scientific American, 218(3):21–31, March 1968. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.nature.com/scientificamerican/
journal/v218/n3/pdf/scientificamerican0368-21.pdf.

Galison:1989:LSD

[GB89] Peter Galison and Barton Bernstein. In any light: Scientists and the decision to build the superbomb, 1952–1954. Historical Studies in the Physical and Biological Sciences, 19(2):267–347, ???? 1989. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (elec-

tronic). URL http://www.jstor.org/stable/27757627.

Golden:1986:PSA

[GBG^{+}86] W. T. Golden, R. F. Bacher, A. J. Goodpaster, E. R. Piore, I. I. Rabi, J. R. Kilian, and H. A. Bethe. President’s Science Advisory Committee revisited. Science, Technology and Human Values, 11 (2):5–19, ???? 1986. CODEN ???? ISSN 0162-2439 (print),1552-8251 (electronic).

Gehrenbeck:1973:CJD

[Geh73] Richard Keith Gehrenbeck. C. J. Davisson, L. H. Germer, and the Discovery of Electron-Diffraction. Ph.D. thesis, Department of Physics, University of Minnesota, St. Paul, MN, USA, 1973. 369

pp. URL http://search.proquest.com/docview/302665039?
accountid=14677.

Gehrenbeck:1978:EDF

[Geh78] Richard K. Gehrenbeck. Electron diffraction: fifty years ago. Physics Today, 31(1):34–41, January 1978. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.

nobelprize.org/nobel_prizes/physics/laureates/1937/;
http://www.physicstoday.org/resource/1/phtoad/v31/i1/p34_
s1.

Gerrish:1975:LTM

[Ger75] Laina Gerrish. Letters: Too much. Bulletin of the Atomic Scientists, 31(6):5, June 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [BAA^{+}75].

Garwin:2005:HWP

[GG05] Richard L. Garwin and Kurt Gottfried. Hans in war and peace. Physics Today, 58(10):52–57, October 2005. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.

physicstoday.org/resource/1/phtoad/v58/i10/p52_s1.

Gibbs:2005:HAB

[Gib05] W. Wayt Gibbs. Hans Albrecht Bethe, 1906–2005. Scientific American, 292(5):36, May 2005. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v292/n5/full/scientificamerican0505-
36a.html;   http://www.nature.com/scientificamerican/journal/
v292/n5/pdf/scientificamerican0505-36a.pdf.

Girdner:1975:LNA

[Gir75] William Girdner. Letters: No alternative? Bulletin of the Atomic Scientists, 31(6):4–5, June 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [BAA^{+}75].

Gottfried:1984:RSW

[GKB^{+}84] Kurt Gottfried, H. W. Kendall, Hans A. Bethe, Peter A. Clausen, Richard L. Garwin, et al. Reagan’s Star Wars. New York Review of Books, 31(7):47–52, April 26, 1984. ISSN 0028-7504 (print), 1944-

7744 (electronic). URL http://www.nybooks.com/articles/
archives/1984/apr/26/reagans-star-wars/. Excerpts from the

report by the Union of Concerned Scientists on “Space-Based Missile Defense” — often referred to as “Star Wars.”.

Glackin:1976:LCD

[Gla76] James J. Glackin. Letter: Courting destruction. Bulletin of the Atomic Scientists, 32(3):3–4, March 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Dud75].

Goldberger:1968:BRB

[Gol68] M. L. Goldberger. Book review: Perspectives in Modern Physics. Essays in honor of Hans A. Bethe on the occasion of his 60th birthday, July 1966. R. E. Marshak, Ed. Interscience (Wiley), New York, 1966. xii + 673 pp., illus. $19.50. Science, 160(3828):666–667, May 10, 1968. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-

9203 (electronic). URL http://www.sciencemag.org/content/
160/3828/666.2.full.pdf.

Gold:2005:CCL

[Gol05] Lauren Gold. Cornell to celebrate the life of Hans Bethe. Cornell Chronicle, ??(??):??, September 7, 2005. URL http://news.

cornell.edu/stories/2005/09/cornell-will-celebrate-life-
hans-bethe-sept-18.

Goodchild:1980:JRO

[Goo80] Peter Goodchild. J. Robert Oppenheimer: shatterer of worlds. British Broadcasting Corporation, London, UK, 1980. ISBN 0-563-17781-0. 301 pp. LCCN ???? Published in conjunction with the BBC/WGBH television series Oppenheimer, produced by Peter Goodchild and written by Peter Prince. See important corrections [Bet82b, page 53].

Goodchild:1981:JRO

[Goo81] Peter Goodchild. J. Robert Oppenheimer: shatterer of worlds. Houghton-Mifflin, Boston, MA, USA, 1981. ISBN 0-395-30530-6. 301 pp. LCCN ???? Published in conjunction with the BBC/WGBH television series Oppenheimer, produced by Peter Goodchild and written by Peter Prince. See important corrections [Bet82b, page 53].

Goodstein:1999:CHB

[Goo99] Judith Goodstein. A conversation with Hans Bethe. Physics in Perspective (PIP), 1(3):253–281, October 1999. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s000160050022;   http:
//www.springerlink.com/content/k4f0pllveakla47q/.

Goodstein:2001:CFR

[Goo01] Judith R. Goodstein. A conversation with Franco Rasetti. Physics in Perspective (PIP), 3(3):271–313, September 2001. CODEN PH-PEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http:

//link.springer.com/article/10.1007/PL00000534. This in-

terview was conducted 4 February 1982 but was unpublished for 19 years.

Gottfried:1998:BRH

[Got98] Kurt Gottfried. Book review: Hans A. Bethe, Selected Works of Hans A. Bethe: With Commentary. Physics Today, 51(7):65–66, July 1998. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://link.aip.org/link/phtoad/v51/i7/
p65/s1.

Gottfried:2005:SIH

[Got05] Kurt Gottfried. Special issue: Hans Bethe. Physics Today, 58(10): 36–37, October 2005. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Gottfried:2006:HB

[Got06a] Kurt Gottfried. Hans Bethe. In Brown and Lee [BL06a], pages 125–129. ISBN 981-256-609-0, 981-256-610-4 (paperback), 981-277-450-5 (e-book). LCCN QC16.B46 H36 2006. URL http://

www.worldscibooks.com/physics/5989.html;   http://www.
worldscientific.com/worldscibooks/10.1142/5989.

Gottfried:2006:OHB

[Got06b] Kurt Gottfried. Obituary: Hans A. Bethe. In Brown and Lee [BL06a], pages 279–281. ISBN 981-256-609-0, 981-256-610-4 (paperback), 981-277-450-5 (e-book). LCCN QC16.B46 H36 2006.

URL http://www.worldscibooks.com/physics/5989.html;
http://www.worldscientific.com/worldscibooks/10.1142/5989.

Goudsmit:1947:GSA

[Gou47] Samuel A. Goudsmit. German scientists in Army employment: I. The case analyzed. Bulletin of the Atomic Scientists, 3(2):64, 67, February 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Grodzins:1963:AAS

[GR63] Morton Grodzins and Eugene I. Rabinowitch, editors. The Atomic Age: Scientists in National and World Affairs. Articles from the Bulletin of the Atomic Scientists 1945–1962. Basic Books, New York, NY, USA, 1963. xviii + 616 pp. LCCN D842 .B78. With the assistance of Harvey Flaumenhaft and Lois Gradner.

Gottfried:2005:OHB

[GS05] Kurt Gottfried and Edwin E. Salpeter. Obituary: Hans A. Bethe (1906–2005). Nature, 434(7036):970–971, April 20, 2005. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v434/
n7036/full/434970a.html.

Garwin:1983:PBS

[GSB^{+}83] Richard Garwin, Carl Sagan, Hans A. Bethe, E. Margaret Burbidge, Clark R. Chapman, Thomas M. Donahue, Sidney D. Drell, Lee A. DuBridge, Herbert Freedman, Admiral Noel Gayler, Donald M. Hunten, Christopher C. Kraft, Vice-Admiral John Marshall Lee, Franklin A. Long, Carson Mark, James S. Martin, Jr., David Morrison, Philip Morrison, Bruce C. Murray, Gerry Neugebauer, Tobias Owen, Wolfgang K. H. Panofsky, Gordon H. Pettengill, William Pickering, Edward M. Purcell, I. I. Rabi, George W. Rathjens, Glenn T. Seaborg, Eugene M. Shoemaker, John A. Simpson, Edward C. Stone, James Van Allen, Victor F. Weisskopf, Jerome B. Wiesner, Robert R. Wilson, Herbert F. York, et al. Perspective: Ban space weapons. Bulletin of the Atomic Scientists, 39(9):2–3, November 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldhaber:2003:SMP

[GSS^{+}03] Alfred S. Goldhaber, R. Shrock, J. Smith, G. Sterman, P. van Nieuwenhuizen, and W. Weisberger, editors. Symmetry & modern physics: Yang Retirement Symposium: State University of New York, Stony Brook, 21–22 May 1999. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2003. ISBN 981-238-503-7 (hardcover), 981-238-563-0 (paperback), 981-279-508-1 (e-book). LCCN QC174.17.S9 S994

2003. URL http://www.worldscientific.com/doi/abs/10.
1142/9789812795083_0001;   http://www.worldscientific.
com/worldscibooks/10.1142/5349.

Gamow:1938:LER

[GT38] George Gamow and Edward Teller. Letter to editor: The rate of selective thermonuclear reactions. Physical Review, 53(7):608– 609, April 1, 1938. CODEN PHRVAO. ISSN 0031-899X (print),

1536-6065 (electronic). URL http://prola.aps.org/abstract/
PR/v53/i7/p608_1.

Gulbransen:1975:LSE

[Gul75] Earl A. Gulbransen. Letters: Not safe enough. Bulletin of the Atomic Scientists, 31(6):5, June 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [BAA^{+}75, Pal76].

Gunther:1964:SAV

[Gün64] Marian Günther. Separation of angular variables in the Bethe– Salpeter equation for two spin-1/2 particles. Journal of Mathematical Physics, 5(2):188–192, February 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v5/i2/p188_s1.

Guran:2001:AIS

[Gur01] A. (Ardéshir) Guran, editor. Acoustic interactions with submerged elastic structures. Part 2, Propagation, ocean acoustics and scattering. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2001. ISBN 981-02-2964-X. xxiii + 343 pp. LCCN ???? Foreword by Hans A. Bethe.

Garwin:2005:MHB

[GvH05] Richard L. Garwin and Frank von Hippel. In memorium: Hans Bethe (1906–2005). Arms Control Today, 35(3):??, April 2005. CODEN ???? ISSN 0196-125X (print), 1943-5754

(electronic). URL http://www.armscontrol.org/act/2005_04/
Bethe;   http://www.armscontrol.org/print/1771.

Garwin:2006:MHB

[GvH06] Richard L. Garwin and Frank von Hippel. In memoriam: Hans Bethe. In Brown and Lee [BL06a], pages 273–278. ISBN 981-256-609-0, 981-256-610-4 (paperback), 981-277-450-5 (e-book). LCCN QC16.B46 H36 2006. URL http://

www.worldscibooks.com/physics/5989.html;   http://www.
worldscientific.com/worldscibooks/10.1142/5989.

Haberer:1969:PCS

[Hab69] Joseph Haberer. Politics and the Community of Science. Van Nostrand Reinhold, New York, NY, USA, 1969. vi + 337 pp. LCCN Q125 .H23 1969.

Hahn:1958:DF

[Hah58] Otto Hahn. The discovery of fission. Scientific American, 198(??):76–84, February 1958. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

crownedanarchist.com/emc2/discovery_of_fission.doc;   http:
//www.nature.com/scientificamerican/journal/v198/n2/pdf/
scientificamerican0258-76.pdf.

Halpern:2017:QLH

[Hal17] Paul Halpern. The Quantum Labyrinth: How Richard Feynman and John Wheeler Revolutionized Time and Reality. Basic Books, New York, NY, USA, 2017. ISBN 0-465-09758-8 (hardcover), 0-465-09759-6 (e-book). ix + 311 pp. LCCN QC174.12 .H347 2017.

Hammond:1973:BPB

[Ham73] Allen L. Hammond. Breeder program: Bethe Panel calls for reorientation. Science, 182(4118):1236–1237, December 21, 1973. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (elec-

tronic). URL http://www.sciencemag.org/content/182/4118/
1236.full.pdf.

Hansen:1974:PDM

[Han74] Carl J. Hansen, editor. Physics of Dense Matter. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1974. ISBN 90-277-0406-6, 90-277-0407-4 (paperback). LCCN QB806 .P48.

Harper:2001:AGG

[Har01] Eamon Harper. In appreciation: George Gamow: Scientific amateur and polymath. Physics in Perspective (PIP), 3(3):335–372, September 2001. CODEN PHPEF2. ISSN 1422-6944 (print),

1422-6960 (electronic). URL http://adsabs.harvard.edu/abs/
2001PhP.....3..335H;   http://link.springer.com/article/
10.1007/PL00000536;   http://www.springerlink.com/content/
m0htmt9ww9b9jjv1/;   http://www.springerlink.com/openurl.
asp?genre=journal&issn=1422-6944.

Hargittai:2007:BRG

[Har07] István Hargittai. Book review: Gerald E. Brown and Chang-Hwan Lee, Hans Bethe and His Physics. Structural Chemistry, 18 (5):723–724, October 2007. CODEN STCHES. ISSN 1040-0400 (print), 1572-9001 (electronic). URL http://link.springer.

com/article/10.1007/s11224-007-9207-5.

Hargittai:2010:JET

[Har10] István Hargittai. Judging Edward Teller: a closer look at one of the most influential scientists of the Twentieth Century. Prometheus Books, Amherst, NY, USA, 2010. ISBN 1-61614-221-9 (hardcover). 575 pp. LCCN QC16.T37 H37 2010. Afterword by Richard Garwin.

Hawking:2011:DSM

[Haw11] Stephen Hawking, editor. The dreams that stuff is made of: the most astounding papers on quantum physics — and how they shook the scientific world. Running Press, Philadelphia, PA, USA, 2011. ISBN 0-7624-3434-1. xi + 1071 pp. LCCN QC173.98 .D74 2011.

Holloway:1940:CSR

[HB40] M. G. Holloway and H. A. Bethe. Cross section of the reaction N^{15} (p, α)C^{12}. Physical Review, 57(8):747, April 15, 1940. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://prola.aps.org/abstract/PR/v57/i8/p747_1.

Heidmann:1951:DCS

[HB51a] J. Heidmann and H. A. Bethe. Z-dependence of the cross section for photocapture by nuclei. Physical Review, 84(2):274–281, October 15, 1951. CODEN PHRVAO. ISSN 0031-899X (print),

1536-6065 (electronic). URL http://prola.aps.org/abstract/
PR/v84/i2/p274_1.

Hurwitz:1951:NCC

[HB51b] H. Hurwitz, Jr. and H. A. Bethe. Neutron capture cross sections and level density. Physical Review, 81(5):898, March 1, 1951. CO-DEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

URL http://prola.aps.org/abstract/PR/v81/i5/p898_1.

Henssge:1969:ZEZ

[HB69] J. Henssge and H. Bethe. Der Zeitfaktor in der Entwicklung zervikaler Osteochondrosen. (German) [The time factor in the development of cervical osteochondroses]. Zeitschrift für Orthopadie und ihre Grenzgebiete, 107(1):62–75, ???? 1969. ISSN 0044-3220.

Holt:2006:HBN

[HB06] Jeremy Holt and Gerald E. Brown. Hans Bethe and the nuclear many-body problem. In Brown and Lee [BL06a], pages 201–237. ISBN 981-256-609-0, 981-256-610-4 (paperback), 981-277-450-5 (e-book). LCCN QC16.B46 H36 2006. URL http://

www.worldscibooks.com/physics/5989.html;   http://www.
worldscientific.com/worldscibooks/10.1142/5989.

Heisenberg:1971:QFP

[HBD71] Werner Heisenberg, Erich Bagge, and H. Peter Dürr, editors. Quanten und Felder; physikalische und philosophische Betrachtungen zum 70. Geburtstag von Werner Heisenberg. (German) [Quanta and fields: physical and philosophical views. On the 70th Birthday of Werner Heisenberg]. Friedrich Vieweg und Sohn, Braunschweig, Germany, 1971. 366 pp. LCCN QC174.45 .Q35.

Hoddeson:1987:DQM

[HBE87] Lillian Hoddeson, Gordon Baym, and Michael Eckert. The development of the quantum-mechanical electron theory of metals: 1928–1933. Reviews of Modern Physics, 59(1):287–327, January 1987. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.59.287;   http://rmp.aps.org/abstract/
RMP/v59/i1/p287_1.

Huang:1970:GFB

[HD70] Justin C. Huang and Brian DeFacio. On the Green’s functions for the Bethe–Salpeter equation. Journal of Mathematical Physics, 11 (2):715–724, February 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v11/i2/p715_s1. See errata [HD72].

Huang:1972:EGF

[HD72] Justin C. Huang and Brian DeFacio. Errata: “On Green’s Functions to the Bethe–Salpeter Equation”, [J. Math. Phys. 11, 715 (1970)]. Journal of Mathematical Physics, 13(3):434, March 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v13/i3/p434_s1. See [HD70].

Heisenberg:1934:CTG

[Hei34] Werner Heisenberg. Considérations théoriques générales sure la structure du noyau. (French) [General theoretical considerations of the structure of the nucleus]. In Cockcroft et al. [CCJ^{+}34], pages 316–?? LCCN ???? Publiés par la commission administrative de l’institut.

Henry:1994:BRM

[Hen94] John Henry. Book reviews: Michael Hunter (ed.), Letters and Papers of Robert Boyle: A Guide to the Manuscripts and Microfilm. Collections from the Royal Society. Bethesda, Maryland: University Publications of America, 1992. Pp. xlix + 90. ISBN 1-55655-217-3. No price given. Peter Jones (ed.), Sir Isaac Newton: A Catalogue of Manuscripts and Papers Collected and Published on Microfilm by Chadwyck-Healey. Cambridge: Chadwyck-Healey, 1991. Pp. xi + 148. ISBN 0-85964-226-7. £50.00 (provided free with the microfilms). British Journal for the History of Science, 27(1):115–116, March 1994. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.jstor.

org/stable/4027586.

Hentschel:2014:CFW

[HH14] Klaus Hentschel and Dieter Hoffmann, editors. Carl Friedrich von Weizsäcker: Physik — Philosophie — Friedensforschung: Leopoldina-Symposium, vom 20. bis 22. Juni 2012 in Halle (Saale): mit 110 Abbildungen und 1 Tabelle, volume 63 of Acta Historica Leopoldina. Deutsche Akademie der Naturforscher Leopoldina — Nationale Akademie der Wissenschafte, Halle (Saale), Germany, 2014. ISBN 3-8047-3244-5. ISSN 0001-5857. LCCN ????

Hormozdiari:1973:GFB

[HHD73] Iraj Hormozdiari, Justin C. Huang, and Brian DeFacio. On the Green’s functions for the Bethe–Salpeter equation. II. Journal of Mathematical Physics, 14(5):563–568, May 1973. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v14/i5/
p563_s1.

Howes:1999:TDS

[HHW99] Ruth (Ruth Hege) Howes, Caroline L. Herzenberg, and Ellen C. Weaver. Their day in the sun: women of the Manhattan Project. Labor and social change. Temple University Press, Philadelphia, PA, USA, 1999. ISBN 1-56639-719-7 (hardcover), 1-59213-192-1 (paperback), 0-585-38881-4 (e-book). viii + 264 pp. LCCN QC773.3.U5 H68 1999.

Hiltzik:2015:BSE

[Hil15] Michael A. Hiltzik. Big Science: Ernest Lawrence, the Cyclotron, and the Birth of the Military–Industrial Complex. Simon and Schuster, New York, NY, USA, 2015. ISBN 1-4516-7575-5 (hardcover), 1-4516-7576-3 (paperback), 1-4516-7603-4 (e-book). x + 512 pp. LCCN QC787.C8 H55 2015.

Hoffman:1937:SDE

[HLB37] J. G. Hoffman, M. Stanley Livingston, and H. A. Bethe. Some direct evidence on the magnetic moment of the neutron. Physical Review, 51(3):214–215, February 1, 1937. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.

aps.org/abstract/PR/v51/i3/p214_1.

Hoier:1983:ESF

[HM83] R. Høier and K. Marthinsen. Effective structure factors in manybeam X-ray diffraction — use of the second Bethe approximation. Acta crystallographica. Section A, 39(6):854–860, November 1983. CODEN ACACBN. ISSN 0567-7394 (print), 1600-8596 (electronic).

Hoddeson:1981:OHT

[Hod81] Lillian Hoddeson. Oral history transcript — Hans Albrecht Bethe. Web document., April 29, 1981. URL http://www.aip.org/

history/ohilist/4505.html.

Hoddeson:2013:NFM

[Hod13] Lillian Hoddeson. Book review: Nuclear Forces: The Making of the Physicist Hans Bethe, Silvan S. Schweber, Harvard U. Press, Cambridge, MA, 2012. $35.00 (579 pp.). ISBN 978-0-674-06587-1. Physics Today, 66(4):50–51, April 2013. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Hoffmann:1996:HBI

[Hof96] Dieter Hoffmann. Hans Bethe — ein Interview zum 90. Geburtstag. (German) [Hans Bethe — an interview on his 90th birthday]. Physik in unserer Zeit, 27(5):214–217, 1996. CODEN PHUZAH. ISSN 0031-9252.

Hoiberg:2001:YNL

[Hoi01] Dale Hoiberg, editor. 100 years with Nobel laureates. Encyclopaedia Britannica (India) and I.K. International, New Delhi, India, 2001. ISBN 81-88237-00-0. xiii + 1098 pp. LCCN AS911.N9 A15 2001.

Holton:1984:MPU

[Hol84] Gerald Holton. The migration of physicists to the United States. Bulletin of the Atomic Scientists, 40(4):18–24, April 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Holbrow:2003:ACC

[Hol03] Charles H. Holbrow. In appreciation: Charles C. Lauritsen: a reasonable man in an unreasonable world. Physics in Perspective (PIP), 5(4):419–472, December 2003. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s00016-003-0178-7. See also

[McC20].

Houtermans:1930:NAQ

[Hou30] Fritz G. Houtermans. Neuere Arbeiten über Quantentheorie des Atomskerns. (German) [New work on the quantum theory of the atomic nucleus]. Ergebnisse der Exakten Naturwissenschaften, 9 (??):123–221, ???? 1930. CODEN EENAA3. ISSN 0367-0325.

Hoyle:1946:SEH

[Hoy46] Fred Hoyle. The synthesis of the elements from hydrogen. Monthly Notices of the Royal Astronomical Society, 106(5):343– 383, ???? 1946. CODEN MNRAA4. ISSN 0035-8711 (print),

1365-2966 (electronic). URL http://adsabs.harvard.edu/abs/
1946MNRAS.106..343H.

Hoyle:1954:NRO

[Hoy54] Fred Hoyle. On nuclear reactions occurring in very hot stars. I. The synthesis of elements from carbon to nickel. Astrophysical Journal. Supplement Series, 1(??):121–146, September 1954. CO-DEN APJSA2. ISSN 0067-0049 (print), 1538-4365 (electronic).

URL http://labs.adsabs.harvard.edu/ui/abs/1954ApJS....

1..121H. Reprinted in the centennial edition Astrophysical Journal 525, 571 (1999) with a modern commentary by David Arnett.

Harper:1997:GGS

[HPA97] Eamon Harper, William Carleton Parke, and G. D. (George David) Anderson, editors. The George Gamow Symposium: sponsored by the George Washington University and the Carnegie Institution of Washington, 12 April 1997 [i.e., 1996], volume 129 of Astronomical Society of the Pacific conference series. Astronomical Society of the Pacific, San Francisco, CA, USA, 1997. ISBN 1-886733-49-X.

LCCN QC16.G37 G46 1996. URL http://adsabs.harvard.edu/
abs/1997ASPC..129.....H.

Hou:2004:ERS

[HSY04] B. Y. Hou, R. Sasaki, and W.-L. Yang. Eigenvalues of Ruijsenaars– Schneider models associated with A_{n−1} root system in Bethe ansatz formalism. Journal of Mathematical Physics, 45(2):559– 575, February 2004. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v45/i2/p559_s1.

Hubisz:2014:BRN

[Hub14] John L. Hubisz. Book review: Nuclear Forces: The Making of the Physicist Hans Bethe, by Silvan S. Schweber and published by Harvard University Press, Cambridge, MA (2012), $35.00, pp. ix + 579, hardback, ISBN: 978-0674-06587-1. The Physics Teacher, 52 (4):255, April 2014. CODEN PHTEAH. ISSN 0031-921X (print), 1943-4928 (electronic).

Hufbauer:1981:ATS

[Huf81] Karl Hufbauer. Astronomers take up the stellar-energy problem, 1917–1920. Historical Studies in the Physical Sciences, 11(2):277– 303, ???? 1981. CODEN HSPSAS. ISSN 0073-2672. URL http:

//www.jstor.org/stable/27757481.

Hufbauer:1982:BRBb

[Huf82] Karl Hufbauer. Book review: Hans Bethe: Prophet of Energy by Jeremy Bernstein. Isis, 73(2):322, June 1982. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/231736.

Hufbauer:2006:SSE

[Huf06] Karl Hufbauer. Stellar structure and evolution, 1924–1939. Journal for the History of Astronomy, 37(2):203–227, May 2006. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556 (electronic).

Inokuti:1978:AIC

[IIT78] Mitio Inokuti, Yukikazu Itikawa, and James E. Turner. Addenda: Inelastic collisions of fast charged particles with atoms and molecules — the Bethe theory revisited. Reviews of Modern Physics, 50(1):23–35, January 1978. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.50.23;   http://
rmp.aps.org/abstract/RMP/v50/i1/p23_1. See [Ino71].

Ince:1992:BRR

[Inc92] Martin Ince. Book review: The Road from Los Alamos, by Hans Bethe. New Scientist, ??(??):442–??, ???? 1992. CODEN NWS-CAL. ISSN 0262-4079 (print), 1364-8500 (electronic).

Inokuti:1971:ICF

[Ino71] Mitio Inokuti. Inelastic collisions of fast charged particles with atoms and molecules — the Bethe theory revisited. Reviews of Modern Physics, 43(3):297–347, July 1971. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.43.297;
 http://rmp.aps.org/abstract/RMP/v43/i3/p297_1. See ad-

dendum [IIT78].

Ioffe:2005:HBG

[Iof05] B. L. Ioffe. Hans Bethe and the global energy problems. arXiv.org,

??(??):??, July 6, 2005. URL http://arxiv.org/abs/physics/
0507043.

Ioffe:2006:HBG

[Iof06] Boris L. Ioffe. Hans Bethe and the global energy problems. In Brown and Lee [BL06a], pages 263–271. ISBN 981-256-609-0, 981-256-610-4 (paperback), 981-277-450-5 (e-book). LCCN QC16.B46

H36 2006. URL http://www.worldscibooks.com/physics/
5989.html;   http://www.worldscientific.com/worldscibooks/
10.1142/5989.

Irion:2006:TPL

[Iri06] Robert Irion. A towering physicist’s legacy faces a threatening future. Science, 313(5783):39–40, July 7, 2006. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/313/5783/39.full.pdf.

Jackiw:2009:HBM

[Jac09] Roman Jackiw. Hans Bethe, my teacher. Physics in Perspective (PIP), 11(1):98–103, March 2009. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://www.

springerlink.com/content/707k21m7328t0860/.

Johnson:1978:CSC

[JB78] M. B. Johnson and H. A. Bethe. Correlations and self-consistency in pion scattering (I). Nuclear Physics A, 305(2):418–460, August 7, 1978. CODEN NUPABL. ISSN 0375-9474 (print), 1873-1554

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0375947478903500.

Jeans:1924:AMS

[Jea24] J. H. Jeans. The ages and masses of the stars. Nature, 114(2875): 828–829, December 1924. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Jehle:1956:BRS

[Jeh56] Herbert Jehle. Book review: Silvan S. Schweber, Hans A. Bethe, and Frederic de Hoffmann, Mesons and Fields. Volume I: Fields. American Journal of Physics, 24(7):534–??, October 1956. CO-DEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

URL http://ajp.aapt.org/resource/1/ajpias/v24/i7/p534_
s1.

Janes:1966:NTA

[JLBF66] G. S. Janes, R. H. Levy, H. A. Bethe, and B. T. Feld. New type of accelerator for heavy ions. Physical Review, 145(3):925–952, May 20, 1966. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065

(electronic). URL http://prola.aps.org/abstract/PR/v145/
i3/p925_1.

Juretschke:1986:PPE

[JMWB86] H. J. Juretschke, A. F. Moodie, H. K. Wagenfeld, and H. A. Bethe. Paul P. Ewald. Physics Today, 39(5):101–104, May 1986. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL

http://link.aip.org/link/phtoad/v39/i5/p101/s1.

Jungk:1958:BTT

[Jun58] Robert Jungk. Brighter than a thousand suns: a personal history of the atomic scientists. Harcourt Brace, New York, NY, USA, 1958. 369 pp. LCCN QC773 .J813. Translated by James Cleugh.

Jungk:1982:BTT

[Jun82] Robert Jungk. Brighter than a thousand suns: a personal history of the atomic scientists. Penguin, London, UK and New York, NY, USA, 1982. ISBN 0-14-020667-1 (paperback). 329 pp.

Jurco:1989:QIM

[Jur89] Branislav Jurčo. On quantum integrable models related to nonlinear quantum optics. An algebraic Bethe ansatz approach. Journal of Mathematical Physics, 30(8):1739–1743, August 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v30/i8/
p1739_s1.

Kaiser:2005:CBL

[Kai05a] David Kaiser. Climbing Bethe’s ladder: Feynman diagrams and the many-body problem. In DTA [Kai05b], pages 230–243. ISBN 0-226-42266-6, 0-226-42267-4 (paperback). LCCN QC794.6.F4

K35 2005. URL http://www.loc.gov/catdir/bios/uchi051/
2004023335.html;   http://www.loc.gov/catdir/enhancements/
fy0617/2004023335-d.html;   http://www.loc.gov/catdir/
toc/ecip051/2004023335.html.

Kaiser:2005:DTA

[Kai05b] David Kaiser. Drawing theories apart: the dispersion of Feynman diagrams in postwar physics. University of Chicago Press, Chicago, IL, USA, 2005. ISBN 0-226-42266-6, 0-226-42267-4 (paperback). xix + 469 pp. LCCN QC794.6.F4 K35 2005.

URL http://www.loc.gov/catdir/bios/uchi051/2004023335.
html;   http://www.loc.gov/catdir/enhancements/fy0617/
2004023335-d.html;   http://www.loc.gov/catdir/toc/ecip051/
2004023335.html.

Konopinski:1938:TEF

[KB38] E. J. Konopinski and H. A. Bethe. The theory of excitation functions on the basis of the many-body model. Physical Review, 54(2):130–138, July 15, 1938. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.aps.

org/abstract/PR/v54/i2/p130_1.

Keam:1966:PCG

[Kea66] R. F. Keam. Properties of a causal Green’s function for the Bethe– Salpeter equation. Journal of Mathematical Physics, 7(12):2196– 2210, December 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v7/i12/p2196_s1.

Keam:1968:RBS

[Kea68] R. F. Keam. Reduction of the Bethe–Salpeter equation for a zeromass system. Journal of Mathematical Physics, 9(9):1462–1467, September 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v9/i9/p1462_s1.

Keam:1969:SBS

[Kea69] R. F. Keam. Solution of the Bethe–Salpeter equation with a square-well potential. Journal of Mathematical Physics, 10(4): 594–603, April 1969. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v10/i4/p594_s1.

Keam:1970:BSE

[Kea70] R. F. Keam. Bethe–Salpeter equation and the Goldstein problem. Journal of Mathematical Physics, 11(2):394–403, February 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v11/i2/p394_s1.

Keam:1971:PTB

[Kea71] R. F. Keam. Perturbation treatment of the Bethe–Salpeter equation for tightly bound fermion-antifermion systems. Journal of Mathematical Physics, 12(3):515–528, March 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v12/i3/
p515_s1.

Kenworthy:1954:DHB

[Ken54] E. W. Kenworthy. The drama of the hydrogen bomb — and Dr. Oppenheimer’s key role: Security case focuses attention on disputes that preceded first successful test of H-bomb at Pacific Proving Ground. New York Times, ??(??):E5, April 18, 1954. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL https://search.proquest.com/hnpnewyorktimes/
docview/113104238/.

Kuhn:1967:SHQ

[KHFA67] Thomas S. Kuhn, John L. Heilbron, Paul Forman, and Lini Allen. Sources for history of quantum physics: an inventory and report, volume 68 of Memoirs of the American Philosophical Society. American Philosophical Society, Philadelphia, PA, USA, 1967. ix + 176 pp. LCCN QC174.1 .S66.

URL http://www.amphilsoc.org/guides/ahqp/;   http://
www.amphilsoc.org/guides/ahqp/s-t.htm#schrodinger.

Karabach:1998:IBA

[KHM98] Michael Karabach, Kun Hu, and Gerhard Müller. Introduction to the Bethe Ansatz. II. Computers in Physics, 12(6):565–573, November 1998. CODEN CPHYE2. ISSN 0894-1866 (print),

1558-4208 (electronic). URL https://aip.scitation.org/doi/
10.1063/1.168740.

Killian:1977:SSE

[Kil77] James Rhyne Killian, Jr. Sputnik, scientists, and Eisenhower: a memoir of the First Special Assistant to the President for Science and Technology. MIT Press, Cambridge, MA, USA, 1977. ISBN 0-262-11066-0. xix + 315 pp. LCCN Q143.K42 A37.

Kulish:2001:COB

[KM01] P. P. Kulish and N. Manojlović. Creation operators and Bethe vectors of the osp(1|2) Gaudin model. Journal of Mathematical Physics, 42(10):4757–4778, October 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kivel:1957:RHA

[KMB57] B. Kivel, H. Mayer, and H. Bethe. Radiation from hot air. Part I. Theory of nitric oxide absorption. Annals of Physics, 2(1): 57–80, July 1957. CODEN APNYA6. ISSN 0003-4916 (print),

1096-035X (electronic). URL http://www.sciencedirect.com/
science/article/pii/0003491657900350.

Karabach:1997:IBA

[KMGT97] Michael Karabach, Gerhard Müller, Harvey Gould, and Jan Tobochnik. Introduction to the Bethe Ansatz. I. Computers in Physics, 11(1):36–??, January 1997. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic). URL https://aip.

scitation.org/doi/10.1063/1.4822511.

Krishnamurthy:1996:PSD

[KMN^{+}96] Ramesh S. (Ramesh Saligrama) Krishnamurthy, Clifford Mead, Mary Jo Nye, Sean C. Goodlett, and Marvin E. Kirk, editors. The Pauling symposium: a discourse on the art of biography: proceedings of the Conference on the Life and Work of Linus Pauling (1901–1994): a Discourse on the Art of Biography: February 28–March 2, 1995, Corvallis, Oregon. Special Collections, Oregon State University Libraries, Corvallis, OR, USA, 1996. ISBN 0-9629082-2-3. LCCN QD22.P35 C651 1995.

Kursunoglu:1998:ENE

[KMP98] Behram Kursunoğlu, Stephan L. Mintz, and Arnold Perlmutter, editors. Environment and nuclear energy: Proceedings of an International Conference on Environment and Nuclear Energy, held October 26–29, 1997 in Washington, DC. Plenum Press, New York, NY, USA; London, UK, 1998. ISBN 0-306-45921-3.

LCCN TK9006 .E54 1998. URL http://www.loc.gov/catdir/
enhancements/fy0821/98041249-d.html;   http://www.loc.
gov/catdir/enhancements/fy0821/98041249-t.html.

Konopinski:1946:IAN

[KMT46] Emil Konopinski, Cloyd Marvin, Jr., and Edward Teller. Ignition of the atmosphere with nuclear bombs. Report LA-602, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, August 14, 1946. Declassified 1974.

Kragh:2013:BRS

[Kra13] Helge Kragh. Book review: Silvan S. Schweber, Nuclear Forces: The Making of the Physicist Hans Bethe. Isis, 104(3):635–636, September 2013. CODEN ISISA4. ISSN 0021-1753 (print),

1545-6994 (electronic). URL http://www.jstor.org/stable/
pdfplus/10.1086/674500.pdf.

Kragh:2016:SSE

[Kra16] Helge Kragh. The source of solar energy, ca. 1840–1910: From meteoric hypothesis to radioactive speculations. European Physical Journal H, 41(4–5):365–394, November 2016. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467 (electronic). URL http://

link.springer.com/accesspage/article/10.1140/epjh/e2016-
70045-7.

Kragh:2017:NPS

[Kra17] Helge Kragh. The Nobel Prize system and the astronomical sciences. Journal for the History of Astronomy, 48(3):257–280, August 2017. CODEN JHSAA2. ISSN 0021-8286 (print), 1753-8556

(electronic). URL http://journals.sagepub.com/doi/full/
10.1177/0021828617721574.

Kleint:2005:WHB

[KRW05] Christian Kleint, Helmut Rechenberg, and Gerald Wiemers, editors. Werner Heisenberg, 1901–1976: Beiträge, Berichte, Briefe: Festschrift zu seinem 100. Geburtstag. (German) [Werner Heisenberg, 1901–1976: Contributions, Reports, and Letters: Celebration of his 100th Birthday], volume 62 of Abhandlungen der Sächsischen Akademie der Wissenschaften zu Leipzig, Mathematisch-Naturwissenschaftliche Klasse; Aufsatzsammlung. Verlag der Sächsischen Akademie der Wissenschaften zu Leipzig, Leipzig, Germany, 2005. ISBN 3-7776-1402-5. ISSN 0365-6470. 424 pp. LCCN QC16.H35 W474 2005. URL http://www.loc.

gov/catdir/toc/fy0613/2006402966.html.

Kuo:1987:WFS

[KS87] T. T. S. (Thomas Tzu Szu) Kuo and J. Speth, editors. Windsurfing the Fermi Sea: proceedings of the International Conference and Symposium on Unified Concepts of Many-Body Problems held at the State University of New York, Stony Brook, September 4–6, 1986. North-Holland Publishing Co., Amsterdam, The Netherlands, 1987. ISBN 0-444-87064-4 (set), 0-444-87065-2 (v. 1), 0-444-87066-0 (v. 2). LCCN QC174.17.P7 I56 1986.

Kutzelnigg:2005:HBLb

[Kut05a] Werner Kutzelnigg. Hans Bethe (1906–2005) and Ligand Field Theory. Angewandte Chemie, International Edition, 44(25):3800– 3801, June 20, 2005. CODEN ACIEF5. ISSN 1433-7851 (print), 1521-3773 (electronic).

Kutzelnigg:2005:HBLa

[Kut05b] Werner Kutzelnigg. Hans Bethe (1906–2005) und die Ligandenfeldtheorie. (German) [Hans Bethe (1906–2005) and Ligand Field Theory]. Angewandte Chemie, 117(25):3866–3867, June 20, 2005. CODEN ANCEAD. ISSN 0044-8249 (print), 1521-3757 (electronic).

Kipphardt:1969:MJR

[KW69] Heinar Kipphardt and Joseph Wiseman. In the matter of J. Robert Oppenheimer [Sound recording]. Caedmon TRS 336, ????, 1969. LCCN ????

Kato:2003:DCF

[KW03] Go Kato and Miki Wadati. A direct calculation of the free energy from the Bethe ansatz equation for the Heisenberg model. Journal of Mathematical Physics, 44(9):4189–4200, September 2003. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Kyriakopoulos:1972:AEB

[Kyr72] E. Kyriakopoulos. Algebraic equations for Bethe–Salpeter and Coulomb Green’s functions. Journal of Mathematical Physics, 13 (11):1729–1735, November 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v13/i11/p1729_s1.

Korepin:2002:ISV

[KZJ02] V. Korepin and P. Zinn-Justin. Inhomogeneous six-vertex model with domain wall boundary conditions and Bethe ansatz. Journal of Mathematical Physics, 43(6):3261–3267, June 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Lai:1974:LGN

[Lai74] C. K. Lai. Lattice gas with nearest neighbor interaction in one dimension with arbitrary statistics. Journal of Mathematical Physics, 15(10):1675–1676, October 1974. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v15/i10/p1675_s1.

Lanouette:2000:BRB

[Lan00] William Lanouette. Book review: In The Shadow of the Bomb: Oppenheimer, Bethe, and the Moral Responsibility of the Scientist, by S. S. Schweber. The Wilson Quarterly (1976–2012), 24(4):138, Autumn 2000. ISSN 0363-3276. URL http://www.jstor.org/

stable/40260132.

Livingston:1937:NPC

[LB37] M. Stanley Livingston and H. A. Bethe. Nuclear physics C. Nuclear dynamics, experimental. Reviews of Modern Physics, 9(3):245–390, July 1, 1937. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.9.245;   http://
rmp.aps.org/abstract/RMP/v9/i3/p245_1.

Levinger:1950:DTN

[LB50] J. S. Levinger and H. A. Bethe. Dipole transitions in the nuclear photo-effect. Physical Review, 78(2):115–129, April 15, 1950. CO-DEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

URL http://prola.aps.org/abstract/PR/v78/i2/p115_1.

Levinger:1952:NYN

[LB52] J. S. Levinger and H. A. Bethe. Neutron yield from the nuclear photoeffect. Physical Review, 85(4):577–581, February 15, 1952. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065

(electronic). URL http://prola.aps.org/abstract/PR/v85/
i4/p577_1.

Leon:1962:NMA

[LB62] M. Leon and H. A. Bethe. Negative meson absorption in liquid hydrogen. Physical Review, 127(2):636–647, July 15, 1962. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://prola.aps.org/abstract/PR/v127/i2/p636_1.

Lee:2007:HAB

[LB07] Sabine Lee and Gerry E. Brown. Hans Albrecht Bethe: 2 July 1906–6 March 2005. Biographical Memoirs of Fellows of the Royal Society, 53:2–20, December 1, 2007. CODEN BMFRA3. ISBN 0-85403-648-2. ISSN 0080-4606 (print), 1748-8494 (electronic). URL

http://www.jstor.org/stable/20461364.

Lewis:1975:RAP

[LBC^{+}75] H. W. Lewis, R. J. Budnitz, A. W. Castleman, D. E. Dorfan, F. C. Finlayson, R. L. Garwin, L. C. Hebel, S. M. Keeny, Jr., R. A. Muller, T. B. Taylor, G. F. Smoot, F. von Hippel, Hans Bethe, W. K. H. Panofsky, and V. F. Weisskopf (APS Study Group Participants and APS Council Review Committee). Report to the American Physical Society by the study group on light-water reactor safety. Reviews of Modern Physics, 47(S1):S1–S123, January 1, 1975. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
47.S1;   http://rmp.aps.org/abstract/RMP/v47/iS1/pS1_1.

Liu:1971:ARS

[LC71] C. J. Liu and Yutze Chow. An algebraic realization of spin-wave theory. Journal of Mathematical Physics, 12(10):2144–2150, October 1, 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v12/i10/p2144_s1.

Lee:2005:HBP

[Lee05] Sabine Lee. Hans A. Bethe publications. Nuclear Physics A, 762 (1–2):13–49, November 28, 2005. CODEN NUPABL. ISSN 0375-9474 (print), 1873-1554 (electronic).

Lee:2007:BPC

[Lee07] Sabine Lee. The Bethe–Peierls correspondence. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2007. ISBN 981-277-135-2. xi + 506 pp. LCCN

QC16.B46 A4 2007eb. URL http://site.ebrary.com/lib/
yale/Doc?id=10255467;   http://www.loc.gov/catdir/toc/
fy0805/2008274437.html;   http://www.worldscientific.com/
worldscibooks/10.1142/6595.

Lee:2007:SRP

[Lee09] Sabine Lee, editor. Sir Rudolf Peierls: selected private and scientific correspondence. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2007–2009. ISBN 981-256-503-5 (vol. 1), 981-279-706-8 (vol. 2). xiv + 866 (volume 1), xii + 1108 (volume 2) pp. LCCN QC16.P375 A4 2007. URL http:

//wspc-prod.literatumonline.com/worldscibooks/10.1142/
6790;   http://www.worldscientific.com/worldscibooks/10.
1142/5941.

Lee:2011:NSF

[Lee11] Sabine Lee, editor. From nuclei to stars: Festschrift in honor of Gerald E. Brown. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2011. ISBN 981-4329-08-8, 981-4329-88-6 (e-book). xiii + 487 pp. LCCN QC776 2011. URL

http://ebooks.worldscinet.com/ISBN/9789814329880/.

Lerche:1971:KDT

[Ler71] Ian Lerche. Kinematic dynamo theory: The Dyson equation and the large-scale field; the Bethe–Salpeter equation and the fluctuation intensity. Journal of Mathematical Physics, 12(8):1538– 1547, August 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v12/i8/p1538_s1.

Links:1999:BAS

[LFK99] Jon Links, Angela Foerster, and Michael Karowski. Bethe ansatz solution of a closed spin 1 XXZ Heisenberg chain with quantum algebra symmetry. Journal of Mathematical Physics, 40(2):726– 735, February 1999. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Logan:1978:FDP

[LG78] Jonothan L. Logan and Samuel A. Goudsmit. The first digit phenomenon. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 122(4):193–197, August 18, 1978. CODEN PAPCAA. ISSN 0003-049X (print),

2326-9243 (electronic). URL http://links.jstor.org/sici?
sici=0003-049X(19780818)122%3A4%3C193%3ATFDP%3E2.0.CO%
3B2-C;   http://www.jstor.org/stable/986530. This paper

contains derivations of both Stigler’s Law and Benford’s Law, and receives strong criticism in [Rai85]. This paper contains an important historical note that is recorded in entry [Ben38].

Liu:1972:PCA

[Liu72] C. J. Liu. Perturbation calculation of an algebraic realization of spin waves. Journal of Mathematical Physics, 13(2):248–252, February 1972. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v13/i2/p248_s1.

Lowen:2013:NFM

[Low13] Rebecca S. Lowen. Book review: Nuclear Forces: The Making of the Physicist Hans Bethe. The Journal of American History, 100 (1):252–253, 2013. ISSN 0021-8723 (print), 1945-2314 (electronic).

URL http://jah.oxfordjournals.org/content/100/1/252.1.
full.pdf+html;   http://jah.oxfordjournals.org/content/
100/1/252.1.short.

Lamb:1947:FSH

[LR47] Willis E. Lamb and Robert C. Retherford. Fine structure of the hydrogen atom by a microwave method. Physical Review (2), 72 (3):241–243, August 1947. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://link.aps.org/doi/

10.1103/PhysRev.72.241.

Lewis:1971:APT

[LW71] Richard S. Lewis and Jane Wilson, editors. Alamogordo plus twenty-five years; the impact of atomic energy on science, technology, and world politics. Viking Press, New York, NY, USA, 1971. ISBN 0-670-11151-1. vi + 281 pp. LCCN QC792 .A43 1971. With Eugene Rabinowitch.

Li:2001:CBA

[LYS01] Guang-Liang Li, Rui-Hong Yue, and Kang-Jie Shi. Coordinate Bethe ansatz for the one-dimensional SU(n) Hubbard model with open boundary conditions. Journal of Mathematical Physics, 42 (6):2466–2476, June 2001. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Mermin:2006:HBC

[MA06] N. David Mermin and Neil W. Ashcroft. Hans Bethe’s contributions to solid-state physics. In Brown and Lee [BL06a], pages 189–200. ISBN 981-256-609-0, 981-256-610-4 (paperback), 981-277-450-5 (e-book). LCCN QC16.B46 H36 2006. URL http://

www.worldscibooks.com/physics/5989.html;   http://www.
worldscientific.com/worldscibooks/10.1142/5989.

Maddox:1983:WWH

[Mad83] John Maddox. Where was Hoyle? It is not in bad taste to suggest another 1983 Nobel prize, sadly overlooked. Nature, 305(5837):750, October 27, 1983. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL https://www.nature.com/nature/
journal/v305/n5937/pdf/305750b0.pdf.

Maddox:1986:HBS

[Mad86] John Maddox. Hans Bethe on solar neutrinos. Nature, 320(6064): 677, April 24, 1986. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v320/n6064/pdf/320677a0.pdf.

Mainland:2003:STB

[Mai03] G. B. Mainland. Solving the two-body, bound-state Bethe– Salpeter equation. Journal of Computational Physics, 192 (1):21–35, November 20, 2003. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999103003450.

Mainland:2004:RBC

[Mai04] G. B. Mainland. The role of boundary conditions in solving finite-energy, two-body, bound-state Bethe–Salpeter equations. Journal of Computational Physics, 197(2):610–623, July 1, 2004. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999103006545.

Marshak:1970:HBR

[Mar70] R. E. Marshak. Hans Bethe. (Russian). Uspekhi Fizicheskikh Nauk, 102(2):281, October 1970. CODEN UFNAAG. ISSN 0042-1294 (print), 1996-6652 (electronic). URL http://ufn.ru/ru/ articles/1970/10/h/. From a Life of Physics, Evening Lectures the International Centre of Theoretical Physics in Trieste, Vienna, IAIA, 1969 [BDH^{+}89]. Translated by V. K. Ignatovich.

Marateck:2008:LEA

[Mar08] Samuel L. Marateck. Letter to the Editor: Alpher, Bethe, Gamow. Physics Today, 61(9):11–12, September 2008. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.

physicstoday.org/resource/1/phtoad/v61/i9/p11_s1.

Matthews:1998:GGR

[Mat98] Robert Matthews. Giants go on the record. New Scientist, ??(??): ??, May 9, 1998. CODEN NWSCAL. ISSN 0262-4079 (print),

1364-8500 (electronic). URL https://www.newscientist.com/
article/mg15821336-300-giants-go-on-the-record/.

Mawson:1978:LCF

[Maw78] Colin A. Mawson. Letter: Capacity figures compared. Bulletin of the Atomic Scientists, 34(3):3, March 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Fes77].

Marshak:1940:GTF

[MB40] Robert E. Marshak and Hans A. Bethe. The generalized Thomas– Fermi method as applied to stars. Astrophysical Journal, 91(??):

329–343, ???? 1940. CODEN ASJOAB. ISSN 0004-637X (print), 1538-4357 (electronic).

Marshak:1941:SSE

[MB41] Robert E. Marshak and Hans A. Bethe. The sources of stellar energy. Transactions of the New York Academy of Sciences, Series II, 3(3 (series II)):74–??, January 1941. CODEN TNYAAE. ISSN 0028-7113 (print), 1939-2966 (electronic).

Marshak:1947:TMH

[MB47] R. E. Marshak and H. A. Bethe. On the two-meson hypothesis. Physical Review, 72(6):506–509, September 15, 1947. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://prola.aps.org/abstract/PR/v72/i6/p506_1.

Maximon:1952:DCS

[MB52] L. C. Maximon and H. A. Bethe. Differential cross section for Bremsstrahlung and pair production. Physical Review, 87(1):156, July 1, 1952. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://prola.aps.org/abstract/PR/
v87/i1/p156_1.

Marshak:1966:PMP

[MB66] R. E. Marshak and J. Warren Blaker, editors. Perspectives in modern physics. Essays in honor of Hans A. Bethe on the occasion of his 60th birthday, July, 1966. Interscience Publishers, New York, NY, USA, 1966. xii + 673 pp. LCCN QC774.B4 M3.

Majumdar:1982:TMB

[MB82] Chanchal K. Majumdar and S. Banerjee. The three magnon bound-state equation in one dimension. Journal of Mathematical Physics, 23(1):71–73, January 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:

//jmp.aip.org/resource/1/jmapaq/v23/i1/p71_s1.

McNamara:1985:RRN

[MB85] Robert S. McNamara and Hans A. Bethe. Reducing the risk of nuclear war: Geneva can be a giant step toward a more secure twenty-first century. Atlantic Monthly, 256(1): 43–51, July 1985. ISSN 1072-7825 (print), 2151-9463 (elec-

tronic). URL http://connection.ebscohost.com/c/articles/
9583518/reducing-risk-nuclear-war.

McNamara:1986:RRN

[MB86] R. S. McNamara and H. A. Bethe. Reducing the risk of nuclear war. Bulletin of Peace Proposals, 17(2):121–130, ???? 1986. ISSN 0007-5035. Reprinted from Atlantic Monthly, July, 1985.

McCoy:2020:NRH

[McC20] Alfred W. McCoy. Nuclear reaction: How an article about the Hbomb landed Scientific American in the middle of the Red Scare. Scientific American, 323(3):73, September 2020. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). See also [Hol03].

McDonald:1962:SCI

[McD62] Donald McDonald. Science as citizen: [interview with Hans Bethe]. Bulletin of the Atomic Scientists, 18(6):25–28, June 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

McKinney:1986:LTB

[McK86] Thomas A. McKinney. Letter: The test ban and reliability. Bulletin of the Atomic Scientists, 42(6):61–62, June/July 1986. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Menzel:1927:SSE

[Men27] D. H. Menzel. The source of solar energy. Science, 65(1688): 431–438, May 1927. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Merzbacher:1968:BRH

[Mer68] Eugen Merzbacher. Book review: Hans A. Bethe, R. E. Marshak, J. W. Blaker, Editors, Perspectives in Modern Physics: Essays in Honor of Hans A. Bethe. Physics Today, 21(1):127–??, January 1968. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Morse:1972:NPM

[MHFW72] Philip M. (Philip McCord) Morse, Bernard T. Held, Herman Feshbach, and Richard Wilson, editors. Nuclear, particle and many body physics. Academic Press, New York, NY, USA, 1972. ISBN 0-12-508201-0 (vol. 1), 0-323-14303-2 (e-book). xxi + 693 pp. LCCN QC173 .N8837 1972. URL http://lccn.loc.gov/

79183529;   http://www.sciencedirect.com/science/book/

9780125082013. Two volume tribute to Amos de-Shalit (1926– 1969).

Mada:2006:ECB

[MIT06] Jun Mada, Makoto Idzumi, and Tetsuji Tokihiro. The exact correspondence between conserved quantities of a periodic box-ball system and string solutions of the Bethe ansatz equations. Journal of Mathematical Physics, 47(5):053507, May 2006. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v47/i5/
p053507_s1.

Malone:1975:NSM

[MJB75] R. C. Malone, M. B. Johnson, and H. A. Bethe. Neutron star models with realistic high-density equations of state. Astrophysical Journal, 199(??):741–748, August 1, 1975. CODEN ASJOAB. ISSN 0004-637X (print), 1538-4357 (electronic).

Molinari:1975:ETB

[MJBA75] A. Molinari, M. B. Johnson, H. A. Bethe, and W. M. Albevier. Effective two-body interaction in simple nuclear spectra. Nuclear Physics A, 239(1):45–73, February 1975. CODEN NUPABL. ISSN 0375-9474 (print), 1873-1554 (electronic). URL http://www.

sciencedirect.com/science/article/pii/037594747591132X.

Metropolis:1987:NDP

[MKR87] N. (Nicholas) Metropolis, Donald M. Kerr, and Gian-Carlo Rota, editors. New Directions in Physics: The Los Alamos 40th Anniversary Volume. Academic Press, New York, NY, USA, 1987. ISBN 0-12-492155-8. xii + 292 pp. LCCN QC44 .N49

1987. URL http://www.osti.gov/energycitations/product.
biblio.jsp?osti_id=6120718&query_id=0.

Margoliash:1983:CSFa

[ML83] D. J. Margoliash and P. W. Langhoff. Computational studies of first-born scattering cross sections I. Spectral properties of Bethe surfaces. Journal of Computational Physics, 49(1):44– 66, January 1983. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999183901146.

Mladjenovic:1998:DYN

[Mla98] Milorad Mladjenović. The Defining Years in Nuclear Physics, 1932–1960s. IOP Publishing, Bristol, UK, 1998. ISBN 0-7503-0472-3 (hardcover). xx + 441 pp. LCCN QC773 .M54 1998.

Manojlovic:2007:ABA

[MN07] Nenad Manojlović and Zoltán Nagy. Algebraic Bethe ansatz for the elliptic quantum group E_{τ,η}(A_{2}^{(2)}). Journal of Mathematical Physics, 48(12):123515, December 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:

//jmp.aip.org/resource/1/jmapaq/v48/i12/p123515_s1.

Moler:1989:LCR

[Mol89] Robert B. Moler. Letter: Circular reasoning? Bulletin of the Atomic Scientists, 45(5):47, June 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bet89c].

Montgomery:1949:CRP

[Mon49] Donald Joseph X. Montgomery, editor. Cosmic Ray Physics. Princeton University Press, Princeton, NJ, USA, 1949. ISBN 1-258-43907-7. viii + 370 pp. LCCN QC485 .M6. Based on lectures given by Marcel Schein at Princeton University.

Morland:2002:BKC

[Mor02] Howard Morland. The 1979 Bethe–Kidder correspondence. Web

page, October 2002. URL http://www.fas.org/sgp/eprint/
bethe-kidder.html.

Moynihan:1997:SMC

[Moy97] Daniel Patrick Moynihan. Senator Moynihan’s comments on Bethe letter. From the Congressional Record Online pages S5887– S5889, DOCID:cr17jn97-209., 1997. URL http://www.fas.org/

bethecr.htm.

Mainland:1991:DCS

[MP91] G. Bruce Mainland and L. Alayne Parson. Derivation of the coordinate system that separates the Bethe–Salpeter equation for the Wick–Cutkosky model. Journal of Mathematical Physics, 32 (4):1072–1075, April 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v32/i4/p1072_s1.

Molinari:1985:NSM

[MR85] A. Molinari and R. A. Ricci, editors. From nuclei to stars: a meeting in nuclear physics and astrophysics exploring the path opened by H. A. Bethe, volume 91 of Proceedings of the International School of Physics “Enrico Fermi”. North-Holland Publishing Co., Amsterdam, The Netherlands, 1985. ISBN 0-444-86988-3 (U.S.). LCCN QC770 .I645 1984.

Morton-Smith:2015:O

[MS15] Tom Morton-Smith. Oppenheimer. Oberon Books Ltd., London, UK, 2015. ISBN 1-78319-198-8. ???? pp. LCCN

???? URL http://stratford-upon-avon-theatre.blogspot.
com/2015/01/rsc-oppenheimer-review-stratford-upon.html.

Mermin:2005:MHB

[MTG05] David Mermin, Saul Teukolsky, and Kurt Gottfried. Memories of Hans Bethe. Cornell Chronicle, ??(??):??, ???? 2005.

URL http://web.archive.org/web/20121008171439/http://
www.news.cornell.edu/Chronicle/05/3.10.05/Bethe_memories.
html;   http://www.news.cornell.edu/Chronicle/05/3.10.
05/Bethe_memories.html.

Masters:1946:OWN

[MW46] Dexter Masters and Katharine Way, editors. One world or none: a report to the public on the full meaning of the atomic bomb. Whittlesey House, McGraw-Hill Book Co., Inc., New York, NY, USA, 1946. x + 79 pp. LCCN UF767 .M3 1946a. Foreword by Niels Bohr. Introduction by Arthur H. Compton. See also reprint [MW07].

Masters:1947:OWN

[MW47] Dexter Masters and Katharine Way, editors. One world or none. Latimer House, London, UK, 1947. 159 pp. LCCN UF767 .M3 1947.

Masters:2007:OWN

[MW07] Dexter Masters and Katharine Way, editors. One world or none: a report to the public on the full meaning of the atomic bomb. New Press, New York, NY, USA, 2007. ISBN 1-59558-227-4 (hardcover). xx + 220 pp. LCCN UG1282.A8 O54 2007. URL http:/

/thenewpress.com/index.php?option=com_title&task=view_
title&metaproductid=1703;   http://www.loc.gov/catdir/

toc/ecip0718/2007020838.htm. Foreword by Niels Bohr. Intro-

duction by Arthur H. Compton. Reprint of [MW46].

McLaughlin:2002:BSI

[MWBB02] G. C. McLaughlin, R. A. M. J. Wijers, G. E. Brown, and H. A. Bethe. Broad and shifted iron-group emission lines in gamma-ray bursts as tests of the hypernova scenario. Astrophysical Journal, 567(1):454–462, ???? 2002. CODEN ASJOAB. ISSN 0004-637X (print), 1538-4357 (electronic).

Myra:1988:NTS

[Myr88] Eric S. Myra. Neutrino transport in stellar collapse. Physics Reports, 163(1–3):127–136, 1988. CODEN PRPLCM. ISSN 0370-1573 (print), 1873-6270 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0370157388900397.

Mosteller:1967:DSR

[MYZ67] Frederick Mosteller, Cleo Youtz, and Douglas Zahn. The distribution of sums of rounded percentages. Demography, 4(2):850– 858, June 1967. CODEN ???? ISSN 0070-3370 (print), 1533-

7790 (electronic). URL http://muse.jhu.edu/journals/dem/
;   http://www.biomedsearch.com/nih/Distribution-Sums-
Rounded-Percentages/21318695.html;   http://www.jstor.

org/stable/2060324. See [DF79] for further work.

Nomoto:1988:PEM

[NaH88] Ken’ichi Nomoto and Masa aki Hashimoto. Presupernova evolution of massive stars. Physics Reports, 163(1–3):13–36, 1988. CODEN PRPLCM. ISSN 0370-1573 (print), 1873-6270

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0370157388900324.

Nakanishi:1963:ISE

[Nak63a] Noboru Nakanishi. Invariant solutions of the exact Bethe–Salpeter equation in the general-mass case. Journal of Mathematical Physics, 4(10):1229–1235, October 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:

//jmp.aip.org/resource/1/jmapaq/v4/i10/p1229_s1.

Nakanishi:1963:RDD

[Nak63b] Noboru Nakanishi. Remarks on the double dispersion approach to the Bethe–Salpeter equation. Journal of Mathematical Physics, 4(10):1235–1240, October 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v4/i10/p1235_s1.

Nakanishi:1966:PPV

[Nak66] Noboru Nakanishi. Poles of the proper vertex function in the Bethe–Salpeter formalism. Journal of Mathematical Physics, 7(4): 698–701, April 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v7/i4/p698_s1.

Nakanishi:1971:VSS

[Nak71] Noboru Nakanishi. Vector-scalar sector solutions to the spinor– spinor Bethe–Salpeter equation. Journal of Mathematical Physics, 12(8):1578–1582, August 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v12/i8/p1578_s1.

Navarro:2006:EAD

[Nav06] Jaume Navarro. Early attempts to detect the neutrino at the Cavendish Laboratory. Physics in Perspective (PIP), 8(1):64–82, March 2006. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-

6960 (electronic). URL http://link.springer.com/article/
10.1007/s00016-005-0249-z.

Navarro:2014:BRS

[Nav14] Jaume Navarro. Book reviews: Silvan S. Schweber, Nuclear Forces: The Making of the Physicist Hans Bethe. Cambridge, MA and London: Harvard University Press, 2012. Pp. viii + 579. ISBN 978-0-674-06587-1. £25.95 (hardback). British Journal for the History of Science, 47(3):579–580, September 2014. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

Nemeth:1968:STF

[NB68] Judit Nemeth and Hans A. Bethe. A simple Thomas–Fermi calculation for semi-infinite nuclei. Nuclear Physics A, 116(2):241– 255, August 5, 1968. CODEN NUPABL. ISSN 0375-9474 (print),

1873-1554 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0375947468906593.

Nordheim:1934:TEM

[NBP^{+}34] L. Nordheim, H. A. Bethe, R. Peierls, A. Sommerfeld, L. Brillouin, A. W. Maue, R. H. Fowler, and H. Jones. La théorie des électrons dans les métaux. (French) [The theory of electrons in metals]. Helvetica Physica Acta, 7(Supplement 2):1–87, 1934. CO-DEN HPACAK. ISSN 0018-0238 (print), 2297-1971 (electronic).

Negele:2005:HBT

[Neg05] John W. Negele. Hans Bethe and the theory of nuclear matter. Physics Today, 58(10):58–61, October 2005. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Negele:2006:HBT

[Neg06] John W. Negele. Hans Bethe and the theory of nuclear matter. In Brown and Lee [BL06a], pages 165–173. ISBN 981-256-609-0, 981-256-610-4 (paperback), 981-277-450-5 (e-book). LCCN QC16.B46

H36 2006. URL http://www.worldscibooks.com/physics/
5989.html;   http://www.worldscientific.com/worldscibooks/
10.1142/5989.

Novick:1976:LD

[Nov76] Sheldon Novick. Letter: A disclaimer. Bulletin of the Atomic Scientists, 32(3):4, March 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [vHWD76].

Nussinov:1966:APB

[NR66] S. Nussinov and J. Rosner. Absorptive parts and the Bethe– Salpeter equation for forward scattering. Journal of Mathematical Physics, 7(9):1670–1679, September 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v7/i9/p1670_s1.

Nityananda:2015:SRM

[NS15] Rajaram Nityananda and Arnold Sommerfeld. Some reminiscences of my teaching career. Resonance, 20(5):472–475, May 2015. CO-DEN RESOFE. ISSN 0971-8044 (print), 0973-712X (electronic).

Oppenheimer:19xx:JRO

[OBxx] J. Robert Oppenheimer and Niels Bohr. J. Robert Oppenheimer papers (1799–1980) (bulk 1947–1967) and Niels Henrik David Bohr papers. 19xx. ???? pp. LCCN 0456G; Vault 0202A; Mi-

crofilm 16,646-1P. URL http://hdl.loc.gov/loc.mss/eadmss.
ms998007.

Olwell:2001:BRB

[Olw01] Russell B. Olwell. Book review: In the Shadow of the Bomb: Oppenheimer, Bethe, and the Moral Responsibility of the Scientist. Technology and Culture, 42(4):823–825, October 2001. ISSN 0040-165X (print), 1097-3729 (electronic). URL http:/

/muse.jhu.edu/journals/technology_and_culture/v042/42.
4olwell.html.

ORaifeartaigh:2018:ILE

[OM18] Cormac O’Raifeartaigh and Simon Mitton. Interrogating the legend of Einstein’s ”biggest blunder”. Physics in Perspective (PIP), 20(4):318–341, December 2018. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Ong:1986:CFS

[Ong86] Michael K. Ong. A closed form solution of the s-wave Bethe– Goldstone equation with an infinite repulsive core. Journal of Mathematical Physics, 27(4):1154–1158, April 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v27/i4/
p1154_s1.

Ong:1988:SWB

[Ong88] Michael K. Ong. The solution of the s-wave Bethe–Goldstone equation with a standard hard core potential. Journal of Mathematical Physics, 29(6):1477–1485, June 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:

//jmp.aip.org/resource/1/jmapaq/v29/i6/p1477_s1.

Oppenheimer:1946:ICAb

[Opp46] J. R. Oppenheimer. International control of atomic energy. Foreign Affairs, ??(??):??, ???? 1946. CODEN FRNAA3. ISSN 0015-7120.

Oppenheimer:1966:PMP

[Opp66] J. R. Oppenheimer. Perspectives in modern physics. In Marshak and Blaker [MB66], page ?? LCCN QC774.B4 M3. With the assistance of J. Warren Blaker.

Orbach:2005:TDH

[Orb05] Raymond L. Orbach. Tribute to Dr. Hans A. Bethe by Dr. Raymond L. Orbach, Director, Office of Science. Web document,

March 8, 2005. URL http://science.energy.gov/news/in-
the-news/2005/03-08-05/.

Orear:2004:EFM

[Ore04] Jay Orear, editor. Enrico Fermi — The Master Scientist. The Internet-First University Press, Ithaca, NY, USA, 2004. ISBN ???? iv + 165 pp. LCCN ???? URL http://dspace.library.

cornell.edu/handle/1813/62.

Palladino:1976:LDZ

[Pal76] N. J. Palladino. Letter: Defends zirconium. Bulletin of the Atomic Scientists, 32(3):5, March 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Gul75].

Palevsky:2000:AFD

[Pal00] Mary Palevsky. Atomic Fragments: a Daughter’s Questions. University of California Press, Berkeley, CA, USA, 2000. ISBN 0-520-22055-2 (hardcover), 0-250-22055-2 (hardcover). xiv + 289 pp. LCCN QC773.3.U5 P35 2000. URL http://catdir.loc.gov/

catdir/bios/ucal052/99087422.html;   http://catdir.loc.
gov/catdir/description/ucal042/99087422.html.

Panofsky:2005:EAM

[Pan05] Wolfgang K. H. Panofsky. An episode among the many contributions of Hans Bethe to physics and society. American Physical Society Forum on Physics and Society Newsletter, April

2005. URL http://www.aps.org/units/fps/newsletters/
2005/april/articles.html#panofsky.

Pandharipande:1973:VMD

[PB73] V. R. Pandharipande and H. A. Bethe. Variational method for dense systems. Physical Review C (Nuclear Physics), 7(4):1312– 1328, April 1, 1973. CODEN PRVCAN. ISSN 0556-2813 (print), 1089-490X, 1538-4497.

Petschek:1997:JCM

[PBB97] Albert G. Petschek, George I. Bell, and Hans A. Bethe. J. Carson Mark: Obituary. Physics Today, 50(10):124, October 1997. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

URL http://link.aip.org/link/phtoad/v50/i10/p124/s1.

Pais:2006:JRO

[PC06] Abraham Pais and Robert P. Crease. J. Robert Oppenheimer: a life. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2006. ISBN 0-19-516673-6. xxii + 353 + 16 pp. LCCN QC16.O62 P35 2006. URL http:

//www.loc.gov/catdir/enhancements/fy0636/2005002173-d.
html;   http://www.loc.gov/catdir/enhancements/fy0723/
2005002173-b.html;   http://www.loc.gov/catdir/toc/ecip056/
2005002173.html.

Pang:1990:NBW

[PcPZ90] Gen Di Pang, Fu cho Pu, and Bao Heng Zhao. Norms of Bethe wave functions for the nonlinear Schrödinger model of spin-1/2 particles. Journal of Mathematical Physics, 31(10):2497–2500, October 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i10/p2497_s1.

Peierls:1957:BRF

[Pei57] R. E. Peierls. Book review: Field theory and meson practice. Mesons and Fields, Vol. 1. Fields. By Prof. Silvan Schweber, Prof. Hans. A. Bethe, and Frederic de Hoffmann. Pp. xvi + 449. Vol 2. Mesons. By Prof. Hans. A. Bethe and Frederic de Hoffmann. Pp. xiii + 446. (Evanston, Ill.: Row, Peterson and Company, 1955). Nature, 179(4562):711–712, April 6, 1957. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL https://

ui.adsabs.harvard.edu/#abs/1957Natur.179..711P.

Peierls:1985:BPR

[Pei85] Sir Rudolf E. (Rudolf Ernst) Peierls. Bird of Passage: Recollections of a Physicist. Princeton legacy library. Princeton University Press, Princeton, NJ, USA, 1985. ISBN 0-691-08390-8, 0-691-02416-2 (paperback), 0-691-60220-4, 1-4008-5461-X (ebook). xii + 350 + 12 pp. LCCN QC16.P375 A32 1985.

URL http://site.ebrary.com/id/10897402;   https://www.
jstor.org/stable/j.ctt7ztn4b.

Peierls:1991:NT

[Pei91a] Rudolf Peierls. The nuclear threat: A proposal. New York Review of Books, 38(18):??, November 7, 1991. ISSN 0028-7504 (print), 1944-7744 (electronic). URL http://www.

nybooks.com/articles/archives/1991/nov/07/the-nuclear-

threat-2/. See reply [BGM91b].

Peierls:1991:NTP

[Pei91b] Rudolf Peierls. The nuclear threat: A proposal. New York Review of Books, 38(18):??, November 7, 1991. ISSN 0028-7504 (print), 1944-7744 (electronic). URL http://www.

nybooks.com/articles/archives/1991/nov/07/the-nuclear-

threat-2/. See reply [BGM91b].

Penzias:1979:OE

[Pen79] Arno A. Penzias. The origin of the elements. Reviews of Modern Physics, 51(3):425–431, July 1979. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.51.425;
 http://rmp.aps.org/abstract/RMP/v51/i3/p425_1.

Peplow:2005:HB

[Pep05] Mark Peplow. Hans Bethe. Nature News, March 8, 2005. ISSN

???? URL http://www.nature.com/news/2005/050308/full/
news050307-7.html.

Percus:1963:MBP

[Per63] Jerome K. (Jerome Kenneth) Percus, editor. The many-body problem: proceedings. Interscience Publishers, New York, NY, USA, 1963. LCCN QC174.5 .S93 1957.

Peruggi:1984:PMHa

[Per84a] Fulvio Peruggi. Probability measures and Hamiltonian models on Bethe lattices. I. Properties and construction of MRT probability measures. Journal of Mathematical Physics, 25(11):3303–3315, November 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v25/i11/p3303_s1.

Peruggi:1984:PMHb

[Per84b] Fulvio Peruggi. Probability measures and Hamiltonian models on Bethe lattices. II. The solution of thermal and configurational problems. Journal of Mathematical Physics, 25(11):3316–3323, November 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v25/i11/p3316_s1.

Pevsner:1978:CEI

[Pev78] Charles D. Pevsner. Cornell Engineer interview: Hans A. Bethe. Cornell Engineer, 44(??):14–20, November 1978. CODEN CN-LEAC. ISSN 0010-8790.

Polanyi:2019:WMD

[Pol19] John Polanyi. We must do more to prevent nuclear war. Times Higher Education, ??(??):??, December 9, 2019. URL https://

www.timeshighereducation.com/features/we-must-do-more-
prevent-nuclear-war.

Porter:1997:BBS

[Por97] L. E. Porter. Bethe–Bloch stopping power parameters for polystyrene, kapton, and mylar. International Journal of Quantum Chemistry, 65(5):997–1003, December 5, 1997. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). URL http://

www3.interscience.wiley.com/cgi-bin/abstract?ID=42831;
 http://www3.interscience.wiley.com/cgi-bin/fulltext?ID=

42831&PLACEBO=IE.pdf. Special Issue: Proceedings of the International Symposium on Atomic, Molecular, and Condensed Matter Theory and Computational Methods. Issue Edited by Per-Olov Löwdin, Yngve Öhrn, John R. Sabin, Michael C. Zerner.

Porter:1998:BBS

[Por98] L. E. Porter. Bethe–Bloch stopping-power parameters for GaAs and ZnSe. International Journal of Quantum Chemistry, 70(4–5):919–924, ???? 1998. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). URL http://

www3.interscience.wiley.com/cgi-bin/abstract?ID=75069;
 http://www3.interscience.wiley.com/cgi-bin/fulltext?ID=
75069&PLACEBO=IE.pdf.

Porter:1999:MBB

[Por99] L. E. Porter. Modified Bethe–Bloch stopping power parameters for kapton. International Journal of Quantum Chemistry, 75(4–5):943–950, November 15, 1999. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). URL http://

www3.interscience.wiley.com/cgi-bin/abstract/66004988/
START;   http://www3.interscience.wiley.com/cgi-bin/fulltext?
ID=66004988&PLACEBO=IE.pdf.

Porter:2003:FOP

[Por03] L. E. Porter. Further observations of projectile-z dependence in target parameters of modified Bethe–Bloch theory. International Journal of Quantum Chemistry, 95(4–5):504–511, ???? 2003. CO-DEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Porter:2004:BEC

[Por04a] L. E. Porter. The Barkas-effect correction to Bethe–Bloch stopping power. Advances in Quantum Chemistry, 46 (part 2):91–119, 2004. CODEN AQCHA9. ISSN 0065-3276. URL http://www.

sciencedirect.com/science/article/pii/S0065327604460043.

Porter:2004:SOT

[Por04b] L. E. Porter. Studies of observed trends in values of Bethe–Bloch parameters extracted from stopping power measurements. International Journal of Quantum Chemistry, 100(6):973–980, ???? 2004. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Penney:1944:SHB

[PPvNF44] W. G. Penney, Rudolf E. Peierls, John von Neumann, and Klaus Fuchs. Shock hydrodynamics and blast waves. Report AECD 2860, U.S. Atomic Energy Commission, Technical Information Division, Oak Ridge, TN, USA, October 28, 1944. 60 pp. URL http:/

/hdl.handle.net/2027/mdp.39015086444083. Notes written by

J. O. Hirschfelder. Declassified 22 April 1949. This report is based on LADC-549.

Povill:1997:UFC

[PR97] Àngels Povill and Jaime Rubio. Using Full-CI algorithms in Bethe–Goldstone-type expansions of the correlation energy. International Journal of Quantum Chemistry, 61(1):35–43, January 5, 1997. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-

461X (electronic). URL http://www3.interscience.wiley.com/
cgi-bin/abstract?ID=42383;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?ID=42383&PLACEBO=IE.pdf.

Primakoff:1956:BRH

[Pri56] H. Primakoff. Book review: Hans A. Bethe and Philip Morrison, Elementary Nuclear Theory. Physics Today, 9(11):37–??, November 1956. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://link.aip.org/link/phtoad/v9/i11/
p37/s3.

Percival:1969:APM

[PS69] Ian C. Percival and H. Henry Stroke. Atomic physicists meet for Arnold Sommerfeld centennial. Physics Today, 22(2):99–107, February 1969. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Rabi:19xx:RPB

[Rabxx] I. I. (Isidor Isaac) Rabi. I. I. Rabi papers, 1899–1989 (bulk 1945– 1968). 19xx. ???? pp. LCCN 0639F; LCA (Unprocessed ac. 22,752); Vault 0202A; Oversize 6:10. URL http://hdl.loc.gov/

loc.mss/eadmss.ms998009.

Raimi:1985:FDP

[Rai85] R. Raimi. The first digit phenomenon again. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 129(2):211–219, June 1985. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http:// www.jstor.org/stable/986989. This paper contains strong criticism of a derivation of Benford’s Law [LG78].

Randall:1948:MBS

[Ran48] C. A. Randall. A mnemonic for Bethe’s solar-energy reactions. American Journal of Physics, 16(1):56, January 1948. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL

http://link.aip.org/link/ajpias/v16/i1/p56/s2.

Ray:1997:BAS

[Ray97] Subhankar Ray. Bethe ansatz study for ground state of Fateev Zamolodchikov model. Journal of Mathematical Physics, 38(3): 1524–1558, March 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rose:1937:ENS

[RB37a] M. E. Rose and H. A. Bethe. Erratum: Nuclear Spins and Magnetic Moments in the Hartree Model (Phys. Rev. 51, 2051937). Physical Review, 51(11):993, June 1, 1937. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.aps.

org/abstract/PR/v51/i11/p993_2. See [RB37b].

Rose:1937:NSM

[RB37b] M. E. Rose and H. A. Bethe. Nuclear spins and magnetic moments in the Hartree model. Physical Review, 51(3):205–213, February 1, 1937. CODEN PHRVAO. ISSN 0031-899X (print),

1536-6065 (electronic). URL http://prola.aps.org/abstract/

PR/v51/i3/p205_1. See erratum [RB37a].

Rose:1939:APE

[RB39] M. E. Rose and H. A. Bethe. On the absence of polarization in electron scattering. Physical Review, 55(3):277–289, February 1, 1939. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://prola.aps.org/abstract/PR/
v55/i3/p277_1.

Rabinowitch:1967:JRO

[RB67a] Eugene Rabinowitch and Hans Bethe. J. Robert Oppenheimer, 1904–1967. Bulletin of the Atomic Scientists, 23(8):2–6, October 1967. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rajaraman:1967:TBP

[RB67b] R. Rajaraman and H. A. Bethe. Three-body problem in nuclear matter. Reviews of Modern Physics, 39(4):745–770, October 1, 1967. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.39.745;   http://rmp.aps.org/abstract/
RMP/v39/i4/p745_1.

Ronay:2002:TEH

[RB02] Maria Ronay and Hans A. Bethe. Teller ‘exiled’ himself for science. Physics Today, 55(3):11–12, March 2002. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.org/resource/1/phtoad/v55/i3/p11_s2.

Rutherford:1930:RRS

[RCE30] Sir Ernest Rutherford, James Chadwick, and C. D. (Charles Drummond) Ellis. Radiations from radioactive substances. Cambridge University Press, Cambridge, UK, 1930. xi + 1 + 588 pp. LCCN QC721 .R94.

Redei:2005:JNS

[Réd05] Miklós Rédei. John von Neumann selected letters, volume 27 of History of mathematics. American Mathematical Society, Providence, RI, USA, 2005. ISBN 0-8218-3776-1 (hardcover). ISSN 0899-2428. xxv + 301 pp. LCCN QA29.V66 A4 2005.

Ridenour:1950:HB

[Rid50] Louis N. Ridenour. The hydrogen bomb. Scientific American, 182(3):11–15, March 1950. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.nature.com/scientificamerican/journal/v182/n3/pdf/
scientificamerican0350-11.pdf. See also Parts II [Bet50c] and

III [Bac50].

Righetti:2006:ABG

[Rig06] Pier Giorgio Righetti. The Alpher, Bethe, Gamow of isoelectric focusing, the alpha-Centaury of electrokinetic methodologies. Part I. Electrophoresis, 27(5–6):923–938, March 5–6, 2006. CODEN ELCTDN. ISSN 0173-0835 (print), 1522-2683 (electronic).

Righetti:2007:ABG

[Rig07] Pier Giorgio Righetti. The Alpher, Bethe and Gamow of IEF, the alpha-Centaury of electrokinetic methodologies. Part II: Immobilized pH gradients. Electrophoresis, 28(4):545–555, February 4, 2007. CODEN ELCTDN. ISSN 0173-0835 (print), 1522-2683 (electronic).

Robbiano:2010:IBL

[RJ10] María Robbiano and Raúl Jiménez. Improved bounds for the Laplacian energy of Bethe trees. Linear Algebra and Its Applications, 432(9):2222–2229, April 15, 2010. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rojo:2011:LGC

[RJ11] Oscar Rojo and Raúl D. Jiménez. Line graph of combinations of generalized Bethe trees: Eigenvalues and energy. Linear Algebra and Its Applications, 435(10):2402–2419, November 15, 2011. CO-DEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rojo:2006:TBA

[RM06] Oscar Rojo and Luis Medina. Tight bounds on the algebraic connectivity of Bethe trees. Linear Algebra and Its Applications, 418 (2–3):840–853, October 15, 2006. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rojo:2009:SGB

[RM09] Oscar Rojo and Luis Medina. Spectra of generalized Bethe trees attached to a path. Linear Algebra and Its Applications, 430(1): 483–503, January 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rodgers:1950:TMR

[RMR^{+}50] Douglas Rodgers, Roger Muir, Eleanor Roosevelt, Albert Einstein, David Eli Lilienthal, J. Robert Oppenheimer, Brien McMahon, Hans A. (Hans Albrecht) Bethe, Harry Winne, Allan Kline, and Detlev W. (Detlev Wulf) Bronk. Today with Mrs. Roosevelt. NBC Television Network, United States, February 12, 1950. LCCN FSA 1920 (arch neg); FSA 1922 (mag trk); FCB 4264 (ref print); FTA 3029 (pos trk); FTA 3028 (masterpos); FCB 4367 (ref print, copy 2); FSA 1921 (neg trk). Seven tape reels.

Read:1963:EFS

[ROB63] A. L. Read, J. Orear, and H. A. Bethe. Exact form for scattering amplitude in Regge pole theory. Il Nuovo Cimento (10), 29(??): 1051–1058, ???? 1963. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (electronic).

Rojo:2007:SGO

[Roj07] Oscar Rojo. The spectra of a graph obtained from copies of a generalized Bethe tree. Linear Algebra and Its Applications, 420 (2–3):490–507, January 15, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rojo:2008:SWG

[Roj08] Oscar Rojo. Spectra of weighted generalized Bethe trees joined at the root. Linear Algebra and Its Applications, 428(11–12):2961– 2979, June 1, 2008. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rojo:2009:SCG

[Roj09] Oscar Rojo. Spectra of copies of a generalized Bethe tree attached to any graph. Linear Algebra and Its Applications, 431(5–7):863– 882, August 1, 2009. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rosner:1966:EMP

[Ros66] Jonathan L. Rosner. Exchange of massive particles in the Bethe– Salpeter equation. Journal of Mathematical Physics, 7(5):875–884, May 1966. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v7/i5/p875_s1.

Rossing:1993:AAP

[Ros93] Thomas D. Rossing. American Association of Physics Teachers 1993 Oersted Medalist: Hans Bethe. American Journal of Physics, 61(11):971–??, November 1993. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://link.aip.org/

link/ajpias/v61/i11/p971/s1.

Rojo:2007:EFE

[RR07] Oscar Rojo and María Robbiano. An explicit formula for eigenvalues of Bethe trees and upper bounds on the largest eigenvalue of any tree. Linear Algebra and Its Applications, 427(1):138–150, November 1, 2007. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rees:1982:SSC

[RS82] Martin J. Rees and Ray J. Stoneham, editors. Supernovae: A survey of current research. Proceedings of the NATO Advanced Study Institute held at Cambridge. 29 June–10 July, 1981, volume 90 of NATO advanced study institutes series C: mathematical and physical sciences. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1982. ISBN 90-277-1442-8. LCCN QB843.S95 N37 1981.

Ray:2005:BAS

[RS05] Subhankar Ray and J. Shamanna. A Bethe Ansatz study of free energy and excitation spectrum for even spin Fateev–Zamolodchikov model. Journal of Mathematical Physics, 46(4):043301, April 2005. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v46/i4/p043301_s1.

Rigden:2008:BNG

[RS08] John S. Rigden and Roger H. Stuewer. Book notes [Gerald E. Brown and Chang-Hwan Lee, Editors, Hans Bethe and His Physics, (Singapore: World Scientific, 2006)]. Physics in Perspective (PIP), 10(1):123–125, March 2008. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s00016-007-0379-6.

Runnels:1967:PTB

[Run67] L. K. Runnels. Phase transition of a Bethe lattice gas of hard molecules. Journal of Mathematical Physics, 8(10):2081–2087, October 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v8/i10/p2081_s1.

Reines:1988:OSN

[RV88] F. Reines and J. Vandervelde. Observation of SN1987A by neutrino light. Physics Reports, 163(1–3):137–140, 1988. CO-DEN PRPLCM. ISSN 0370-1573 (print), 1873-6270 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0370157388900403.

Rand:1973:CPS

[RYBB73] R. E. Rand, M. R. Yearian, H. A. Bethe, and C. D. Buchanan. Comments on the present status of elastic and inelastic magnetic electron-deuteron scattering. Physical Review D (Particles and Fields), 8(??):3229–3232, November 1, 1973. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM).

Sakharov:198x:SDI

[SAB^{+}8x] Andrei Sakharov, General James A. Abrahamson, George Ball, George P. Shultz, Secretary of State, Hans A. Bethe, James R. Schlesinger, John Bardeen, John Kogut, President of MIT, Representative Bob Dornan, Senator George Aiken, Senator Malcolm Wallop, Senator Dan Quayle, several former Secretaries of Defense, Strategic Defense Initiative Office, and Zellman Warhaft. Strategic Defense Initiative. Found in USGPO Congressional Record document list; publication year and status unknown, 198x.

Saenger:1967:BSE

[Sae67] Robert M. Saenger. Bethe–Salpeter equation in momentum space. Journal of Mathematical Physics, 8(12):2366–2375, December 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v8/i12/p2366_s1.

Salpeter:1952:NRS

[Sal52] Edwin E. Salpeter. Nuclear reactions in stars without hydrogen. Astrophysical Journal, 115(??):326–328, ???? 1952. CO-DEN ASJOAB. ISSN 0004-637X (print), 1538-4357 (electronic).

URL http://adsabs.harvard.edu/abs/1952ApJ...115..326S.

Reprinted in the centennial edition Ap. J. 525C, 450 (1999) with a modern commentary by André Maeder; also excerpted, with commentary, in Lang, Kenneth R. & Owen Gingerich, eds., A Source Book in Astronomy & Astrophysics, 1900–1975 Harvard Univ. Press, Cambridge, MA, 1979, 349–352.

Salpeter:2002:GLB

[Sal02] Edwin E. Salpeter. A generalist looks back. Annual Review of Astronomy and Astrophysics, 40(1):1–25, September 2002. CODEN ARAAAJ. ISSN 0066-4146 (print), 1545-4282 (electronic).

Salpeter:2005:HBS

[Sal05] Edwin E. Salpeter. Hans Bethe and societal issues. American Physical Society Forum on Physics and Society Newsletter, April

2005. URL http://www.aps.org/units/fps/newsletters/
2005/april/articles.html#salpeter.

Salpeter:2006:MSY

[Sal06] Edwin E. Salpeter. My sixty years with Hans Bethe. In Brown and Lee [BL06a], pages 115–123. ISBN 981-256-609-0, 981-256-610-4 (paperback), 981-277-450-5 (e-book). LCCN QC16.B46 H36 2006.

URL http://www.worldscibooks.com/physics/5989.html;
http://www.worldscientific.com/worldscibooks/10.1142/5989.

Sartori:2001:BRS

[Sar01] Leo Sartori. Book review: S. S. Schweber, In the Shadow of the Bomb: Oppenheimer, Bethe, and the Moral Responsibility of the Scientist. American Journal of Physics, 69(3):394–??, March 2001. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

URL http://link.aip.org/link/ajpias/v69/i3/p394/s2.

Sato:1963:NCB

[Sat63] Iwao Sato. Normalization condition for the Bethe–Salpeter wavefunction and a formal solution to the Bethe–Salpeter equation. Journal of Mathematical Physics, 4(1):24–35, January 1963. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/
i1/p24_s1.

Sauberman:1975:LCL

[Sau75] Nat. H. Sauberman. Letters: Credibility lost. Bulletin of the Atomic Scientists, 31(6):5, June 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [BAA^{+}75].

Sazdjian:1987:CTP

[Saz87] H. Sazdjian. The connection of two-particle relativistic quantum mechanics with the Bethe–Salpeter equation. Journal of Mathematical Physics, 28(11):2618–2638, November 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v28/i11/
p2618_s1.

Sommerfeld:1933:EMG

[SB33] Arnold Sommerfeld and Hans Bethe. Elektronentheorie der Metalle. (German) [Electron theory of metals]. In Aufbau Der Zusammenhängenden Materie. (German) [Structure of Connected Matter], Handbuch der Physik, pages 333–622. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc.,

1933. URL http://link.springer.com/chapter/10.1007/978-

3-642-91116-3_3. See also English translation [SB53].

Sommerfeld:1934:EMGc

[SB34] Arnold Sommerfeld and Hans Bethe. Elektronentheorie der Metalle. (German) [Electron theory of metals]. In Handbuch der Physik, volume 24, page ?? Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1934.

Seitz:1947:HCD

[SB47] Frederick Seitz and Hans A. Bethe. How close is the danger? In Masters and Way [MW47], pages 92–101. LCCN UF767 .M3 1947.

Salpeter:1951:REB

[SB51] E. E. Salpeter and H. A. Bethe. A relativistic equation for boundstate problems. Physical Review (2), 84(6):1232–1242, December 15, 1951. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://prola.aps.org/abstract/PR/
v84/i6/p1232_1.

Sommerfeld:1953:ETM

[SB53] Arnold Sommerfeld and Hans A. Bethe. Electron theory of metals. II. Electrons in the periodic potential field. Report AEC-TR-2055, United States Atomic Energy Commission, Technical Information Service, Oak Ridge, TN, USA, October 15, 1953. English translation of [SB33].

Shepley:1954:HBM

[SB54] James R. Shepley and Clay Blair, Jr. The hydrogen bomb: the men, the menace, the mechanism. Jarrolds, London, UK, 1954. 216 pp. LCCN UF767 .S4 1955. This book is seriously flawed; see the lengthy now-declassified rebuttal in [Bet82b], as well as the comments in a short review [All56], and a longer paper [Tel55].

Shepley:1955:WKB

[SB55a] James R. Shepley and Clay Blair, Jr. Die Wasserstoffbombe: der Konflikt, die Bedrohung, die Konstruktion. (German) [The hydrogen bomb: the conflict, the threat construction]. Steingrüben Verl., Stuttgart, Germany, 1955. 288 pp. LCCN ???? Translation of [SB54] to German by Hans Dieter Müller. This book is seriously flawed; see the lengthy now-declassified rebuttal in [Bet82b], as well as the comments in a short review [All56].

Shepley:1955:GSA

[SB55b] James R. Shepley and Clay Blair, Jr. La guerre secrète autour de la superbombe: la conception, la compétition, le risque mondial. (French) [The hydrogen bomb: conception, competition, and world risk]. Plon, Paris, France, 1955. x + 272 pp. LCCN ???? Translation of [SB54] to French by Anne Frauger. Preface by Yves Rocard. This book is seriously flawed; see the lengthy now-declassified rebuttal in [Bet82b], as well as the comments in a short review [All56].

Suh:1959:RMD

[SB59] K. S. Suh and H. A. Bethe. Recoil momentum distribution in electron pair production. Physical Review, 115(3):672–677, August 1, 1959. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065

(electronic). URL http://prola.aps.org/abstract/PR/v115/
i3/p672_1.

Schumacher:1961:UPT

[SB61] Clifford R. Schumacher and Hans A. Bethe. Usefulness of polarized targets and the polarization transfer tensor in reconstruction of the nucleon-nucleon scattering matrix. Physical Review (2), 121 (5):1534–1541, March 1, 1961. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.aps.

org/abstract/PR/v121/i5/p1534_1.

Siemens:1967:SHN

[SB67a] Philip J. Siemens and H. A. Bethe. Shape of heavy nuclei. Physical Review Letters, 18(17):704–706, April 24, 1967. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Sommerfeld:1967:MGG

[SB67b] Arnold Sommerfeld and Hans Bethe. Das Metall als Ganzes. (German) [The metal as a whole]. In Elektronentheorie der Metalle.

(German) [Electron theory of metals], page ?? Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc.,

1967. URL http://link.springer.com/chapter/10.1007/978-
3-642-95002-5_3.

Sommerfeld:1967:HFE

[SB67c] Arnold Sommerfeld and Hans Bethe. Die Hypothese der freien Elektronen. (German) [The hypothesis of free electrons]. In Elektronentheorie der Metalle. (German) [Electron theory of metals], page ?? Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1967. URL http://link.springer.

com/chapter/10.1007/978-3-642-95002-5_1.

Sommerfeld:1967:EIP

[SB67d] Arnold Sommerfeld and Hans Bethe. Elektronen im periodischen Potentialfeld. (German) [Electrons in the periodic potential field]. In Elektronentheorie der Metalle. (German) [Electron theory of metals], Heidelberger Taschenbücher, pages 36–253. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc.,

1967. URL http://link.springer.com/chapter/10.1007/978-
3-642-95002-5_2.

Sommerfeld:1967:EMG

[SB67e] Arnold Sommerfeld and Hans A. (Hans Albrecht) Bethe. Elektronentheorie der Metalle. (German) [Electron theory of metals], volume 19 of Heidelberger Taschenbücher. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1967. vii + 290 pp. LCCN QD171 .B45.

Shepley:1971:HBM

[SB71] James R. Shepley and Clay Blair, Jr. The hydrogen bomb: the men, the menace, the mechanism. Jarrolds, London, UK, 1971. ISBN 0-8371-5235-6. vii + 244 pp. LCCN UF767 .S4 1955. This book is seriously flawed; see the lengthy now-declassified rebuttal in [Bet82b], as well as the comments in a short review [All56].

Schmidt:1976:FNP

[SB76] Fred Henry Schmidt and David Bodansky. The fight over nuclear power: the energy controversy. Albion, San Francisco, CA, USA, 1976. ISBN 0-87843-002-4. xiii + 154 pp. LCCN ???? Foreword by Hans A. Bethe.

Seitz:1986:CTA

[SB86] F. Seitz and H. A. Bethe. A Chernobyl-type accident can’t happen here. New York Times, page 27, May 17, 1986. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Shukla:1997:NCB

[SBB^{+}97] P. K. Shukla, R. Bingham, H. A. Bethe, J. M. Dawson, and L. Stenflo. Nonlinear coupling between intense neutrino fluxes and dense magnetoplasmas. Physica Scripta, 55(1):96–98, ???? 1997. CO-DEN PHSTBO. ISSN 0031-8949 (print), 1402-4896 (electronic).

Su:1994:NNP

[SBDB94] J. J. Su, R. Bingham, J. M. Dawson, and H. A. Bethe. Nonlinear neutrino plasmas interactions. Physica Scripta, T52(??):132–134, ???? 1994. CODEN PHSTBO. ISSN 0281-1847.

Su:1995:NNP

[SBDB95] J. J. Su, R. Bingham, J. M. Dawson, and H. A. Bethe. Nonlinear neutrino plasmas interactions. Physica Scripta, T52(??):132–134, ???? 1995. CODEN PHSTBO. ISSN 0031-8949 (print), 1402-4896 (electronic).

Schweber:1955:MFF

[SBdH55] S. S. (Silvan S.) Schweber, Hans A. (Hans Albrecht) Bethe, and F. (Frederic) de Hoffmann. Mesons and fields. I. Fields, volume 1. Row, Peterson, Evanston, IL, USA, 1955. xvi + 449 pp. LCCN QC173 .S288.

Bethe:1956:MF

[SBdH56] Silvan S. Schweber, Hans A. Bethe, and Frederic de Hoffmann. Mesons and fields. Row, Peterson, Evanston, IL, USA, 1956. xvi + 449 pp. LCCN QC173 .S288.

Seitz:1986:RCA

[SBR^{+}86] Frederick Seitz, Hans Bethe, Dixy Lee Ray, Miro M. Todorovich, Robert K. Adair, Bernard L. Cohen, Thomas J. Connolly, Herbert Goldstein, Herbert Kouts, Leon Lidofski, John McCarthy, Eugene Wigner, and Edwin L. Zebroski. Review of the Chernobyl accident, and comparison of Soviet with U.S. reactors, by a Select Panel of Scientists and Engineers for Secure Energy (SE_{2} ). Report, Select Panel for Post-Chernobyl Safety Review, 570 Seventh Avenue, Room 1007, New York, NY 10018, USA, August 29, 1986. 64–71

pp. URL http://www.energytruth.com/Strike_Back_Or_Die.
pdf.

Szilard:1960:SWSb

[SBSW60] Leo Szilard, Hans A. (Hans Albrecht) Bethe, Lewis L. Strauss, and Jerome B. (Jerome Bert) Wiesner, editors. Small world: Szilard, Bethe, Strauss, Wiesner [Motion picture]. Columbia Broadcasting System, ????, 1960. LCCN ???? 28 minute 16mm film.

Schachenmeier:1932:WVT

[Sch32] R. Schachenmeier. Wellenmechanische Vorstudien zu einer Theorie der Supraleitung. (German) [Wave-mechanical investigation toward a theory of superconductivity]. Zeitschrift für Physik, 74(7– 8):503–546, July 1932. CODEN ZEPYAA. ISSN ???? URL http:/

/www.springerlink.com/content/g83170r826861868/. See also

[Bet34c, Sch34].

Schachenmeier:1934:TSE

[Sch34] R. Schachenmeier. Zur Theorie der Supraleitung. Entgegnung an Herru Bethe. (German) [On the theory of superconductivity. Reply to Mr. Bethe]. Zeitschrift für Physik, 90(9–10):680–692, September 1934. CODEN ZEPYAA. ISSN ???? URL http://www.

springerlink.com/content/mv32p632q1gh37p7/. See [Bet34c].

Schwinger:1958:SPQ

[Sch58] Julian Schwinger, editor. Selected Papers on Quantum Electrodynamics. Dover books on engineering and engineering physics. Dover, New York, NY, USA, 1958. ISBN 0-486-60444-6. xvii + 424 pp. LCCN QC680 .S35.

Schulze:1967:SHB

[Sch67a] G. E. R. Schulze. A. Sommerfeld und H. Bethe: Elektronentheorie der Metalle. Springer-Verlag, Berlin-Heidelberg-New York 1967. 290 Seiten mit 60 Abbildungen, Format 13,5 × 20,5 cm. Preis: DM 10,80. (German) [A. Sommerfeld and H. Bethe, Electron theory of metals, etc.]. Berichte der Bunsen-Gesellschaft für Physikalische Chemie, 71(6), August 1967. CODEN BBPCAX. ISSN 0005-9021.

Schulze:1967:BAS

[Sch67b] G. E. R. Schulze. Buchbesprechung: Arnold Sommerfeld und Hans Bethe: Elektronentheorie der Metalle. Springer-Verlag, Berlin-Heidelberg-New York 1967. 290 Seiten mit 60 Abbildungen, Format 13,5 × 20,5 cm. Preis: DM 10,80. (German) [Book review: Arnold Sommerfeld and Hans Bethe, Electron theory of metals, etc.]. Berichte der Bunsen-Gesellschaft für Physikalische Chemie, 71(6):641, August 1967. CODEN BBPCAX. ISSN 0005-

9021. URL http://onlinelibrary.wiley.com/doi/10.1002/
bbpc.19670710625/full.

Schmitt:1968:BRS

[Sch68] C. B. Schmitt. Book review: Sixteenth and Seventeenth Centuries: A Catalogue of Sixteenth-Century Printed Books in the National Library of Medicine. Compiled by Richard J. Durling. Bethesda, Maryland. U.S. Department of Health, Education, and Welfare. Public Health Service, National Library of Medicine. 1967. Pp. xii + 698. $5.00. British Journal for the History of Science, 4(2): 187–188, December 1968. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL http://www.jstor.org/

stable/4025160.

Schmidt:1978:BRH

[Sch78] P. C. Schmidt. Book review: H. A. Bethe and E. Salpeter: Quantum Mechanics of One- and Two-Electron atoms. Plenum/Rosetta, New York 1977. 370 Seiten, Preis: US$8.95. Berichte der Bunsen-Gesellschaft für Physikalische Chemie, 82(7), July 1978. CODEN BBPCAX. ISSN 0005-9021.

Schweber:1995:CBP

[Sch95] Silvan S. Schweber. Crossing boundaries: Perspectives from the life of H. A. Bethe. In Weingart [Wei95], pages 54–84. ISBN 3-7890-4009-6. LCCN Q175.55 .G74 1995.

Schweber:1996:WBL

[Sch96] Sylvan S. Schweber. Writing the biography of a living scientist: Hans Bethe. In Krishnamurthy et al. [KMN^{+}96], pages 159–196. ISBN 0-9629082-2-3. LCCN QD22.P35 C651 1995. URL http://

oregonstate.edu/Dept/Special_Collections/subpages/ahp/
1995symposium/schweber.html;   http://osulibrary.oregonstate.
edu/specialcollections/events/1995paulingconference/video-
s4-1-schweber.html.

Schwarzschild:2000:BHS

[Sch00a] Bertram Schwarzschild. Bohr–Heisenberg Symposium marks Broadway opening of Copenhagen. Physics Today, 53(5):51–52, May 2000. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://link.aip.org/link/phtoad/v53/i5/
p51/s1.

Schweber:2000:SBB

[Sch00b] S. S. (Silvan S.) Schweber. In the shadow of the bomb: Bethe, Oppenheimer, and the moral responsibility of the scientist. Princeton series in physics. Princeton University Press, Princeton, NJ, USA, 2000. ISBN 0-691-04989-0 (hardcover). xviii + 260 + 8 pp. LCCN QC774.O56 S32 2000.

Schwinger:2003:SPQ

[Sch03] Julian Schwinger. Selected Papers on Quantum Electrodynamics. Dover, New York, NY, USA, 2003. ISBN 0-486-60444-6. xvii + 424 pp. LCCN QC680 .S35. Reprint of [Sch58] with ISBN.

Schweber:2005:HBS

[Sch05a] Silvan S. Schweber. Hans Bethe: Superlative scientist, incomparable mentor, humanitarian, friend. Physics Today, 58(5):12–13, May 2005. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Schweber:2005:HT

[Sch05b] Silvan S. Schweber. “The happy thirties”. Physics Today, 58(10): 38–43, October 2005. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Schweber:2006:HT

[Sch06] Silvan S. Schweber. “The Happy Thirties”. In Brown and Lee [BL06a], pages 131–145. ISBN 981-256-609-0, 981-256-610-4 (paperback), 981-277-450-5 (e-book). LCCN QC16.B46 H36 2006.

URL http://www.worldscibooks.com/physics/5989.html;
http://www.worldscientific.com/worldscibooks/10.1142/5989.

Schweber:2008:HB

[Sch08] Silvan Schweber. Hans Bethe. In Noretta Koertge, editor, The New Dictionary of Scientific Biography, pages 269–276. Scribner’s, New York, NY, USA, 2008. ISBN 0-684-31321-9. LCCN Q141 .N45 2008.

Schweber:2012:NFM

[Sch12] S. S. (Silvan S.) Schweber. Nuclear forces: the making of the physicist Hans Bethe. Harvard University Press, Cambridge, MA, USA, 2012. ISBN 0-674-06587-5 (hardcover). viii + 579 + 12 pp. LCCN QC774.B4 S39 2012. £25.95.

Schaaf:2014:WBN

[Sch14a] Michael Schaaf. Weizsäcker, Bethe und der Nobelpreis. (German) [Weizsäcker, Bethe and the Nobel Prize]. In Hentschel and Hoffmann [HH14], pages 145–158. ISBN 3-8047-3244-5. ISSN 0001-5857. LCCN ????

Schweber:2014:WBH

[Sch14b] Silvan S. Schweber. Writing the biography of Hans Bethe: Contextual history and Paul Forman. Physics in Perspective (PIP), 16(2):179–217, June 2014. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.

com/article/10.1007/s00016-014-0136-6.

Schutz:2015:BWM

[Sch15] Gunter M. Schütz. The Bethe Wavefunction, by Michel Gaudin, Scope: monograph. Level: postgraduate, early career researcher, researcher, scientist, specialist. Contemporary Physics, 56(2):239– 240, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Scott:1994:CFB

[Sco94] D. R. D. Scott. Classical functional Bethe ansatz for SL(N): Separation of variables for the magnetic chain. Journal of Mathematical Physics, 35(11):5831–5843, November 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v35/i11/p5831_s1.

Segre:1953:ENPb

[Seg53] Emilio Segrè, editor. Experimental Nuclear Physics, volume 2. Wiley, New York, NY, USA, 1953. viii + 600 pp. LCCN QC173 .S313 1953.

Segre:1953:ENPa

[Seg59a] Emilio Segrè, editor. Experimental Nuclear Physics, volume 1. Wiley, New York, NY, USA, 1953–1959. x + 789 pp. LCCN QC173 .S313 1953.

Segre:1959:ENP

[Seg59b] Emilio Segrè, editor. Experimental Nuclear Physics, volume 3. Wiley, New York, NY, USA, 1959. ix + 811 pp. LCCN QC173 .S313 1959.

Segre:1985:HPR

[Seg85] Emilio Segrè. Historical perspective: Refugee scientists and nuclear energy. Annals of the New York Academy of Sciences, 452(1):xv– xix, 1985. CODEN ANYAA9. ISBN 0-89766-298-9, 0-89766-299-7 (paperback). ISSN 0077-8923 (print), 1749-6632 (electronic). Sixth International Conference on Collective Phenomena: reports from the Moscow Refusnik Seminar / edited by Inga Fischer-Hjalmars and Joel L. Lebowitz. Contributions from the Moscow Refusnik Seminar and from two International Conferences on Collective Phenomena, one held in Stockholm, Sweden, 1–2 December 1983, and the other in Tel Aviv, Israel, 31 May–1 June 1984.

Seidel:2001:BRBb

[Sei01] Robert W. Seidel. Book review: In the Shadow of the Bomb: Oppenheimer, Bethe, and the Moral Responsibility of the Scientist by S. S. Schweber. Isis, 92(2):422–423, June 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/3080703.

Selbin:1977:RUT

[Sel77] Joel Selbin. Rebuttal: Unreal thinking about energy. Bulletin of the Atomic Scientists, 33(7):54–55, September 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bet77e, Bet77d].

Serneels:1973:DGB

[SG73] R. Serneels and R. Gevers. A discussion and generalization of Bethe’s second-order approximation. Physica Status Solidi. B, Basic Research, 55(1):335–346, January 1, 1973. CODEN PSSBBD. ISSN 0370-1972 (print), 1521-3951 (electronic).

Sherwin:1975:WDA

[She75] Martin J. Sherwin. A world destroyed: the atomic bomb and the Grand Alliance. Alfred A. Knopf, New York, NY, USA, 1975. ISBN 0-394-49794-5. xvi + 315 + xi pp. LCCN D753 .S48 1975. US$10.00.

Skrypnyk:2007:QIS

[Skr07] T. Skrypnyk. Quantum integrable systems, non-skew-symmetric r-matrices and algebraic Bethe ansatz. Journal of Mathematical Physics, 48(2):023506, February 2007. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:

//jmp.aip.org/resource/1/jmapaq/v48/i2/p023506_s1.

Skrypnyk:2009:NSS

[Skr09] T. Skrypnyk. Non-skew-symmetric classical r-matrices, algebraic Bethe ansatz, and Bardeen–Cooper–Schrieffer-type integrable systems. Journal of Mathematical Physics, 50(3):033504, March 2009. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v50/i3/p033504_s1.

Swift:1964:SBS

[SL64] Arthur R. Swift and Benjamin W. Lee. Singular Bethe–Salpeter scattering amplitudes. Journal of Mathematical Physics, 5(7):908– 921, July 1964. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v5/i7/p908_s1.

Smith:1975:LWN

[Smi75] Anthony Wayne Smith. Letters: We need restraint. Bulletin of the Atomic Scientists, 31(6):5, June 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [BAA^{+}75].

Shukla:1997:SSN

[SSB^{+}97a] P. K. Shukla, L. Stenflo, R. Bingham, H. A. Bethe, I. M. Dawson, and J. T. Mendonça. Stimulated scattering of neutrinos by electron-cyclotron-harmonic and convection modes in magnetoplasmas. Physics Letters A, 226(6):375–377, March 3, 1997. CODEN PYLAAG. ISSN 0375-9601 (print), 1873-2429

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0375960196009498.

Shukla:1997:NGR

[SSB^{+}97b] P. K. Shukla, L. Stenflo, R. Bingham, H. A. Bethe, J. M. Dawson, and J. T. Mendonça. Nonlinear generation of radiation by intense neutrino fluxes in super dense magnetized plasmas. Physics Letters A, 224(4–5):239–242, January 13, 1997. CODEN PYLAAG. ISSN 0375-9601 (print), 1873-2429

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S037596019600816X.

Shukla:1997:NPN

[SSB^{+}97c] P. K. Shukla, L. Stenflo, R. Bingham, H. A. Bethe, J. M. Dawson, and J. T. Mendonça. Nonlinear propagation of neutrinos in a strongly magnetized medium. Physics Letters A, 230(5–6):353– 357, June 23, 1997. CODEN PYLAAG. ISSN 0375-9601 (print),

1873-2429 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0375960197002764.

Shukla:1997:GMF

[SSB^{+}97d] P. K. Shukla, L. Stenflo, R. Binghani, H. A. Bethe, J. M. Dawson, and J. T. Mendonça. Generation of magnetic fields by nonuniform neutrino beams. Physics Letters A, 233(3):181–183, August 25, 1997. CODEN PYLAAG. ISSN 0375-9601 (print),

1873-2429 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0375960197004507.

Shukla:1998:NEA

[SSB^{+}98] P. K. Shukla, L. Stenflo, R. Bingham, H. A. Bethe, J. M. Dawson, and J. T. Mendonça. The neutrino electron accelerator. Physics of Plasmas, 5(1):1–3, 1998. CODEN PHPAEN. ISSN 1070-664X (print), 1089-7674 (electronic), 1527-2419.

Shukla:1999:PCN

[SSB^{+}99a] P. K. Shukla, L. O. Silva, H. A. Bethe, R. Bingham, J. M. Dawson, L. Stenflo, J. T. Mendonça, and S. Daihed. The physics of collective neutrino plasma interactions. Plasma Physics and Controlled Fusion, 41(3A):A699–A707, March 1999. CODEN PPCFET. ISSN 0741-3335 (print), 1361-6587 (electronic).

Shukla:1999:NBB

[SSB^{+}99b] P. K. Shukla, L. Stenflo, H. A. Bethe, R. Bingham, J. M. Dawson, L. O. Silva, and J. T. Mendonça. Neutrinos as building blocks of the universe. In J. Buechner, Sir Ian Axford, E. Marsch, and V. Vasyliunas, editors, Plasma Astrophysics and Space Physics, pages 731–737. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1999.

Sobel:1975:SWC

[SSBB75] M. I. Sobel, P. J. Siemens, J. P. Bondorf, and H. A. Bethe. Shock waves in colliding nuclei. Nuclear Physics A, 251(3):502–529, October 27, 1975. CODEN NUPABL. ISSN 0375-9474 (print),

1873-1554 (electronic). URL http://www.sciencedirect.com/
science/article/pii/037594747590545X.

Shapiro:1986:HMA

[ST86] Stuart L. (Stuart Louis) Shapiro and Saul A. (Saul Arno) Teukolsky, editors. Highlights of modern astrophysics: concepts and controversies. Wiley, New York, NY, USA, 1986. ISBN 0-471-82421-6. LCCN QB461 .H54 1986. US$35.95 (est.). Derived from a symposium held at Cornell University in October 1984, in honor of E. S. Salpeter.

Strauss:1963:MDa

[Str63] Lewis L. Strauss. Men and decisions. Doubleday, Garden City, NY, USA, 1963. viii + 468 + 16 pp. LCCN E741 .S78.

Szilard:1949:DSB

[STS^{+}49] Leo Szilard, The Franck Committee, Frederick Seitz, Hans Bethe, Harold C. Urey, Harrison Brown, Irving Langmuir, Vannevar Bush, and Leslie R. Groves. Did the Soviet bomb come sooner than expected? The atomic scientists: The wartime administrators. Bulletin of the Atomic Scientists, 5(10):262–264, October 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Stuewer:1979:NPR

[Stu79] Roger H. Stuewer, editor. Nuclear physics in retrospect: proceedings of a symposium on the 1930s. University of Minnesota Press, Minneapolis, MN, USA, 1979. ISBN 0-8166-0869-5. LCCN QC773 .S95 1977.

Stuewer:2018:AIN

[Stu18] Roger H. Stuewer. The Age of Innocence: Nuclear Physics Between the First and Second World Wars. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2018. ISBN 0-19-186658-X, 0-19-882787-3 (hardback), 0-19-256290-8 (e-book). xv + 484 pp. LCCN QC773 .S78 2018.

Sullivan:1975:EDV

[Sul75] Walter Sullivan. Experts doubt view that atom blast could end all life. New York Times, ??(??):??, November 23, 1975. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL http://search.proquest.com/hnpnewyorktimes/

docview/120557618/. This story comments on the recentlydeclassified report [Bre53] on whether a hydrogen bomb explosion could ignite the atmosphere or oceans, or blow the atmosphere into space. See also [Dud75].

Sutherland:2003:BAN

[Sut03] Bill Sutherland. Bethe’s Ansatz: now and then. In Goldhaber et al. [GSS^{+}03], pages 125–141. ISBN 981-238-503-7 (hardcover), 981-238-563-0 (paperback), 981-279-508-1 (e-book). LCCN

QC174.17.S9 S994 2003. URL http://www.worldscientific.
com/doi/pdf/10.1142/9789812795083_0009.

Teller:1982:HBH

[TB82] Edward Teller and Hans A. Bethe. Hydrogen bomb history. Science, 218(4579):1270, December 24, 1982. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

jstor.org/stable/1690175. Response to [Bro82].

Teller:1984:FND

[TB84] Edward Teller and Hans Bethe. Face-off on nuclear defense. Technology Review (M.I.T.), 84(3):48–49, April 1984. CODEN TEREAU. ISSN 0040-1692. Excerpts from a debate over Star Wars weaponry in November 1983 at the Kennedy School of Government at Harvard.

Tsytovich:1998:CNP

[TBDB98] V. N. Tsytovich, R. Bingham, J. M. Dawson, and H. A. Bethe. Collective neutrino plasma interactions. Astroparticle Physics, 8(4): 297–307, ???? 1998. CODEN APHYEE. ISSN 0927-6505 (print), 1873-2852 (electronic).

Teller:1955:WMP

[Tel55] Edward Teller. The work of many people. Science, 121(3139): 267–275, February 25, 1955. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.jstor.org/

stable/1681529.

Teller:1966:TQ

[Tel66] E. Teller. On a theory of quasars. In Marshak and Blaker [MB66], pages 449–462. LCCN ????

Teller:1987:SND

[Tel87] Edward Teller. Scientists and national defense, an open letter to Hans Bethe. Policy Review, 39(??):18–23, Winter 1987. CODEN ???? ISSN 1080-6857.

Thouless:1999:BRT

[Tho99] David Thouless, F.R.S. Book review: Two pioneers of modern theoretical physics: H. A. Bethe, Selected works of Hans A. Bethe, with commentary. World Scientific Publishing Company, 1997. Pp. viii + 605, £47.00 (hardback). ISBN 981-02-2876-7. R. H. Dalitz and Sir Rudolf Peierls, Selected scientific papers of Sir Rudolf Peierls, with commentary. World Scientific Publishing Company and Imperial College Press, 1997. Pp. xxiii + 805, £67.00 (hardback). ISBN 981-02-2692-6. Notes and Records of the Royal Society of London, 53(2):282–288, May 22, 1999. CODEN NOREAY. ISSN 0035-9149 (print), 1743-0178 (electronic). URL http://www.

jstor.org/stable/532214.

Tomozawa:1983:NTB

[Tom83] Yukio Tomozawa. Normalization of the three-body Bethe–Salpeter wavefunction for protons. Journal of Mathematical Physics, 24 (2):369–374, February 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v24/i2/p369_s1.

Tucker:2005:OHB

[Tuc05] Anthony Tucker. [obituary:] Hans Bethe. The Guardian, ?? (??):??, March 8, 2005. ISSN 0261-3077 (print), 1756-3224 (elec-

tronic). URL https://www.theguardian.com/news/2005/mar/
08/guardianobituaries.nuclear.

Turner:2008:PCA

[Tur08] Michael S. Turner. From αβγ to precision cosmology: The amazing legacy of a wrong paper. Physics Today, 61(12):8–9, December 2008. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://www.physicstoday.org/resource/1/

PHTOAD/v61/i12. This is a retrospective of influence the famous paper [ABG48] on the origin of the chemical elements, and the Big Bang theory of the evolution of the Universe. See also comments [AWCT09]. The correct theory of the origin of the elements appears in [BBFH57].

Ulam:1966:TD

[Ula66] Stanisław M. Ulam. Thermonuclear devices. In Marshak and Blaker [MB66], pages 593–601. LCCN ????

Unser:2000:WPM

[Uns00] Margit Unser, editor. Wolfgang Pauli and modern physics = Wolfgang Pauli und die moderne Physik. ETH-Bibliothek, Zürich, Switzerland, 2000. ISBN 3-9521386-2-2. 124 pp. LCCN QC16.P37 W635 2000.

Usdin:2009:RRR

[Usd09] Steven T. Usdin. The Rosenberg ring revealed: Industrialscale conventional and nuclear espionage. Journal of Cold War Studies, 11(3):91–143, Summer 2009. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.

jhu.edu/article/268392;   http://muse.jhu.edu/article/
268392/pdf.

VanVleck:1932:TVP

[Van32] John H. Van Vleck. Theory of the variations in paramagnetic anisotropy among different salts of the iron group. Physical Review (2), 41(2):208, July 1932. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://link.aps.org/doi/

10.1103/PhysRev.41.208;   http://prola.aps.org/abstract/

PR/v41/i2/p208_1. This paper applies the Crystal Field Theory of earlier work [Bet29b, Bec29] to chemistry.

VanDeMark:2003:PKN

[Van03] Brian VanDeMark. Pandora’s keepers: nine men and the atomic bomb. Little, Brown and Co., Boston, MA, USA, 2003. ISBN 0-316-73833-6. xii + 399 + 16 pp. LCCN QC774.A2 V36 2003. URL

http://www.loc.gov/catdir/toc/fy045/2002043646.html.

vonBaeyer:1992:BRH

[vB92] Hans C. von Baeyer. Book review: Hans A. Bethe, The Road from Los Alamos. American Journal of Physics, 60(1):91–??, January 1992. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

vonderLage:1947:MOE

[vdLB47] Fred C. von der Lage and H. A. Bethe. A method for obtaining electronic eigenfunctions and eigenvalues in solids with an application to sodium. Physical Review, 71(9):612–622, May 1, 1947. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://prola.aps.org/abstract/PR/v71/i9/p612_1.

vonHippel:1976:LTC

[vHWD76] Frank von Hippel, Robert H. Williams, and H. C. Dudley. Letter: Taxes credulity. Bulletin of the Atomic Scientists, 32(1):2–3, January 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Dud75] and comment [Nov76].

Vosko:1960:ASB

[Vos60] S. H. Vosko. Approximate solutions of the Bethe–Salpeter equation. Journal of Mathematical Physics, 1(6):505–515, June 1960. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v1/i6/p505_s1.

vonWeizsacker:1935:TKG

[vW35] C. F. von Weizsäcker. Zur Theorie der Kernmassen. (German) [On the theory of nuclear mass]. Zeitschrift für Physik, 96(??):431–458, ???? 1935. CODEN ZEPYAA. ISSN 0044-3328.

vonWeizsacker:1938:EISa

[vW38a] Carl Friedrich von Weizsäcker. Über Elementumwandlungen in Innern der Sterne. (German) [On element transformations in stellar interiors]. Physikalische Zeitschrift, 38(??):176–191, ???? 1938. CODEN PHZTAO. ISSN 0369-982X. URL http:/

/hdl.handle.net/2027/mdp.39015076063125;   http://www.
nobelprize.org/nobel_prizes/physics/articles/fusion/.

vonWeizsacker:1938:EISb

[vW38b] Carl Friedrich von Weizsäcker. Über Elementumwandlungen in Innern der Sterne. II. (German) [On element transformations in stellar interiors. II]. Physikalische Zeitschrift, 39(17–18):633–646, ???? 1938. CODEN PHZTAO. ISSN 0369-982X. URL http:/

/hdl.handle.net/2027/mdp.39015013742401;   http://www.
nobelprize.org/nobel_prizes/physics/articles/fusion/.

vonWeizsacker:1988:BGC

[vW88] Carl Friedrich von Weizsäcker. Bewußtseinswandel. (German) [Change of consciousness]. C. Hanser, Munich, West Germany, second edition, 1988. ISBN 3-446-14649-0. 476 pp. LCCN ????

vonWeizsacker:1991:BGC

[vW91] Carl Friedrich von Weizsäcker. Bewußtseinswandel. (German) [Change of consciousness]. Deutscher Taschenbuch Verlag, Munich, West Germany, 1991. ISBN 3-423-11388-X. 476 pp. LCCN ????

Weigel:1994:BSE

[WA94] H. Weigel and R. Alkofer. The Bethe–Salpeter equation for mesons as quark — anti-quark bound states in a soliton background. Computer Physics Communications, 82(1):57–73, August 1994. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0010465594901317.

Walker:2015:BRB

[Wal15a] Mark Walker. Book review: Émigré Biographies: Anja Skaar Jacobsen, Léon Rosenfeld: Physics, Philosophy, and Politics in the Twentieth Century. Hackinsack: World Scientific, 2012. xii + 354 pp., illus., index. ISBN 978-981-4307-81-9. $87.00 (hardcover).; Christoph Laucht, Elemental Germans: Klaus Fuchs, Rudolf Peierls and the Making of British Nuclear Culture 1939– 1959. London: Palgrave Macmillan, 2012. xiv + 274 pp., index. ISBN 978-0-230-35487-6. $75.00 (hardcover).; Mary Jo Nye, Michael Polanyi and His Generation: Origins of the Social Construction of Science. Chicago: University of Chicago Press, 2011. xi + 405 pp., illus., index. ISBN 978-0-226-61063-4. $50.00 (hardcover).; Silvan S. Schweber, Nuclear Forces: The Making of the Physicist Hans Bethe. Cambridge, MA: Harvard University Press, 2012. viii + 579 pp., illus., index. ISBN 978-0-674-06587-1. $35.00 (hardcover). Historical Studies in the Natural Sciences, 45(1): 206–216, February 2015. CODEN ???? ISSN 1939-1811 (print),

1939-182X (electronic). URL http://www.jstor.org/stable/
10.1525/hsns.2015.45.1.206.

Walker:2015:BRE

[Wal15b] Mark Walker. Book review: Émigré Biographies: Anja Skaar Jacobsen, Léon Rosenfeld: Physics, Philosophy, and Politics in the Twentieth Century. Hackinsack: World Scientific, 2012. xii + 354 pp., illus., index. ISBN 978-981-4307-81-9. $87.00 (hardcover).; Christoph Laucht, Elemental Germans: Klaus Fuchs, Rudolf Peierls and the Making of British Nuclear Culture 1939– 1959. London: Palgrave Macmillan, 2012. xiv + 274 pp., index. ISBN 978-0-230-35487-6. $75.00 (hardcover).; Mary Jo Nye, Michael Polanyi and His Generation: Origins of the Social Construction of Science. Chicago: University of Chicago Press, 2011. xi + 405 pp., illus., index. ISBN 978-0-226-61063-4. $50.00 (hardcover).; Silvan S. Schweber, Nuclear Forces: The Making of the Physicist Hans Bethe. Cambridge, MA: Harvard University Press, 2012. viii + 579 pp., illus., index. ISBN 978-0-674-06587-1. $35.00 (hardcover). Historical Studies in the Natural Sciences, 45(1): 206–216, February 2015. CODEN ???? ISSN 1939-1811 (print),

1939-182X (electronic). URL http://www.jstor.org/stable/
10.1525/hsns.2015.45.1.206.

Walske:1951:AFS

[WB51] M. C. Walske and H. A. Bethe. Asymptotic formula for stopping power of K-electrons. Physical Review, 83(2):457–458, July 15, 1951. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://prola.aps.org/abstract/PR/
v83/i2/p457_2.

Weisskopf:1964:FAD

[WB64] V. F. Weisskopf and H. A. Bethe. Fermi Award defended. New York Times, page 28, March 23, 1964. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Weber:1973:RWS

[Web73] Robert L. Weber, editor. A Random Walk In Science. The Institute of Physics, Bristol and London, 1973. ISBN 0-85498-027-X. xvii + 206 pp. LCCN Q167.W42. Compiled by R. L. Weber, edited by E. Mendoza, with a foreword by William Cooper.

Weisskopf:1975:LCE

[Wei75] Victor Weisskopf. Letters: A clerical error. Bulletin of the Atomic Scientists, 31(4):3, April 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reports incorrect inclusion of Weisskopf’s name in [BAA^{+}75].

Weisskopf:1991:JIP

[Wei91] Victor Frederick Weisskopf. The Joy of Insight: Passions of a Physicist. Alfred P. Sloan Foundation series. Basic Books, New York, NY, USA, 1991. ISBN 0-465-03678-3, 0-465-03677-5. xiv + 336 + 8 pp. LCCN QC16.W516 A3 1991.

Weingart:1995:GWC

[Wei95] Peter Weingart, editor. Grenzüberschreitungen in der Wissenschaft = Crossing boundaries in science, volume 1 of Interdisziplinäre Studien = Interdisciplinary studies / ZIF. Nomos, Baden-Baden, Germany, 1995. ISBN 3-7890-4009-6. 218 pp. LCCN Q175.55 .G74 1995.

Watanabe:2011:NGP

[WF11] Y. Watanabe and K. Fukumizu. New graph polynomials from the Bethe approximation of the Ising Partition Function. Combinatorics, Probability and Computing, 20(2):299–320, March 2011. CODEN CPCOFG. ISSN 0963-5483 (print), 1469-2163 (electronic).

Wesley:1968:BWM

[WG68] James Paul Wesley and Alex E. S. Green. The Bethe–Weizsäcker mass formula and Lennard-Jones N-N potentials. American Journal of Physics, 36(12):1093–??, December 1968. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://

link.aip.org/link/ajpias/v36/i12/p1093/s1.

Wellerstein:2017:SSH

[WG17] Alex Wellerstein and Edward Geist. The secret of the Soviet hydrogen bomb. Physics Today, 70(4):40–47, April 2017. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Wiser:1991:APS

[WGO^{+}91] Christine Zurbach Wiser, Donald Goldsmith, David Oyster, Richard Chamberlain, and Hans A. (Hans Albrecht) Bethe. The astronomers. part 5: Stardust. WETA-TV, May 13, 1991. Two one–hour videocassettes.

Weiner:1972:EHN

[WH72] Charles Weiner and Elspeth Hart, editors. Exploring the history of nuclear physics: proceedings of the American Institute of Physics– American Academy of Arts and Sciences conferences on the history of nuclear physics, 1967 and 1969, volume 7 of AIP conference proceedings. American Institute of Physics, Woodbury, NY, USA, 1972. ISSN 0094-243X, 1551-7616, 1935-0465. LCCN QC173 .E88 1972.

Whitaker:1995:BBB

[Whi95] M. A. B. Whitaker. Bethe, Bell and de Broglie. American Journal of Physics, 63(10):873, October 1995. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://link.

aip.org/link/ajpias/v63/i10/p873/s1.

Wiesler:1981:DBH

[Wie81] Andreas Wiesler. Derivation of “Bethe’s hypothesis” from the quantum inverse scattering transformation for the nonlinear Schrödinger equation. Journal of Mathematical Physics, 22(4): 750–753, April 1981. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v22/i4/p750_s1.

Wiescher:2014:CFW

[Wie14] Michael Wiescher. Carl Friedrich von Weizsäcker und der Bethe– Weizsäcker-Zyklus. (German) [Carl Friedrich von Weizsäcker and the Bethe–Weizsäcker cycle]. In Hentschel and Hoffmann [HH14], pages 117–144. ISBN 3-8047-3244-5. ISSN 0001-5857. LCCN ????

Wilkins:1973:EBD

[Wil73] J. Ernest Wilkins, Jr. The eigenvalues of the Bethe differential system. SIAM Journal on Mathematical Analysis, 4(1):62–77, February 1973. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Weekes:1936:MDS

[WLB36] D. F. Weekes, M. Stanley Livingston, and H. A. Bethe. A method for the determination of the selective absorption regions of slow neutrons. Physical Review, 49(6):471–473, March 15, 1936. CO-DEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

URL http://prola.aps.org/abstract/PR/v49/i6/p471_1.

Wilson:1988:CCC

[WM88] James R. Wilson and Ronald W. Mayle. Convection in core collapse supernovae. Physics Reports, 163(1–3):63–78, 1988. CO-DEN PRPLCM. ISSN 0370-1573 (print), 1873-6270 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0370157388900361.

Ward:2015:RWF

[WM15] Tim Ward and Domenic Mastrippolito. Race for the world’s first atomic bomb: A thousand days of fear. [Los Alamos National] Laboratory produced documentary on the Manhat-

tan Project., 2015. URL http://www.governmentattic.org/
11docs/NNSAfilmVideoList_2014.pdf;   https://www.youtube.

com/watch?v=xwpgmEvlRpM. Written by Michael Wilson. The hour-long program features interviews with some of those who worked on the Project Hans Bethe, Norman Ramsey, Phil Morrison, Edward Teller and others. The program profiles the creation of the Manhattan Project the obstacles that were faced, life in the “secret city,’ and events that led up to and followed the Trinity Test.

Wu:1976:RSS

[WPBW76] C. S. Wu, W. K. H. Panofsky, H. Bethe, and V. F. Weisskopf. Reactor safety study. Physics Today, 29(7):15–??, July 1976. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

URL http://link.aip.org/link/phtoad/v29/i7/p15/s1.

Wright:2010:HBD

[Wri10] James Wright. Happy birthday Dr. Bethe. Web page, July

2, 2010. URL http://www.fas.org/blog/nutshell/2010/07/

happy-birthday-dr-bethe/. Posthumous greetings on Bethe’s 104th anniversary.

Woosley:1988:PMS

[WW88] S. E. Woosley and T. A. Weaver. Presupernova models: Sensitivity to convective algorithm and Coulomb corrections. Physics Reports, 163(1–3):79–94, 1988. CODEN PRPLCM. ISSN 0370-1573 (print),

1873-6270 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0370157388900373.

Yang:2013:SPI

[Yan13] Chen Ning Yang, editor. Selected papers II, with commentaries. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2013. ISBN 981-4449-00-8 (hardcover), 981-4449-01-6 (paperback), 981-4449-02-4 (e-book). x +

34 pp. LCCN QC21.3. URL http://www.worldscientific.com/
worldscibooks/10.1142/8640.

Yonas:1985:CSW

[YB85] G. Yonas and H. A. Bethe. Can Star Wars make us safe? Science Digest, 93(9):30–??, September 1985. CODEN ???? ISSN 0036-8296.

Yang:2006:BH

[YG06] Chen Ning Yang and M.-L. Ge. Bethe’s hypothesis. In Brown and Lee [BL06a], pages 185–187. ISBN 981-256-609-0, 981-256-610-4 (paperback), 981-277-450-5 (e-book). LCCN QC16.B46 H36 2006.

URL http://www.worldscibooks.com/physics/5989.html;
http://www.worldscientific.com/worldscibooks/10.1142/5989.

Ying:2002:EBA

[YLGC02] Zu-Jian Ying, You-Quan Li, Shi-Jian Gu, and Bin Chen. Extension of Bethe ansatz to multiple occupancies for one-dimensional SU(4) fermions with δ-function interaction. Journal of Mathematical Physics, 43(12):5977–5986, December 2002. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

York:1989:AOT

[Yor89] Herbert F. (Herbert Frank) York. The advisors: Oppenheimer, Teller, and the superbomb. Stanford nuclear age series. Stanford University Press, Stanford, CA, USA, 1989. ISBN 0-8047-1713-3, 0-8047-1714-1 (paperback). xiv + 201 pp. LCCN UG1282.A8 Y67

1989. URL http://www.loc.gov/catdir/description/cam023/
88062671.html;   http://www.loc.gov/catdir/toc/cam027/
88062671.html.

Yuille:2001:DLA

[Yui01] Alan Yuille. A double-loop algorithm to minimize the Bethe free energy. Lecture Notes in Computer Science, 2134:3–??, 2001. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer-ny.com/link/service/series/
0558/bibs/2134/21340003.htm;   http://link.springer-ny.
com/link/service/series/0558/papers/2134/21340003.pdf.

Zhou:2003:SMB

[ZLGM03] Huan-Qiang Zhou, Jon Links, Mark D. Gould, and Ross H. McKenzie. Solvable models of Bose–Einstein condensates: a new algebraic Bethe ansatz scheme. Journal of Mathematical Physics, 44(10):4690–4701, October 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

ZumWinkel:1975:ESW

[Zum75] D. Zum Winkel. On the exact solutions of the Wick-rotated fermion-antifermion Bethe–Salpeter equation. Journal of Mathematical Physics, 16(1):93–95, January 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v16/i1/p93_s1.